A Complete Bibliography of Publications in the Journal of Statistical Computation and Simulation: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

10 June 2022 Version 2.00

Title word cross-reference

(0, 1) [ZZXS17]. (m/n) [AMBP17]. 2^{K} [Lu19]. 2 × 2 [IS13]. 3 [SYL^{+}14]. $38.66 [Sha19f]. $39.95 [Sha19i]. ^{18} [ZB13]. A [DBC12]. AP(q) [TTWT19]. AR(p) [LM13]. B [DA16]. X̄ [GVTGH^{+}18, MQR18]. BINAR(1) [KSJ16]. C_{Npk} [RAK16b]. C_{pk} [AWJ^{+}13, Cha17, PWW15]. C_{pl} [HCY16]. C_{pm} [Per10]. C_{pmk} [Per10, WL10]. C_{pu} [HCY16]. D [DBC12, JP17, LJV^{+}18, LY15, MJ16, PT14, Özk19, XY11]. E [Han17, LS14]. F [MY13, NFFM12]. G [NCO12, NCO15, SP16, AY15, LZZ^{+}15]. GARCH(1, 1) [BV16, ZS16b]. gh [BT16b]. I [MY13]. INAR(1) [IRNB18]. J [JK14]. k [AB11, AMB12, KLK15, LXZ11, PS14, XY11, YP17, Yoo13]. l_{1} [FP15b, YZ14]. L_{2} [LS14]. M [NAA17, SRP11, AMBP17, ASS11]. N [AR16b, TGL12a, AMBP17, ASS11, BH19, CC19]. ν [ASBM19]. P [OO12, DYX15, KR13, Kel16, LXZ11, MP15, Moi17b, MMP12, SKX^{+}18].

P (X > Y) [Gen13]. P (Y < X) [KDG17, NKP14]. r [Özk19, XY11]. R = P (Y < X) [SKK12]. R = P [Y < X] [KR15b]. R^{2} [MH16]. ρ [SGZM14]. S [EE15, GB18, KL13b]. S_{pk} [HCY16]. T [ATON18, AO12, BCY16, Ho12, HY16, LK17, NMPR14, WCC13, WL15b, Yos18]. T^{2} [AO12, FKS10, FHSC14, NMS18, SFS15]. τ [SGZM14]. U [AvR15, CVL18]. V [HOR17]. φ [VK14]. X [Rav19, SWZ15]. X(Y) [JR13]. X^{−} [DH14, Gau10, NTG13]. Y [GGM18]. Y (X) [JR13]. z [LLB12].

-bar [Rav19]. -Bayesian [Han17]. -divergence [GB18]. -estimation [NAA17]. -estimators [EE15]. -generalized [KR13]. -link [NMPR14]. -linked [GGM18]. -location-scale [XHEM17]. -means [YP17]. -measures [VK14]. -mixture [TGL12a]. -nearest-neighbour [KLK15]. -optimal [DBC12, JP17, LJV^{+}18, LY15, MJ16, PT14, SP16]. -ordered [AR16b]. -out-of- [AMBP17, ASS11]. -pooled [LLB12]. -prior [AY15]. -process [JK14]. -quantile [SRP11]. -record [AMB12]. -records [AB11]. -robust [DA16]. -statistics [AvR15, CVL18]. -step [BH19]. -Student [AO12]. -unit [PS14]. -value [DYX15, LXZ11, MP15, Moi17b, OO12]. -values [Kel16, MMP12, SKX^{+}18].

2 [ABJR13a, ABJR13b]. 2013 [KS15].

3 [Sha15e].

4 [Sha15c, Sha19i, Sha19f].

5 [Sha15i].

8 [Sha15d]. 80 [Pak11].

978 [Sha15d, Sha15c, Sha15e, Sha15i, Sha19i, Sha19f]. 978-0-521-19676-5 [Sha15i]. 978-1-118-11777-4 [Sha15c]. 978-1-138-04960-4 [Sha19i]. 978-1-4200-9336-0 [Sha15i]. 978-1-4398-0021-8 [Sha15d]. 978-1-4522-1648-3 [Sha15e]. 978-981-4616-32-4 [Sha19f].

a-posteriori [MAV17]. ABC [TPM17]. abilitie [GdCCDS18]. ability [Hir11]. abrupt [LMSX16]. absolute [CP12, CKP15, NHA18, ZZ15b]. abundant [LWT^{+}17]. Accelerated [DYT10, AS15a, Ano14c, CZX18, DT13, DW18, HW17, Ism10, IA10, Ism14, Kan17, LHLB19, LXL11, WH19, YCD15, ZS16a]. Accelerating [PF16a]. acceleration [Ism10, ZS16a]. accept [Bot11]. accept-reject [Bot11]. Acceptance [AW17, AS12, AJA11a, ABJR13a, ABJR13b, ABA16, DG19, NS17, VC15, YCA15]. accounts [Wes16]. Accuracy [GDR12, LH14, RNSR19]. Accurate [PGF19, SKX^{+}18, SW18]. Accurately [ES11]. across [Kuk18, Moi17a]. active [SZJW19]. actuarial [GG16]. acyclic [TZSP19]. Adaptation [ÖK19]. Adaptive [CLH14, CZX18, CLYX18, GHRAM13, Han10, LLQ^{+}16, XSF17, YZX18, ZC13, ZAK13, AH19a, AH19b, BS16, Bot11, CJ13, EDASM17, FWZ^{+}15, FKS10, Haq19b, Hwa11, KL13b, LL13, MPB19, NKB19, Pam19, QKUH18, QRH19, SMBS10, SHP12, TCHS19, WMDO11, WyT17, XX15, YCXN14, YS19, ZS16a, ZWCK16, Zie11]. adding [LDB10]. Additive [HZL16, KE10, LLBL14, PL18, QX12, XLW10, YHWS18, ZCZ17, ZLZZ18]. Addressing [NMPR14]. adjacent [DHP14]. Adjusted [Ali14, ASB14, BTR16, SZS16, TK14]. adjusting [DGLV17]. adjustment [Moi17b, SXTJ17]. adjustments [CS16, RS17]. ADMM [SWXJ18]. advanced [ACG^{+}16]. affects [Sha19f]. after [LCZ18, Say12]. against [AK16, AD10b, AB14, Kus11, Li14, MAEP14, SNG12, Yam19]. Age [Fuk11, CCM12, Yu11]. Age-period-cohort [Fuk11]. age-specific [Yu11]. aggregate [PJ15]. aggregated [HWWZ16]. Aggregating [MM15]. aggregation [Kus11]. Aging [HLN^{+}15]. aided [CKP16]. air [DRC^{+}16]. alarm [HP11]. alert [AD16b]. Algorithm [Sta10, BH19, BSS17, Bor17, BOG15, Bot11, CC16a, CH15, DHP14, DN13, DLZ19, FFP16, FM15a, FZ18, GBdL16, GGdC17, GM16, HSR15, Hof12, HY16, HH15b, JKLR11, KM14, Kim18, KSLN^{+}18, KL18b, KC16, LLXY17, MBL15, McF16, PD13, RB18, SP16, San12, Seo11, SL15, SWXJ18, SZJW19, SC16, SZWL19, SWLZ15, SZM17, SS15, TPM17, TKÖ19, VVNTVD^{+}17, WMDO11, WH14, WTJW17, WT18, WyT17, Wu12, XHYX14, XLW10, XHEM17, YZ14, YLL17, YXZ19, YW14, ZCW^{+}17]. algorithm-based [Bor17, RB18, VVNTVD^{+}17]. Algorithmic [HM17, LJV^{+}18, WC14]. algorithms [AKU11, GHRAM13, Hol19, HBC11, KK15, LMSX16, LWT^{+}17, LL18b, NHGS14, PM10, SL17, SD15, Sha18b, SKTC11, TW14, TNS14, TJK13, Tu19, ZF16]. Ali [Ano14c]. alignment [FMK15]. allocation [ARB13, BD17, BS16, GB17b, HW17, KGA12, SPS19, VM19, WL14]. Almon [ÖK17]. almost [Wu16a]. alpha [FS15b]. alternating [BAB15, CR18]. alternative [AM13, BCO13, CNP19, GG16, KR15a, LR18a, Mug16, VAW15, WN13, dLHT17]. alternatives [AD10b, AB14, TCM11]. Aly [Ano14c]. ambiguity [LL19]. amendment [LN13]. American [FM19, JK14]. AMMI [PRMM12]. among [Xie14]. amputation [SLV18]. ANA [CLS^{+}19]. analog [PA15]. analyse [EKE^{+}18]. analyses [GDPH12, KKL^{+}15, Sha19u]. Analysing [Par12, She14a, SKJ17, Tso15]. Analysis [AYR16, CB10, Cox13, DK17, GK16, GWH14, HK16b, HAC16, JG16, KK14, PM10, RSD14, VFLR10, ASM17, ARY17, AZS15, ACG^{+}16, AB13, AW14, AHJ16, BTD18, BBL13, Car16, CW16, CL13, CL15, CK14, CLL10, CB11b, Da 15, DW19a, DM16, DB10, DAB11, DM12, DFT17, Fal16, FM19, Fan17, FMZ18, FDGD16, FAV18, FPG13, FAT19, FM15b, HKK^{+}16, HSJ18, JY13, JSM13, JTL18, KKY15, KE10, KEW13, LKM^{+}15, LCLP15, LSA^{+}15, LN13, LLV^{+}14, LJ18, LLC17, Lin16, LPS12, LV17, Mal16, MN15, DMV17, NLOT19, OMY12, PBSZ13, PJ15, PRMM12, PB15, RZ13a, RdSF16, Rin12, RTM18, RRDU13, RCL15, SCL^{+}18, SB15, Sha10f, Sha14f, Sha15f, Sha19b, Sha19d, Sha19h, Sha19o, She12, SE11, SA16, SU11, TNJ17, TA11, TKK19, UG10, VKK14, WL15b, WL15c, WYZK16, WZS17]. analysis [WC17, WT18, Wei12, Wil15, WWB18, XHYX14, XLW10, XYZ19, XWMA18, YWL18, YBAA15, ZP14, ZCZ17, ZY19, ZY15, ZB13, dCOC16]. analysis-based [LJ18]. Analytical [Tsa18, BS12, HM17, LPS13]. analyzing [YQ19]. anchovy [PSY18]. and/or [FHO15]. annealing [SGZM14, Woo10]. ANOM [MY13]. anomalous [DF14]. ANOVA [ÇŞ18a, CL15, GGSS19, MY13, NFFM14]. ANOVA- [MY13]. appearing [AS15b, Ten19]. applicability [Wes16]. applicable [Gau11]. Application [THG15, AAG^{+}19, BCL18, BS13, Bor17, CCF19, Con10, COS11, Erd13, FW15, HOC^{+}19, HLVRS18, HLN^{+}15, Jer13, JWWdL16, KAJB19, Kia10, KMS17, KBM19, LLP^{+}14, Mån13, Mug16, Nou10, Nou19, OS14, PF16b, RB17, SHLT17, Sha18c, SC12, TMG18, TKK19, WDCK15, WSC18, YWL18, ZZXS17]. applications [AHH17, Aci18, ACN^{+}17, AACR18, CLS^{+}19, CTTS19, CGdS14, COP14, COS14, CAO^{+}17, CYRO18, DLS18, DSVY14, FI17, FKM13, FZ18, GDCO11, GDCO18, GLB17, Haq19a, Hay15, HM19, JY13, JKP19, JL16, KR15a, MS11, MFD16, OY15, OWLP16, Özk19, SDS16, Sha11e, Sha13c, Sha15c, TCA^{+}18, UM14, VP16, WSLX17, Woo10, Wu16c, Yos18, ZF16]. applied [JK14, NJ18, RY13, WM12, Sha19d, Sha19e]. Applying [Bru19, ASS11]. Approach [Sha19j, AE17, ABA12, BT16b, BMK14, CR18, CLC17, CM19, CB11b, DHP14, DK17, DI11, FHSC14, FP17, FNRCM17, FS15a, FAT19, Fuk16, GYV^{+}13, GVW17, GdCCDS18, GV18, GGSS19, HH17, HBT12, Ho16, HH15a, HSJ18, HS13, HNPB18, JMM^{+}17, JYML13, LDCL17, LJV^{+}18, LWZ17, LXZ11, LLBL14, LLWY15, LZZ^{+}15, MA17, MT17, Mug16, NS18, NLOT19, OSdVM13, PC11, PWW15, PC10, QX12, QMZ15, RAB14, RWCD17, RB17, RIY18, Saf13, SS12a, SCL^{+}18, SFS15, SA16, Şir19, SR16, SKJ17, TK18, TGL12a, Tsi15, UA16, WXF11, WL16, WYZK16, WH17, WHX^{+}17, Wu11, WFC^{+}18, YA16a, YAEU13, ZP14, ZZ15a, ZAK13, dSdS17]. approaches [BTD18, BAB15, BBC10, CB19, CNP19, SGGC10, WBAS15, XLB12]. Approximate [Are12, BV15b, CP12, HHC15, KP15, QKY16, RNA17, SAM13b, CP14, Chi10, KK15, KL12, KL13a, PF16a, RCL15, XZY13, You14]. Approximating [Bak18, Len11, SLD16]. Approximation [WN18, Alw17, AGA18, CK14, EFGMD13, GMR15, HN13, JG10, LLC17, LPS12, MG17, MBL15, McC14, McL14]. Approximations [Kel16, GP15a, JSM13, SB12a, Ye16, GI17]. Arc [TGL12b]. Archimedean [GHH17, Hof12, SU18]. architecture [BR16]. area [CCS12, Con10, HM13, LXL17, MS19, QMZ15, SRP11, TS14]. areas [NTC11, NKZ19]. ARFIMA [Tsa10b]. Arieh [Sha19f]. arising [AR16b]. arithmetic [DTZZ12, MS15]. ARL [CC11b, GWX14, HP11]. ARL-unbiased [CC11b, GWX14, HP11]. arm [SKX^{+}18]. ARMA [FP11a, LL11, MMK10]. ARMS [ZWCK16]. array [TP15a]. array-based [TP15a]. arrays [Jou15b]. ARTICLE [Ano14c]. artificial [LCN^{+}17]. ascent [FH11, Kia10]. ascertainment [RSD14]. Asian [Meh15, MS15]. AsiaSim [Zha17]. AsiaSim/SCS [Zha17]. aspect [Sha19f]. aspects [HR12, SZM17]. assess [MMP12, NdCOP15, Wu11]. Assessing [EXH16, PC10, ZA12b]. assessment [Car19, GZT14, LB11, LR18a, WC14, XBL18, Yuc17]. assignable [NMS18, NTG13, SSMF18]. assisted [PV19]. associated [BAKZ16, DB10, GB17a, KC14, LMFMA15, Lüt15, NdCOP15, SAT16, ZAK13]. association [Ali12, BS12, BS14, CL13, PZY^{+}14, XZY13]. assumption [AW14, Aus18]. assumptions [KKY15, WCK11]. assured [LH14]. Asymmetric [APF18, AR16c, BOG15, dlCHBCD19, KK14, KC16, LK17, Lem12, MN19, MKW16, TZA10, dABS16]. asymptotes [DHP14]. Asymptotic [BCJG12, NJdC14, RS17, WPXL14, AP17, CRV15, DTZZ12, McC14, RA17, Shu12, WL10, YZX18]. Asymptotically [DY16, WN11a, AKS^{+}15]. asynchronous [SYL^{+}14]. attributable [IS13]. attribute [Gov17, LC10, TD19]. attributed [FAT19]. attributes [DG19]. augmented [MK12]. augSIMEX [ZY19]. autocorrelated [AÖ16, WL19]. autocorrelation [GFS15, GVTGH^{+}18, HKKL17, Jin15, MSK14]. autocorrelations [LM13]. autocovariances [TGL12b]. Automatic [CJ13, LKM^{+}15, CH15, Ten19]. autoregression [FP17, YWL18]. autoregressions [Akn13]. autoregressive [AHAM15, AAH19, BCCN18, BH18, CL11, CLH14, CTTS19, Haj19, Jou15a, KL17b, KL17a, LL15b, LJL15, MAV17, MN19, MS17, NSMFR15, PB19, Rai12, SM15, SL15, THG15, YPL13, ZB10, ZS16b]. AutumnSim [Zha17]. auxiliary [AH19b, ARQ19, HK16a, HK18a, KTJ18, LAuaS^{+}15, MSS18, MKSN18]. Average [NdCOP15, Adr18, AAH19, BCCN18, CHB18, Haq14, HBMAO15, HL15, Jou15a, LPJ14, LZGW14, LCZ18, MMR16, PB15, RS14, TKJ13, TK16, Wu16b]. averaged [FP15b]. averages [Zha16]. averaging [FW15, Moi17a, Tsi15]. avoid [Sha19f]. aware [PSS15].

B [MT17]. B-spline [MT17]. backcross [NPJ14]. backward [BK17a]. Bain [GHMR19]. Balanced [LPSR16, CP14, CB11b, LZ10, PO14, YML14]. bands [MC14]. bandwidth [Bag11, Bai16, BC19, BAKZ16, DL19, JW18, JP14, KC14, WZ19, XY16, YBAA15]. Bangladesh [IGR13]. bank [Erd13]. bankruptcy [Erd13, JL16]. bar [Rav19]. barley [PRMM12]. barrier [KS17a]. Bartlett [LCM12]. based [AF17, AJFB14, AMB12, ARY15, AG19, AKS^{+}15, ADRA15, AAS18, AO11, AOR13, ABP16, AB16b, AS15c, AKAW15, AJA11a, AJA11b, AWJ^{+}13, ABA16, AÖ13, AF12, BH19, Bak14, BZ16a, BT16a, BU17, BBA15, DS18, BMP12, BRY17, BS14, Bor17, BBM18, BCO13, CS18b, CJ13, CC11a, Cha17, CR18, CPK19, CC16b, CYL17, CMX17, Cor13, CBG16, DBC12, DS11, DL19, DF14, DW19b, DAB11, DB11, DK10, DF19, DC15, DW18, DA13, EDASM17, EG18, Fan19, FF14, FP11a, FDGD16, FGHRM12, FAT19, Fuk16, GGdC18, GL16, GZT14, GZL18, GLLO14, GP19, GTB14, GAB14, GBCS16, GRPP10, GEC19, GWX14, GV18, GGSS19, HT12, HK18b, HC17, HMP17, HH15a, HCY16, HMM13, Hua16, HKKL17, HLL18, IRNB18, JK14, KK14, KL17a, KM17, KN15, KN16, Kiz17, KL18b]. based [KBL^{+}15, LF16, LP16a, LL10, LL11, LK17, LPK18, Lee19, LSA^{+}15, LW12, LYQ^{+}15, LWZ17, LXL17, LJ18, Li18, LWS19, LBN^{+}19, LY13, LLN13, LSA16, LP16b, LSF^{+}17, LGWZ18, Llo10, Lu14, MAV17, MMR16, Man15, MLCL18, Med16, Mug16, NK15, NTZ19, NB13b, NB14, NB15a, NB16, NS17, NLHD12, NKB19, NA11c, NA13, NP16, Nou17b, OvP10, PB13b, PBSZ13, PA14, PK17, Par17, PSS15, PS19b, PC10, PS14, PJR15, Rav19, RL15, RWCD17, RB18, SJN15, SP16, SAB15, SMM19, SM11, SGGC10, Seo11, SBAA14, SBS14, Sha16, Shi15, SK17, SC19, SA15, Shu12, SYL^{+}14, SDWL17, SLLZ18, SZWL19, SKM14, SZ16, SU18, TdSC19, TK18, TTZS15, TP15a, TdC18, VSG^{+}18, VGTGCFC17, VVNTVD^{+}17, WDCK15, WGC14, WL15c, WBGJ15, WBG15, WL16, WH17, WKJ^{+}19, WyT17, WJ19, Wu10, WY11, WWCL11]. based [WLL12, WC14, WL14, Wu16a, WWB18, XY16, XYZ19, XBL18, YWZ18, ZA11, ZA12a, Zam15, ZMKAV19, ZCL19, ZLQ^{+}17, ZCW^{+}17, ZX14, ZZ15a, ZS16b, ZGW14, ZB13, ZNW19]. Baseline [CCM12, DDD17]. Basic [Sha14a]. Basic(R) [OO12]. basis [DDZ13, MMK14]. Basu [PK16b]. batch [JY14]. Bate [CDG^{+}15]. bathtub [ASH16, KTSR17, UGMK13, WWCL11]. bathtub-shaped [ASH16]. Baumgartner [Mur12, Mur15]. Bayes [AHAA10, Ak19, CS11, GZT14, GP15b, HBFSGD11, Ism10, JG16, LLN13, NKZ19, PK11, PK16b, QMZ15, RASR16, SSK13, SAM13b, TS14, UG10, UGMK13, XZY13, XLMX19, ZZXS17, ZWDM19]. Bayesian [Lia14, AHAH14, EB19, AJM11, AMB12, AY14, AY15, Ali15, ACG^{+}16, Ami11, AHJ16, ABA12, BL19, BAKZ16, CRV15, CCZJ17, CW16, CS17, CQJ12, CTC17, CYL17, CK14, CF17, CB11b, DA14, DP15, DH14, DD12, DRLP14, DW19b, DAB11, DW17, DFT17, EKBO16, FW13, FW15, FDGD16, FPG13, FTS10, Fu16, GdCCDS18, GGM18, GMR15, GI17, GHMR19, Gun18, GIDB15, Han17, HBT12, HLVRS18, HZL16, HY16, HW12, HC15, HLN^{+}15, Hua16, IGR13, IP14, JP17, JMM^{+}17, JYML13, JTL18, Kal14, KLK17, KAR13, Kiz17, KK15, KSLN^{+}18, KK11, LDCL17, LB11, LC19, LW14, LW17, LY18, LY13, LHB11, LHB13, LPSR16, LZZ^{+}15, LL18b, Lu19, LLT12, MJ16, MKW16, MT17, MGR15, MWL14, NKP14, NTC11, NY16, NS18, PJ15, PMM18, PS19a, PWW15, PF16a, RZ12, RdSF16, RIY18, RRDU13, SB15, SPS19]. Bayesian [SN17, SF12, SBAA14, SBS14, Sha16, Sha14b, Sha15i, Sha19b, SK18, SR16, SAK14, SXTJ17, TPM17, TNJ17, TTWT19, TGL12a, TNM17, TCLY14, Tsi15, WL15b, WL15c, Wan18a, WWW15, WFC^{+}18, YCD15, YL14, Yu15, YZWM19, YW14, Zha14, ZL15, ZAK13]. beauty [Sha19i]. Behaviour [CI15, LMRW17]. behavioural [Sha11c]. Behrens [PA15, RS15, WRN18]. belonging [GVW17]. Ben [Sha19f]. Ben-Naim [Sha19f]. benchmarking [NTC11]. benefit [DRAR19]. Berger [LLB12]. Bernoulli [Bak18, LZ11]. Bernstein [SU18, YZWM19]. best [BDKM11, Che14, GMG13, RRB10]. Beta [BCCN18, AB10, ASB14, BSSC10, BSS17, BBT13, CMC13, CGdSO13, CNO13, Cor13, CSPO14, CF17, CNS12, CNQ14, FBV18, FP11b, JP17, LY15, NCO11, PCN14, RPN15, Shi15, TNS14, ZZXS17]. beta-binomial [ZZXS17]. beta-inflated [FBV18]. beta-modified [NCO11]. between [Are12, BF12, BD17, CC11b, HH17, IGR13, JMM^{+}17, Jin15, KLKK19, KL13a, LZ10, Mal16, Mar14a, Pak10, Pak11, Per10, PRMM12, PP15, TD13, dB15]. between-group [Are12, KLKK19]. Bi [FWZ^{+}15]. Bi-level [FWZ^{+}15]. Bias [ABGM18, AAVG16, BT14, CG15, MD18, NB15b, SY15, TN16, WN11b, AB16a, FGHRM12, LS19, LPS13, SD15, WJ19]. Bias-corrected [MD18]. Bias-correction [SY15]. Bias-variance [NB15b]. biased [FH15, TKÖ19, TP13, XZ19]. bicompositional [Ber12]. bimodal [AR16c, ROCH16, dABS16]. Bimodality [VSG^{+}18]. BINAR [SKJ17]. binary [Ahm16, AI16, BS16, BAB15, CB10, CTC17, DL13, Fan17, FHS12, GSL^{+}14, GMR15, HSC16, dlCHBCD19, KOM19, LS19, LLM16, LYLM17, LR18a, Lu19, MHA10, NMPR14, NJdC14, NdCOP15, PMM18, PQ14, RNA17, RNSR19, RSD14, SH16, SKX^{+}18, SBD10, SJ10, SR11, Woo10, ZB10, ZS18]. binned [WW16a]. binomial [Ame12, AH16, ÇK19, FI17, Fu16, GGdC17, GY16, HY14, KS17c, Li15a, LLY19, LLB12, Mån13, Mog11, Özk19, PA15, PRNG18, RZ12, RZ13a, SP11, TK14, Wel16, You14, YGX14, YXZ16, ZPL16, ZZXS17, ZHB18]. bioassay [DS10a]. biological [AP19]. biometric [KAWA12]. Biostatistics [Sha14c, Sha14e]. Birnbaum [AJA11a, BZ14, LDCL17, Lem11a, LCM12, Lem12, Lem13, LMFMA15, Lem16, LX16, MLCL18, NGXZ14, PJR15, SN17, XYZ19]. Bivariate [AE17, FI17, Muh16, dALNCdATdC11, PBSZ13, RPFOMGRM17, AJFB14, CMX17, CF17, DN13, FOC14, Fam12, GB15, GB17a, GGAM13, GV18, HP15, HEB13, HN11, Kiz18, LP13, LLGP17, LLN13, LXL11, LC10, MB16, PNN17, PK16b, RSD14, SK11, Tso15, UY12, WGC14, WBGJ15, WSLX17, WZ19, WN12]. BL [DGW10]. black [SF12, Gül10]. black-box [SF12]. BLIEs [SRAO11]. blind [GGM18]. Block [ADA18, CB11b, DS10b, Juh16, PK16b]. blocks [HM17]. BLUEs [SRAO11]. Board [Ano10, Ano11, Ano12, Ano13, Ano14a, Ano15b]. Boca [Sha19i]. Book [Sha14d, Sha15f, Sha15b, Sha15d, Sha15g, Sha15c, Sha15e, Sha15h, Sha15i, Sha19c, Sha19d, Sha19e, Sha19i, Sha19f, Sha19g, Sha19h, Sha19j, Sha10a, Sha10h, Sha10j, Sha10k, Sha10l, Sha11a]. Books [Sha10b, Sha10c, Sha10d, Sha10f, Sha10e, Sha10g, Sha11b, Sha13a, Sha19k, Sha19l, Sha19n, Sha19m, Sha10i, Sha11d, Sha11e, Sha13b, Sha14c, Sha14b, Sha14d, Sha14a, Sha14e, Sha14f, Sha15f, Sha15g, Sha15h, Sha15i, Sha15j, Sha19p, Sha19q]. Boole [HCY16]. Boos [LLB12]. Boost [YZ15]. boosting [LLQ^{+}16, Yua15]. Bootstrap [BBC10, BD12, DBL10, FP11a, FP15a, KL18a, LLP^{+}14, PB19, PV19, RAK16b, Rin12, SG18, SGB13, WJ19, ZS16b, AMBP17, ASS11, BC19, Bru19, Chi10, DS10b, FP17, HA10, JCW19, Jou15a, KL10, LP10, LVB18, MG17, PSK14, RMS14, THR17, THR18, XMC^{+}14, YPAC11]. Bootstrap-assisted [PV19]. Bootstrap-based [FP11a, WJ19, ZS16b]. Bootstrapping [BMM15, HA10, MK12, SH10]. both [JP17]. bottom [JKLR11]. bottom-up [JKLR11]. boundary [AI12, SYL^{+}14, Wel16]. bounded [Bot11, IP14]. bounds [Hay15, JR16, SGZM14]. box [SF12, CWM17, Dia10]. boxplot [CPP^{+}19]. branching [GGM18]. breadth [NB15b]. breadth-depth [NB15b]. break [MSK14]. breaking [AR10]. breaks [JZZH18, MV14, RL15]. bridge [BF12, TTWT19]. bridge-randomized [TTWT19]. broken [Car16]. broken-stick [Car16]. Brownian [BF12, FS15a, Guo17, KSM16, XZZ15]. Buckley [WCK11]. budget [WH19]. building [MC13, Sha15e]. Burr [AJM11, ACN^{+}17, ABJR13a, ABJR13b, BAS17, CGdS14, CYRO18, DA16, NK15, PS16, POCdP13, SSMF18, SWZ15, SEAEM13]. Burr-type [SWZ15, ABJR13a, ABJR13b]. burst [ASM17]. BUSDM [JKLR11]. business [CCS12, Sha13a].

CAC [Jer13]. calculating [CC16a]. calculus [JK14]. Calibrated [XLMX19, BMM15]. Calibrating [AL15, FGV14]. calibration [ARQ19, BBL13, HBT12, JR13, Jer13, Özg18, RB17]. can [DRAR19]. cancer [Bor17, DLS18, She11]. candidate [TZSP19]. Cannabis [CW16]. Canonical [CAC17, AY13, Mal16, SE11, SA16]. capability [AWJ^{+}13, BU17, Cha17, GV16, HCY16, HAB12, NS17, NAA18, PWW15, Per10, SZS14, Wei12, WL10]. capacity [MCW17]. capture [ABM17, HHC15, Moh17, YP10, YPAC11]. capture-mark-recapture [YP10, YPAC11]. capture-recapture [ABM17, HHC15, Moh17]. CAR [BL19]. card [FMK15]. Care [Sha19j]. Carlo [AB16a, AB13, AB14, BZF18, Bør15, CDG^{+}15, Car16, CK14, CCK13, FP15b, GP15a, HH17, HLVRS18, How13, KE10, LLC17, LZZ^{+}15, LED16, Meh15, Ned11, NA11b, NA11a, OSN17, PP10, PO14, Saf13, SJZ17, TPM17, UG10, WJ17, YW14]. Case [XZY13, AJFB14, AAS18, AO12, DFT17, EKE^{+}18, FMZ18, GDPH12, IS13, Kan17, LCB13, LCB14, PM10, PRMM12, RAB14, RNA17, TNM17, WFC^{+}18, YPAC11]. case-cohort [FMZ18]. Case-control [XZY13, EKE^{+}18, GDPH12, IS13, Kan17, RNA17]. cases [JP17]. categorical [FAT19, GSL^{+}14, GBCS16, JM10, PA14, Sha14f, THCZ18, Wei11]. categories [Alt19]. Cauchy [AACR18, GFS15, TS17, Zha14]. Cauchy-type [TS17]. Causal [TK16, LM13, Sha18b]. causality [LS11, TD13]. causes [AH18b, DK17, NMS18, NTG13, SSMF18, TZSP19]. cdf [BF12, CWM17, JZD18]. cellwise [VVW11]. CEM [BBV17]. Censored [KBS11, AHAH14, AMB12, AYR16, ARY16, ARY17, AB16b, AR16b, AY18, AAY19, BS13, BZ16a, BBA15, BBW17, BMP12, ÇŞ18a, Che14, CYL17, Che18, CHT16, CBG16, DA14, DVA15, DG16, DNMT18, DK17, DW17, DPK11, EDASM17, FM15b, GTB14, GAB14, GP15b, Gun18, GWX14, GIDB15, HC17, HSC11, HY16, Ili15, Ism14, KS16, KA13, KM12, KK13, KVK15, KDG17, LDCL17, Lee11, LY13, LH12, LLN13, LXL11, MSS14, MTSR19, NTZ19, NB15a, NB16, NLHD12, Nou17b, OCPC12, PB13a, PB13b, PZZ19, PS16, PP15, PPK16, PP19, PS14, RAB16, SBAA14, SBS14, Sha16, sS11, sS12b, She12, She14b, She14c, She15a, She15b, SK16, STW15, SEAEM13, SZ16, Su16, TCLY14, VK14, WYW12, WH14, WDY18, Wu10, WY11, WWCL11, WLL12, WC14, WFC^{+}18, XHYX14, XZ19, YCD15, ZLW16]. censoring [AHA15, AHH17, AS15a, ARY15, ADA18, AG19, AVK15, BDKM11, BB12, BAS17, CM17, CPK19, DBC12, DN13, DYT10, DT13, DC19, DKG16, ER19, HSR15, HW12, HW17, Ism10, IA10, IAGEK11, KTSR17, KR18, KP18, Lia14, LHB11, LHB13, LCB13, LCB14, LHLB19, PB12, PGF19, RT14, SKK12, SBK13, She10, sS16, sS19, SAM13b, SAK14, TN16, TP15b, WY19, WH19, XYZ19, YCXN14, ZS16a, Ano14c]. censorship [Zar17, Lia11]. central [Par12, PPGT19]. certain [AS15b, Ten19]. cervical [Bor17]. Chain [LED16, CDG^{+}15, CCK13, FP15b, LZZ^{+}15, MCW17, UG10, ZB13]. chains [DRB17, DS10b]. chances [RSD14]. change [CB11a, CK14, Dog15, EFGMD13, JZZH18, Kel16, LDB10, LLC17, RG10, SR16, Sun11, VBL17, Wu16c, ZGW14]. change-point [Kel16, LLC17, SR16, Wu16c, ZGW14]. change-points [EFGMD13]. changed [LN13]. changepoint [ACG^{+}16, dSdS17]. changes [SHST13, SB13, TC10, Yam19]. characteristic [BP15, JGPF17, KKY15, LYQ^{+}15, PL16b, SRG11, Sha15d, WN18]. characteristic-based [LYQ^{+}15]. characteristics [GWX14, Kiz18, NB16, WPC15]. characterization [OSN17]. characterizations [NA11c]. chart [AM13, ASBM19, AKJ16, AAJ16, ASA^{+}17, AH18a, CS13, CHB18, Cha17, FKS10, FHSC14, GVTGH^{+}18, Gau10, Gau11, Haq14, HK16a, HK18a, HWA^{+}14, HOR17, HCT19, HLL18, LPJ14, LAuaS^{+}15, LMSX16, LP10, MQR18, NMS18, NTG13, OWKC15, Rav19, SSMF18, SFS15, SLLZ18, SKTC11, TCHS19, TPM17, VGTGCFC17, YZL17]. charts [AF17, AH19a, AH19b, AP15, AO12, CM10, Cha17, CC11b, CKPS11, Dog15, GWX14, HBMAO15, HM18a, Haq19b, HP11, HL15, HSJ18, KPM16, KP19, KL13b, Li18, LWS19, MM13a, MMR16, MLCL18, MH17, NTC^{+}19, PC11, RA14, Rak16a, RNH19, SHST13]. checking [AB18, ZLQ^{+}17]. checks [CYPGGM16]. Chen [KTSR17]. Chi [Yoo13, CM14, CF10, SLD16]. chi-square [CM14, SLD16]. Chi-squared [Yoo13, CF10]. child [DRAR19]. Chile [PSY18]. Choice [GGG19, Ahm16, AS15b, GGM18, LLB12, Zie11]. choices [ASS11, KC14, PJ15]. Cholesky [BKR17]. Choosing [Pam19, RRB10, PP15, Li14]. circles [ASM19]. circuits [XBL18]. circular [AHM13, DFPT16a, DFPT16b, FDGD16, KS17b, LP13, MC16, ZA12b]. circular-linear [MC16]. claim [GY16]. class [AYR16, AS15c, BPH12, BSC14, CF17, DP15, GTB14, KF16, Lin14, MBKW19, MMP12, OY15, OWLP16, Özk19, RN18, SM18, XY11, Zie11]. Classes [RS14]. classic [Sha10h]. Classical [Gun18, NKP14, IGR13, LGWZ18, PWW15]. Classification [JKM16, LL19, CVL18, HZL16, JK19, KOM19, LAGM11, NdCOP15, SKTC11, YL14]. classified [CC17a]. classifier [DW19b, RZ12, TK18]. classifiers [MM15]. classifying [MMP19]. clinical [BD17, DS11, LN13, Sha14e, YYW15]. clipped [CP12]. cloning [AWH19, GC17b, MRBR15]. Closed [KL12, BV16, HCY16, LR18b, MCW17]. Closed-form [KL12, BV16, LR18b]. closed-loop [MCW17]. closeness [AB11, BDKM11]. cloud [LYL17]. clumped [RGNM13]. cluster [AL18, KOM19, KL18b, LCN^{+}17, MT13, QKUH18, QRH19, TK18, VHV^{+}16, YS19]. clustered [Ali14, CCF19, DL13, FMV^{+}19, HBL14, IPK10, QKUH18, QRH19, SMM19, She15b, Su16, VHV^{+}16]. Clustering [CVL18, KBL^{+}15, MMP12, AF12, CP14, CH15, CPP^{+}19, DLS18, DF19, HZSA19, Ism16, KLK15, MMP19, MA17, NG16, PSS15, RWCD17, SP16, SC16, SZS16, TJK13, VVNTVD^{+}17, WH17]. clustering-based [SP16]. clusters [AL18, LLP^{+}14, SPS19]. CO [Sha19f, DS18]. co-dependence [DS18]. coarsened [BBV17]. code [UM14]. codes [VP16]. coding [DTZZ12]. coefficient [BKJ16, CS13, CT14, Hay15, HM13, HLN^{+}15, LYZ11, LM18a, LCZ18, LZZM18, LLB12, MCZ17, MFR^{+}18, MMK14, NO10, NX18b, NTC^{+}19, NdCOP15, PB15, QMZ16, RNA17, SCW16, TCM11, WL15a, XZD15, YLG15, Zie11, vdAvA15]. coefficients [LL17, LXZ11, LLM16, LM18b, NJdC14, SCW16, SW18, XLB12, Xu17a]. COGARCH [MRBR15]. cohort [FMZ18, Fuk11, Yu11]. cointegrated [AHC15, HAC16]. cointegration [BD12, GHJC10, LLS13, LL15a]. cold [JG16]. cold-standby [JG16]. Collective [HBFSGD11]. collinearity [JR18]. column [Sta10]. COM [ASA^{+}17]. combat [BKR17, BAT11, LLV^{+}14, RN18]. combination [ATH12, GBCS16, LN13, WL14, Yan10]. combination-based [GBCS16]. combinations [AJFB14, GBCS16, MC13, WM15]. Combinatorial [Sha19o, VS15]. Combined [CM10, DYX15, Saz19, YXZ19]. Combining [JvBF13, Özk12]. Comedian [SS12a]. command [SZS14]. Comment [AH16, Lia14]. Comments [Lia11]. common [BD12, LW14, SY15, TK15, XLB12]. communication [PSS15]. Community [TD19]. compact [SD15]. comparative [AY13, BCT16, CGSTG18, CYPGGM16, DP15, DC13, MMK10, PMM18, PF16b, RKV17, SNGMRC16, SK13, TD14, TNJ17, ZL11]. compare [CHQ17, JCW19, NJdC14, RNSR19]. compared [MP15]. Comparing [BAB15, DGK12, Jin15, MMP15, PQ19, PM10, Wil10, WEHCC14, vdAvA15, CCP12, CB19, GMG13, HM17, KKE17, LL13, MMP19, MG17, Sha15a]. Comparison [EKE^{+}18, HTZ^{+}16, JMM^{+}17, MY13, NGXZ14, PV17b, PAFPM12, Rai12, dESM15, TJK13, TK14, Tu19, TRC^{+}18, ASM^{+}11, AB14, AGM15, BAJN14, BZF18, BD17, BCO13, Car19, CQJ12, CLC19, CNP19, DDDD10, DS10a, DS10b, FGH14, GHRAM13, IGR13, KB18, KEW13, KL12, KL13a, LXL17, MS18b, Mur12, NA11b, NA11a, NNB14, OJRACL18, PP10, PS19a, PRMM12, PQ14, PB15, RY13, RM19, RDC10, SJN15, SRP11, SEGMA19, ST13, SLSW15, SY17a, SBK13, SK17, SBMF18, Tsa11, UGMK13, VKK14, WLL12, Wu16b, XLMX19, YP10, YL15]. Comparisons [DL13, Fam12, Hwa11, YS11, Ahm19, GP15a, KLK17, KP15, Li15b, LWT^{+}17, RRB10, RBHSL11, RAB16, TQP10, Wil15, Zha15]. compatibility [GB15]. compensation [LJ18]. competing [AE17, AYR16, ARY16, ARY17, Bør15, CLC17, DG16, DK17, FMZ18, GK16, HU14, HH15a, HW17, Ili15, Lee17, MSS14, WYW12]. competitive [TMG18]. complementary [BZAZ15, RYS11, TLRB14]. complete [AE17, CLS^{+}19, GTB14, GAB14, JvBF13, LL16, PM10, SH16]. complete-data [JvBF13]. complex [CM16, Kum15, Lu14, RCL15, ZCZ15]. Complexity [HC10, MPB19, Pam19, SZS14, Zie11]. component [AHJ16, Car16, EXH16, Fuk16, HC17, LCLP15, LZ10, VKK14, WZS17, YBAA15, ZLQ^{+}17, vGK17]. component-based [ZLQ^{+}17]. components [AL18, CCF19, Car16, Cha16, Fuk11, LGW16, SR11, YXZ19]. components-combined [YXZ19]. composite [BTR16, CS16, CS17, HC15, LYQ^{+}15, LLQ^{+}16, XZ19, YZX18, ZL16]. composition [FP15b]. compositional [Sha19d]. compound [AÖ13, Che19, OWLP16]. comprehensive [LX14]. compression [DTZZ12]. Compromise [ARB13]. Comput [Pak11]. Computation [AA16, Nom14, PNN17, Sha19p, SHST13, WYX17, BF12, BZAZ15, CRV15, JL16, KK15, LZGW14, OO12, PGF19, PF16a, PRNG18, Sha15i, TNJ17, WJ17, ZHB18]. Computational [EFGMD13, SZM17, WLL12, WC14, DL13, FNRCM17, FGHRM12, HTZ^{+}16, JCW19, RB18, Sha14d]. Computationally [HRR^{+}17, VAW15]. computations [MSA12]. compute [CP14]. Computer [CKP16, DP13, Kum15, LH14, Sha19i]. Computer-aided [CKP16]. Computing [SGTKBL15, SLD16, ACG^{+}16, LWT^{+}17, RGNM13, XHEM17, YS13, YQ19]. concave [BF12]. concentration [Chr15]. concentrations [PAFPM12]. concentric [ASM19]. concept [MKL13]. conceptual [KÖ19]. concomitants [AJFB14]. concordance [Mar14a]. Concurrent [AD16a]. Condition [SGG18, NX18b]. Conditional [GDPH12, HCT19, JK14, ZZ15b, AHAM15, Bai16, CL15, CI15, Che18, CWM17, DFPT16a, ES11, GB15, KC16, KÖ19, LC18, MKW16, MARS19, PB19, SM15, SL15, THR18, Tsi15, UM14, Yua15]. conditionally [Ahm16, AR16b, BPH12]. conditioned [Kib12]. conditioning [JR17]. conditions [Ahm19, Kel16, LSF^{+}17]. conduct [NB18]. Confidence [AKS^{+}15, HK18b, Hay15, MADASAM11, WN11c, WN12, WN13, YXZ16, AR10, AB15, Bia15, BBC10, Bur14, CSJ17, Ili15, Ili16, KKE17, KL13a, LL17, LZ11, LLB12, Moh17, MC14, RAK16b, Rin12, RNA17, SG18, Tar12, Tsa10a, WM12, WY11, Zar17, Zha15, ZMW13, vdAvA15]. conjugate [AL15, DD12, WL15c, ZWDM19]. consecutive [HM17]. consideration [AYJ11, WW19, WCW15]. Considering [HL15, NdCOP15]. Consistency [SWW19, BDKM11, NB16]. Consistent [NB15a, LW12, LZZW17, NB13a]. constancy [YK15]. constant [AS15a, AKS^{+}15, ER19, IAGEK11, LHLB19, QX12]. constant-partially [AS15a]. constant-stress [ER19, IAGEK11]. constants [Lia10]. Constrained [DSP15, CMCH12, GHMR19, HL15, LMRW17, SML19]. constraint [AGA18, LPY18]. constraints [MSS18, ZY15]. construct [LJV^{+}18]. constructing [Gau10, KM14]. Construction [Bia15, HCY16, LZ11, BK17a, HM17, Kia12]. consumption [KKL^{+}15, YFT10]. contaminants [BG11a]. contamination [CC17a]. context [MWL14]. contingency [Alt19, BS14, Lin14, NBB15, NKZ19]. continuation [SWXJ18, SZJW19]. Continuous [SLM16, AHJ16, BK17b, Fan19, JSM13, KNM^{+}15, LL18b, San12, Sha15d, Tso15, WZ13]. contrasts [KKM16]. contribution [SGTKBL15]. contributions [SJZ17, SG18]. Control [AF17, Li18, LWS19, NFFM12, RA14, Tsa11, ASBM19, AP15, AO12, AKJ16, AAJ16, ASA^{+}17, CS13, CM10, Cha17, CC11b, CKPS11, DH14, Dog15, EKE^{+}18, FKS10, FHSC14, GVTGH^{+}18, Gau10, Gau11, GDPH12, GMG13, GWX14, Haq14, HBMAO15, HK16a, HK18a, HOR17, HP11, HSJ18, IS13, Kan17, KPM16, KP19, KL13b, LAuaS^{+}15, Li15b, LMSX16, LP10, MM13a, MMR16, MLCL18, MQR18, Moi17b, MH17, NMS18, NTC^{+}19, NTG13, OWKC15, PC11, Rak16a, Rav19, RNH19, RNA17, SSMF18, SFS15, Sha19g, SOH13, SHST13, SZS14, SKTC11, TPM17, VGTGCFC17, XZY13, YZL17]. Control-variate [Tsa11]. controlled [BBL13, FS15b, JMM^{+}17]. controlling [LPJ14]. controversy [CM14]. conventional [IJL18]. Convergence [DRB17, CLS^{+}19, FM19, JL16, VHV^{+}16]. convex [CPP^{+}19, QX12, SWXJ18]. convolutional [GP19]. Conway [COBH11, SY19]. coordinate [Kim18, Nom14, SP16, TW14, Wu13, YZ14]. coordinate-wise [Kim18]. coordinates [OvP10]. copula [BS14, CR18, JWWdL16, KK14, Med16, NS18, Su16, SU18, UY12, WWW15]. copula-based [BS14, CR18, KK14]. copulas [CS18b, GHH17, Hof12, DDJ16, MS11, PBSZ13, PNN17, Yos18]. copy [RRDU13]. correct [AP17]. Corrected [CCC10, MD18, TN16]. correcting [NP16, ZA11]. Correction [Ano19, ABGM18, AI12, BT14, EE15, dlCHBCD19, Rai12, SY15, Wes16, Wil11]. corrections [AB16a, LCM12, WJ19]. correlated [ADRA15, BK17a, BAB15, Dia10, KC16, LYLM17, NTG13, PQ14, RAN11, RBK16, Xie14, YQ19, ZC13]. Correlation [HN11, LJ18, ZCZ17, AY13, FRL17, GGG19, Ho16, HM19, JW18, KOM19, LLGP17, LP17, LCZ18, LLM16, MM13b, Mal16, SA16, TCM11, Wes16, WD16, WN12, XLB12, YR15, ZZ15a]. correlation-based [ZZ15a]. correlation-network [KOM19]. correlation-type [WD16]. correlations [PQ14, SE11, THR18, WRP18]. correspondence [FAT19, Rin12]. corresponding [Cha17]. Corrigendum [Ano15a, Ano16a, Ano16b, Ano17a, Ano18a]. Cosine [PSS15]. cost [ARB13, DB10, GB17b, KGA12, Sha19b, UA16, VM19, WW12]. cost-effectiveness [Sha19b, WW12]. count [ATF19, ABM17, BG11b, CF15, FMV^{+}19, FPG13, HK16b, HTZ^{+}16, KNM^{+}15, KC14, KF16, LS14, LWS19, MH16, OJRACL18, RIY18, SMM19, SM18, TGL12a, Tso16, WJ19, dSdS17]. counting [BBM18, IRNB18]. counts [LWT^{+}17, Wei12, YWL18]. coupling [MH12]. Covariance [Lem11a, AvR15, Cha15, CLC19, CB11b, Fal16, FWF16, LZZW17, MPB19, Pam19, PGT11, QMZ16, SH19, WLSZ18, Wes16, Wil15, XWZS15, Xu17b]. covariances [APF18]. covariate [DSP15, Hua16, JSM13, KA13, MFD16]. covariates [CB10, DDD17, FVB13, GM16, Kal14, KSLN^{+}18, LYZ11, Seo11, SXTJ17, YD16, ZY19, ZFZQ18]. Coverage [AB10, AKS^{+}15, Jon16, Tak17]. Cox [Aus18, CCM12, CWM17, OS14, sS11, TD14, TNJ17, YHWS18, YD16, ZF11]. CRC [Sha19i]. credible [RZ12]. credit [CKP15]. cricket [PDS16]. Crisis [Juh16]. criteria [AKU11, DS11, GHJC10, HZ14, Kal14, LDB10, LN13, Lin14, LT16, MMK10, Mar14a, MMK14, PSK14]. criterion [LYQ^{+}15, LLQ^{+}16, XY16]. criterion-based [XY16]. Critical [LCZ18, RR13b]. Cross [AÖ16, HN13, Bai16, FF14, TS14, VAW15]. cross-sectional [TS14]. cross-trained [FF14]. Cross-validation [HN13, VAW15]. crossed [GI17]. crossing [EB19, PQ19]. crossover [CB19]. CTP} [OJRACL18]. Cubic [GLB17, RCL15]. Cumulant [KK12]. Cumulative [MCBPF16, PK17, Lee17, MMR16, MH17, RASR16, XLZ18, Zar19, ZMW13]. Cumulative/dynamic [MCBPF16]. Cure [FVB13, FOC14, GBdL16, GGdC17, GGdC18, LVB18, OS14, PB17, RBC^{+}15]. current [PL18, SK11]. current-status [SK11]. Curtailed [CYL17]. curvature [Wil15]. curve [AF12, MCBPF16, Sha15d]. curves [BCO13, CHQ17, ODBT15, SSD17, Sha15d]. CUSUM [AH19a, AH19b, HM18a, Haq19b, HCT19, Lee19, MM13a, SHST13]. cycle [FR15]. cycle-equitable [FR15]. cycles [Juh16]. cyclical [AA11]. cylindrical [LP16a].

D [SYL^{+}14]. Data [JY13, MRBR15, Sha18a, Sha19q, AE17, AHAH14, ASM^{+}11, AS15a, ASM17, AR10, AMB12, AYR16, ARY16, ARY17, AAS18, AWH19, ASS11, AS15b, AB16b, AD15, AD16a, ATF19, AP17, ABM17, AR16c, AI16, AY18, AAY19, AH18b, BZF18, BBW17, BCT16, BAB15, BMP12, BK17b, Bor17, CCF19, CLDB16, CGA10, CGAP19, Che14, CL14, Che18, CM19, CHT16, CWM17, COS11, CYPGGM16, CBG16, DS11, DRAR19, DSP15, DRC^{+}16, DP15, DG16, DTZZ12, DL13, DNMT18, DS14, DK17, DLS18, DB10, DC13, DFT17, EXH16, FW15, Fan17, FHS12, FMV^{+}19, FPG13, FAT19, FTS10, FHO15, Fuk16, GYV^{+}13, GTB14, GEC19, GDPH12, GC17b, Gun18, GGAM13, GWH14, GIDB15, HBL14, HU14, HP15, HK18b, HK16b, HC17, How13, HH15a, HLN^{+}15, HWWZ16, Hua16, dlCHBCD19, HNPB18, IPK10, Ili15, Ism14, JYML13]. data [JSM13, JWWdL16, JvBF13, JTL18, Kan17, KKY15, KNM^{+}15, KAJB19, KA13, KC14, KM12, KVK15, LH18, LDCL17, LP13, LP16a, LL18a, LDB10, Lee11, LAGM11, LWS19, LNC17, Lin14, LT16, LXL11, LLM16, LYLM17, LC18, Lue19, LLT12, MB13, MSS14, MH16, MC16, MHA10, MBL15, Mei11, MT13, MTS14, MPB19, Moi17a, MR17, NG16, NTZ19, NB15a, NB16, NMPR14, NTG13, NJdC14, NMOV18, NA11d, Nou17b, OJRACL18, OCPC12, OS14, OMY12, PB13a, PB13b, PQ19, PS19a, PS19b, PZY^{+}14, PP15, PPK16, PP19, PF16b, PL18, PS14, PV19, QLW16, QMZ16, RZ12, RZ13a, Rak16a, RY13, ROCH16, RdSF16, RAJ16, RL15, RWCD17, RT19, RSD14, RIY18, RCL15, SNGMRC16, SS12a, SEGMA19, SB15, SPS19, SMM19, SLA17, STL16, Sha14f, Sha19k, Sha19d, Sha19h, Sha19o, Sha19u]. data [Sha18c, sS11, SK11, She11, sS12b, She12, She14a, She14b, She14c, She15a, She15b, SOH13, SK17, SC19, SHP12, SK16, Shu12, SH19, SBMF18, SEAEM13, SZWL19, SWLZ15, Su16, SZM17, SS13, SBD10, SJ10, TX14, TS14, TGL12a, Tso15, Tso16, TKK19, UGMK13, VFLR10, VKK14, VHV^{+}16, VK14, WYW12, WH14, WL15a, WW16a, WWW17, WM12, WN11a, WDY18, Woo10, Wu13, WFC^{+}18, XHYX14, XZ19, YK15, YL15, YCD15, YWL18, YWZ18, YQ19, Yu11, YXZ19, Yuc17, YA16b, ZX12, ZL11, ZP14, ZS16a, ZLW16, ZCW^{+}17, ZY19, ZS18, ZFZQ18, ZF16, Sha15d, Sha15e]. data-dependent [ASS11, AS15b]. Data-driven [JY13, Sha18a, ASM17, SNGMRC16, SPS19]. Data-oriented [Sha19q]. datasets [CP14, DW19b, MZZ^{+}19]. DataSifter [MZZ^{+}19]. de-grouped [CM19]. deal [RKV17]. dealing [ASM^{+}11, KS16]. Decision [Sha19r, GMG13, LLV^{+}14, MES19, NSMFR15]. Decomposition [CJS10, AP12, BKR17, CL15, LS11]. decompositions [Fuk11]. deconvolving [Lüt15]. decreasing [JG10]. Deepwater [JKP19]. defined [HL13]. definite [HLVRS18, QX12]. deflation [BBM18, LWS19]. Degradation [WZS17, AH18b, DW18, PBSZ13, RPFOMGRM17, SK18, WGC14, WBGJ15]. dementia [Yu11]. denoising [CC12]. densities [Bai16, Bot11, CYRO18, DFPT16a, LL13, THR17, THR18]. Density [LSA^{+}15, QPQ18, Ame12, BBW17, CH15, DFPT16b, DI11, GZL18, GV18, HLVRS18, HWWZ16, JY14, JP11, JP14, KAJB19, KL17a, LL10, LWZ17, DMV17, NJ18, SNGMRC16, SJN15, Tsa18, VVNTVD^{+}17, WZ19, YBAA15]. density-based [SJN15]. departures [JKLR11, RIY18]. Dependence [Gha16, DS18, Da 15, DM16, Gha19, KAJB19, KK14, Wei11]. Dependency [GBCS16, BR16, DDDD10]. dependent [ASS11, AS15b, AR16b, CL15, CGP15, DP13, Fan19, HSC16, IRNB18, JCW19, KA13, LWS19, MB13, MMP19, PV19, PL15, SWW19, SRK13, SB13, Wei12, Yan10, YD16, ZX12]. depth [DS11, DC15, NB15b, PS19b, SK17, SC19, Wil10]. depth-based [DS11, DC15, SK17]. derivation [Hat16]. derivative [Tsa18]. derivatives [Bai16]. derived [LPV13, SLM16]. descent [TW14, Wu13, YZ14]. description [AP19]. Design [DT13, KL13b, San12, ARQ19, BD17, CC11b, DG19, ER19, FMZ18, FKS10, FHSC14, Gov17, Hir11, HCT19, HSJ18, JP17, JR16, JR17, KP19, LSL10, LXL17, LPS13, MJ16, NMS18, NTG13, PT14, RNA17, RNSR19, SSMF18, SFS15, Sha19g, SJ10, VM19, WC17, Wil15, Zha15]. designing [AS12, ABJR13a, ABJR13b, BU17, LMSX16, RNH19]. designs [CKP16, CB19, DBC12, FR15, HM17, KM14, KPM16, Kia12, KEW13, LF16, LJV^{+}18, LY15, Lu19, SP16, Sta10, TdSC19, TPM17, VFLR10, Woo10, XMC^{+}14, YS18]. destructive [GBdL16, GGdC17, PB17]. detect [Aus18, DHP14, JY13]. Detecting [LGWZ18, MV14, LM16, Noo19, PC11]. Detection [AHM13, LW15, AKU11, CCF19, DC13, HMM13, Jur12, MT13, Par17, PMP14, SS12a, SDWL17, SO10, TD19, WLCL18, YWZ18, YTNT14, ZGW14, dSdS17]. detector [SHLT17]. determinant [GGG19]. determination [IP14, JK14, Su16, TZSP19, TKÖ19]. determining [SA16, Sta10]. deterministic [DTZZ12]. detrended [CLL10, Lin16]. Developing [AWJ^{+}13, BU17, Mån13]. Development [HWA^{+}14]. deviance [Adr18, YL14]. deviances [RM19]. deviation [CP12, Haq14, Wil11, ZZ15b]. deviations [NHA18]. device [CCF19]. Diagnostic [VRC13, AB18, BSS17, BCO13, DGK12, FAV18, NJdC14, PS19a, RNA17, RNSR19, WZ13]. Diagnostics [CL11, CA19, AAG^{+}19, AAAQ19, BMK14, DR17, FLB15, GLLO14, LJL15, dCOC16]. diagonality [Xu17b]. dichotomous [SCL^{+}18]. Dickey [MK12]. dictionary [LMRW17]. Difference [Lu14, ADRA15, AD12, BF12, DA13, Kla15, KL13a, PB15, RZ12, Wu16a]. Difference-based [Lu14, ADRA15, DA13, Wu16a]. differences [Per10]. different [AKU11, DAP15, Han17, HTZ^{+}16, LWT^{+}17, LPS13, MQR18, NA11a, PC10, RdSF16, TRC^{+}18]. differential [JSM13, NHWT14, OSdVM13, WS16]. differentially [LM16]. diffusion [DSE16, DF14, NHWT14, SYL^{+}14]. diffusion-type [DSE16]. diffusions [PF16a]. digital [XBL18]. dilation [Zar19]. Dimension [DKY17, Par11, ZLZZ18, Adr18, Ahm19, MT13, XLZ18, Yoo13, dCFOM12]. dimensional [AvR15, AAG^{+}19, AB13, AAY19, BTD18, CCF19, CAC17, Che18, CLC19, DLS18, DW19b, DDD17, HM18b, HH15b, JSM13, KOM19, LKM^{+}15, LL16, LY18, LLBL14, LC18, MCZ17, MWL14, PZZ19, Par17, QLW16, RY13, RT19, SS12a, SZJW19, SE11, SH10, TS16, WLCL18, WLSZ18, Wu13, XX15, Xu17a, Xu17b, YHWS18, YWZ18, YXZ19, ZFZQ18]. dimensionality [Fan19, SA16]. dimensions [DR17, SWXJ18]. directed [TZSP19]. direction [YK15, ZWCK16]. directional [KK14]. directions [Lue19]. Dirichlet [Ber12, HZSA19, HBC11, SPS19, YZWM19]. disaster [JKP19]. Discover [Sha19f]. discovery [LPJ14, Li15b, PZY^{+}14]. discrepancies [HH17]. Discrepancy [TA11]. Discrete [KC14, TKK19, AÖ13, GB15, GB18, GDCO11, JO11, Lee17, MB13, MA17, MC16, QKY16, TZSP19, WCK11, ZAK13]. Discrete-time [TKK19]. discretely [XZZ15, Zha16]. discriminant [GOC18, HZL16, OMY12, RY13, SE11, Shu12, SAM13a]. discriminate [BSS15]. Discriminating [Pak10, dB15, Pak11]. discrimination [JWWdL16, PSV11, SLSW15]. disease [SZS16]. dispersed [HK16b, SMM19, SKJ17, Tso16]. dispersion [AM13, CW16, CNS12, DFT17, Haq14, PA15, SP11, dSSCR19]. dissimilar [AD16a]. Distance [Käm13, AP12, AGA18, BMK14, GL16, Jin15, Kim18, LF16, LGWZ18, SG18, WZ18, YR15, ZX14, ZZ15a]. distance-based [GL16, LGWZ18]. distances [WY11]. distinctness [Yuc17]. distortion [ZCZ17, ZLZZ18, ZLL19]. distributed [DH14, DRLP14, DGW10, GWX14, LCN^{+}17, Lüt15, McF16, ÖK17, Özk19, TK16]. Distribution [MMR16, SBK13, Yan10, AE11, AHAH14, AHA15, AHH17, AAR15, Aci18, AJM11, AMB12, ARY15, AYR16, ARY17, ADA18, AG19, AKA^{+}16, AAS18, Akg19, AZS15, ASH16, AL15, AO12, AY15, Ali15, AB16b, ATON18, AAVG16, AVK15, AJA11b, ABJR13a, ABJR13b, AKJ16, ASA^{+}17, BAJN14, BCY16, Bak14, BDKM11, BS13, BZ14, BZAZ15, BZ16a, BSSC10, BSdMC11, BB12, BET16, BT16b, Ber12, BBT13, BMP12, Car16, CMC13, CM17, CC16a, Che11, CM19, CMX17, Coi13, CLAH17, CL10, COS11, CGdSO13, CNO13, COL14, COP14, COS14, COP15, CBG16, DA14, DTRB11, DN13, DD12, DAP15, DNMT18, DYT10, DA16, DB10, DAB11, DB11, DDZ13, DC19, DPK11, DKG16, EDASM17, ER19, EG18, FH15, FM15a, Fre12, FM15b, GGdC18, Gen13, GY16, GAB14, GSW17]. distribution [GdSCO14, GDCO11, GDCO18, GdCCDS18, GA15, GWH14, Han17, Haq17, HK16b, Hay15, HN11, HC17, Ho12, HOR17, Hua11, Ili16, Ism10, IAGEK11, Ism14, JZD18, JGPF17, KR13, KAT15, KTSR17, KAR13, KLK14, KN15, KN16, Kiz18, KK12, KM12, KK13, KVK15, KDG17, KR15a, KS17c, KR15b, LLGP17, LPK18, LRV17, Lem11c, LMFMA15, LX16, LW17, Lia14, LHB11, LH12, LLN13, LHB13, LL19, LR18b, MJ16, MTSR19, MD18, MARS19, Muh16, MADASAM11, NKP14, NK15, NCO11, NCO12, NSR14, NB13a, NB13b, NB14, NB15a, NB16, NLHD12, Noo19, NA11c, NNB14, NP16, NJ18, OJRACL18, OWLP16, ÖI10, PS16, POCdP13, Par17, PPK16, PF16b, PS14, PJR15, PK11, QKY16, ROCH16, RAB16, RT14, RASR16, RB12, RN14, RPN15, SJN15, SEGMA19, dESM15, SKK12, SGG13]. distribution [SN17, SBS14, SSD17, Sha18c, Shi15, SK16, Shu12, SK18, SSK13, STW15, SEAEM13, SC12, SAT16, SU18, TCA^{+}18, Tak17, TLRB14, TP15b, Tsa18, TKB16, Tza11, TP13, UG10, VSG^{+}18, WCC13, WW16a, WSLX17, Wan18b, WY19, WM12, Wu10, WL10, WWCL11, WLL12, WZX13, WC14, YL18, YCXN14, YXZ16, Zha14, ZHB18, ZWDM19, ZGW14, dABS16, BMM14]. Distribution-free [MMR16, SBK13, Par17, ZGW14]. distributions [EB19, AJFB14, ARY16, AOH16, ACN^{+}17, AACR18, AR16c, AGM15, AGA18, AJA11a, ABA16, AD12, BCJG12, BSS15, BK16, BSC14, BCY^{+}17, CCZJ17, Che14, CI15, CYL17, CC17b, CdC11, Cor13, CAM16, CAO^{+}17, CF17, CBG16, DSCS19, DYA19, FDGD16, FMB16, FGV14, GLLO14, GTB14, GB15, GOC18, GEC19, GLB17, GGAM10, HEB13, HH17, HBFSGD11, JO11, JWWdL16, Kla15, KP18, LK17, Len11, LW14, LBN^{+}19, LY13, LHLB19, Lon12, MAV17, MLCL18, MG17, MQR18, NCO15, NTZ19, NGXZ14, Nou17b, NFFM12, OY15, OS14, Pak10, Pak11, PB13a, PPGT19, PK16b, QPQ18, RAK16b, RS14, SB12a, SRG11, SDS16, SBK13, SWZ15, SG15, SAM13b, SAM13a, TS17, TZC^{+}16, TNS14, TdC18, TCLY14, Wel16, XHEM17, YZL17, dB15]. diurnal [FFP16]. divergence [Ali12, Ali14, BMPZ14, DR17, DI11, GB15, GB18, KL17a, VK14]. divergence-type [BMPZ14]. divergences [LL10]. diverging [NKB19]. divisions [Kha12]. DNA [FMK15, RRDU13, TNM17]. DNA/RNA [FMK15]. do [PCMMA13]. does [SRP11]. Domain [LLV^{+}14, RRV13]. Domain-specific [LLV^{+}14]. dominance [KAWA12, NX18a, NX18b]. Donoho [VVW11]. dose [CC11a, DS10a, DHP14, WDCK15]. dose-response [DHP14]. Double [Cha15, FHSC14, ABA16, CT16, DW17, DG19, FM19, HK18a, Lia10, MS18a, NKB19, PM10, sS16, sS19, WPXL14]. double-censored [DW17]. Double-objective [FHSC14]. double-truncation [sS16]. doubly [ASM19, DSCS19, McC14, sS11, sS12b, She12, She15b, TKJ13, YL15, ZL11]. Downton [HN11, LLN13]. DP [BCLM17]. DP-Lasso [BCLM17]. drawing [JvBF13]. drift [KSM16, MKL13]. drifts [ZST15]. driven [ASM17, AWH19, JY13, KBM19, KSM16, NB15b, SNGMRC16, SPS19, Sha18a]. dropout [NLK11]. dropouts [QZZ16, ZFZQ18]. drought [MMP12]. drug [AAG^{+}19]. dual [CM16, LT16, LPS13, SZJW19]. dual-frame [LPS13]. dual-inflated [LT16]. dual-record [CM16]. dubious [MSS18]. dummy [HA13]. duration [LED16, MKW16, PB19, She10, sS16, sS19]. durations [DW17, FFP16]. Dynamic [RB17, dSdS17, BZF18, BKJ16, FMK15, JTL18, LTJB18, MCBPF16, PSS15, RASR16, SRK13, SS13, SBD10, SJ10, THR18, VBL17, YQ19, ZS18]. dynamically [Saz19]. dynamics [FF14, SYL^{+}14]. dynamics-based [FF14]. Dysfunction [Sha19j].

E-Bayesian [Kiz17]. each [NO10, NX18b]. early [LBN^{+}19]. EBLUP [PC10]. econometric [SZM17]. Economic [FKS10, KPM16, NMS18, NTG13, SSMF18, TPM17, FHSC14, HL15, KP19, LMSX16, Moi17a, SFS15]. economic-statistical [TPM17]. ecosystem [PSY18]. Edge [OCT18]. edges [SHLT17]. Editor [Kru12]. Editorial [Ahm15, GIP15, RP18, ZT14, Ano10, Ano11, Ano12, Ano13, Ano14a, Ano15b]. Effect [ASBM19, GVTGH^{+}18, HBMAO15, LDB10, AL18, CCM12, LH18, LR18a, MC14, RASR16, RBC^{+}15, WDCK15]. Effective [MC14, AVG18, CC11a, DS10a, PHCS11, WDCK15]. Effectiveness [MC13, LLV^{+}14, Sha19b, WW12]. Effects [TC10, ATF19, AL18, Car19, CCMGA14, DM12, FBV18, FHO15, GC17b, GI17, HBL14, HTZ^{+}16, Hua16, KE10, KLKK19, KEW13, KP15, LC19, LLWY15, LZ10, MBL15, NFFM14, PO14, PMP14, RRDU13, SH16, She15b, SJ10, TKJ13, TK16, VHV^{+}16, XLW10, ZC13, ZB13]. Efficiency [ADRA15, MA10, Nou17a, RN18, GB18, JR16, Mar10, YZX18, ZL16]. Efficient [AHAM15, BF12, Chr15, DGLV17, JZD18, JLM15, LR18b, PJ15, SMBS10, TCLY14, WL16, AM13, BU17, DSVY14, Fre16, HRR^{+}17, HSJ18, HM17, KM14, PSKC18, PSK14, SR11, VNM14, WT18, WYX17, YW14]. eigenmaps [Fan19]. elastic [MPB19, Pam19, PK16a, XX15]. elastic-net [XX15]. electric [DLZ19]. electrical [KBL^{+}15]. electronic [MZZ^{+}19]. elementary [SHLT17]. Elements [Sha18b]. elliptical [AJFB14, BMPZ14, CL11, DGW10, MFP17, ZA12b]. EM-algorithm [BSS17]. EM-Type [SL17]. Embedding [SF12]. Empirical [Ak19, CS16, Che19, Ciu13, JGPF17, Wei11, Zha14, Zha16, AGM15, AGA18, BMM15, BZ16b, BMP12, CS11, CLC19, FW13, GHRAM13, GV18, Ho16, LPS13, LPV13, Med16, PL16b, RDC10, SJN15, Say12, TW14, XZY13, YS13, ZWDM19, ZY15]. encryption [DTZZ12]. Encyclopedia [Sha11c]. endogeneity [RKV17]. endogenous [HA13]. Endorsement [VC15]. endpoints [SKX^{+}18]. Energy [LX14, SHLT17, SYL^{+}14]. energy-based [SYL^{+}14]. engineering [Sha19g]. Engle [LL15a]. Engraulis [PSY18]. Enhanced [AH19a, JR16, JR17]. enough [CC19]. ensemble [MKL13]. ensembles [ZF11]. Entropy [AOH16, BRY17, KL18a, KR16, LK17, LPK18, MA10, Nou10, NA11a, NN13, NNB14, NP16, NJ18, PK17, RASR16, ZA11, ZA12a, Zam15, Zar19]. entropy-based [LK17]. environment [LS11, LYL17, MKL13, XWMA18]. environmental [Sha10j]. epidemic [EKBO16, MT17]. equal [RRB10, RBHSL11]. equality [DVA15, IPK10, LBN^{+}19, LXZ11, LLM16, LLB12, PA15, PGT11, SW18, SK17, TPW17]. equalizers [Käm13]. Equation [BH19, Con10, GHMR19, Gül10, JY13, LK13, SLM16, ZP14]. Equation-solving [BH19]. equations [Kha12, KMS17, NHWT14, SZWL19, TKG18, Wes16, ZPL16]. equilibrium [Alw17]. equitable [FR15]. equivalence [LZ10, Wel16]. equivalent [TGL12b]. Equivariant [TK15]. ER-Boost [YZ15]. Erratum [Ano14b, Ano15c, Ano15d, Ano15e, Ano17b, Ano18b, BMM14, Pak11]. erroneous [NdCOP15]. error [AHM13, AL18, CCZJ17, Car19, CSJ15, ÇŞ18a, CB10, GGSS19, HBMAO15, IJL18, KLKK19, LYZ11, LJ18, MG17, MFP14, MY13, PF16a, Rai12, STL16, Şir19, SGB13, TdC18, Tsa10a, TRC^{+}18, WL14, YPL13, ZR14, ZA12b, ZY19]. error-prone [LYZ11]. Errors [GG11, ADRA15, BZF18, CL11, FW15, FHO15, GZB13, GM16, HAC16, HHC15, LH18, Lem12, LL15b, Lüt15, LK13, Moi17b, NAA18, NFFM12, OSN17, PO14, QZZ16, RAN11, RBK16, RIY18, SWW19, SD15, Sha10a, SCW16, THG15, TX14, Wu11, ZA12b, ZCZ17, ZLZZ18, ZLL19, ZS16b]. Errors-in-variables [GG11, RIY18, SD15, TX14]. Essentials [Sha19j]. estimate [CC19, Coi13, SU18, TKJ13, WH14]. estimated [AB10, CM10, Cha17, MM13a, NO10, NX18b, TCHS19]. estimates [AAVG16, BT14, CCM12, DL19, DS10a, DHP14, FRL17, GDR12, GDPH12, JP11, Jon16, NA11a, RASR16, WN11a]. Estimating [AS15a, Bur14, CDG^{+}15, GD19, GC17b, LM18b, NAA18, PWW15, Yu11, AOH19, BAS17, CR18, Che11, FFP16, GBdL16, GM16, JMM^{+}17, LED16, NB14, SNGMRC16, STS14, STW15, SZWL19, TK14, Tu19, Wes16, ZP14, ZPL16]. Estimation [ARY15, AHC15, AO11, AOR13, AWH19, AA11, AAY19, BB12, CM17, CSJ17, CCZ13a, CCZ13b, DNMT18, DDZ13, EDASM17, FNRCM17, FAV18, FPP18, Gen13, GOC18, GGAM10, GGAM13, Haq17, HSR15, HWWZ16, HKST17, KTSR17, KR18, KN15, KN16, KVK15, KDG17, KSM16, KR15b, LZZM18, LPS13, MES19, NK15, OvP10, ÖK18, PS16, Per10, PL18, SF13, sS11, sS12b, sS16, sS19, SEAEM13, SAM13a, TP15b, UY12, Wan18b, YP10, ZS16a, ZB10, Adr18, AJM11, AG19, AMAMS12, AHAM15, Akg19, Akn13, AOH16, Ak19, Ali15, AB16b, AI12, AR16a, Ami11, ABM17, BR16, BL19, Bag11, Bak14, BZ14, BZ16a, BC19, BSS17, BAT11, BBW17, BT16b, BG11b, BAZ19, BBC10, BS12, BS14, Bor17, Bør15, BEBG14, BV15a, BH18, CG15, CB11a, CH19, CMCH12, ÇŞ18a, CS17, CJ13]. estimation [CCS12, CQJ12, CM16, CA12, CCK13, CWM17, Chr15, DD15, DN13, DD12, DAP15, DFPT16a, DFPT16b, DRLP14, DW17, DFT17, DPK11, DSVY14, EG18, FHS12, FGHRM12, FTS10, GGdC17, GG11, GTB14, GAB14, GRPP10, GHRAM13, GHH17, GMR15, HN11, HKK^{+}16, HM13, HLVRS18, HAH14, Hua11, HW12, HLN^{+}15, HC10, IJL18, Ism10, JW11, JW18, JZD18, JP14, KSÖG11, KAJB19, KC14, KL17a, Kim18, KB18, KF16, KM12, KK13, KTJ18, LH18, LP13, LSL10, LL10, LS14, LOK16, Lem11b, Lem11c, LW17, LWZ17, LXL17, LYL17, Li18, LLN13, LSA16, LLBL14, LSF^{+}17, Llo10, Lue19, MA10, MM13b, MAV17, MN19, MRBR15, MCBPF16, MD18, MT17, MS19, MAAM10, MADASAM11, NKP14, NB13a, NB13b, NB15a, NB16, NLHD12, NS18, NLK11, NAA17, PB17, PNN17, PMP14, PL16b, PK11, PK16b, QMZ15]. estimation [QKUH18, QRH19, RGNM13, RT14, RYS11, RKV17, RB18, RL14, SP11, SRP11, SM15, SA12, dESM15, SKK12, SGGC10, SGG13, SL15, SGÇ18, Sha16, She10, SK11, SY15, SCW16, SA15, SK18, SH19, Şir19, SB13, Stå16, SLM16, SR16, SWLZ15, SAM13b, SS13, SKJ17, SJ10, SGB13, TCA^{+}18, THG15, TTWT19, TP15a, TS14, TK15, TC10, Tsa10b, TKB16, VBL17, VNM14, WZ13, WCK11, WGC14, WBGJ15, WBG15, WL15a, WY19, WZ19, Wes16, Wu10, WWW15, WZ18, WHX19, XZ19, YPAC11, Yos18, YXZ16, YD16, YA16b, Zha14, ZZ17, Zie11, ZAK13, dCOC16, dABS16, dCFOM12]. estimations [AHH17, BAJN14, FWF16, Han17, HHC15, Kiz17, SAK14]. estimator [AÖ16, AKS^{+}15, AD10a, Ame12, BH19, BV16, BAKZ16, CTX18, CC12, DK10, DA13, FMV^{+}19, Fre16, GFS15, HMP17, HY14, LW12, LZZW17, LGW16, Lu14, Mån13, MKSN18, NO10, NX18a, NX18b, Nou10, Nou19, ÖK17, ÖKD17, ÖK18, Özg18, Özk12, Özk19, PP10, PL15, QRH19, RBK16, RA17, RN18, STL16, SKM14, TKG18, TKJ13, WCK11, WW12, Wu16a, WYX17, XY11, YS19, ZA11]. estimators [Adk12, ADRA15, AKV17, AY18, BZF18, BDKM11, BRY17, Bru15, ÇK19, Cha15, CS11, CLC19, CCC10, DA16, DI11, EE15, FGH14, GVTGH^{+}18, GP15a, GSW17, GB18, HA13, Hwa11, Kib12, KR16, LF16, LPS12, LP16b, Lon12, LR18b, LK13, MFR^{+}18, Mar10, MPB19, NHA18, NN13, NP16, NJ18, Özk12, PSKC18, Pam19, PC11, PP10, PV17b, QYX17, SMBS10, SWW19, Saz19, SSK13, TKÖ19, Tza11, VVW11, Wil11, WLL12, YP10, YL15, YS19, ZA12a, Zam15, ZL11, ZCZ15, ZWDM19, ZBG18, ZMW13]. ethical [BD17]. evaluate [Hir11, LLP^{+}14]. Evaluating [HA13, HSC11, LCLP15, LC10, Mar10, dLHT17]. Evaluation [BL19, BAJN14, GV16, HBC11, SFS15, vGK17, DL19, GSL^{+}14, HL13, LX14, LH14, LPS13, TD14, TGL12a]. event [HNPB18, KM14, MBL15, PV17a, RCL15, Wu13, YYG16]. event-related [KM14]. events [ASM17, CR18, CC11b, CMX17]. every [Sha19f]. everyone [Sha19s]. evidence [MV14, OSN17]. Evidential [DSE16]. evolution [MFD16]. evolutionary [San12]. Evolving [MBG17]. EWMA [AP15, ASA^{+}17, AH18a, CM10, HSJ18, HLL18, KP19, TCHS19]. Exact [BZ16a, Cha17, Ili16, Lia14, LHB11, LHB13, LYLM17, Llo10, MSS14, PPGT19, Sha15a, SZ16, Zha11, AP17, BBV17, CBG16, DG19, Fre12, Ili15, LLB12, WT18]. examination [Car16]. examinee [Hir11]. Examining [MHA10, GYV^{+}13]. example [YAEU13]. examples [Sha19d]. excavations [ASM17]. exceedance [CHB18]. Excellent [Sha14e, Sha14f]. Excel(R) [OO12]. excess [MH16]. excesses [GRPP10]. exchange [SP16]. Exhaustive [KLK15, LR18a]. Existence [NB16, EFGMD13]. exogeneity [KM17]. exogenous [AHC15]. expansion [Shu12]. expansions [MMK14]. expectation [Cha17, JK14, MBL15, MARS19, SWLZ15]. expectation-maximization [MBL15, SWLZ15]. expectations [WJ17]. expected [GC17a, MP15, PRNG18]. expectile [YZ15]. Expedient [Tso16]. experiment [Moh17]. experimental [SP16]. experiments [CB11b, NPJ14, San12, Sha19g]. experts [Li14]. explicit [Zör15]. exploration [FH11]. Exponential [JG10, ZLW16, AHA15, AMB12, AYR16, ARY16, ARY17, ADA18, BAJN14, Bak14, BDKM11, BZAZ15, BCJG12, BSSC10, BSS15, CCMGA14, Che14, CYL17, CCZ13a, CCZ13b, Coi13, COL14, CBG16, DN13, DAP15, DNMT18, DB10, DB11, DC19, DKG16, EDASM17, FM15b, GZL18, GVW17, GGA18, GMG13, GWH14, HN11, HBFSGD11, Ili15, JKM16, KAR13, KN15, Kiz18, KK13, KDG17, KL17c, LK17, LY13, Lia14, LHB11, LH12, LLN13, LHB13, LCB14, MSS14, MJ16, MKSN18, NA11c, OvP10, Pak10, Pak11, Rak16a, RG10, RAB16, RB12, RPN15, RS17, SB12a, SKK12, SBAA14, SBS14, SSK13, SZ16, TCLY14, TKB16, WSLX17, WY19, WY11, WC14, Wu16b, YL18, YS19, CM17]. exponential-generalized [GWH14]. exponential-geometric [BZAZ15]. exponential-Poisson [BAJN14, BSS15]. exponential-Weibull [COL14]. exponentiality [AD10b, AB14, DS14, NA11a, NA11c, NA11d, PB13b, TMG18]. exponentially [DH14, GWX14, Haq14, HBMAO15, HL15, LPJ14, MMR16, McF16, PB17]. exponentiated [AHA15, CC17b, COS11, CGdSO13, GTB14, GG16, MTSR19, NCO15, NTZ19, OCPC12, RT14, RB12]. expressed [LM16]. expression [Cha17, OSdVM13, Tsa18]. Extended [AP19, FMK15, RBK16, ASH16, CGdS14, CSPO14, DNMT18, GGAM10, GV12, Kuk19, PZY^{+}14, dESM15, TKB16, ZS16a]. Extending [FR15, Sha19j]. extension [Aci18, CF10, Li14, LLXY17, NFFM14, RKV17, UGMK13, VP16, YLL17]. Extensions [Kia10, GDPH12, MGR15]. extra [HK16b, Mog11]. extraction [CSJ15]. extrapolation [WZ19]. extremal [MR17]. extreme [AO11, AB16b, CH14, FP14, Gha16, Gha19, LS14, MS18a, Pak10, Pak11, PV17b, PF16b, WY11, YCXN14]. extreme-value [FP14]. extremes [CS18b, LdNdSF18].

F [LZ10]. F-test [LZ10]. Factor [SGG18, ABGM18, AB13, DW19a, EE15, GGG19, HKK^{+}16, Ism10, MS15, NLOT19, XZY13]. factorial [Fan17, Lu19, Sta10, VFLR10]. factors [Wil11, XLMX19, ZS16a]. failure [ARY16, Ano14c, AAGV12, AH18b, COP14, DK17, DKG16, GP15b, IA10, JG10, Kan17, KAR13, KDG17, Lee17, Sha18c, SK18, SRK13, SEAEM13, WY19, YCD15]. failure-rate [SRK13]. failure-step [Ano14c, IA10]. failures [AVK15, LBN^{+}19]. FAIR [Sha18a]. faithfulness [AW14]. fallible [RZ12]. false [LPJ14, Li15b, XX15]. familial [SR11]. familial-longitudinal [SR11]. families [Ten19, ZX14, ZBG18]. family [ACN^{+}17, AACR18, AD12, BMPZ14, BCY^{+}17, CCMGA14, Che11, CTX18, CdC11, CAM16, CAO^{+}17, GHJC10, GVW17, GHH17, GDPH12, KL17b, KP18, Lüt15, MP16, NJR13, NCO15, ODBT15, PSKC18, RS17, TZC^{+}16, XHEM17]. Fan [LL13]. FARIMA [RMS14, SB12b]. Farlie [UY12]. Fast [CTC17, DP13, FMV^{+}19, JW18, MCZ17, MMT16, YQ19, DLZ19, FM15a, Ism16, KM14, Kim18, VAW15, YLL17]. fatigue [BSC14]. FDR [Hwa11]. feasible [ÖKD17]. Feature [YHWS18, Che18, LWZ17, LY18, MCZ17]. feed [dB15]. feed-forward [dB15]. Fence [NPJ14]. Fiducial [LX16, GGSS19, YL18]. fiducial-based [GGSS19]. fields [ABP16, APF18]. filling [Jou15b]. filter [CC16b, YTNT14, dSdS17]. filtering [FG13, Ned11]. finance [Sha13a]. financial [Ahm16, CGAP19, CL16, FFP16, GC17a, Gus15, HM19, MV14, Ral17, YTNT14]. finding [DI11]. Finite [AM19, RRV13, SA12, Wil11, Ak19, CHTZ14, DS10a, LC19, LZZ^{+}15, Lu19, MAV17, MA17, MKSN18, NY16, PSKC18, Wes16, YWL18]. Finite-population [SA12, Lu19]. finite-range [YWL18]. finite-sample [Wes16]. first [CL11, DKG16, HP11, KL17b, KDG17, NSMFR15, SGTKBL15, SEAEM13, SC12, TP15a, WY19]. first- [TP15a]. first-failure [DKG16, SEAEM13]. first-failure-censored [KDG17]. first-order [CL11, HP11, KL17b, NSMFR15, SGTKBL15]. Fisher [Chr15, DBC12, PB12, PA15, RS15, WRN18, Yan10]. fit [AE11, AOH16, Ali14, AS15b, BPH12, BET16, BMP12, DS15, DC19, FH16, GEV18, JV14, KK12, LK17, LPK18, Li11, LLY19, MA10, NA13, NP16, Nou17b, Nou19, PB13a, QHB15, RDC10, SJN15, Ten19, TA11, UY12, VK14, WD16, YS18, ZA11, ZX14, Zör15]. Fitting [AR16c, FS10, GI17, Kan17, WW16a, ASM19, CM19, HRR^{+}17, HTZ^{+}16, LWT^{+}17, Mar14b, MMT16, Rai12, TGL12a, WFC^{+}18, ZC13]. five [NA11a]. Fixed [NFFM14, AHAA10, LZ11, PMP14]. FL [Sha19i]. flatness [TS16]. Flexible [BK17b, Bør15, HU14, XBL18, CW16, CF17, GGdC18, GHJC10, LLWY15, MMT16, PB12, ROCH16, WTJW17, WCW15, WFC^{+}18]. fluctuation [CLL10, Lin16]. focusing [MMK10]. fold [LDB10]. fold-change [LDB10]. folded [CC16a]. following [GSL^{+}14, HU14]. Food [LSL10]. footpaths [LF16]. forecast [FW13, THR17]. Forecasting [AI16, BS19, MMK10, Moi17a, DLZ19, GC17a, Her11, KL18b, LL11, RWCD17, Sha19a, YAEU13]. forecasts [Li14, LL15b]. forensic [OSN17, Sha11e]. Forest [CAC17, HBL14]. forests [CCF19]. form [AD16b, BV16, KL12, LR18b, SK11]. forms [GBCS16]. formula [Cor13, Lem11a]. forward [dB15]. foundations [Sha18b]. four [CB19, NSR14]. four-parameter [NSR14]. four-sequence [CB19]. fourth [RR13b]. fraction [CLDB16, FOC14]. Fractional [CGA10, FBV18, KSM16, Mar10, NHWT14, RL14, RRV13, SF13, Sta10, XZZ15]. frailty [AÖ13, FRL17, HP15, WXF11]. frame [LPS13]. frames [Moi17a]. framework [ABA12]. Francia [MP15]. Fréchet [GG16, NTZ19]. free [CC17a, Che18, LWZ17, MMR16, Par17, SGGC10, SBK13, SYL^{+}14, ZGW14]. freemium [CLC17]. frequencies [Zör15]. Frequency [Tak17, AA11, CGA10, CGAP19, CMX17, FPG13, LLBL14, RRV13]. frequency- [RRV13]. frequentist [JMM^{+}17, LB11, Yao15]. fresh [MSS18]. Frobenius [HM19]. frog [YPAC11]. Frontiers [Sha10a]. Full [HKK^{+}16]. Fuller [MK12]. Fully [LY18, MHA10]. function [ASBM19, ARB13, Ami11, AD16b, BP15, BMP12, CCZ13a, CCZ13b, FGH14, FPP18, GOC18, Haq17, HZL16, HL15, HWWZ16, IP14, JW18, JY14, JGPF17, LPK18, DMV17, NO10, OO12, ÖI10, Par11, RY13, She10, SK11, sS16, sS19, Shu12, SSK13, SAM13b, SAM13a, SU18, UA16, VM19, YZWM19, Zar17, ZP14, ZZXS17, ZHB18, ZWDM19, ZMW13]. Functional [KK18, LWW18, MG15, MS18b, AS15c, BCT16, BS19, HN13, JSM13, KM14, LAGM11, Mal16, Sha19a, SOH13, WKJ^{+}19, YBAA15, ZHH19, Zie11]. functional-coefficient [Zie11]. functions [AÖ13, AA16, BAKZ16, CR18, CPK19, CH15, Han17, KAWA12, LPV13, McL14, NJ18, PQ19, PL16b, RASR16, RCL15, SRG11, VVNTVD^{+}17, WN18, YBAA15, ZAK13]. fund [MV14]. Fused [CKP15, KK18]. future [AVK15, DDD17, RAJ16, VDBA14]. Fuzzy [GB17b, Saf13, JY14, MBG17, WTJW17, WyT17, YAEU13].

Gaining [RZ13b]. Gains [Bai16]. GAMLSS [CNL17]. Gamma [HP15, AAAQ19, AD12, BSS15, COBH11, CMX17, COS11, COP14, COP15, DTRB11, FKM13, GY16, GG17, HOC^{+}19, Hat16, Ili16, Kla15, LR18b, Nad10, NB15a, OCPC12, Özk19, PB17, PK11, QAA18, RB12, SB12a, Tza11, WBGJ15, WM12, ZWDM19, vdAvA15, VM19]. gamma-exponentiated [RB12]. gamma-linear [COP14]. gamma-Lomax [COP15]. GARCH [CH19, DGW10, FNRCM17, HKL17, JGPF17, LL11, Lee19, MRBR15, TC10, THR17, ZH12]. gas [Alw17]. Gaussian [AOH16, ABP16, Ami11, BB12, CL15, DY16, DW18, Fal16, GDCO18, GV12, JK19, JWWdL16, KR13, KAT15, KKE17, Kib12, KK12, LL11, LP10, MA10, MC13, NB13a, NB14, NLOT19, NJ18, PL15, RL14, SC12, TD14, TNJ17, VAW15, XSF17, Zha16, ZB10]. Gaussianity [LW15]. GEE [SMM19]. GEE-based [SMM19]. GEEs [PA14]. gene [PZY^{+}14, XWMA18]. gene-environment [XWMA18]. General [CNO13, CSPO14, NCO11, NCO12, NA13, STS14, Ahm19, Alw17, AJA11b, BPH12, BH19, GTB14, GA15, Her11, KL18a, KL12, Lem11b, MFP17, PK17, Sha10b, Tso16, WL19, WLL12, XHEM17, YLL17, ZLL19]. generalization [ATON18, Dia10, PL15, SN17]. generalizations [PF16b]. Generalized [AAH19, CCMGA14, CGP15, DKG16, GA15, KK15, LX14, LL17, QRH19, SSD17, Tsa10a, WS16, YL18, YS19, YFT10, AE11, AAR15, Aci18, ARY15, AG19, ADRA15, AHAA10, ACN^{+}17, AKAW15, BAJN14, BS13, BSSC10, BA15, Bor17, COBH11, CW16, CCZ13a, CCZ13b, CT14, COS11, CdC11, Cor13, CAO^{+}17, DSP15, DN13, EDASM17, ER19, EG18, FAV18, Fuk16, GdSCO14, GdCCDS18, Gun18, GV12, GWH14, GG17, HTZ^{+}16, Ili15, Kal14, KR13, KN15, Kiz18, KK13, KDG17, KR15b, LW12, LW17, LM18a, LXZ11, LZZM18, LV17, MSS14, MD18, NKB19, NFFM12, OY15, OCPC12, ODBT15, ÖKD17, Özk19, Pak10, Pak11, PB17, PA14, PPGT19, QMZ16, RPN15, RZ13b, SM15, SMM19, SL15, Sha16, SLM16, SZ16, SB11, TJK13, TKJ13, Tu19, VRC13, VSG^{+}18, VNM14, WM15, WCC13, WW16a]. generalized [WH17, Wan18b, WY19, Wes16, ZP14, ZPL16, ZHB18, ZST15]. generalized-least-squares [Fuk16]. generalized-order [BA15]. generally [CHB18]. generated [HOC^{+}19, ZB10, Zör15]. Generating [Ber12, OS14, SLV18, BP15, HBC11, McF16, McL14, SP16]. Generation [YD16, AD15, AD16a, PSS15, dABS16]. genes [LM16]. Genetic [AKU11, RB18, SD15, TKÖ19, VVNTVD^{+}17]. genetics [CVL18]. genome [XZY13]. genome-wide [XZY13]. Geodesic [HLVRS18]. geographic [AVG18]. geographical [FGH14]. geographically [WL17]. geometric [AA16, BZAZ15, BSdMC11, BBT13, BBM18, CW16, KAT15, NCO15, ÖI10, Pak10, Pak11, Sha19o, TLRB14, UA16, WSLX17, XZZ15, AKA^{+}16]. geometric-exponential [WSLX17]. GEV [GOC18]. Gibbs [BTR16, CI15, CK14, DRB17, DSVY14, KK11, SDC12]. given [BP15, JO12]. Global [Da 15, BCO13, CTX18, SZS16, WYZK16, WC17, XG11]. Globally [Gau11]. GMM [BZF18]. GOF [DY16]. goft [GEV18]. gold [RZ12]. Gompertz [Bor17, GAB14, Ism10, KA13, LTJB18]. good [Sha11e]. Goodman [vdAvA15]. Goodness [AE11, Nou17b, PB13a, QHB15, VK14, ZA11, ZX14, AOH16, Ali14, AS15b, BMP12, DS15, DC19, FH16, GEV18, JV14, KK12, LK17, LPK18, MA10, NA13, Nou19, RDC10, SJN15, Ten19, UY12, WD16, YS18, Zör15]. Goodness-of-fit [Nou17b, PB13a, QHB15, VK14, ZA11, ZX14, AOH16, Ali14, AS15b, BMP12, DS15, DC19, JV14, KK12, LK17, LPK18, MA10, NA13, RDC10, SJN15, Ten19, UY12, WD16, YS18, Zör15]. GQL [SS13, SKJ17]. Gra [LV17]. graded [FTS10]. gradient [BOG15, HSJ18, LLQ^{+}16, Seo11, Yua15]. gradient-based [Seo11]. Gram [Kia10, LWW18]. Granger [LL15a]. graph [FR15, GDR12, LCN^{+}17]. Graphical [SEGMA19, HT12, MC13, SH19, VAW15]. graphs [OCT18, Sha15e, TZSP19]. green [YPAC11]. Griddy [DRB17]. Group [ZX12, Are12, AJA11b, AYJ11, AWJ^{+}13, ABJR13a, ABJR13b, BS16, CA19, DW19a, FWZ^{+}15, KLKK19, LOK16, NJ13, NS17, PMP14, RR13a, YXZ16]. group-level [NJ13]. grouped [CM19, LYQ^{+}15]. grouping [DW19b]. groups [CC17a, Jin15, LLM16, SGÇ18, TQP10, WEHCC14, Wil15]. growth [FHO15, Sha10c]. Gumbel [UY12, PF16b].

H [Ano14c]. Haghighi [TCA^{+}18]. half [Aci18, ARY15, AG19, BS13, CSPO14, CAM16, ER19, GA15, MD18, RT14, Wan18b]. half-logistic [BS13, CAM16, RT14]. half-normal [Aci18, ARY15, AG19, CSPO14, ER19, GA15, MD18, Wan18b]. Hammerstein [LJ18]. Hamming [AP12]. Handbook [Sha15c]. handle [Ism16]. Handling [DS18, NMOV18]. hard [HR12]. harmonics [LLBL14]. Hastings [Lia10]. Hawkes [KB18]. hazard [AAGV12, Bag11, BCL18, BC19, BBA15, CPK19, CHQ17, Kiz17, Kuk19, PQ19, RCL15, SBK13, UGMK13]. Hazards [CLC17, Aus18, CZX18, DW17, HU14, OS14, PL18, sS11, She11, She14b, WDY18]. HDDA [SML19, MZZ^{+}19]. Health [Sha19j, Con10, DSVY14, MZZ^{+}19, Sha19h, Sha19r]. heavy [CG15, CCZJ17, DYA19, LPSR16, SL17, TdC18]. heavy-tailed [CG15, CCZJ17, DYA19, SL17, TdC18]. hedge [MV14]. Hellinger [BMK14, WZ18]. Heston [CDG^{+}15, FM19]. heterogeneity [Bør15, LL19, ZPL16]. heterogeneous [ABM17, GGSS19, KLKK19, LS14, LC18, RPFOMGRM17]. Heteroscedastic [DYA19, BPH12, CCZJ17, CCC10, EFGMD13, FHO15, HHC15, JLM15, SCW16, TdC18, VFLR10]. heteroscedasticity [ASM19, Are12, DDDD10, KKE17, KL18a, KL10, Wu16b, XMC^{+}14]. heteroskedastic [Ahm16, CT16, CNP19, MFP14, SM15, SL15]. heteroskedasticity [LZZW17, SDWL17]. heteroskedasticity-consistent [LZZW17]. heuristic [DI11, GHRAM13, Ism16]. Hidden [MC16, DLZ19, GB17a, LP13, LL10, Mar14b, TNM17, Wu16c]. Hierarchical [BKJ16, Fu16, NLOT19, TS14, WXF11, ABA12, GHH17, Han17, KNM^{+}15, Kiz17, MWL14, NLK11, SF12, Yu15]. Hierarchically [KBJ16]. hierarchies [CP14]. High [CCF19, CAC17, CLC19, AvR15, AAG^{+}19, AB13, AAY19, BTD18, Che18, DLS18, DDD17, HM18b, HH15b, KOM19, LKM^{+}15, LL16, LY18, LC18, Par17, QLW16, RY13, RT19, SS12a, SHLT17, SWXJ18, SZJW19, SE11, SH19, SH10, TS16, WLCL18, WLSZ18, Wu13, XX15, Xu17a, Xu17b, YWZ18, YXZ19, ZFZQ18]. High-dimensional [CAC17, CLC19, AvR15, AB13, AAY19, BTD18, DDD17, HM18b, HH15b, KOM19, LKM^{+}15, LL16, LY18, Par17, QLW16, RY13, RT19, SZJW19, SE11, TS16, WLCL18, WLSZ18, Wu13, XX15, Xu17a, Xu17b, YWZ18, YXZ19, ZFZQ18]. high-energy [SHLT17]. high-throughput [SH19]. Higher [CRV15, SWZ15, SAT16, DRAR19, DR17, ZS16b]. higher-order [DRAR19]. highlighting [SHLT17]. Hit [ZWCK16]. hitting [SC12]. HMM [RRDU13]. Hoeffding [FS13]. hold [SRP11]. homogeneity [CGSTG18, LYLM17, SK13]. homogeneous [CC17a, CM16, MMP12]. homoscedastic [MBKW19]. homoskedasticity [ES11, HA10]. homotopy [Gül10]. Horizon} [JKP19]. Horvitz [PP10, YS19]. Hosmer [LL18a]. Hotelling [AO12]. households [IGR13]. human [LMRW17]. Hurdle [CLAH17]. Hurst [SF13]. Hybrid [HBM16, Sun11, AHA15, AB16b, AVK15, CBG16, DPK11, Haq17, Ili15, Ism14, JP14, LY13, Lia14, LHB11, LH12, LHB13, LHLB19, LXL11, MSS14, Meh15, Pam19, PS16, PB12, Sha16, SK16, SZ16, TP15b, ZS16a]. hybrid-censored [AB16b, PS16]. hybridized [MPB19]. hyper [LY18, dSSCR19]. hyper-LASSO [LY18]. hyper-Poisson [dSSCR19]. hypercube [LLXY17, YLL17]. hypotheses [CRV15, GHMR19, Hwa11, MS17, QYX17, Say12, TS16]. Hypothesis [PBM16, dSSCR19, ASS11, AKV17, CNL17, DGK12, MFP17, Mur15, NJdC14, TPW17, Sha19j]. hysteretic [CT16, CTTS19].

IBM [Li14]. ideas [Sha10d]. identical [AR16b]. Identifiability [LS19]. Identification [CC11a, FR15, KC16, LJ18, Shi15, WL17, XZZ15]. identifiers [DS11]. identify [Car16]. Identifying [Dog15, PSY18]. identity [FHSC14, WLSZ18]. if [CM14]. IFR [AD10b]. ignorability [Yuc17]. ignorable [LLWY15, NJdC14]. ignorably [Kal14]. II [Ano14c, AHAH14, AHA15, AHH17, ARY15, ARY16, AG19, AB16b, AR16b, BDKM11, BS13, BZ16a, BB12, BBA15, ÇŞ18a, CM17, CPK19, CYL17, CBG16, DBC12, DNMT18, DT13, DC19, EDASM17, ER19, GAB14, GP15b, GWX14, HW17, IA10, KR18, KM12, KK13, KP18, NTZ19, NB15a, NB16, NLHD12, Nou17b, PB13a, PV17a, RT14, SKK12, SFS15, SBAA14, SBS14, Sha16, SBK13, SOH13, SK16, SAK14, TN16, Wu10, WY11, WLL12, WC14, WFC^{+}18, XYZ19, YCXN14]. II-censored [WLL12]. III [CGdS14]. ill [Kib12]. ill-conditioned [Kib12]. illustration [DRAR19]. imaging [KM14]. imbalanced [dlCHBCD19]. Immaculating [STL16]. Impact [FRL17, Yuc17, GZB13]. impacts [RL15]. Implementation [OO12, LZZ^{+}15, SS15, VNM14]. implementations [HTZ^{+}16]. implemented [JV14]. Implementing [WH14, Sha18a, VC15]. implication [OO12]. implications [GSL^{+}14, THR18]. importance [GZL18, WL16, WWB18, XSF17]. imprecise [HSC11]. improper [OS14]. improve [Wes16]. Improved [ABA16, BV15a, ÇK19, FP11b, HM18a, JP11, KP18, Lem11b, Lem11c, LX14, LL15b, MFP17, NAA17, SGG13, SY17b, WM15, WRP18, ZHH19, AH16, BZ14, BAT11, CHQ17, Haq14, MKSN18, WW19, WyT17, YGX14, ZCW^{+}17]. Improvement [BR16]. Improvements [GB18, DM16, RL15]. Improving [BSS17, CMCH12]. Imputation [OMY12, ASM^{+}11, AHJ16, Bru19, DS11, DG16, EKE^{+}18, GSL^{+}14, GG17, KKY15, SBMF18, WDCK15, XHYX14, Yuc17, vGK17]. imputations [JvBF13]. imputing [MSS18]. INAR [BV15a, BVRI16, BBM18, SM18, WJ19]. incidence [Lee17, Yu11]. inclusion [LB11, LN13]. Income [IGR13]. incompatibility [GB15]. incompatible [CI15]. incomplete [AP17, DSP15, HN11, KKY15, LP13, LNC17, OMY12]. inconsistent [STL16]. incorporates [MTS14]. increasing [Sha18c]. incremental [WW12]. independence [LL16, Med16, NBB15, NKZ19, SU18, ZZ15a]. independent [IRNB18, LCLP15, Sha15a, SWZ15, WEHCC14, WPC15]. Index [LX14, AWJ^{+}13, BU17, Bia15, BAZ19, GRPP10, Jer13, LM18b, NS17, PV17b, PWW15, PL18, Wan18a, WL10, WC14, WHX19, YZWM19, ZL15]. indices [CGP15, Cox13, HCY16, HAB12, NAA18, Per10, SGTKBL15, TP15a, XG11]. Indirect [SM15, Car19, GC17b, LGWZ18]. individual [NO10, NX18b, SG18, WL14, YPAC11]. individuals [MA17]. inequalities [HL13]. inequality [FS13, HCY16, ZY15]. Inference [AHA15, AD16b, AD12, ER19, FLB15, FKM13, IS13, Lee17, LHLB19, PS14, Say12, ABP16, Ame12, BS13, BZ16b, CCZJ17, CL14, CF17, CNS12, CNQ14, CNL17, CNP19, DSE16, DA14, DSP15, DF19, EKBO16, FP11b, FP14, GGM18, GC17b, GA15, GV12, GIDB15, Haq19a, HC17, HS13, Ism14, JK17, JYML13, KAT15, KP18, LRV17, LMFMA15, Lem16, LYZ11, LX16, LZZW17, LM18a, LZZM18, LVB18, MB16, Man15, MSS14, MKW16, PF16a, SN17, SBAA14, Sha10i, Sha18b, Sha19c, Sha19e, Sha19o, SK16, SWZ15, SZWL19, SZ16, SAT16, SB11, TK16, TP13, WL15c, WT18, WSC18, Wes16, YL18, YCXN14, Yuc17, Sha15i]. Inferences [AAS18, YML14, ZS18, ATH12, BBA15, CPK19, CT14, GZB13, Gun18, Lu19, SY17b, SR11, WRP18, WWCL11]. Inferential [SRK13, GB17a, GLB17, SSD17]. inferiority [Llo10]. infinite [DHP14]. inflated [AWH19, DSCS19, FI17, FBV18, HK18b, KNM^{+}15, KL17a, KR15a, Li11, LLY19, LT16, LWT^{+}17, MH16, MGR15, PCN14, PCMMA13, SB15, SMM19, SY19, ZCL19]. Inflation [SGG18, BBM18, FS15b, GZT14, LWS19]. Influence [AAAQ19, LJL15, XLW10, DFT17, Ema16, IJL18, JTL18, dCOC16]. Influential [ZH12, CA19, LGWZ18]. Information [Sha19j, AH19b, AKU11, BS12, Cha17, DBC12, GHJC10, HK16a, HK18a, HKK^{+}16, HSC11, Kuk18, Lin14, MMK10, MFD16, MPB19, MKSN18, NA13, Nou19, Pam19, PB12, PSK14, PK17, Sha10h, XY16]. informative [KLK14, PS19a, SSK13]. inheritance [SS15]. initial [DI11]. INLA [TD14]. innovation [JGPF17]. innovations [DGW10, Haj19, KL17b, MN19, TZA10]. innovative [LZZ^{+}15]. input [DMV17, WL16]. inputs [CJS10, CL15]. insight [RZ13b]. inspection [WW19, WCW15]. instrumental [Adk12]. insurance [Sha13a, TKK19]. integer [HKL17, KL17b, KBM19, Lee19]. integer-valued [HKL17, KL17b, KBM19, Lee19]. Integrated [GI17, GMR15, HKST17, LPK18, Tsa18]. Integration [CKPS11, SKTC11, CGA10, Mar10, RRV13]. inter [KN15, KN16, NK15]. inter-record [KN15, KN16, NK15]. interaction [FR15, KKM16, XWMA18]. interactions [CKP16]. intercept [MJ16, RKV17]. intermittent [UM14]. interpretation [AAH19]. Interval [Bak14, BBW17, BEBG14, KBS11, RYS11, SP11, She10, WY19, WWW17, Wu10, AYR16, AH16, Are12, BZ16a, BAS17, Che14, DG16, DTZZ12, DYT10, DT13, DK17, GIDB15, HCY16, HWWZ16, KA13, KL18b, MADASAM11, NY16, dALNCdATdC11, She12, She14b, She15a, SWLZ15, Tsa10a, WYW12, WM12, WDY18, WFC^{+}18, WH19, XHYX14, YP10, YCD15, YGX14, YXZ16, ZMW13]. interval- [KA13]. Interval-Censored [KBS11, BBW17, Che14, DG16, GIDB15, She14b, She15a, WYW12, WDY18, WFC^{+}18, XHYX14, YCD15]. Interval-valued [WWW17, dALNCdATdC11, SWLZ15]. intervals [AR10, AB10, BA15, Bia15, BBC10, Bur14, DD12, Haq19b, HK18b, Hat16, Ili15, Ili16, KKE17, KL12, KL13a, LZ11, Moh17, PO14, PB19, QKY16, RAK16b, RAJ16, RNA17, RMS14, SFS15, SG18, Tar12, WM15, WN11c, WN12, WN13, WY11, XHEM17, You14, Zar17, Zha15, vdAvA15]. intervened [KS17c]. intraclass [XLB12]. intractable [HH17, Lia10]. Introduction [PL16a, Zha17, Sha19g, Sha19h]. invariant [BDKM11, CC12, DDJ16, NB13b, NB14]. Inverse [SDS16, AOH16, Ami11, BB12, DY16, DW18, GTB14, GDCO18, GV12, JK19, JZD18, KAT15, KKE17, KLK14, KK12, LP10, MA10, Moh17, Muh16, NB13a, NB14, NJ18, RAK16b, SH19, SC12, SAM13b, SAM13a, SAK14]. inversely [AR10]. inverted [DKG16, KK13, KDG17, MTSR19, SSK13, TCA^{+}18, WY19]. investigation [JYML13, Say12]. investing [FS15b]. investment [MV14]. involving [CKP16]. irregular [How13]. Irregularities [JR13]. irregularity [LX14]. IRT [ABA12]. ISBN [Sha15d, Sha15c, Sha15e, Sha15i, Sha19i, Sha19f]. ISBN-13 [Sha15d, Sha15c, Sha15e, Sha15i]. Ismail [Ano14c]. isomorphic [Sta10]. issue [Zha17]. issues [DS18, GB17a, GLB17]. Istanbul [YAEU13]. item [AB13, CCK13, GdCCDS18, MTS14, TA11]. iterated [NHGS14]. iterative [CP14, FFP16, HY16, MM13b, MSA12, ZZ15a]. IV [GB17a].

J [Pak11]. jackknifed [ÖK19]. Jacobi [MM13b, SB12a]. Jacobi-type [MM13b]. James [AKS^{+}15, WCK11]. Jeffreys [Ho12]. job [WTJW17]. Johnson [MG15]. Joint [AH18b, MBL15, PM10, QMZ16, YYG16, Bai16, BAKZ16, HZSA19, HNPB18, IJL18, KNM^{+}15, MA17, PQ19, PRMM12, She10, sS16, sS19, SZ16, WZX13, XZY13]. jointly [SBAA14]. Jolly [HS13, YPAC11]. judgement [Fre16]. judgment [Haq19a]. jump [LZZ^{+}15]. jumps [MS15]. Jun [Sha19i].

Kaplan [CC11a, Lee11, SKM14]. kappa [NJdC14, NdCOP15, RNA17]. Katz [KL17b]. Keep [Her11]. Kendall [SGZM14, SU18]. Kernel [How13, KAJB19, Ame12, ABA12, BBW17, BAKZ16, DL19, DW19b, GL16, GZL18, JW11, JP14, KC14, KPK^{+}13, LWZ17, TNS14, WZ19, ZAK13]. kernels [Lüt15]. key [PSS15]. kind [AKJ16]. Knot [SESY13, JKP19]. know [Sha19s]. known [JP17, MQR18, TCHS19, TdC18, WLL12]. Kolmogorov [Fre12, OO12, SY17a]. Korean [LLP^{+}14]. KPSS [KE10]. Kruskal [vdAvA15]. Kullback [DR17, NA13, Nou19]. Kumaraswamy [CM17, COS14, GdSCO14, KN16, Lem11c, NKP14, NCO12, POCdP13]. Kumaraswamy- [NCO12]. kurtosis [Bur14].

Label [Yao15]. labelled [BG11a, ZB13]. labelling [YL14]. labels [YL14]. laboratory [DS11]. lack [Li11, LLY19]. lack-of-fit [Li11, LLY19]. lag [DRLP14, ÖK17]. Lagrangian [HLVRS18]. Laguerre [SB12a]. lambda [WW16a]. Lanchester [KMS17]. landings [PSY18]. Langevin [SLM16]. Laplace [GI17, AOH16, BZ16a, BOG15, DYA19, GMR15, NP16, TZA10]. Laplacian [Fan19]. large [Ahm19, CC19, Cha15, FWF16, HKK^{+}16, Ism16, JKP19, LL18a, LCN^{+}17, LXL17, PF16a, RGNM13, RWCD17, SHLT17, ZCW^{+}17, ZF16]. large-scale [Ism16, JKP19, LCN^{+}17, RGNM13, SHLT17]. largely [PS19a]. largest [KS16]. Laspeyres [Bia15]. Lasso [AY14, FWZ^{+}15, Wu13, AP19, CZX18, DW19a, HM18b, KK18, YW14, FP17, Fu16, LY18, LPSR16, WKJ^{+}19, BCLM17, NKB19]. Lasso-mixed [AY14]. lasso-type [AP19]. Latent [RBC^{+}15, ARY16, Lin14, Sha10c, VNM14, YYW15]. latin [LLXY17, YLL17]. lattice [Alw17, SM11]. law [MA10, WL15c, Zör15]. laws [HCY16, RS14, SA15]. lead [BS16]. lead-in [BS16]. lean [Sha19g]. Learning [TZSP19, Cha16, DW19b, KKL^{+}15, LMRW17, LL18b, Sha18b, Tu19, YZ14]. Least [AR16a, RA17, ATF19, CKP15, Fuk11, Fuk16, JR16, KS16, LJ18, Mal16, MSA12, NHA18, RBK16, SESY13, SML19]. least-square [LJ18]. least-squares [SESY13]. Least-trimmed [RA17]. left [CHT16, DN13, FM15b, KA13, She10, She14b, She14c, She15a]. left-censored [FM15b, KA13]. left-truncated [CHT16, She14b, She14c, She15a]. Lehmann [ODBT15]. Leibler [DR17, NA13, Nou19]. Lemeshow [LL18a]. length [FH15, LZGW14, SHST13, TGL12b, WN11c, XZ19]. length-biased [FH15, XZ19]. lengths [CCP12, DRLP14, Moi17a]. Leone [Akg19, BCY^{+}17, Gen13, Sha18c]. leptokurtic [AD12]. level [ATF19, DRAR19, FWZ^{+}15, KPM16, MS19, NJ13, SNGMRC16, Sta10, TC10]. levels [Kuk18]. Levenberg [GVW17]. Levene [Mar11]. leverages [ZR14]. Libby [RPN15]. library [OO12]. life [AS15a, Ano14c, AJA11a, AJA11b, AKJ16, ABA16, BSC14, DYT10, DT13, HW17, Ism10, IA10, IAGEK11, Ism14, LSL10, LHLB19, LXL11, LSF^{+}17, NB16, WGC14, WBGJ15, WBG15, WSLX17, WH19, ZS16a, Lia11, Sha19f]. lifetime [AHH17, ASH16, AS12, BCL18, CLDB16, COS11, COL14, EXH16, GdSCO14, GIDB15, HC17, HSC11, LHLB19, NJR13, NSR14, Nou17b, PB17, PS14, ROCH16, RN14, WC14, Zar17]. lifetimes [CSJ17]. Likelihood [ABP16, BG11a, GTB14, HC17, Ism14, ZCL19, ZGW14, AAVG16, BTR16, BZ16a, BSS15, BS19, BZ16b, Bor17, BOG15, CH19, CS16, CS17, Che19, Ciu13, CL10, CCC10, DL19, DFPT16b, DGK12, EFGMD13, FHS12, FP11b, FP14, GAB14, GV18, HKK^{+}16, Ho16, HC10, JW11, Kha12, KB18, LP13, Lem11b, Li18, LZ10, MM13b, MB16, MMP15, MD18, MFP14, NBB15, NLK11, NLOT19, PCN14, RGNM13, RL14, SJN15, Saz19, SGG13, SL15, SY17b, SZS16, SDWL17, SLLZ18, SZ16, SAK14, SB11, TW14, Tsa10b, Tso16, VNM14, WXF11, WPXL14, XZ19, YS13, Yos18, Zha16, ZST15, ZY15]. Likelihood-based [ABP16, ZCL19, BZ16a, DL19, SDWL17]. likelihoods [GVW17, LPV13]. Likert [MTS14]. Likert-type [MTS14]. limit [PPK16]. limited [WH19, ZST15]. limits [DH14, Gau10, KP15, NHGS14]. Lindley [AAR15, AAS18, Ali15, GSW17, GDCO11, HBFSGD11, OY15, QKY16, dESM15, SAT16]. line [FF14, FG13, QLW16]. Linear [Jou15a, LS11, TKB16, ZY15, AÖ16, AJFB14, Ali12, AR16a, AKAW15, ATH12, AB18, AGA18, AW17, AL18, AH18b, Bag11, BCY16, BDKM11, CL11, CCMGA14, CT14, COP14, CNP19, DSP15, DK10, FS10, GG17, HL13, HN13, HY16, HXT15, HTZ^{+}16, HKST17, HLL18, JR13, Kal14, KAK16, Kib12, KF16, KL12, Kuk18, KÖ19, LC19, LW12, LP16b, LGW16, LWW18, Man15, MC16, Moi17b, MS19, MMP12, MC14, NX18b, NKB19, ÖK18, ÖK19, Özk19, PS18, PZ10, PMP14, PAFPM12, PL18, QMZ16, RY13, RKV17, RB17, RAN11, RZ13b, SGG18, SRP11, SWW19, ST13, SD15, Say12, Sha10b, sS12b, Shu12, SU11, SS13, SB11, THG15, TTWT19, Tsa10a, VRC13, VNM14, WM15, WL15b, WL15a, WW16b, Wan18a, WL19, WKJ^{+}19, WBAS15, WWW17, Wel16, Wu16a, XY11, Yu15, ZR14]. linear [ZC13, ZZ17, ZST15, dCFOM12]. linear-by-linear [Ali12]. linearity [Li15a]. linearly [MM13b, MN15]. lines [HP11, Wil10]. lineups [PDS16]. linex [IP14, NO10, ASBM19]. Link [Sha19f, NMPR14, QHB15, YZWM19]. linked [GGM18]. links [dlCHBCD19]. Liu [AD10a, ADRA15, KAA18, Kib12, Mån13, QAA18, Şir19, TKG18, Wu16a]. Liu-type [AD10a]. LMARS [AP19]. load [DLZ19]. Local [AGM15, Bag11, CWM17, Ema16, JO12, KBS11, XZ19, Alt19, BAKZ16, CL16, DFPT16b, DKY17, FMK15, JTL18, WMDO11]. Locality [LMRW17]. Locally [Che14, SNG12, CJ13]. location [Che11, GA15, KL18a, LMSX16, LCB13, LHLB19, LP16b, MADASAM11, NB13b, NB14, PB13a, PL16b, Saz19, SRAO11, SC19, Wil15, WZX13, Wu16b, WZ18, XHEM17, Zha14, ZX14, ZBG18]. location-based [LP16b]. location-scale [Che11, KL18a, PB13a, ZX14]. location-shifted [WZ18]. locations [DC15, PS19b, SK17]. Log [TNJ17, BCJG12, BCY^{+}17, CMC13, CAO^{+}17, DPK11, GRPP10, GGAM13, HK16b, KF16, LDCL17, Lem12, LCB13, LHLB19, Lon12, MMP12, OWLP16, OCPC12, PP15, QHB15, RAK16b, SH19, TD14, VP16, XHEM17, ZZXS17, ZBG18, dCOC16]. log-Birnbaum [LDCL17, Lem12]. log-excesses [GRPP10]. log-exponential [BCJG12]. log-exponentiated [OCPC12]. Log-Gaussian [TNJ17, TD14]. log-linear [KF16, MMP12]. log-link [QHB15]. log-location-scale [LCB13, LHLB19, ZBG18]. log-logistic [BCY^{+}17, CAO^{+}17, OWLP16, RAK16b, dCOC16]. log-normal [CMC13, DPK11, GGAM13, Lon12, SH19]. log-odd [dCOC16]. log-rank [PP15]. log-skew-normal [HK16b]. log-symmetric [VP16]. Logarithmic [AZS15, KR15a, MP16]. logistic [AL18, BS13, BAB15, BCY^{+}17, CL13, CAM16, CAO^{+}17, CA19, EKE^{+}18, FH16, GI17, HMP17, HAH14, KK18, KSLN^{+}18, LJV^{+}18, LYQ^{+}15, LY18, LSA16, MJ16, MS18b, MADASAM11, OWLP16, PK16a, PT14, RAK16b, RT14, SCL^{+}18, SL17, SJ10, TK18, WZ13, dCOC16]. logit [DGLV17, GC17b, GMR15, SBD10, ZS18]. loglogistic [GIDB15]. lognormal [BET16, DDZ13, HK18b, LW17, NB13b, SRG11, STW15, WM12, WN18, WN13]. Lomax [AZS15, COP15, GGAM10, HSR15]. Lomax-Logarithmic [AZS15]. Long [CGAP19, AA11, BBC10, Bor17, COBH11, FPP18, Her11, JK17, Kuk19, Kus11, MAEP14, PL15, RL14, RTM18, Sha10k, SB12b]. long-memory [AA11, RL14, SB12b]. long-range [PL15]. long-term [Bor17, COBH11, Kuk19]. longest [Bak18]. Longitudinal [HLN^{+}15, UM14, AI16, BG11b, CWM17, DFT17, FW15, Fan17, Hua16, HNPB18, LH18, LLT12, Mar14b, MBL15, NMOV18, PQ19, QMZ16, QZZ16, Sha10c, SJ10, SR11, TX14, WL15a, YLG15, YYG16, ZP14, ZPL16, ZFZQ18]. Looking [CC19]. loop [MCW17]. Lorenz [SSD17]. loss [ASBM19, AYJ11, Han17, HL15, IP14, NO10, RASR16, WW19, WCW15, YZL17, ZZXS17, ZWDM19]. losses [KC16, NdCOP15]. lot [ABJR13a, ABJR13b, NS17]. low [BTD18, JSM13, SH10]. low-dimensional [JSM13]. low-sample [SH10]. lower [Gha19, KS17a, KL17c, WEHCC14]. LR [BD12]. LRT [MMP12]. LSTAR [LS11]. Ltd [Sha19f].

M [Ano14c]. machine [KOM19, KKL^{+}15, LYL17, Tu19, HZL16]. machines [CKPS11, Erd13]. macro [Ahm16, BBV17, LV17]. macro-financial [Ahm16]. Macros [NB18]. magnetic [KM14]. magnitude [CMX17]. Mahalanobis [LGWZ18, SG18]. maintenance [AD16b]. majorant [BF12]. majorants [JR17]. majorization [YZ14]. make [ATH12]. making [Sha19r, WRP18]. Malliavin [JK14]. malnutrition [DRAR19]. Management [Sha19j]. Mann [AS15c]. MANOVA [KL10, RR13a]. manufacturer [MCW17]. MAR [JvBF13]. Marginal [VNM14, AI16, FHS12, GDPH12, PQ14, RPFOMGRM17, TK14]. marginal-likelihood [FHS12]. marginally [AI16]. marginals [AD15, JO12, PNN17]. mark [YP10, YPAC11]. marked [HEB13]. market [Cox13, YAEU13, YTNT14]. markets [CGAP19, GC17a]. Markov [CDG^{+}15, CCK13, DRB17, DLZ19, FP15b, Haj19, LP13, LL10, LZZ^{+}15, LED16, Mar14b, MC16, TNM17, UG10, Wu16c, ZB13]. Markov-chain-based [ZB13]. Markowitz [CLC19]. Marquardt [GVW17]. MARS [AP19]. Marshall [ASH16, DNMT18, GGAM10, LLGP17, MS11, dESM15]. Marshall-Olkin [DNMT18]. masked [AH18b, WYW12]. masking [WYW12]. mass [BAKZ16, ZB13]. massive [CP14]. matching [BBV17]. mathematical [LAGM11]. mathematics [Sha19i]. MATLAB [Wan15]. matrices [AvR15, Cha15, FWF16, HLVRS18, HM19, PGT11, Wes16, YBAA15]. matrix [APF18, CA12, CLC19, GGG19, Ho16, Lem11a, LZZW17, MM13b, McC14, VS15, WLSZ18, Xu17b]. matrix-valued [APF18]. max [CS17]. max-stable [CS17]. maximal [BF12, WD16, YR15]. maximization [BOG15, GVW17, MBL15, SWLZ15]. Maximum [LP13, Lee11, MM13b, MAV17, RGNM13, SO10, Tsa10b, Yos18, AAVG16, CH19, CCC10, HKK^{+}16, Ho16, HR12, JW11, Jin15, KB18, Lem11b, LR18b, MB16, MD18, PPK16, RL14, Saz19, SGG13, SL15, SAK14, WDCK15, WRN18]. Maximum-likelihood [RGNM13, Lem11b, SGG13]. maximum/minimum [MB16]. Maxwell [COBH11, SY19, HOR17, KM12, KVK15, TP15b]. MCMC [BMK14, DR17, GGM18, Hol19, PD13, TD14]. MDEWMA [AM13]. MDIC [MMK10, MK12]. MDS [AAJ16]. Mean [FH15, FMB16, AF17, AH19a, AH19b, AHM13, Ahm19, AO11, AOR13, AOH19, Ak19, AWJ^{+}13, AH18a, AA16, CL16, FBV18, Fre12, Fre16, Gau11, Haq14, HK16a, HM18a, HK18a, HK18b, IJL18, JZZH18, KTJ18, LAuaS^{+}15, LP17, LW14, LLBL14, MMR16, NY16, OWKC15, PSKC18, PJR15, QMZ16, RR13a, SGTKBL15, SRP11, SGÇ18, SRAO11, STS14, SLLZ18, Sun11, TS16, TK15, WN11c, Yam19, YL15, YS19, Zar17, ZL11, ZMW13]. mean-covariance [QMZ16]. Mean-shift [FMB16]. mean-variance [PJR15]. means [KKE17, KL13a, PA15, PQ14, RRB10, RBHSL11, TPW17, WM12, WN11b, YP17]. measles [YWL18]. Measure [LP16b, AB11, BS12, NdCOP15, Wil15, WWB18]. Measurement [SZS14, CCZJ17, CB10, GZB13, GM16, HBMAO15, HAC16, HHC15, IJL18, Lüt15, MFP14, NAA18, PF16a, QZZ16, STL16, Şir19, TdC18, Tsa10a, Wu11, ZR14, ZCZ17, ZLZZ18, ZLL19, ZY19]. measurements [DH14, LC10, MTS14, MC14]. measures [BS14, Da 15, DM16, GZL18, GB15, HTZ^{+}16, IJL18, MH16, PC10, QKUH18, TA11, VK14, Wei11]. Measuring [Jon16, Sha19s]. mechanism [JY13, LPY18]. median [AOR13, CC16b, HMP17, HCT19, MAEP14, MAAM10, RS15, TCHS19, WDCK15, YZL17, ZZ15b]. mediation [SCL^{+}18]. mediators [SCL^{+}18]. medical [CCF19, Sha19b]. medicine [Sha19r]. medium [Her11]. Mehl [MG15]. Meier [CC11a, Lee11, SKM14]. membership [GC17b]. memetic [WTJW17]. memory [AA11, BBC10, CGAP19, FPP18, JK17, KKL^{+}15, Kus11, RL14, Sha10k, SB12b]. Merging [SLA17, CC17a]. Merton [Guo17]. meshes [SHLT17]. meta [PB15, WT18]. meta-analysis [PB15]. method [AOH19, ATH12, BZ14, BCT16, BCO13, CHTZ14, Che11, Coi13, Dia10, EZ12, Fan19, FR15, GL16, GGM18, Gül10, GA15, JL16, KKM16, Kia10, LX14, LYL17, LJ18, LLP^{+}14, LGWZ18, Mån13, MKL13, MADASAM11, NB13a, NB14, NHWT14, OvP10, ST13, Sha19a, TPM17, Tso15, Tso16, WyT17, XY16, YP10, YPAC11, YL14, ZCW^{+}17, dLHT17]. Methodology [Jer13, BS19, FGHRM12, JP14, MSS18, PS18, RAN11, Sha10l]. methods [AY13, BP15, Car19, CMCH12, CGP15, CHT16, DS11, DL13, DAP15, GSL^{+}14, GD19, HT12, dlCHBCD19, IGR13, JMM^{+}17, Jou15a, KF16, KEW13, KK11, LW14, LXL17, LYLM17, LED16, MS18b, MMT16, NPJ14, ÖK19, PQ14, PB15, RY13, RKV17, RNSR19, SNGMRC16, dESM15, SJZ17, Sha19h, SBMF18, TCA^{+}18, TNJ17, Tsa11, TRC^{+}18, VKK14, WN18, WRP18, Xie14, YS13, Zha16, ZNW19]. Metropolis [KK15, Lia10, VNM14, ZWCK16]. MGARCH [CQJ12]. MHDR [BH19]. microarray [LDB10, PZY^{+}14, RY13]. microRNA [SH19]. Microsoft [OO12]. microwave [RB17]. middle [ARY17]. midpoint [AH16, YGX14]. military [KMS17]. mines [ASM17]. Minimal [WD16]. Minimal-maximal [WD16]. minimization [PM10]. minimizing [DRAR19, LM18b, YL14]. Minimum [Adr18, WZ18, CC11a, CTX18, DI11, GB18, Kim18, LL19, MB16]. mining [PZY^{+}14]. Minitab [Sha19e]. minorants [JR17]. misclassification [NMPR14, RYS11, SXTJ17, WL14, ZY19]. misclassified [CB10, RZ13a]. Mises [Chr15, CL10, FM15a]. mismeasured [Hua16]. Missing [VKK14, ASM^{+}11, CH19, Ciu13, CYPGGM16, DK17, DM12, FVB13, FTS10, FHO15, HKK^{+}16, HLN^{+}15, Hua16, JY13, Jon16, Kal14, KAR13, Lin14, LLWY15, NJdC14, NMOV18, PS19a, SLV18, Seo11, Shu12, SZWL19, SZM17, WL15b, YL15, ZL11]. missingness [JM10, UM14]. mission [WWB18]. misspecification [LCLP15]. misspecified [ES11, FRL17, NX18b, XWZS15]. misusing [Sha19f]. mitigating [JR18]. Mitochondrial [Sha19j]. Mixed [HBL14, AY14, AD15, AD16a, ATF19, AB18, BCY16, BV15b, CL11, CB10, DFT17, FHS12, FBV18, FHO15, GC17b, GMR15, HTZ^{+}16, Hua16, HKST17, IRNB18, JWWdL16, KP15, KÖ19, LC19, LW12, LLWY15, MBL15, Mog11, MS19, MC14, ÖK18, ÖK19, PS18, SRP11, SL17, SH16, SBD10, SB11, VHV^{+}16, WL15b, WSLX17, WBAS15, XLW10, ZR14, ZC13, ZY19, dCFOM12]. mixed-effect [MC14]. Mixed-effects [HBL14, ATF19, HTZ^{+}16, Hua16, LC19, LLWY15, MBL15, VHV^{+}16, XLW10, ZC13]. Mixing [Hol19]. Mixture [BCY16, GZB13, PBM16, AM19, AJM11, Akn13, Cha16, CCK13, DP15, Fal16, GDR12, GY16, GYV^{+}13, GOC18, HZSA19, HEB13, JM10, LL11, LC19, LNC17, LZZ^{+}15, MA17, PJR15, SPS19, SL15, SAM13b, SAM13a, TS17, TGL12a, Tsa18, WPXL14, WH14, WW16b, WYX17, WZ18, YL14, Yao15, ZF16]. mixtures [DYA19, FS10, FLB15, FMB16, GLLO14, GdCCDS18, JR18, LDCL17, MAV17, MARS19]. MLEs [CBG16, TN16]. mobile [DRC^{+}16]. modal [YLG15]. mode [vGK17]. Model [CL13, Che18, CYPGGM16, FP15b, HZ14, HAH14, KÖ19, LP16a, Lin14, LT16, LSA16, LZZ^{+}15, MMK14, PBM16, RT19, SM11, XWZS15, AHAH14, AHA15, AP19, AG19, AHAM15, Ahm16, Akg19, AAH19, AL15, Alt19, Alw17, Ame12, AS12, AAAQ19, AKAW15, ATF19, AAGV12, Aus18, AAY19, ABA12, BR16, BL19, BV16, BCL18, BAS17, BAZ19, BBL13, Bor17, Bør15, CG15, CB11a, COBH11, CLDB16, CCZJ17, CJS10, CW16, CL14, CT16, CC17b, CTTS19, CA12, CCK13, CH14, Con10, CGdS14, CSPO14, CYRO18, DW17, DF19, DW18, DBL10, DDD17, EKBO16, FOC14, FM19, FW15, FI17, FHS12, FF14, FRL17, FBV18, FMB16, FKM13, FTS10, GBdL16, GGdC17, GGdC18, GK16, GY16, GB17a, GdSCO14, GGA18, GdCCDS18, GG16, Guo17, GV12, GG17, GIDB15, Haj19, HZSA19, HRR^{+}17, HBFSGD11, Hir11]. model [HR12, HSC11, HW12, HY14, HS13, HKST17, HHC15, IRNB18, Ili15, Ism14, JP17, JY14, JYML13, Juh16, JKLR11, Kal14, KNM^{+}15, Kha12, KA13, Kib12, KLKK19, Kiz17, KBM19, KL18b, KP15, Kuk19, LJV^{+}18, LKM^{+}15, LL11, LCLP15, LS19, Lem11b, Lem12, Lem13, LS11, LWZ17, LXL17, LJ18, LCZ18, LTJB18, LJL15, LZ10, Mån13, MSS14, MJ16, MC13, MC16, MS15, MFP17, MTS14, MPB19, Moi17a, MG15, NMS18, NX18a, NX18b, NTC11, OWLP16, OCPC12, ÖK18, PSV11, PB17, Par11, PO14, PRMM12, PD13, PL18, PT14, QAA18, QMZ16, RGNM13, Rai12, RYS11, RBC^{+}15, SSMF18, Saf13, SRP11, SWW19, dSSCR19, SPS19, SMM19, SGGC10, STL16, sS11, sS12b, SCW16, SY17b, Shi15, SM18, SYL^{+}14, SH19, SRK13, TK18, TX14, TTWT19, TKG18, TdC18, TGL12a, Tsa10a, Tsi15, VBL17, WXF11]. model [WYW12, WZS17, WC17, Wan18a, WBAS15, WWW17, Wu16a, Wu16c, WYX17, WZ18, XYZ19, XZY13, XZD15, XLMX19, YCD15, YYW15, YHWS18, YZWM19, ZZXS17, ZLQ^{+}17, ZF11, ZB13, dCOC16]. Model-averaged [FP15b]. Model-based [LP16a, SM11, AKAW15, DF19, LXL17]. Model-free [Che18, LWZ17, SGGC10]. modeling [AH18b, Sha15c, Sha15i, Sha19m]. Modelling [DRC^{+}16, GY16, HEB13, KSJ16, MCW17, MFD16, MR17, WCC13, ZPL16, BAB15, CS18b, GHJC10, GG16, HNPB18, JY13, LS14, LLV^{+}14, MBG17, MMK14, MBL15, MWL14, Mug16, PK16a, Par17, PQ19, RPFOMGRM17, SPS19, Sha15i, Sha19n, Stå16, YZWM19, dCOC16]. models [AM19, AHM13, AÖ16, AHC15, AD10a, ADRA15, AWH19, AAH19, AB16a, AACR18, ABGM18, ATON18, AR16a, ASB14, AB18, AI16, AÖ13, AL18, AY18, BKR17, BT14, BCY16, BZF18, BKJ16, BG11a, BSS17, BBA15, BMAW14, BCCN18, BG11b, BK17b, BEBG14, BV15a, BH18, BV15b, CL11, CDG^{+}15, CSJ15, CH19, CCMGA14, ÇK19, CCM12, Cha16, CQJ12, CPK19, CLH14, CZX18, CT14, CWM17, Ciu13, CAO^{+}17, CA19, CNL17, CCC10, DRAR19, DSP15, DP15, DY16, DHP14, DRLP14, DGW10, DP13, DFT17, DA13, Ema16, FH16, Fam12, FW13, FW15, FP15a, FP14, FLB15, FNRCM17, FAV18, FPP18, FVB13, FHO15, GLLO14, GDR12, GYV^{+}13, Gha19, GB18, GM16, GC17b, GMR15, GI17, GHMR19, HP15, HOC^{+}19, HM13, HMP17, HA10, HA13, HAC16, HY16, HLN^{+}15, HTZ^{+}16, Hua16]. models [HKL17, HKKL17, JM10, JKP19, JvBF13, Jou15a, JTL18, JLM15, Kal14, Kan17, KKY15, KK14, KL17a, KL18a, KF16, KSLN^{+}18, KL12, KP15, Kum15, KÖ19, LH18, LDCL17, LP13, LY15, LL10, LC19, Lee19, Lem11a, LCM12, Li11, LYZ11, LW12, Li15a, LM18a, LLY19, Lia10, LNC17, LdNdSF18, Lin14, LSA16, LLWY15, LZZ^{+}15, LGW16, LWT^{+}17, LZZM18, LVB18, LM18b, LLT12, LK13, MCZ17, MMP19, MBKW19, Man15, MMK10, MRBR15, MH16, Mar14b, MC13, MBL15, MS17, ME15, MFP14, MKW16, MT17, MGR15, Mog11, Moi17b, MS19, MMP12, MC14, NJR13, dALNCdATdC11, NKB19, NLK11, NHA18, OJRACL18, OS14, ÖK17, ÖK19, Özk19, Pam19, PS18, PV17a, PZ10, PMP14, PAFPM12, PB19, QMZ16, QHB15, RKV17, RS17, RAN11, RBK16, RA17, RN18, RIY18, RZ13b, SM15, SL17, SH16, Say12, Seo11]. models [SL15, SF12, Sha10b, She14a, She15b, SK18, Şir19, SB12b, SC12, Su16, SZM17, SS13, SO10, SBD10, SJ10, SR11, SB11, SGB13, THG15, THCZ18, TNM17, TZA10, TC10, Tsi15, UGMK13, UM14, VHV^{+}16, VRC13, VP16, VNM14, VAW15, WM15, WDCK15, WL15b, WL15a, WL17, WDY18, WZX13, WHX19, XLW10, YK15, YLG15, YYG16, Yao15, YML14, YPL13, Yu15, YD16, YA16b, ZRH15, ZR14, ZMKAV19, ZCL19, ZH12, ZC13, ZZ17, ZLL19, ZL15, ZB10, ZS18, ZS16b, ZF16, Zie11, dCFOM12, dLHT17, vGK17, Sha15i]. moderate [SHP12]. Modern [MB16, Sha19t]. modification [LSA^{+}15]. modification-based [LSA^{+}15]. Modified [AY18, JZZH18, KR16, MFP14, Mur12, PCN14, PT14, SRAO11, VVNTVD^{+}17, AO11, AAH19, AH16, AAVG16, ASA^{+}17, Bot11, CC16b, CLAH17, COS14, CB13, DS14, DS15, DYX15, GSW17, GD19, HMP17, HNPB18, KS17a, Kia10, KS17c, LN13, McF16, NCO11, Saz19, SK11, UG10, UGMK13, Wel16, YS19, YGX14, ZA11]. moment [CLS^{+}19, GSW17, McL14, RR13b, Yoo13]. moment-generating [McL14]. Moments [NJ18, BS13, BZAZ15, Cor13, NP16, PGF19, SAT16, WLL12, ZA11]. monitor [AM13, Haq14]. Monitoring [HKL17, KAK16, PV17a, Rak16a, WL19, AP15, AKAW15, AH18a, CMX17, CKPS11, DRC^{+}16, FAT19, Gau11, GWX14, HK16a, HM18a, HK18a, HP11, HLL18, KP19, LAuaS^{+}15, Li18, MMR16, MH17, NTC^{+}19, OWKC15, QLW16, RA14, SLLZ18, SKTC11, WW16b, YA16a, YZL17, ZST15]. monotone [JM10, MMT16, Shu12]. Monotonic [MH12, SGB13, Yam19, ZNW19]. monotonicity [BDKM11, Wu12]. Monte [LED16, AB16a, AB13, AB14, BZF18, Bør15, CDG^{+}15, Car16, CK14, CCK13, FP15b, GP15a, HH17, HLVRS18, KE10, LLC17, LZZ^{+}15, Meh15, Ned11, NA11b, NA11a, OSN17, PP10, PO14, Saf13, SJZ17, TPM17, UG10, WJ17, YW14]. Moran [LLN13]. Morgan [HCY16]. Morgenstern [UY12]. morphology [LAGM11]. Morris [FR15]. motion [FS15a, Guo17, KSM16, XZZ15]. motor [TKK19]. movement [LCN^{+}17]. Moving [AP15, AAH19, BT16a, BCCN18, CHB18, Haq14, HBMAO15, HL15, Jou15a, LPJ14, LCZ18, MMR16, Zha16]. moving-average [Jou15a]. mRMR [BCT16]. MSE [NX18b]. MSV [CQJ12]. MTM [ZBG18]. Muller [Dia10]. Multi [Cha14, DRAR19, KGA12, RAB14, UA16, WTJW17, XX15, AR10, DW19b, GB17b, PS19b, RWCD17, SC19, WyT17, WHX19]. multi-dimensional [DW19b]. multi-index [WHX19]. Multi-level [DRAR19]. Multi-objective [KGA12, RAB14, UA16, WTJW17, GB17b, WyT17]. multi-sample [AR10, PS19b, SC19]. multi-stage [RWCD17]. Multi-step [Cha14, XX15]. multiattribute [HWA^{+}14]. multicollinear [YA16b]. multicollinearity [AAH19, BKR17, BAT11, Jur12, RN18]. multicomponent [AG19, Akg19, Gun18]. multidimensional [Bai16, Cox13, FTS10, KC16, QX12]. multifactor [TQP10]. multigamma [MH12]. multilevel [AL18, GM16, GI17, Yuc17]. multimodal [AB10]. multimodality [NB18]. multinomial [AB15, HSC16, HAH14, RGNM13]. multipath [LPY18]. Multiple [Ahm19, AHJ16, DG16, GG17, WLCL18, YXZ19, AKU11, BKR17, BL19, BG11a, BCLM17, CSJ15, CA19, DS11, DDDD10, DYX15, DW19b, ER19, EKE^{+}18, FAT19, GC17b, Jin15, KLK17, KS17b, KK15, KL17c, Li15b, LLM16, LWW18, NMS18, NTG13, NJdC14, QYX17, SSMF18, SGG18, SA12, SFS15, SLA17, sS12b, SBMF18, SU11, THCZ18, TKJ13, TK16, WDCK15, Wil15, WRP18, WPC15, Wu16b, Xie14, XSF17, Yam19, YZ15, YYG16, Yuc17, Zha15, vGK17]. multiple-imputation [vGK17]. multiple-membership [GC17b]. multiple-point [KK15]. multiplicative [LLBL14]. multiplicity [Hwa11]. multipliers [JGPF17]. multiply [FM15b, SBS14, WY11, WLL12, WC14]. Multiresolution [MN15]. multisample [Mur12]. multistate [DDD17]. multivariable [ZCW^{+}17]. Multivariate [CF10, CS18b, KS17b, Lem13, LMFMA15, MARS19, PJR15, Sha11d, SC19, AJFB14, ABP16, APF18, ARB13, AAG^{+}19, AD15, AD16a, AI16, BCY16, BAKZ16, BS12, BS14, CCF19, CLDB16, CA12, CKPS11, DS11, DSCS19, Dog15, DC13, DC15, DC16, GSL^{+}14, GBCS16, GV16, GB17b, HT12, Haq19b, HZ14, HSJ18, JV14, KGA12, LNC17, LLWY15, Lue19, MM13a, MLCL18, MA17, MV14, MFP14, MFP17, Mur15, NTC^{+}19, Nom14, OvP10, PC11, PJ15, PS19b, PK16b, PL15, RAB14, RWCD17, SSMF18, SLSW15, SLV18, SFS15, SA15, SLD16, SR16, SKTC11, TD13, Tsa18, Tsa10b, UA16, VM19, WL15b, Wan15, WBG15, Wan18a, Yuc17, ZZ17]. Multivariate-multiple [KS17b]. municipal [YFT10]. mutation [WyT17]. Mutiple [GMG13]. Mutual [BS12]. MVRB [GHRAM13]. MWM [ZBG18]. myopia [FW15].

Nadarajah [TCA^{+}18]. Naim [Sha19f]. Naive [JL16]. Nakagami [SGG13]. natural [BS16, WL15c]. NBUE [AB14]. near [GB15]. nearest [KLK15]. necessary [BMPZ14]. needs [Con10, Sha19s]. negative [ÇK19, CLYX18, FI17, FDGD16, Fu16, GGdC17, GY16, HY14, KAJB19, KS17c, Li15a, LLY19, Mån13, MS17, Mog11, PA15, SP11, You14, YXZ16]. neighbour [KLK15]. neighbourhood [RdSF16]. neoteric [TdSC19]. Nested [GI17, TW14, CNL17, GMR15, HZSA19, Hof12, Kan17, XMC^{+}14]. Nesting [AB13]. net [MPB19, Pam19, PK16a, XX15]. net-type [PK16a]. Network [LPY18, AP19, KOM19, KBL^{+}15, SZS14]. networking [PSS15]. networks [FAT19, GP19, LH14, LPY18, LL18b, VC15, YFT10, dB15]. neural [GP19, VC15, YFT10, dB15]. neuro [JY14]. neuro-fuzzy [JY14]. Neyman [LL13, PSV11]. node [LH14]. Noise [LLBL14, GP19, RL14, WHX^{+}17]. noise-based [GP19]. noisy [RT19]. Non [CNL17, GGM18, KLK14, LH18, Med16, EB19, AO12, AR16b, BCJG12, BEBG14, CHTZ14, DFT17, FDGD16, GSL^{+}14, GYV^{+}13, HBT12, HY16, Hua16, IJL18, JM10, JG16, JYML13, Kal14, Kel16, KSJ16, LW15, LL15b, LLWY15, Llo10, MS17, MSS18, NJ13, NTG13, NLOT19, OSdVM13, PC11, PO14, PS19a, PPGT19, RAB14, RRB10, SRP11, SWXJ18, SSK13, Sta10, SKJ17, TK16, WBGJ15, WYZK16, WRN18, Wil10, WN11a, WN13, WFC^{+}18, XY16, Xu17b, Yuc17]. non-Bayesian [HBT12]. non-binary [GSL^{+}14]. non-central [PPGT19]. non-conventional [IJL18]. non-convex [SWXJ18]. non-crossing [EB19]. non-distinctness [Yuc17]. non-Gaussian [NLOT19]. non-Gaussianity [LW15]. non-identical [AR16b]. non-ignorability [Yuc17]. non-ignorable [LLWY15]. non-ignorably [Kal14]. non-inferiority [Llo10]. Non-informative [KLK14, PS19a, SSK13]. non-isomorphic [Sta10]. non-iterative [HY16]. non-monotone [JM10]. non-negative [FDGD16, MS17]. Non-nested [CNL17]. non-normal [BEBG14, DFT17, GYV^{+}13, Hua16, JYML13, LL15b, NTG13, PO14, RRB10, WN11a]. non-normality [AO12, Xu17b]. non-normally [TK16]. Non-parametric [GGM18, Med16, OSdVM13, SRP11, WYZK16, WRN18, Wil10, WN13, WFC^{+}18, XY16]. Non-penalty [LH18]. non-random [PC11]. non-regular [BCJG12]. non-repairable [JG16]. non-response [IJL18, MSS18, NJ13, RAB14]. non-standard [Kel16]. non-stationary [CHTZ14, KSJ16, SKJ17, WBGJ15]. nonconformities [KP19]. nonconvex [LOK16, SZJW19]. nonignorable [FTS10]. nonignorably [SZWL19]. noninferiority [Wel16]. Noninformative [KLKK19]. Nonlinear [ZLL19, BPH12, BV15b, Ciu13, CCC10, EKBO16, FG13, GLLO14, GOC18, GV16, Haj19, KMS17, LCM12, LS11, MBKW19, MBL15, Par12, RAB14, RS17, SR16, WBG15, XLW10]. nonlinear-multivariate [WBG15]. nonnormality [Bur14, GGSS19, KEW13, RBHSL11]. Nonparametric [AR10, BS14, CB19, DM12, Ned11, PS19b, PL16b, SK11, She12, She15a, TPW17, VDBA14, YZ15, ZCZ15, AGM15, BC19, CYPGGM16, DS11, DS10a, DFPT16a, DC15, DC16, KB18, LMSX16, Lu14, MG17, NY16, NS18, NJ18, SK17, WSC18, WBAS15, WWW15, Xu17a, YZWM19]. nonresponse [JvBF13, VM19]. nonstationary [MKL13]. normal [Aci18, ARY15, AG19, AL15, AB16b, AD15, AAVG16, ABA12, BMM14, BEBG14, CMC13, ÇŞ18a, CC16a, CSPO14, Dia10, DYA19, DC13, DFT17, DPK11, ER19, FLB15, FMB16, FAV18, GSL^{+}14, GLLO14, GYV^{+}13, GEC19, GA15, GGAM13, HK16b, HL13, Hay15, Hua16, JYML13, JTL18, KLK17, LW14, LL15b, LXZ11, Lon12, MAV17, MB16, MD18, MARS19, NTG13, Nom14, PO14, PJR15, RRB10, SL15, Sha18c, SH19, SG15, Sun11, TK15, Tsa18, VSG^{+}18, Wan18b, Wil11, WN11a, WN12, WZX13, XZD15, ZMKAV19, ZA12b, dABS16]. normality [AO12, BRY17, CF10, GV18, Han10, HT12, JV14, KR16, MP15, Mei11, Nou10, NA11b, Nou17a, PZ10, RDC10, RA17, RIY18, SEGMA19, SNG12, Wan15, Xu17b, YS11, ZA12a]. normalizing [Lia10]. normally [Lüt15, McF16, TK16]. northern [PSY18]. note [BMM14, DFPT16a, GGdC17, Kla15, Kru12, LRV17, LPV13, TS17, WL10, Wu12, ZL16]. novel [AP17, FAT19, Ism16, LLXY17, WHX^{+}17, Wu11, YAEU13]. Novick [RPN15]. nowcasting [MFD16]. np [HWA^{+}14]. NSGA [SFS15]. NSGA-II [SFS15]. nuisance [CRV15, JP17, Mug16]. null [CRV15, Hwa11, QYX17]. Number [SGG18, AL18, Cha16, GY16, HKK^{+}16, Hwa11, NKB19, RRDU13, SPS19, dABS16]. numbers [GB17b, McF16, YD16]. numerical [CGP15, DS11, LPS12, LR18a, Mur12, WT18].

O [ZB13]. O-labelled [ZB13]. obfuscation [MZZ^{+}19]. Objective [KLK17, LM13, FHSC14, GB17b, KGA12, LW17, QMZ15, RAB14, UA16, WTJW17, WyT17]. objects [LP17, MWL14]. observational [TJK13, TKJ13, TK14]. observations [CC17a, CA19, DM12, JO11, KS16, RAB16, ZH12]. observed [PF16a, XZZ15, Zha16]. Obtaining [RMS14, Dia10]. occlusion [LLP^{+}14]. occupation [DDD17]. odd [AACR18, BCY^{+}17, CAO^{+}17, dCOC16, ACN^{+}17]. odds [Alt19, WFC^{+}18, ZRH15]. offences [CW16]. oil [HOC^{+}19]. Olkin [ASH16, DNMT18, GGAM10, LLGP17, MS11, dESM15]. On-line [FG13, QLW16]. One [NTC^{+}19, RZ13a, AS15c, AW17, BEBG14, ÇŞ18a, DHP14, DS15, DC16, GGSS19, KLKK19, KP15, LZ10, PO14, PHCS11, PCMMA13, SB15, SK11, Wu16b, YML14, YCA15]. one-regressor [DHP14]. one-sample [DS15, DC16]. One-sided [NTC^{+}19, AW17, KP15, PHCS11, YCA15]. one-stage [Wu16b]. One-way [RZ13a, BEBG14, ÇŞ18a, GGSS19, KLKK19, LZ10, PO14, YML14]. online [Akn13, Gau11, SGGC10, YL14]. only [Mug16]. open [HS13, Sha18a]. operating [KKY15, LYQ^{+}15, Sha15d]. operational [HBFSGD11]. Operations [Sha13b]. operator [CKP15]. Optimal [AH19b, Alw17, ABJR13a, ABJR13b, DA16, DB10, DB11, DW18, DG19, Hua11, HW17, IA10, KS17a, LY13, LXL11, PDS16, TKÖ19, WH19, YBAA15, ASM19, ATH12, CTX18, DBC12, ER19, Hir11, JP17, Kia12, LJV^{+}18, LSL10, LY15, LLB12, MJ16, Pam19, PT14, RNH19, SP16, SNG12, WN11a, WL14, Ano14c]. optimality [WPXL14]. optimization [CMCH12, CLC19, GB17b, KGA12, LOK16, LL18b, San12, WH14, WyT17, ZCW^{+}17]. optimizations [Sha11b]. Optimized [IJL18, LLXY17]. Optimum [IAGEK11, VM19, GB17b, KGA12]. Option [Guo17, FM19, MS15, SJZ17]. options [JK14, Meh15]. order [AJFB14, AHAA10, BS13, BZAZ15, BT16a, BA15, CRV15, CG15, CL11, CLH14, DRAR19, FP11a, FP15a, FP17, FGHRM12, Gha16, GRPP10, GHMR19, HP11, JKM16, KL17b, LL10, Llo10, MMP15, MES19, NSMFR15, PGF19, Rai12, SGTKBL15, SAB15, SBS14, SWZ15, SAT16, TP15a, UY12, VDBA14, WSC18, Zar19, ZS16b]. Ordered [Haq19a, JK19, AR16b, GYV^{+}13, GBCS16, SAB15, TCM11, THCZ18]. ordering [Kiz18]. ordinal [Alt19, DP15, FH16, SXTJ17, Wan18a]. oriented [MN15, Sha19l, Sha19p, Sha19q]. Ornstein [HKST17, Zha11]. orthogonal [JP11, Jou15b, TP15a, ZCW^{+}17]. orthogonalization [LWW18]. other [BAB15, DRB17, MZZ^{+}19, MP15, OJRACL18, WLL12]. Outcome [BS16, GG17, TK16, YYG16]. Outcome-adaptive [BS16]. outcomes [BS16, CB10, JM10, Lu19, MA17]. Outlier [CC17a, DC13, Jur12, Par17, PMP14, SS12a, AKU11, CCF19, DS11, HMM13, KR18, SDWL17, YWZ18]. Outlier-free [CC17a]. Outlier-resistant [Par17]. outliers [AHM13, FMB16, KE10, KL17c, LGWZ18, Noo19, RBHSL11, SB12a, SC12, SO10, SB11, WLCL18, ZLW16]. output [VNM14, WC17]. outputs [DP13]. over- [OJRACL18]. over-dispersed [HK16b, SKJ17]. over-dispersion [PA15]. overdispersed [KNM^{+}15, PCMMA13]. overdispersion [FMV^{+}19, RIY18]. overestimations [KC16]. overlapped [AAG^{+}19]. overview [LZGW14]. ozone [PAFPM12].

P [ACG^{+}16]. P-splines [ACG^{+}16]. package [GEV18, Wan15, YS13, ZY19]. pair [Alw17, Dia10, JW18]. pair-approximation [Alw17]. paired [AOH16, AP17, LLM16, RNSR19, SHP12]. pairs [HA10, HM17]. Pairwise [FHS12, Lue19, RBHSL11, RR13a, RM19, FMK15, KKM16, LPV13, RRB10, Zha15]. Pairwise- [FHS12]. Palm [NHGS14]. panel [AB16a, BZF18, Fuk16, JTL18, Mei11, SS13, SBD10, YK15]. paper [BMM14]. Paperback [Sha19i, Sha19f]. parabolic [GB17b]. parachute [GZT14]. Parallel [SA16, AHH17, HC17, Kiz18, PS14, WZS17]. parallelism [TS16]. Parallelizing [PRNG18]. Parameter [ABA12, DPK11, EG18, KBM19, NB13b, NLHD12, RT14, SWLZ15, XZZ15, Adk12, ADA18, AP15, AWH19, ASH16, AY15, AS15b, AKV17, BT14, Bak14, BB12, CG15, CMCH12, CCMGA14, ÇK19, CCM12, CTX18, CA12, Chr15, Coi13, DHP14, DAB11, DB11, FWF16, FRL17, GZB13, GSW17, GRPP10, GHH17, GA15, Han17, HM18b, HAH14, HY14, HKL17, Ili16, JP17, JKM16, KC16, KR15b, Mar10, Mug16, NSR14, NB13a, NB14, NB15a, NB16, ÖK17, ÖK18, PV17a, PSK14, PK16a, Par17, PS14, PK11, RL14, SP11, SF13, San12, dSSCR19, SGG13, STL16, SSK13, SJ10, Ten19, Tza11, WY11, WWCL11, WLL12, WC14, YK15, ZZXS17, ZWDM19]. Parameter-driven [KBM19, AWH19]. parameterized [PV17b]. parameters [AE11, AS15a, AB16b, BZ14, BAS17, CRV15, CDG^{+}15, CM10, CM17, Che11, DN13, DA16, DA13, GAB14, GM16, HCT19, Ism10, JKM16, JK19, KSM16, LGW16, MM13a, MD18, MQR18, MES19, MADASAM11, NB13a, NB14, NB15a, NKB19, PA15, PL16b, QAA18, RYS11, Saz19, SRAO11, STW15, SAM13b, SAK14, TCHS19, TKÖ19, TK14, WM15, WH14, Wan18b, Wu16b, ZS16a, ZBG18, ZY15]. Parametric [CT14, Chi10, XMC^{+}14, BT16a, BCLM17, Bør15, CWM17, GC17a, GRPP10, GGM18, Haq19a, JK17, KL10, Li11, Med16, ME15, OSdVM13, Saf13, SRP11, SK18, WYZK16, WRN18, Wil10, WN13, WFC^{+}18, WHX19, XY16, Xie14, YPAC11, YA16b, ZS18]. parametrization [Lem11b]. Pareto [AS15a, AS12, AKJ16, CC17b, DAB11, GB17a, Han17, HSC11, Hua11, IAGEK11, Lon12, Nad10, Noo19, PNN17, RASR16, SSD17, Wu10, WLL12]. parsimonious [DF19]. part [AKS^{+}15, NX18a, NX18b]. Partial [SGÇ18, BCO13, Fuk11, LM13, Mal16, NHWT14, QX12, RAN11, SML19, Tsa18, WL15a, WWW17, Wu16a, XZ19]. partially [AS15a, AR16a, Ano14c, BCJG12, Haq19a, Ism10, IA10, Ism14, LGW16, PL18, QMZ16, SWW19, Wan18a, ZZ17]. particle [LL18b, Ned11, SHLT17, WH14, WyT17, YTNT14, dSdS17]. partition [WYW12]. Partitioning [LCN^{+}17, RZ13b]. past [AAGV12, Li14, Zar17]. path [CP12]. Pattern [JM10]. Pattern-mixture [JM10]. patterns [FFP16, MN15, PC11, PSY18, SBMF18]. PDF [JZD18]. Pearson [ASB14, CM14, RS17]. Pena [KAA18]. penalization [CP12]. Penalized [Ali12, YPL13, HS13, HC10, KBJ16, LOK16, MT17, NHA18, SWXJ18, SZJW19, SDWL17, SESY13, TTWT19, VAW15, Wu13, XWMA18, YZ14, YA16b]. penalty [KK18, LH18]. pepper [WHX^{+}17]. percentile [HK18b]. percentiles [HP11, LP10, SHST13]. perfect [MH12]. perform [BBV17]. Performance [DDDD10, dlCHBCD19, LF16, LLS13, MFR^{+}18, Moh17, WM12, Wu16a, CS17, DS11, DFPT16b, GVTGH^{+}18, HA13, HSC11, LS14, LK13, MM13a, MQR18, NO10, QAA18, RRV13, Tar12, Wu11, WC14, ZBG18]. Performances [Kal14, YS18]. period [CB19, Fuk11]. periodic [BCL18, BH18, MN19]. permanent [McC14]. permutation [GBCS16, KEW13, YXZ19]. Permuting [AP17]. Perron [HM19]. personality [MTS14]. perspectives [SLA17]. perturbation [Gül10, TTWT19]. perturbations [RL15]. perturbed [DRB17]. PFC [XLZ18]. pharmacokinetic [CL14]. Phase [SOH13, EKE^{+}18, PV17a, AKAW15, YYW15]. phased [WWB18]. phenomenon [PSV11]. Phylogenetic [AP12]. physiology [WSC18]. Piecewise [AGA18, SB12b, WDY18, GGA18, Wel16]. Pitman [AB11, BDKM11, RAB16]. pivotal [MADASAM11]. pivotals [Llo10]. plan [AR10, AS12, AJA11b, AWJ^{+}13, ABJR13a, ABJR13b, ABA16, NS17, WW19, WH19, YCA15]. Planning [LCB13, LCB14, Ano14c, IA10]. plans [AJA11a, AYJ11, AW17, CYL17, DYT10, DT13, DW18, DG19, Gov17, IAGEK11, LY13, Lia14, LHB11, LHB13, LCB13, TCLY14]. plant [WSC18]. play [Li14]. plot [SGTKBL15, Sta10]. plots [KK12]. plotting [Han10]. PLSR [AAG^{+}19]. plug [FFP16]. plug-in [FFP16]. PMSE [NX18a]. Point [ARY16, Stå16, ACG^{+}16, BT14, BTR16, BZ16a, HEB13, Kel16, KK15, Lem11c, Len11, LLC17, MN15, RG10, RS15, SR16, Wu16c, YP10, YTNT14, ZGW14]. points [EFGMD13, JZZH18, LGWZ18, Wil15]. Poisson [AD15, ASA^{+}17, AÖ13, BK17a, BAJN14, BSS15, BV15a, BVRI16, COBH11, ÇK19, CW16, Che19, Chi10, DSCS19, Fre12, GBdL16, GDCO18, GWH14, HEB13, HK16b, HBFSGD11, HR12, HKL17, KP19, Kha12, KSJ16, KL17a, KL13a, Lee19, Li11, LWT^{+}17, OWLP16, ÖI10, Özk19, PB17, QKY16, RYS11, dSSCR19, SMM19, Sha15a, SH19, TZC^{+}16, Wei12, ZHB18, ZWDM19, ZST15]. Poisson-binomial [ZHB18]. Poisson-generalized [COBH11]. Poisson-reciprocal [GDCO18]. Poisson-X [TZC^{+}16]. Poissons [SY19]. polar [Nom14]. policies [AD16b]. pollution [DRC^{+}16]. polygon [SHLT17]. polylogarithm [GGdC18]. polynomial [AD16b, Coi13, JP11, KSM16, MMT16, SB12a]. polynomials [MM15]. polytomous [GYV^{+}13, WZ13]. pool [DHP14, Kuk18]. pool-adjacent-violators-algorithm [DHP14]. Pooled [Tza11, LLB12, NKZ19]. pooling [Rav19]. popular [ABA16]. population [AOR13, AOH19, Ak19, BMM15, CM16, GD19, IJL18, KTJ18, LXL17, Lu19, Moh17, MKSN18, NY16, PSKC18, QKUH18, QRH19, SA12, SRAO11, STS14, SG15, YP10, YPAC11, YS19]. populations [HS13, JKM16, JK19, LXZ11, RRB10, SBAA14, Stå16, SZ16, TK15, WY11]. PORT [FGHRM12, GHRAM13]. portfolio [CLC19, THR18]. portmanteau [AB18, ZHH19]. position [Han10, HCT19]. positive [AKS^{+}15, CLYX18, HLVRS18, NX18a, NX18b, PV17b, Sha19u]. positive-part [NX18a, NX18b]. positives [XX15]. possibilistic [MBG17, WH17]. possible [JZZH18, RSD14]. post [Fre16, LSA16, Sun11]. post-change [Sun11]. post-stratification [Fre16]. posterior [FM15a, HH17, LRV17, WJ17]. posteriori [LR18b, MAV17]. potentially [CI15]. Power [DC15, Li15b, SY17a, SK17, Ali12, Ali14, AACR18, AL18, Aus18, BMPZ14, BMP12, Coi13, CLAH17, DC19, DI11, GGA18, JV14, KL17a, LL18a, LL10, LK17, MY13, NNB14, RDC10, RS14, SJN15, SDS16, TKB16, TRC^{+}18, WL15c, Yam19, ZCL19, ZZXS17, dABS16]. power-divergence [Ali12, Ali14]. power-log [ZZXS17]. power-normal [dABS16]. powerful [LL15a, PHCS11, WPC15]. pp [Pak11, Sha19i, Sha19f]. Practical [DFPT16b, Sha19j, Jon16]. practice [GSL^{+}14, POCdP13]. Practices [Sha19j]. pre [CC17a, Llo10, NO10, WW16a]. pre-binned [WW16a]. pre-classified [CC17a]. pre-estimation [Llo10]. pre-test [NO10]. Precedence [LBN^{+}19]. precise [CRV15]. precision [LC10, WW12]. predator [Alw17]. predict [PAFPM12]. predicting [WDCK15, Yua15]. Prediction [AJFB14, AVK15, BA15, Erd13, RAJ16, SAB15, AHAH14, AMB12, AHAA10, AB10, CF15, DD12, DNMT18, DDD17, FNRCM17, Hat16, HW12, KTSR17, KR18, KN16, LX14, Lia11, NK15, NY16, PS16, PB19, PK11, RMS14, SBS14, Sha16, TD14, VDBA14, WL15c, XHEM17, XLZ18, YFT10]. Predictive [Gov17, FGV14, HBT12, KÖ19, MPB19]. predictor [GP15b, Her11, LY15]. predictors [ASM17, ARY16, LB11, RAB16, THCZ18]. Preface [AV19]. Preliminary [BPH12, Gau10]. premium [HBFSGD11]. presence [ASM19, Bor17, CRV15, CH19, GK16, GGSS19, JK17, JZZH18, KR18, KA13, LL15b, MSK14, NJdC14, PA15, RL15, SBMF18, VM19, Wu11]. present [ASM17, Mug16]. prespecified [HU14]. Press [Sha19i, Sha19f]. pretest [LSA16, SLA17]. prevalences [LYLM17]. prevalent [Yu11]. prey [Alw17]. prey-predator [Alw17]. price [Bia15]. Pricing [MS15, FM19, Guo17, JK14, Meh15, SJZ17]. primal [SZJW19]. Principal [Fuk16, JSM13, LGW16, AHJ16, Car16, Fuk11, VKK14, YBAA15]. Principal-component-based [Fuk16]. Principles [Sha19j, Sha18a]. prior [AL15, AY15, ABA12, BMM14, DD12, Tak17, WL15c, YZWM19, ZWDM19]. priors [Ho12, KLK14, KLKK19, LW17, LY18, LM13, LZZ^{+}15, SSK13]. probabilistic [KL18b, DMV17, RWCD17]. probabilities [AB15, Bak18, GP15a, Gou11, HL13, Hay15, LB11, LLB12, Nom14, OvP10, SLD16, TK14, YL14]. Probability [SS15, AKS^{+}15, BMM15, BAKZ16, CH15, DD15, GB15, HWWZ16, JY14, JL16, LZ11, MQR18, NJ18, OWKC15, ÖI10, Sha10d, Sha19f, VVNTVD^{+}17, XBL18]. probability-proportional-to-size [BMM15]. probable [ZA12b]. probit [Adk12, AB16a, HA13, LS19, NMPR14, SBD10, Wan18a, YYW15]. problem [CLC19, MN19, NS17, PA15, PHCS11, QX12, RAB14, RGNM13, RS15, SH10, WRN18, WPC15]. problems [BTD18, Che19, CVL18, Mur12, NSMFR15, PD13, Wu16c]. procedure [CHQ17, FH11, GMG13, JCW19, KB18, Kuk18, LL18a, DDJ16, PB17, Rav19, SLV18, STS14, TP15a, WC14, WL14, You14]. procedures [DS18, DB11, DBL10, FKM13, Hwa11, Li15b, Rai12, SSD17, Sha15a, TP15b, UY12, WZ13, Wu16b, ZMW13, vGK17]. Proceedings [KS15]. Process [Wei12, AM13, AH19a, AH19b, Ami11, AYJ11, AWJ^{+}13, AÖ13, AH18a, AA16, BT14, BKJ16, BU17, BVRI16, BBM18, CW16, Cha17, Che19, CKPS11, DW18, DLZ19, Gau11, GV16, HZSA19, Haq14, HK16a, HM18a, HK18a, HCY16, HCT19, HAB12, Jer13, JK14, KAT15, KS17a, KSJ16, KL17b, KB18, LAuaS^{+}15, LSF^{+}17, Med16, MH17, NS17, NAA18, OWKC15, PWW15, Per10, RA14, SOH13, SC16, SLLZ18, WGC14, WL15c, WBGJ15, WW19, WL10, Wu11, YA16a, YZWM19, ZMKAV19]. processes [ACG^{+}16, BTR16, BK17a, BBC10, CS17, CL15, DSE16, GGM18, HEB13, HR12, JY14, MAV17, Moi17a, NAA18, PBSZ13, PV17a, Per10, PL15, RPFOMGRM17, RMS14, SF13, SM11, Sha13c, SB13, SLM16, SKTC11, TD14, TNJ17, Tsa10b, WBG15, Wei12, XSF17, YYG16, Ye16, Zha11, ZST15, ZB10]. product [BS13, BZAZ15, MKSN18, Nad10, PGF19]. production [FF14, HOC^{+}19]. products [CLC17, PBSZ13, VS15, WC14]. profiles [AKAW15, AW17, GV16, HLL18, KAK16, Li18, WW16b, WL19]. program [GHMR19]. programming [FMK15, QX12, RAB14, Sha19l, UA16]. programming-oriented [Sha19l]. programs [KKL^{+}15]. progressive [AHAH14, AHA15, AHH17, AS15a, ARY15, AG19, AB16b, BB12, BBA15, BAS17, CM17, CPK19, DBC12, DYT10, DT13, DC19, DKG16, ER19, HSR15, HW12, HW17, KTSR17, Lia14, LHB11, LH12, LHB13, PGF19, RT14, SKK12, SBK13, SEAEM13, SAK14, TN16, TP15b, WY19, Wu10, WWCL11, WH19, XYZ19, YCXN14]. progressive-censored [AHAH14]. progressive-stress [AHAH14, AHA15]. progressively [AMB12, AYR16, ARY16, AR16b, BS13, CBG16, DNMT18, EDASM17, GTB14, GAB14, GP15b, HC17, Ism14, KM12, KK13, KDG17, MTSR19, NLHD12, PB13a, PS14, RAB16, SK16, STW15, ZS16a]. Projection [AF12, CWZ18]. Projection-based [AF12]. prone [LYZ11]. propensity [TJK13, Tu19, TRC^{+}18]. Properties [GDCO18, TQP10, AHH17, AACR18, AB10, AR16b, BG11b, BS12, Bru15, CGdS14, CSPO14, CAO^{+}17, CYRO18, CLL10, DAP15, GDCO11, Kib12, KMS17, LMFMA15, OWLP16, SWW19, SDS16, SLM16, Tak17]. property [HM19]. proportion [GD19, QYX17, SB15, Tso15, YGX14, AH16]. Proportional [She11, She14b, AAGV12, Aus18, BMM15, BBA15, CPK19, DW17, DW18, Kiz17, OS14, sS11, WDY18, ZRH15]. proportions [ATH12, CC19, CCS12, RZ13a]. proposal [DGLV17, SU11]. proposals [TMG18]. protocol [LN13]. prototype [RGNM13]. proxy [NX18a]. Pseudo [CS11, DK17, DDD17, TGL12a, YD16]. Pseudo-Bayes [CS11]. pseudo-empirical [CS11]. pseudo-replicated [TGL12a]. pseudo-value [DDD17]. pseudorandom [McF16]. Pte [Sha19f]. public [DSVY14]. Publishing [Sha19f]. purposes [GG16, SLV18]. puzzles [VS15].

QMLE [BZF18]. QR [RN18]. quadratic [ARB13, GB17b, KGA12, OMY12, QX12, UA16]. quality [QLW16, Rav19, Sha19g, WW19, YZL17]. quantal [DS10a]. quantification [OSN17]. quantifying [HH17]. Quantile [CHT16, FMZ18, HH15a, Hua16, LC18, MSA12, QZZ16, EB19, AR10, AY14, AY15, Cha14, CSJ17, HC15, HXT15, HKKL17, JLM15, KBJ16, KM17, KPK^{+}13, KK11, Kuk18, LLT12, NB13b, PZZ19, SRP11, SB15, She14c, Tar12, TTWT19, Tsi15, WKJ^{+}19, YZX18, Yu15, ZL15, ZL16]. quantiles [AR10, Cor13, Lon12, NB14, NB15a, Shi15, TTZS15, WEHCC14, Yua15]. quantitative [SBMF18]. quantization [FG13, FS15a]. Quasi [Alt19, CH19, SH16, Wu16c, FGHRM12, Kha12, Li18, NBB15, Ned11, SB11, Tsi15]. quasi-Bayesian [Tsi15]. Quasi-complete [SH16]. Quasi-hidden [Wu16c]. quasi-independence [NBB15]. quasi-likelihood [Kha12, Li18, SB11]. Quasi-maximum [CH19]. quasi-Monte [Ned11]. quasi-PORT [FGHRM12]. question [Lin14]. questionnaire [VKK14]. questionnaires [NMOV18].

R [ACG^{+}16, CM14, GEV18, JV14, KKL^{+}15, TNM17, VP16, YS13, ZY19, Sha19d]. Rademacher [MM15]. radiometry [RB17]. Random [MWL14, NJ13, Yua15, dABS16, AHAA10, APF18, AMB15, AR16b, AD16b, AL18, BP15, BA15, Ber12, BEBG14, Bru15, CCF19, CCMGA14, CLS^{+}19, DN13, DYT10, DT13, GP15a, GDR12, GI17, HBL14, HOC^{+}19, HM13, HBC11, Juh16, KLKK19, KP15, KTJ18, LH18, Li14, LL16, Lia11, LZ10, MJ16, McL14, MFD16, MSS18, NFFM14, PSKC18, PC11, PO14, RGNM13, RG10, RKV17, RRDU13, SA12, SH16, SY17a, She15b, WYW12, WL15b, YML14, YD16, Zar17, ZC13, ZWCK16, ZX14, ZB13]. random-clumped [RGNM13]. random-intercept [RKV17]. randomization [AP17, TQP10]. Randomized [RTM18, CS11, CB11b, JMM^{+}17, MES19, Mur15, SM15, TTWT19, TP15a]. randomly [DA14, KVK15, PP19, RS14, Zör15]. randomness [DTZZ12]. range [AP15, PL15, YWL18]. Rank [HMM13, BZ16b, CP14, CWZ18, DD15, DKY17, GLB17, HZ14, HLL18, LM16, MMR16, Man15, NKB19, PP15]. Rank-based [HMM13, HLL18, NKB19]. rank-sum [DD15, MMR16]. Ranked [AMB15, BK16, AMAMS12, AAS18, AO11, AOR13, AOH16, CTX18, EG18, FZ18, GD19, Haq14, HBMAO15, HBM16, Haq17, LBN^{+}19, MFR^{+}18, MA10, MS18a, MMR16, MAAM10, Nou17a, SAB15, SDC12, SY17a, SGÇ18, SRAO11, STS14, SKM14, TdSC19, TKK19, YS18]. ranked-set [FZ18]. ranks [VGTGCFC17]. Rapid [SDWL17, YTNT14]. Rare [QKUH18, QRH19]. Rasch [FHS12]. rate [AAGV12, Bag11, BCL18, BC19, CPK19, COP14, DW18, FVB13, GBdL16, GGdC17, GGdC18, Guo17, JG10, Kiz17, LPJ14, Li15b, LH14, LVB18, MY13, PQ19, RYS11, RBC^{+}15, Sha18c, SRK13, UGMK13]. rates [AB10, AL18, Car19, Sha15a, SBK13, TRC^{+}18, VHV^{+}16]. Ratio [NHGS14, BTR16, BG11a, BSS15, BEBG14, CS16, Che19, CL10, EFGMD13, Gus15, LZ10, Mar11, MMP15, MFP14, MKSN18, Nad10, NBB15, OWKC15, PK17, PCN14, SJN15, SZS16, SLLZ18, WPXL14, WM12, WN11b, WW12, YS19, ZST15, ZGW14]. ratio-based [ZGW14]. ratio-exponential-type [YS19]. ratio-product [MKSN18]. ratio-type [YS19]. rational [OWKC15]. ratios [CSJ17, DGK12, Kuk19, LL17]. Raton [Sha19i]. Rayleigh [AE11, AHAH14, DD12, EG18, FH15, GdSCO14, JZD18, KR15b, MTSR19, NNB14, RAK16b, SJN15]. Re [SCW16]. Re-weighting [SCW16]. real [KAJB19, Sha15c]. real-world [Sha15c]. realistic [HM19, RBHSL11]. reasonable [ASM17]. recapture [ABM17, HHC15, Moh17, YP10, YPAC11]. Receiver [Sha15d, KKY15, LYQ^{+}15]. reciprocal [AJM11, GDCO18]. recognition [LMRW17]. recombinant [PRNG18]. Recommended [RNSR19]. Reconstruction [AAGV12, KAR13]. record [AR10, AMB12, CM16, DB10, DAB11, DB11, KN15, KN16, Kiz17, NK15, RASR16, SAB15, Sha10e]. record-based [DB11]. records [AB11, Bak14, DDZ13, MZZ^{+}19]. recovery [AB10, ABA12]. Recurrence [BS13]. recurrent [CR18, Wu13]. Recursive [Akn13, BZAZ15, GM16, LJ18, RZ13b]. recycling [MCW17]. reduced [FGHRM12, HZ14]. reduced-bias [FGHRM12]. Reducing [MT13, SD15, CP14, XX15]. reduction [Adr18, AAVG16, CG15, DKY17, Fan19, LS19, Par11, WL16, WN11b, XLZ18, Yoo13, ZLZZ18, dCFOM12]. redundancy [LV17]. reference [ABJR13a, ABJR13b, BMM14, YL14]. referring [Wel16]. refined [DY16]. refinement [FS13]. regard [MY13]. regime [Guo17]. region [Wel16]. regions [AB15, CSJ17, LL17, MMP12, Rin12, Sun11]. Regression [DSCS19, LdNdSF18, PM10, SGGC10, THCZ18, AM19, AHM13, AÖ16, EB19, AD10a, ADRA15, AY14, AY15, AACR18, AKU11, AI12, AAAQ19, AR16a, AKAW15, ATF19, ASB14, AGM15, AL18, Aus18, AY18, AAY19, BKR17, BSS17, BZ16b, BAZ19, BK17b, Bor17, COBH11, ÇK19, Cha16, Cha14, CL13, CLC17, CHT16, CT14, CA12, Coi13, CAO^{+}17, CYRO18, CA19, CYPGGM16, DG16, DYA19, DKY17, DK10, DP13, DA13, DDD17, EKE^{+}18, FOC14, FH16, Fam12, FW15, FMZ18, FI17, FBV18, FP11b, FP14, FLB15, FAV18, FPP18, Fu16, GLLO14, GYV^{+}13, GG16, HOC^{+}19, Hat16, HM13, HMP17, HA10, HAH14, HN13, HH15a, HY16, HY14, HC15, HXT15, Hua16, dlCHBCD19, HKKL17, HH15b, JP17, JLM15, KSÖG11, KBS11, KBJ16, KAA18, Kib12, KK14, KS17b, KK18, KPK^{+}13, KSLN^{+}18, KK11]. regression [Kuk18, KÖ19, LDCL17, LY15, Lem11a, Lem11b, LCM12, Lem12, Lem13, LYQ^{+}15, Li15a, LL17, LY18, LNC17, LSA16, LZZ^{+}15, LP16b, LGW16, LWW18, LM18b, Lu14, Lue19, LLT12, Mån13, MMP19, MBKW19, MJ16, MH16, MG17, Mar14b, MFP14, Moi17b, MS18b, MSA12, NO10, NX18a, NX18b, dALNCdATdC11, NHA18, OCPC12, ÖKD17, ÖK18, PB17, Pam19, PZZ19, PV17a, PSK14, PK16a, Par17, PZ10, PRMM12, PMP14, PD13, QAA18, QZZ16, QPQ18, QHB15, RAN11, RBK16, RA17, RN18, RIY18, SGG18, SRP11, SCL^{+}18, SWW19, SB15, dSSCR19, ST13, SD15, SY19, SF12, Sha19u, She11, sS12b, She14b, She14c, SY17b, SWXJ18, SZJW19, SU11, SO10, SGB13, TK18, TTZS15, Tar12, THG15, TX14, TTWT19, VRC13, VP16, WZ13, WL17, WLCL18, WKJ^{+}19, WBAS15, WWW17, Wil10, WS16, WN13, WW12, Wu13, XLW10, XY11, Xu17a]. regression [XWMA18, YZ15, YR15, YPL13, Yu15, YFT10, ZP14, ZCZ15, ZLW16, ZLL19, ZL15, ZL16, ZY15, ZB13, dCOC16]. Regression-based [SGGC10, Hua16]. regression-free [SGGC10]. regression-type [WW12]. regressions [BAB15, CTC17, CNS12, CNQ14, CNP19, FS10, KM17, Lem16, NMPR14, PCN14, XY16, YZX18, ZLZZ18]. regressor [DHP14]. regressors [HA13]. regular [BCJG12]. regularization [FP15b, PK16a, RY13, WS16]. Regularized [CPP^{+}19, LYQ^{+}15, ZRH15, FWF16, MMK14]. regularly [EB19]. Reisensburg [KS15]. reject [Bot11]. rejection [ZWCK16]. Relabelling [ZF16]. related [Ame12, BSS17, KM14, KSLN^{+}18, Sha15h, Sha15j]. relation [Mal16]. Relations [Xie14, BS13]. relationship [PZY^{+}14]. Relative [JO11, FGH14, JMM^{+}17, MMP12, YZX18]. Reliability [AG19, Akg19, CC17b, GZT14, KM12, KK13, LH14, XYZ19, XG11, AAS18, Bak14, CPK19, EXH16, Gun18, GGAM10, GGAM13, Kiz18, LSA^{+}15, LPY18, NTZ19, PNN17, RT14, SK18, SSK13, SWZ15, SAM13b, TMG18, WWB18, XBL18, YML14]. remanufacturing [MCW17]. remarks [Mal16]. Remedying [PSV11]. removals [DYT10, DT13]. remove [WHX^{+}17]. renewal [Ami11, CCZ13a, CCZ13b, JG10]. repair [AD16b, RBC^{+}15]. repairable [JG16]. Reparameterizing [MS11]. repeated [FHO15, HTZ^{+}16, MBL15, MC14, PA14, RG10, YPAC11, ZS18]. Repetitive [NS17, YCA15, AYJ11, AWJ^{+}13, LAuaS^{+}15]. replacement [Moh17]. replicated [CCZJ17, MAAM10, STL16, TGL12a]. replicates [BBL13]. representation [FNRCM17, Hof12]. representative [GL16]. represented [AD12]. Reproducibility [DD15]. reproductive [DFT17]. repulsive [QPQ18]. reputation [PSS15]. Resampling [XLB12, DS18, PC10, ZNW19]. resampling-based [DS18, PC10]. rescaling [Bru19]. Research [Sha19j, MCW17, NJ13, Sha13b, Sha19h, Sha19s]. reservoir [Sha19n, Sha19m]. Residual [Lee19, LSF^{+}17, WGC14, WBGJ15, WBG15, CSJ17, FH15, FRL17, HKKL17, RAJ16, RASR16, Zar19, dCOC16]. Residual-based [Lee19, HKKL17]. Residuals [ZR14, ASB14, MS17, PA14, RS17]. resin [HOC^{+}19]. resistant [Par17]. resonance [KM14]. resources [LYL17]. respondent [NB15b]. respondent-driven [NB15b]. response [CTC17, CS11, CCK13, DHP14, FH11, Fan17, FBV18, FTS10, GdCCDS18, IJL18, LS19, Lue19, MTS14, MSS18, NJ13, PV17a, RAB14, SH16, Sha10l, Sha19u, TA11]. responses [CB10, Ciu13, KBS11, LS14, Özk19, PA14, WL15b, Wan18a, ZPL16]. restoration [CC16b]. restricted [KSÖG11, ÖKD17, RCL15, XY11]. restriction [MES19]. restrictions [JKM16, ÖK18]. results [BCJG12, CL10, CNO13, NCO11, NCO12]. retain [Car16]. RETRACTED [Ano14c]. Retraction [Ano14c]. retransmission [LPY18]. returns [Ral17, THR17, Yos18]. reversed [AAGV12, Kiz17]. reversible [LZZ^{+}15]. Review [Sha14d, Sha15f, Sha15b, Sha15d, Sha15c, Sha15e, Sha19d, Sha19e, Sha19i, Sha19f, Sha19g, Sha19h, OJRACL18, PF16b, TMG18]. reviewed [Sha10a, Sha10b, Sha10c, Sha10d, Sha10f, Sha10e, Sha10g, Sha10h, Sha10i, Sha10j, Sha10k, Sha10l, Sha11b, Sha11a, Sha11d, Sha11e, Sha13b, Sha14c, Sha14b, Sha14a, Sha14e, Sha14f, Sha15f, Sha15b, Sha15d, Sha15g, Sha15c, Sha15e, Sha15h, Sha15i, Sha15j, Sha14d]. Reviews [Sha15g, Sha15h, Sha15i, Sha19j]. revised [BKR17]. rich [EKE^{+}18]. Ridge [RAN11, AÖ16, AY15, AAAQ19, AR16a, AKV17, DK10, Ema16, GGG19, KSÖG11, Kib12, KÖ19, NO10, NAA17, ÖKD17, ÖK19]. ridge-type [Kib12]. right [AK16, AY18, AAY19, BZ16a, BMP12, CC19, CPK19, Che18, CHT16, KA13, Li14, MAEP14, NB15a, NB16, PP15, PPK16, PP19, She10, She14c, sS16, sS19, Stå16, Su16, XZ19]. right- [KA13]. right-censored [AY18, AAY19, BMP12, CHT16, NB15a, PP15, PPK16, PP19, She14c, Su16, XZ19]. right-censoring [sS16]. ringens [PSY18]. Risk [FS15b, NO10, THR18, BR16, DK17, FGH14, GZL18, Gus15, HBFSGD11, IS13, JMM^{+}17, LL11, PB15, RZ12, Sha10f, Sha15j]. risk-related [Sha15j]. risks [AE17, AYR16, ARY16, ARY17, Bør15, CLC17, DG16, FMZ18, GK16, HU14, HH15a, HW17, Ili15, Lee17, MSS14, WYW12]. RNA [FMK15]. Robust [AO12, AB15, AAG^{+}19, AI12, BAZ19, CB11a, ÇŞ18a, DF19, GZL18, IPK10, KL17a, KF16, KL18b, LL10, Lem16, LMSX16, LNC17, MS18a, MLCL18, Mar14b, NKB19, NLK11, PSK14, PK16a, Sha19u, SZWL19, SB11, THR17, THR18, VBL17, Woo10, XWMA18, YLG15, AF17, AM13, AO11, AY13, BV16, CT14, DA16, HMP17, KM17, MCZ17, MBKW19, MSK14, NAA17, PC11, PMP14, PP19, QKUH18, RA17, San12, SGGC10, Sha19a, Tso15, Tso16, TKJ13, Wil15, Wil11, WYX17, YL15, ZL11]. robustness [DC16, MHA10]. ROC [BCO13, MCBPF16, ODBT15, WZ13, Sha15d]. rock [ASM17]. rock-burst [ASM17]. root [CL16, MK12, ZS16b]. roots [GHJC10]. Rounding [GSL^{+}14]. row [Sta10]. row-column [Sta10]. rule [MAEP14, PZY^{+}14, ZZXS17]. rules [CL13, CC11b, JKM16, JK19, MQR18, RA14, Rak16a]. run [Bak18, Her11, LZGW14, SHST13, ZWCK16]. run-length [SHST13]. runs [CC11b, CB13, CLYX18, MQR18, RA14, Rak16a]. Runtime [KKL^{+}15]. rural [IGR13].

Saddlepoint [GP15a, Ye16, AE17, McL14]. sagax [PSY18]. SALSA [WMDO11]. salt [WHX^{+}17]. salt-and-pepper [WHX^{+}17]. Sample [Su16, AF17, AM19, AVG18, AR10, AHAA10, ARQ19, AS15c, AAJ16, ABA12, BA15, BTD18, BG11b, Bru15, CC19, CTC17, CWZ18, Che19, DS10a, DS15, DB10, DC16, EE15, EDASM17, EG18, FKS10, FP14, GB18, GP15b, HW17, IP14, KAK16, KK13, KL17c, Lee11, LCZ18, Lon12, LK13, Mur12, Mur15, PS19b, PP15, PP19, PF16a, PL16b, RAB14, RRB10, RBHSL11, RRV13, Saf13, SGTKBL15, SBAA14, SBS14, Sha16, SC19, SHP12, SH10, SZ16, Tar12, UA16, WSC18, Wes16, Wil11, Wu10, WWCL11, WLL12, WC14, ZBG18]. sampled [AR10]. sampler [CI15, CK14, DSVY14, Lia10, SDC12]. samples [AOR13, BZ16a, BMM15, BBA15, BK16, ÇŞ18a, CYL17, DVA15, FM15b, Gau10, GAB14, GWX14, HN11, LBN^{+}19, LY13, LLN13, NLHD12, NFFM14, RR13a, RAB16, SB12a, SBS14, SKM14, TCLY14, WY11]. sampling [AMAMS12, AAS18, AO11, AOH16, AOH19, Ali14, AS12, AMB15, AJA11a, AJA11b, AYJ11, AWJ^{+}13, ABJR13a, ABJR13b, ABA16, AAJ16, AW17, BU17, BK16, CS11, CYL17, CTX18, DYT10, DRB17, DGLV17, DG19, FM15a, FZ18, GZL18, GD19, Gou11, Gov17, GB17b, Haq14, HBMAO15, HBM16, Haq17, HK18a, Haq19a, Haq19b, HSC16, Hof12, HY16, KGA12, KK11, KTJ18, LF16, LAuaS^{+}15, LLXY17, LY13, Lia14, LHB11, LHB13, LZ11, MFR^{+}18, MA10, MS11, MS18a, MMR16, MH12, Moh17, MAAM10, Ned11, NB15b, NS17, Nou17a, Özg18, PSKC18, QKUH18, QRH19, SMBS10, SAB15, SA12, SFS15, SY17a, SGÇ18, SRAO11, Sha19t, STS14, SG15, TdSC19, TCLY14, TKK19, VM19, VC15, WW19, WCW15, XSF17, YLL17, YCA15, YS18, YS19, ZWCK16]. SAR [LPS12]. sardine [PSY18]. Sardinops [PSY18]. SAS [BBV17, LV17, NB18, Sha15g, UM14]. saturated [LJV^{+}18]. Saunders [AJA11a, BZ14, LDCL17, Lem11a, LCM12, Lem12, Lem13, LMFMA15, Lem16, LX16, MLCL18, NGXZ14, PJR15, SN17, XYZ19]. SCAD [KEW13]. Scale [LDCL17, Che11, FLB15, FMB16, GLLO14, Ism16, JKP19, KL18a, KP18, LCN^{+}17, LCB13, LHLB19, MAV17, MARS19, MADASAM11, PB13a, PS14, PL16b, RGNM13, RM19, SHLT17, Saz19, SGGC10, SL15, SRAO11, WGC14, WLL12, WZX13, XHEM17, ZX14, ZBG18]. scale-mixture [SL15]. scale-mixtures [FMB16, MAV17]. scaler [BMM14]. scales [Mar11]. scaling [Cox13, QX12]. scheduling [WTJW17]. scheme [ADA18, HBM16, PB12, PSS15, SFS15, TN16, TP15b, WCW15, MM13b]. schemes [HBMAO15, KAK16, SA12, YCXN14]. Schmidt [Kia10, LWW18]. Scholes [Gül10]. science [Sha10g, Sha11c, Sha18a, Sha19k, Sha19i]. sciences [Sha19h]. Scientific [Sha19f]. score [AH16, CL10, DY16, LPV13, Mog11, Mug16, SO10, TJK13, Tu19, TRC^{+}18, XZD15, YGX14]. score-based [Mug16]. scoring [CKP15]. Scott [NTC11, PSV11]. screening [BTD18, Che18, DM16, LB11, LWZ17, LC18, MCZ17, PZZ19, She11, WYZK16, WL14, YHWS18, YQ19, ZZ15a]. screenings [CTC17]. SCS [Zha17]. SDE [KSM16]. search [BOG15, FH11, JKLR11, PD13]. Searching [OSdVM13]. seasonal [BCCN18, BVRI16, KE10, RRV13]. Seber [HS13, YPAC11]. second [AKJ16, CG15, FGHRM12, GRPP10, Llo10, TP15a]. second-order [CG15, FGHRM12, GRPP10, TP15a]. sectional [TS14]. segmentation [BCLM17, LP16a, TNM17]. segmented [Mug16]. selected [AR16c]. Selection [PBM16, Bai16, BC19, BCT16, DS18, BAZ19, CCF19, CL13, CLH14, CKP15, DW19a, DP15, DK10, EZ12, FWZ^{+}15, FWF16, FP11a, FP15a, FP17, GL16, Her11, HZ14, HM18b, HAH14, HH15b, JW18, JKP19, JP14, Kal14, KSLN^{+}18, KÖ19, LKM^{+}15, LC19, LYQ^{+}15, LY18, Lin14, LT16, LSA16, LZZ^{+}15, LWW18, MMK14, NKB19, OCT18, PS18, PSK14, PK16a, Par17, PHCS11, RT19, Saf13, SPS19, SESY13, SU11, SJ10, TK18, Ten19, TX14, THR18, WDCK15, WL15a, WL17, WZ19, WZX13, WPC15, XY16, XX15, XWZS15, YLG15, YR15, ZX12, ZFZQ18, ZF11]. selector [BAKZ16]. selectors [DL19]. Self [Rav19, DGLV17, SC16, SZS16]. self-adjusted [SZS16]. self-adjusting [DGLV17]. Self-starting [Rav19]. self-updating [SC16]. Semi [ASM19, BCLM17, GC17a, YP17, DLZ19, GRPP10, HH15a, JK17, LZZM18, ME15, QX12, WHX19, YA16b, ZS18]. semi-competing [HH15a]. semi-definite [QX12]. Semi-doubly [ASM19]. semi-Markov [DLZ19]. Semi-parametric [BCLM17, GC17a, GRPP10, JK17, ME15, WHX19, YA16b, ZS18]. Semi-supervised [YP17]. semi-varying [LZZM18]. semicircle [RS14]. Semiparametric [HS13, She15b, WZ13, YCD15, ZZ17, Zie11, AD10a, ADRA15, AY18, AAY19, DFT17, DA13, Ema16, HZL16, Hua16, Kan17, KC14, Kuk19, PL15, RBK16, RA17, RN18, Seo11, She14a, TX14, WBAS15, Wu13, WYX17, WZ18, XZY13, XZD15]. sensible [MSS18]. sensitive [MZZ^{+}19]. sensitivities [GP15a]. Sensitivity [AW14, BTD18, KKY15, LPS12, ABA12, CL15, CGP15, Da 15, DM16, GDPH12, JY13, LSA^{+}15, DMV17, SGTKBL15, WYZK16, WC17, XG11]. sensitizing [RA14]. sensory [Sha10g]. Separated [MS17, Say12]. separation [SH16]. sequence [Bak18, CB19, LP17, RG10, Sun11, TNM17]. sequences [FMK15]. sequencing [SH19]. Sequential [HAB12, Kum15, Ak19, FS15b, LZ11, OWKC15, SMBS10, SJZ17, SBS14, WPXL14, WL14, ZGW14]. serial [FHO15, LP17, Wei11]. serially [Wei12]. series [AHH17, Ahm16, ABGM18, AB18, AA11, BPH12, BCLM17, BS19, BBM18, CHTZ14, CCP12, CB11a, CL16, Chr15, CF15, CLAH17, EFGMD13, Her11, HZ14, IRNB18, Jin15, JCW19, KL18a, KF16, Kiz18, KBM19, KR15a, LP17, MN19, MC16, MV14, Moi17a, NSMFR15, Par12, PMM18, PSY18, RWCD17, RTM18, SDS16, Sha19a, Sha10k, Sha15i, SM18, SB12b, SR16, SKJ17, SO10, TD13, TS14, TZA10, WZS17, WSLX17, Wei11, WJ19, YK15, YWL18, YAEU13, ZMKAV19, ZCL19, ZNW19, dSdS17]. series-parallel [AHH17]. services [CLC17]. set [AMAMS12, AAS18, AO11, AOR13, AOH16, AMB15, BK16, CJ13, CTX18, EG18, FW15, FZ18, GD19, Haq14, HBMAO15, HBM16, Haq17, LBN^{+}19, MFR^{+}18, MA10, MS18a, MMR16, MWL14, MFD16, MAAM10, Nou17a, SAB15, SY17a, SGÇ18, SRAO11, SZJW19, STS14, SKM14, TdSC19, TKK19, YS18]. SETAR [FP15a]. sets [AKS^{+}15, LL18a, RZ12, SNGMRC16, ZF16]. setting [DLS18, Mar14b]. settings [TS16]. setup [ZS18]. seven [NA11b]. Several [WBAS15, GMG13, KKE17, MMP19, Mar14a, NGXZ14, PV17b, PGT11, PAFPM12, TK15, Wil11]. sex [GGM18]. shape [AP15, CG15, Ili16, Ism16, KTSR17, KP18, NG16, SGG13, SY15, Tza11, WWCL11, WLL12]. shape-scale [KP18]. shaped [ASH16]. shapes [WH17]. Shapiro [GEC19, MP15]. shared [FRL17]. sharpening [BBW17]. shelf [LSL10]. shell [WH17]. Shewhart [CM10]. shift [FMB16, HKL17, LMSX16]. shifted [MP16, WZ18]. shifts [MMR16, OWKC15, TC10, Yam19]. shock [NMS18, SSMF18]. shocks [BG11a]. shop [WTJW17]. short [BZF18, DLZ19, Guo17, Her11, Kuk19]. short-term [Kuk19]. shortfall [GZL18, GC17a]. showing [UGMK13]. Shrinkage [CA12, MAAM10, NHA18, YA16b, Cha15, CKP15, LH18, NX18a, NX18b, SLA17, Tak17, THG15, WN11a]. SICA [SWXJ18, SZJW19].

SICA-penalized [SWXJ18, SZJW19]. sided [AW17, KP15, NTC^{+}19, PO14, PHCS11, Rak16a, WL14, WPC15, YCA15]. sieve [CH14, RMS14]. sigma [Sha19g, Rav19]. sign [DC16, TTZS15]. sign-based [TTZS15]. sign-type [DC16]. signal [CSJ15, CC16b, LZGW14, SLSW15, XBL18]. signals [Bru15]. Signed [BZ16b, Wei11]. Signed-rank [BZ16b]. Significance [DLS18, Adk12, CRV15, LDB10]. Simian [TNM17]. similarities [TD19]. similarity [PSS15]. similarity-based [PSS15]. Simon [LRV17]. simple [AHM13, Ali14, AH18b, BT16b, Bru15, CC16a, Che11, Cor13, Dia10, GSL^{+}14, GK16, Her11, HA13, JP17, KAK16, Kus11, LL18a, LP16b, PSKC18, SY17a, WW16b, Wil11]. simpler [SCL^{+}18]. simplex [DGLV17, OvP10, TPW17]. simplification [SM11]. SimSel [EZ12]. Simul [Pak11]. Simulated [SGZM14, CS16, Woo10]. Simulating [HSC16, How13, HM19, LWT^{+}17, MB13, McL14, PQ14]. Simulation [How13, KR13, Adk12, BK17a, DS18, Bru19, CHTZ14, CDG^{+}15, DF14, DSVY14, EZ12, EKE^{+}18, FM19, FF14, FVB13, FS15a, GI17, HU14, HZ14, HR12, JMM^{+}17, JL16, KTJ18, LF16, LLGP17, LW12, LLV^{+}14, LZ11, Meh15, NHGS14, Özk12, PC11, PO14, RT19, RCL15, SNGMRC16, Saf13, SLV18, Sha15c, SBMF18, TJK13, UG10, Yuc17, Zha11, ZMW13, Sha15c]. simulation-based [LW12]. simulations [Jon16]. Simultaneous [AD15, DW19a, KKE17, LM16, Man15, WL15a, WHX19, Zha15, CSJ17, CT14, LL17, MC14, PS18, TKG18, WM15, WY11, XHEM17]. Simultaneously [CMX17, NJdC14]. Singapore [Sha19f, HLN^{+}15]. Single [CSJ15, BS13, BZAZ15, BAZ19, Dog15, Gov17, LY15, LM18b, Özk12, PGF19, PL18, Wan18a, YZWM19, ZL15]. single-index [BAZ19, LM18b, PL18, Wan18a, YZWM19, ZL15]. Singular [JR17, RTM18]. SIR [EKBO16, SZWL19]. six [Sha19g, Rav19]. size [AVG18, AHAA10, AAJ16, AL18, ABA12, BMM15, BA15, BTD18, CC19, CTC17, CM16, DB10, HW17, IP14, KAK16, Moh17, NFFM14, SH10, Su16, TP13, YP10, YPAC11]. size-biased [TP13]. sized [ES11]. SiZer [LM18a]. sizes [FKS10, PF16a, RRB10, RBHSL11, SHP12, VHV^{+}16]. Skew [XZD15, ATON18, AAVG16, ABA12, BMM14, BK16, ÇŞ18a, DYA19, DC13, FLB15, FMB16, FAV18, GLLO14, GEC19, Haj19, HK16b, JTL18, LS19, MAV17, MARS19, SNG12, Stå16, SG15, TdC18, VSG^{+}18, WZX13, Yos18, ZMKAV19]. skew- [Yos18, ATON18]. skew-generalized-normal [FAV18]. Skew-normal [XZD15, AAVG16, ABA12, BMM14, DC13, GLLO14, JTL18, MAV17, MARS19, SG15, VSG^{+}18, WZX13, ZMKAV19]. skew-normality [SNG12]. skew-probit [LS19]. skew-symmetric [Haj19]. skewed [YZL17]. skewness [AK16]. skip [ABJR13a, ABJR13b]. skip-lot [ABJR13a, ABJR13b]. skipped [WRP18]. SkSP [ABJR13a, ABJR13b]. SkSP-2 [ABJR13a, ABJR13b]. slicing [XLZ18]. sliding [RWCD17]. slightly [Juh16]. slippage [BG11a]. slope [STL16, Tsa10a]. slot [GZT14]. slot-parachute [GZT14]. slowly [FPP18]. Small [BG11b, CCS12, FP14, HM13, LXL17, Lon12, MS19, SRP11, Tar12, ZBG18, Bru15, CTC17, EE15, GB18, HM17, LK13, NTC11, NKZ19, QMZ15, RR13a, SHP12, TS14, HT12]. Small-sample [FP14, Lon12, ZBG18, Bru15, EE15]. Smirnov [Fre12, OO12, SY17a]. Smith [NTC11]. Smooth [BMAW14, DTRB11]. Smoothed [BC19, Lin16, MPB19]. smoothing [CJ13, CWM17, JO11, JO12, Ned11, WMDO11, ZAK13]. smoothly [CP12]. Sobol’ [TP15a, CGP15]. social [FAT19, PSS15]. society [LCN^{+}17]. software [ASM^{+}11, BR16]. solution [KL10, SLM16]. solutions [Yao15]. Solving [NHWT14, VS15, BH19, Kha12]. Some [AB18, Kib12, KMS17, Kiz18, Mal16, MGR15, SJZ17, Adk12, ACN^{+}17, ABA16, CGdS14, DDDD10, DC16, Fam12, GZL18, KL12, KP15, KR15a, LC18, PP10, PNN17, QAA18, QKUH18, SF13, SK13, SG15, TKÖ19, YS18, Zie11, ZMW13]. sometimes [Rav19]. sources [SLA17]. Space [Jou15b, ABP16, CJS10, CSJ15, FNRCM17, HLVRS18, KBM19, San12]. Space-filling [Jou15b]. space-time [ABP16]. Sparse [LOK16, SH19, Cha15, DW19a, DS10b, FWZ^{+}15, FP17, JO11, NBB15, NKZ19, PSK14, SML19, WLCL18]. spatial [BT14, BTR16, BL19, BZF18, CS18b, CWZ18, Con10, DL19, DBL10, FRL17, HLL18, JKP19, JTL18, LP16a, Lia10, MN15, NHWT14, RdSF16, RS15, SZM17, TD14, WS16, XWZS15]. spatial-fractional [NHWT14]. spatial-temporal [DBL10]. spatially [BKJ16, WMDO11]. spatio [ACG^{+}16, DRC^{+}16, VBL17]. spatio-temporal [ACG^{+}16, DRC^{+}16, VBL17]. spatiotemporal [DL19]. Spearman [SGZM14, TCM11]. special [Zha17]. specific [Her11, LLV^{+}14, NAA18, Yu11]. specific-to-general [Her11]. specification [YCA15]. specifications [AW17]. specified [AI16, PQ14, Wil15]. spectra [ZB13]. spectral [HLVRS18, LL13, NHWT14]. spectrally [AAG^{+}19]. spectrum [RTM18]. speed [PF16b]. spheres [APF18]. spiked [HOC^{+}19]. spline [HS13, MT17]. splines [ACG^{+}16, RCL15, SESY13]. split [Sta10]. split-plot [Sta10]. sports [Sha11a]. spread [Len11]. SPSS [Sha15e, Sha15e]. spTDyn [BKJ16]. spurious [Kus11]. square [Alt19, ATF19, CM14, LJ18, SLD16]. squared [CF10, Tsa18, VGTGCFC17, Yoo13]. squares [AR16a, Fuk11, Fuk16, HL15, KS16, Lee19, Mal16, MSA12, RBK16, RA17, SESY13, SML19]. Srivastava [HT12]. stability [Gau11, KBL^{+}15, Li15b]. stabilizer [GP19]. stable [CS17, SA15, Zha11]. stage [AJA11b, BL19, BV15b, CC16b, CHQ17, FMV^{+}19, HNPB18, Kuk18, NAA18, RWCD17, TdSC19, TKG18, Wu16b]. Stahel [VVW11]. standard [Dia10, Kel16, LLP^{+}14, LK13, RZ12, Tsa11, Wil11]. standardized [RR13b]. standardizing [LL18a]. standby [JG16]. starting [Rav19]. starts [SA12]. state [Alw17, CJS10, CSJ15, DSVY14, DDD17, FNRCM17, KBM19, SDC12]. state-space [CJS10, CSJ15, FNRCM17, KBM19]. Statement [Ano14c]. states [RRDU13]. station [DRC^{+}16]. stationarity [BD12, dLHT17]. stationary [CHTZ14, KSJ16, NHGS14, SKJ17, Tsa10b, WBGJ15]. Statist [Pak11]. statistic [AHM13, ASA^{+}17, HMP17, HP11, KAA18, KÖ19, MMR16, Mur15, PPK16, Yan10]. Statistical [ASM17, ARY17, AZS15, Aus18, BBA15, BMP12, BBL13, CPK19, CL14, CLL10, FM15b, GLLO14, GV12, KAT15, LYZ11, RB18, Sha10i, SK16, TP13, WWCL11, YCXN14, ZLQ^{+}17, ABA16, AL18, BS12, DB11, FKS10, FHSC14, GIDB15, HBT12, KPM16, KP19, LLV^{+}14, LCN^{+}17, MZZ^{+}19, NMS18, PAFPM12, RGNM13, RB17, SSMF18, SFS15, Sha19e, Sha19g, SOH13, TPM17, TRC^{+}18, WHX^{+}17, ZCW^{+}17]. statistically [HRR^{+}17, HL15]. statisticians [Sha13a]. Statistics [Sha10a, Sha15h, AJFB14, AvR15, AHAA10, Ali12, Ali14, AR16b, BS13, BZAZ15, BT16a, BA15, CS16, CC11a, CP14, CVL18, ES11, Gha16, Man15, MG15, Mur12, NB13b, NB14, PCN14, PGF19, PPGT19, QHB15, Ral17, SAB15, SB12a, SBS14, Sha10j, Sha11a, Sha11c, Sha11d, Sha11e, Sha14d, Sha14a, Sha19g, Sha19q, SK13, SRK13, SAT16, UY12, VDBA14, WyT17, Sha19j]. status [PL18, SK11]. Steady [SDC12, DSVY14]. Steady-state [SDC12, DSVY14]. steepest [FH11, Kia10]. Stein [AMAMS12, AKS^{+}15, ZWDM19]. Stein-type [AMAMS12]. step [ARQ19, Ano14c, AH18b, BH19, Cha14, Dog15, DW18, GK16, IA10, Ism14, LCB13, LCB14, LXL11, XX15, ZS16a]. step-stress [AH18b, DW18, GK16, Ism14, LCB13, LCB14, LXL11, ZS16a]. stepwise [HH15b, KEW13, QMZ15]. stewardship [Sha18a]. stick [Car16]. Stochastic [LLC17, Sha13c, ARB13, Alw17, AKV17, BH18, CK14, EKBO16, FM19, FW13, HA10, Hof12, Jer13, Jon16, Kiz18, LLQ^{+}16, LH14, MBL15, McC14, MS15, MKW16, ÖK18, RPFOMGRM17, SHLT17, WCC13, WSC18, ZLQ^{+}17]. stock [Cox13, YAEU13, Yos18]. storytelling [Sha18a]. Strategies [Gou11, NMOV18, BAT11, LXL17, LPS13, MV14]. strategy [BCLM17, Kia10, Li14, PZY^{+}14]. stratification [Fre16]. stratified [AM19, ARB13, GB17b, KGA12, KTJ18, Özg18, RAB14, SMBS10, UA16, VM19]. streams [QLW16]. strength [AG19, AAS18, Akg19, Bak14, BBA15, CC17b, Hir11, Jou15b, NTZ19, SWZ15]. strengths [SC16]. Stress [Ho16, NTZ19, AHAH14, AHA15, AG19, AAS18, Akg19, Ano14c, AH18b, Bak14, BBA15, CC17b, DW18, ER19, GK16, Hir11, IA10, IAGEK11, Ism14, LCB13, LCB14, LHLB19, LXL11, SWZ15, ZS16a]. Stress-strength [NTZ19, AG19, AAS18, Akg19, Bak14, BBA15, CC17b, Hir11, SWZ15]. strong [FM19]. strongly [SB13]. Structural [SB13, CB11a, CK14, Con10, GHMR19, JY13, LK13, MV14, MSK14, RL15, TD19, VBL17, Yam19]. structural-change [CK14]. Structure [LL18b, WL17, BR16, BVRI16, BH18, FRL17, Gha16, GHH17, LLGP17, WL15a, XWZS15, YTNT14]. Structured [KPK^{+}13, MM13b, Wes16]. structures [Jin15, RdSF16, ZX12]. Student [Ho12, AO12, HY16, LK17]. Student- [Ho12, HY16, LK17]. Studentized [SO10]. studies [DSVY14, EKE^{+}18, Fan17, KL12, NMOV18, QZZ16, ST13, TJK13, TKJ13, TK14]. Study [GB15, XHYX14, AB16a, AY13, BCT16, DS18, Bør15, Bru19, CGSTG18, CYPGGM16, DS11, DP15, DC13, EKE^{+}18, FGHRM12, FVB13, GI17, HZ14, HM18b, IS13, JMM^{+}17, JV14, KTJ18, LF16, LLGP17, LLP^{+}14, LZ11, MMK10, MSS18, Özk12, PO14, PMM18, PRMM12, RASR16, RKV17, RT19, RCL15, SNGMRC16, SKX^{+}18, SK13, TJK13, Wes16, XZY13, YPAC11, YS13, YYW15, ZL11, ZMW13, HLN^{+}15]. subdiffusive [Guo17]. subdistribution [HU14]. subject [GM16, LH14, RNA17, RNSR19, ZLZZ18]. subordinator [Ye16]. subsamples [AMB15]. Subsampling [WZ19]. Subsampling-extrapolation [WZ19]. subset [GL16, Haq19a, HR12]. subspace [AKV17, KLK15, LLBL14, Par12, SML19]. success [LZ11]. successive [She10, sS16, sS19]. Sufficient [AMBP17, XLZ18, Adr18, NX18b, Yoo13]. suitability [Gau10]. Sum [Nad10, DD15, HL15, MMR16, MH17, SY19]. summary [MG15]. sums [CLS^{+}19, FDGD16, GP15a]. superadditive [KAJB19]. supersaturated [BMAW14, CKP16, KEW13]. supervised [Fan19, YP17]. supplemented [MQR18]. supplier [PHCS11, WPC15]. supply [MCW17]. support [ATF19, Bot11, CKPS11, Erd13, KOM19]. sure [ZZ15a]. surface [FH11, Sha10l, WZ13]. surfaces [ODBT15, Sha19u]. surgery [LED16]. surplus [KS17a]. surrogates [DP13]. survey [ARQ19, AD10b, MSS18, Sha19t, UA16]. surveys [ARB13, CCS12, LXL17, LPS13, Lu14, RAB14, ZCZ15]. Survival [AÖ13, Sha15f, Bor17, CCF19, CLDB16, CR18, FVB13, GWH14, HP15, LKM^{+}15, PQ19, She10, SK11, sS16, sS19, TKK19]. surviving [CLDB16]. survivors [Bor17, COBH11]. Survo [VS15]. SV40 [TNM17]. swarm [LL18b, WH14, WyT17]. Swerdlow [Sha19j]. switching [Haj19, RA14, Yao15, YTNT14]. symbolic [dALNCdATdC11]. Symmetric [JW11, AD12, CCC10, Haj19, MN19, VP16]. Symmetrical [WC17]. symmetry [AE17, Alt19, AK16, BMPZ14, CB13, DC16, MAEP14, PV19]. synthetic [CS13, HK16a, HWA^{+}14, MHA10, NTC^{+}19, RNH19]. synthetic-np [HWA^{+}14]. System [CYRO18, AHH17, BU17, CM16, FF14, Gun18, LLV^{+}14, LTJB18, PS14, RBC^{+}15, SRK13, SZS14, WWB18]. systematic [NJ13, SA12]. systems [AAG^{+}19, BCO13, EXH16, HC17, JG16, Kiz18, KBL^{+}15, Nom14, WZS17, WSLX17, YQ19, ZLQ^{+}17, Sha19g].

table [Alt19]. tables [BS14, KKM16, NBB15, NKZ19]. tagging [YPAC11]. Taguchi [HL15]. Tail [GRPP10, Bak18, GP15a, Gha19, MARS19, MAEP14, Stå16]. tailed [CG15, CCZJ17, DYA19, SL17, TdC18]. tails [LPSR16]. Taylor [Chr15]. TCP [LPY18]. technique [Bør15, dCFOM12]. techniques [GZL18, HRR^{+}17, JY14, MHA10, OMY12, ST13, SLSW15, SF12, SKTC11, VC15]. telephone [LPS13]. tempered [Zha11]. template [WH17]. template-based [WH17]. Temporal [DDD17, ACG^{+}16, BKJ16, DRC^{+}16, DBL10, Kus11, MS19, VBL17]. temporary [TC10]. term [Bor17, COBH11, DLZ19, Kuk19, SGB13]. terms [ÇŞ18a, MMP15, DMV17, YPL13]. tessellations [NHGS14]. Test [BRY17, AE11, Ali12, Ali14, AK16, AP17, ABA16, AH18b, BTR16, BPH12, BT16a, BSS15, CF10, CL16, CWZ18, CB13, CLYX18, DD15, DVA15, DS14, DS15, DYT10, DT13, DGK12, DYX15, DLS18, DW18, ES11, Fre12, GFS15, Gau10, GEC19, Han10, Hir11, HMP17, HW17, HKKL17, IAGEK11, Ism14, JL16, KE10, KM17, KL18a, KEW13, Kus11, LL18a, LK17, LPK18, Lee19, LN13, Li11, LL16, LLY19, LCB13, LCB14, LXL11, LZ10, LLB12, MK12, MP15, MY13, Mog11, MSK14, NO10, NBB15, NKZ19, NJdC14, NdCOP15, NNB14, OWKC15, PHCS11, PP15, PPK16, PP19, QYX17, RNA17, SJN15, SNG12, SY17a, SK13, SHP12, SZS16, SDWL17, SLLZ18, TQP10, WPXL14, Wan15, WRN18, WPC15, WH19, XZD15, Xu17a, YXZ19, ZA11, ZST15]. test-based [SLLZ18]. Testing [Adk12, CNS12, CNP19, Gha19, KAWA12, KL17c, LW14, Li15a, LLM16, Mar14a, Mei11, Mug16, NA11c, NA11d, PB13b, PZ10, PA15, PGT11, PBM16, RG10, RL15, SU18, WLSZ18, Wel16, Xu17b, YK15, Yam19, ZA12a, Zam15, Zar19, ZNW19, Ali12, ASS11, AL18, BG11a, BMPZ14, BV15b, ÇŞ18a, CNQ14, CNL17, DD15, DS14, ER19, EFGMD13, GEV18, GHMR19, GV18, Ho16, HAB12, Hwa11, KKM16, KR16, LDB10, DDJ16, Li15b, LXZ11, LYLM17, LVB18, LZ10, MA10, MS17, MFP17, MAEP14, Mur15, Nou10, NJ18, RR13a, SB12a, dSSCR19, Say12, Sha15a, SK13, SK17, TD13, TCM11, TPW17, WC14, Xie14]. Tests [AvR15, AB14, BET16, CCP12, FH16, LP17, NP16, SC12, TS16, AF17, AE17, AS15a, AOH16, AS15b, AS15c, AD10b, Ano14c, AP12, AB18, AJA11a, AJA11b, AKJ16, BG11a, BMPZ14, BMP12, BD12, CRV15, Che14, Chi10, CGSTG18, CL10, DTRB11, DDDD10, DY16, DGK12, DC19, DC15, DC16, FS15b, GBCS16, HT12, IPK10, Ism10, IA10, JGPF17, JZZH18, JV14, Kel16, KK12, Lee11, LLS13, LL15a, LS11, LM16, LBN^{+}19, LHLB19, Llo10, LR18a, LL13, Mar11, Med16, ME15, MFP14, MY13, MGR15, NB18, NJdC14, NA11b, NA11a, NA13, Nou17a, Nou17b, Nou19, PB13a, PS19a, PS19b, PP15, PV19, RRV13, RNSR19, RDC10, SEGMA19, SM11, She15a, SW18, SK17, SC19, SO10, Ten19, TMG18, TGL12b, VK14, WZ13, WM12, WD16, Xie14, XMC^{+}14, Yan10]. tests [YS11, YS18, Yoo13, ZHH19, ZCL19, ZS16a, ZX14, ZS16b, Zör15]. th [Yoo13]. th-moment [Yoo13]. their [Bru15, GZB13, Kib12, LL17, Lüt15, YZX18]. them [MGR15]. theoretical [SLM16]. theory [AvR15, CCK13, COP14, COS14, GZT14, JG16, POCdP13, Sha10h, TA11]. there [Wil15]. Thompson [PP10, YS19]. three [ATF19, BB12, BS16, Car19, CB19, GMG13, KPM16, KR15b, MWL14, MS18b, NB13a, NB13b, NB14, NB15a, NB16, NLHD12, NAA18, sS19, Tza11, Xie14, vGK17]. three-decision [GMG13]. three-dimensional [MWL14]. three-group [BS16]. three-level [ATF19, KPM16]. three-mode [vGK17]. three-parameter [BB12, KR15b, NB13a, NB13b, NB14, NB15a, NB16, NLHD12]. three-period [CB19]. three-stage [NAA18]. Threshold [JKP19, YWL18, FW13, LLS13, RBC^{+}15]. Thresholding [YWZ18]. Thresholding-based [YWZ18]. throughput [SH19]. thumb [MAEP14]. tightness [LLP^{+}14]. Time [JY14, Sha15i, TZA10, YAEU13, ZMKAV19, Ahm16, ABP16, ABGM18, AB18, AKJ16, AH18b, BPH12, BS19, CHTZ14, CCP12, CB11a, CL16, CC11b, CWM17, CF15, Dog15, DP13, FNRCM17, FPP18, Her11, HZ14, HNPB18, Jin15, JCW19, Kan17, KA13, KL18a, KF16, KBM19, LP17, LZGW14, LdNdSF18, LH14, LPY18, LSF^{+}17, MN19, MC16, MBL15, MV14, MTS14, Moi17a, NSMFR15, PV17a, Par12, PMM18, PSY18, RWCD17, RRV13, RTM18, RCL15, Sha19a, Sha10k, SM18, SK18, SRK13, SB12b, SR16, SC12, SKJ17, SO10, TD13, TS14, TKK19, WGC14, Wei11, WJ19, YK15, YCD15, YYG16, YWL18, YD16, ZNW19, dSdS17, Sha15i]. time-between-events [CC11b]. time-dependent [DP13, KA13, SRK13, YD16]. time-domain [RRV13]. time-scale [WGC14]. Time-series [YAEU13, TS14]. time-to-event [HNPB18, MBL15, PV17a, RCL15, YYG16]. time-to-failure [SK18]. Time-varying [JY14, CWM17, FNRCM17, FPP18, LdNdSF18, LSF^{+}17]. times [ARY16, AAGV12, DAB11, GP15b, KAR13, KN15, KN16, Lee17, NK15, She10, sS16, sS19]. tips [VNM14]. tissue [SYL^{+}14]. Tobit [AY15, ZL15]. Toh [Sha19f]. tolerance [KL12, KP15, PO14, QKY16, You14]. tool [LYL17]. tools [ASM^{+}11, BSS17, VRC13]. topology [KBL^{+}15]. Topp [Akg19, BCY^{+}17, Gen13, Sha18c]. total [NAA18]. track [LX14]. trade [FFP16]. tradeoff [RB18]. tradeoffs [NB15b]. trained [FF14]. trajectory [Alw17]. transfer [Par11]. transform [DGLV17, OvP10]. transformation [CWM17, She14a, She15b]. transformations [GEC19, MP16, WGC14]. transformed [DS14, NA11d]. transition [HEB13, dLHT17]. transitions [MMP12]. translation [CC12]. translation-invariant [CC12]. transmission [KBL^{+}15]. transmuted [GLB17]. transparent [SHLT17]. Treatment [SJ10, CB19, DM12, LR18a, NA13, TKJ13, TK16]. treatments [BD17, GMG13, HM17, MMP15, Sha19b, TKJ13, TK16]. tree [CJ13, Cha14, LLQ^{+}16, MES19, TK18, YPAC11]. tree-based [CJ13]. trees [AP12, PRNG18]. trend [BT16a, FPP18, JK17, ZNW19]. trends [BD12]. trial [BD17, LN13]. trials [Bak18, BS16, JMM^{+}17, Llo10, Sha14e, YYW15]. triangular [KC14]. tricks [VNM14]. trigonometric [FDGD16]. trimmed [AR16a, CC12, RBK16, RA17, XWMA18]. trivariate [HL13]. tropospheric [PAFPM12]. true [Hwa11, QYX17]. truncated [AJA11b, AKJ16, ABA16, CTX18, CHT16, GB17a, GBCS16, sS11, She12, She14a, She14b, She14c, She15a]. Truncation [DDJ16, AOR13, NFFM12, She10, sS16, sS19]. truncation-based [AOR13]. trust [LYL17, PSS15]. trusted [PSS15]. Tsallis [Zar19]. Tuck [Sha19f]. Tukey [BT16b, RNH19]. Tuning [FWF16, AS15b, HM18b, Kum15, PSK14, PK16a, Par17, Ten19]. Tuning-parameter [FWF16]. Turkish [TKK19]. tutorial [TD14]. Tweedie [BK17b]. Twenty20 [PDS16]. twisting [GZL18]. Two [ARQ19, ASS11, CC16b, CWZ18, QYX17, SBS14, Sha11e, TKG18, WSC18, AF17, ADA18, AHAA10, ASH16, AJA11b, AD12, Bak14, BD17, BRY17, BCO13, BV15b, CCMGA14, ÇK19, CR18, CB19, CHQ17, Che19, Chi10, CNP19, DAB11, DB11, EKE^{+}18, FKS10, FMV^{+}19, GSW17, GHRAM13, GOC18, GGM18, HA13, HY14, HNPB18, JP17, JKM16, Jin15, JCW19, Jou15b, KKM16, KL13a, Kuk18, KTJ18, Lee11, LBN^{+}19, LLBL14, LL13, LLB12, Mar11, MMP15, MS15, Mur12, Mur15, NBB15, ÖK17, ÖK18, PO14, PA15, Per10, PP15, PP19, PPGT19, PL16b, PK11, PB15, Rak16a, RNSR19, SBAA14, Sha15a, SKX^{+}18, SBK13, Sta10, SZ16, SAM13b, SAM13a, SR11, TdSC19, TS16, TD13, TMG18, WH14, WZS17, WM12, WEHCC14, WY11, WWCL11, WC14, WL14, WPC15, Zha15]. two-arm [SKX^{+}18]. two-component [WZS17]. two-dimensional [LLBL14]. two-factor [MS15]. two-level [Sta10]. two-parameter [ADA18, ASH16, Bak14, CCMGA14, ÇK19, DAB11, DB11, GSW17, HY14, PK11, WY11, WWCL11, WC14]. two-parameter-weighted [ÖK18]. two-phase [EKE^{+}18]. Two-sample [CWZ18, SBS14, WSC18, AF17, AHAA10, Che19, Lee11, Mur12, Mur15, PP15, PP19, PL16b]. two-sex [GGM18]. two-sided [PO14, Rak16a, WL14, WPC15]. Two-stage [CC16b, TKG18, AJA11b, BV15b, CHQ17, FMV^{+}19, HNPB18, Kuk18, TdSC19]. Two-step [ARQ19]. two-treatment [CB19]. two-way [KKM16, NBB15, Zha15]. Type [ARY16, Ano14c, PS16, SAM13b, AHA15, AHH17, AP19, ARY15, AG19, AMAMS12, AD10a, AB16b, ABJR13a, ABJR13b, BMPZ14, BAS17, Car19, CM17, CPK19, DSE16, DNMT18, DC16, ER19, GWX14, HMP17, HW17, Ili15, Kib12, KP18, LHLB19, MM13b, MSS14, Mar11, MTS14, MKSN18, NK15, PB17, PK16a, RT14, SBAA14, SWZ15, SAK14, TS17, TN16, TP15b, WD16, WW12, WC14, XYZ19, YS19, AHAH14, AJM11, AYR16, AR16b, ABJR13a, ABJR13b, AL18, BDKM11, BS13, BZ16a, BB12, BBA15, BMP12, ÇŞ18a, CYL17, CAM16, CBG16, DBC12, DYT10, DT13, DC19, EDASM17, FM15b, GAB14, GP15b, Ism10, IA10, IAGEK11, Ism14, KR18, KM12, KK13, LCB13, LCB14, LXL11, MY13, Moi17b, NTZ19, NB15a, NB16, NLHD12, Nou17b, PB13a, PB13b, SL17, SKK12, SBS14, Sha16, SBK13]. Type [SK16, SEAEM13, SZ16, TCLY14, TRC^{+}18, Wu10, WY11, WLL12, YCXN14]. Type-2 [SAM13b]. type-I [BAS17, Ili15, LHLB19, MSS14, TP15b, AYR16, BMP12, Ism10, IAGEK11, Ism14, LCB13, LCB14, LXL11, PB13b, SZ16]. Type-I-censored [TCLY14]. Type-II [Ano14c, ARY16, AHA15, AHH17, ARY15, AG19, AB16b, CM17, CPK19, DNMT18, ER19, GWX14, HW17, KP18, SBAA14, SAK14, TN16, XYZ19, AHAH14, AR16b, BS13, BZ16a, BB12, BBA15, CYL17, CBG16, DBC12, DC19, EDASM17, GAB14, GP15b, IA10, KR18, KM12, NTZ19, NB15a, NB16, NLHD12, PB13a, SKK12, SBS14, Sha16, SBK13, SK16, YCXN14]. types [AD16a, GZL18, YS11]. Typical [NSMFR15, GZL18].

Uhlenbeck [HKST17, Zha11]. ultra [Che18, LC18]. ultra-high [Che18]. ultra-high-dimensional [LC18]. ultrahigh [MCZ17, PZZ19, YHWS18]. ultrahigh-dimensional [MCZ17, YHWS18]. unadjusted [TK14]. Unbalanced [VHV^{+}16, BEBG14, CKP16, KLKK19, VFLR10, XMC^{+}14, YML14, Zha15]. unbiased [BDKM11, CTX18, CC11b, GWX14, HP11, Wu16a]. Uncertainty [ABM17, Sha15j, PC10, WJ17]. unconditional [LR18a, LLB12, Sha15a, SKX^{+}18]. under- [KC16]. under-dispersed [SMM19]. underdispersed [OJRACL18]. underground [ASM17]. understand [Sha15e]. undirected [OCT18]. unequal [BK16, CCP12, PA15, RRB10, RBHSL11, Zha15]. unidirectional [SXTJ17]. unified [PS16]. uniform [Tak17, AKA^{+}16]. Uniform-Geometric [AKA^{+}16]. uniformity [Zam15]. unimodal [BBW17, SEGMA19]. unique [Sha10j, Sha10k, Sha10l]. uniqueness [NB16]. Unit [GHJC10, SE11, CL16, MK12, MS19, PS14, ZS16b]. unit-level [MS19]. units [ARY16]. univariate [Bot11, CF10, RDC10]. Universal [BP15, Tso15, Tso16]. unknown [AE11, AWJ^{+}13, Cha16, Fre12, HCT19, JP17, JK19, MQR18, MAEP14, NB13b, NB14, RRDU13]. unobserved [Bør15]. unreliable [LH14]. unreliable-node [LH14]. unreplicated [KKM16]. unrestricted [Özk12, ZMKAV19]. unstable [Alw17, Juh16]. Unsupervised [Cha16, LAGM11]. update [CGSTG18]. Updated [RR13b]. updating [SC16]. upper [JR16, KL17c, RASR16, WEHCC14]. Use [Kha12, DM16, Ho12, RCL15, Sha19f, TCM11, TGL12a]. useful [SB12a]. Using [BAT11, CP14, GYV^{+}13, KSLN^{+}18, LAGM11, TKJ13, Wu11, AHM13, AvR15, AMAMS12, AOR13, AOH16, AY15, AKJ16, AAJ16, ASA^{+}17, AHJ16, Aus18, BCY16, BKJ16, BCLM17, BS19, CCZJ17, CDG^{+}15, CMCH12, CW16, CWZ18, Chr15, DS11, DRAR19, DTZZ12, DS14, DR17, DK17, DSVY14, EXH16, Erd13, FR15, FP15b, FHO15, Fuk11, FMK15, GZL18, GY16, GVW17, GdCCDS18, GG17, HK16a, HK18a, HEB13, HK16b, HN11, HSR15, Hir11, HP11, HWWZ16, HS13, IJL18, IS13, JK14, JY14, JGPF17, KAK16, KKM16, KK14, KTJ18, Kus11, LAuaS^{+}15, LW14, LW17, LZ11, MMR16, MK12, MP15, MBL15, McL14, MSS18, MPB19, MMP12, MKSN18, MADASAM11, NLK11, NA11a, NA11d, Pam19, PF16a, QPQ18, QKUH18, Rak16a, RdSF16, RM19, RMS14, SS12a, SRP11, San12, SB15, SD15, SRAO11, SK18, SSK13]. using [STS14, SEAEM13, Stå16, SWLZ15, Su16, TPM17, TKÖ19, TS14, TNM17, VS15, WCC13, WTJW17, WWW17, Wil15, XZY13, XWMA18, YR15, Yu11, YZWM19, YTNT14, Zar19, ZP14, ZCW^{+}17, ZAK13, dB15]. uterine [Bor17]. utility [GZL18]. utility-based [GZL18]. utilizing [MFD16].

Vacuolating [TNM17]. Validation [ME15, AÖ16, Bai16, HN13, JYML13, VAW15]. Value [LL11, THR18, AB16b, AA16, DYX15, DDD17, FP14, Gha19, LCZ18, LXZ11, MP15, OO12, OSN17, PV17b, PPK16, PF16b, YCXN14, Moi17b]. Value-at-Risk [THR18, LL11]. valued [APF18, HKL17, KL17b, KBM19, Lee19, dALNCdATdC11, SWLZ15, WWW17]. values [AMB12, CH19, CC16a, CH14, DK17, DAB11, DI11, HKK^{+}16, Kel16, KN15, KN16, Kiz17, LS14, LLWY15, MMP12, NK15, PRNG18, RG10, RR13b, RASR16, SAB15, SLV18, SKX^{+}18, Sha10e]. VaR [FGHRM12]. variability [MH17, SLLZ18]. Variable [AYJ11, DK10, PZZ19, TX14, WDCK15, WZX13, YR15, ZFZQ18, ZF11, AAJ16, Bag11, BCT16, DS18, CCF19, CTC17, CNS12, DW19a, DP15, DTZZ12, EZ12, FWZ^{+}15, GA15, GG17, Haq19b, HH15b, KAK16, KSLN^{+}18, LAuaS^{+}15, LC19, LYQ^{+}15, Lia14, LHB11, LHB13, LWW18, NKB19, PS18, SFS15, SU11, VNM14, WL15a, WL17, XX15, YLG15, YCA15, ZX12]. variable-interval [DTZZ12]. variables [Adk12, AHC15, ARQ19, AD16a, AWJ^{+}13, AHJ16, BU17, BP15, CGP15, CLS^{+}19, Dia10, Fan19, FBV18, GSL^{+}14, GG11, GBCS16, HOC^{+}19, HSC16, KTJ18, LC18, MHA10, McL14, MSS18, DMV17, Nad10, NX18a, dALNCdATdC11, PQ14, RG10, RIY18, SD15, SG18, Sun11, TZSP19, TX14, TK16, WW19, WCW15, YYW15]. Variance [SGG18, WCK11, WWB18, ZP14, Are12, AWJ^{+}13, AL18, AA16, BEBG14, Bur14, CTX18, CGSTG18, FKM13, JZZH18, KLKK19, LP17, LZ10, Lu14, MSK14, MKSN18, NB15b, OWKC15, PP10, PJR15, QRH19, RdSF16, SA12, SR11, SGB13, TD13, VGTGCFC17, WL16, YA16a]. Variance-based [WWB18]. variances [DVA15, GGSS19, IPK10, KLK17, KLKK19, RRB10, RBHSL11, SK13, SHST13, TdC18, Zha15]. variants [BL19]. variate [Tsa11]. variates [AD12, Ber12, Chi10]. variation [CS13, Hay15, LXZ11, LM18b, MFR^{+}18, NTC^{+}19, RA14, SW18, WL16]. variational [GDR12]. variations [DRAR19]. various [ASM^{+}11, Moi17a, Sta10, Tu19, YP10, YS11]. varying [EB19, BKJ16, CWM17, FNRCM17, FPP18, Gou11, HLN^{+}15, JY14, LYZ11, LM18a, LdNdSF18, LSF^{+}17, LZZM18, MCZ17, MMK14, PB15, QMZ16, SCW16, WL15a, XZD15, YLG15]. varying-coefficient [HLN^{+}15, LYZ11, MMK14, YLG15]. vector [ATF19, AR16b, CTTS19, CKPS11, Erd13, Jou15a, Kha12, KOM19, LJL15, MB16, PV17a, Rai12, WN11c]. vectors [Ahm19, HBC11, LL16, RR13a, TS16]. version [CLAH17, KR15a, KS17c, NHGS14, RR13b]. versus [FH15, GC17b, PWW15, SBD10]. very [PB12, SH10]. via [ASM^{+}11, AWH19, AP12, AGM15, BMM15, BSS17, BBW17, BZ16b, CQJ12, CL13, CK14, CPP^{+}19, Coi13, DW19a, DKY17, DDD17, FWZ^{+}15, GI17, HZSA19, Jin15, KÖ19, LWW18, MMK14, MM15, MSA12, PSK14, PK16a, SPS19, SDWL17, UG10, VK14, WZ13, Wil10, WEHCC14, Wu13, XLW10, YZ15, ZY15]. view [RS15]. violation [Aus18]. violations [AW14]. violators [DHP14]. virus [TNM17]. Visual [OO12]. visualization [SHLT17]. visualizing [SHLT17]. volatilities [THR17]. volatility [ATON18, BH18, FM19, FW13, JK17, Jer13, MS15, WCC13, Zie11]. volume [HR12]. VP [FHSC14]. vs. [HA10, NJ13, YPAC11]. VSI [NTG13, TPM17]. VSIX [ASBM19]. vulnerability [IGR13].

Wald [FHSC14, HMP17]. Wald-type [HMP17]. Walsh [MM15]. warning [HP11]. Warp [Ral17]. warping [SYL^{+}14]. Wasserstein [AGA18]. water [YFT10]. Wavelet [JK17, SA15, WKJ^{+}19, AI12, Bru15, DF14, LS11, RL15, SR16]. Wavelet-based [SA15, WKJ^{+}19, DF14, RL15]. wavelets [GG11]. way [BEBG14, ÇŞ18a, GGSS19, KKM16, KLKK19, LZ10, NBB15, PO14, RZ13a, YML14, Zha15]. Weibull [Pak11, AP15, AVK15, BSdMC11, BBT13, CGdSO13, CNO13, COL14, COS14, DA14, DYT10, HP11, HW12, Ism14, JG10, KLK14, LLGP17, Muh16, NCO11, NMS18, NB16, NLHD12, OWLP16, Pak10, PK16b, RAJ16, SRG11, SDS16, SY15, SY17b, SK16, SAM13b, SAM13a, SAK14, TLRB14, TP13, UG10, UGMK13, VRC13, WXF11, WH14, WM12, ZS16a, dCOC16]. Weibull-geometric [BSdMC11, BBT13]. Weighted [DAP15, Haq19b, Aci18, Ali15, AB18, BS19, CHB18, CC11a, CLS^{+}19, FH15, GBdL16, GSW17, Haq14, HBMAO15, HH15a, HL15, HXT15, KS16, Lee11, LPJ14, LZZ^{+}15, MMR16, MG17, Med16, NJdC14, NHA18, ÖK18, PB17, RNA17, RS14, Saz19, She14c, SZS16, SLLZ18, WL17, WLL12]. weighting [SCW16]. weights [ARQ19, VVW11]. whether [Sta10]. Whitney [AS15c]. wide [TMG18, XZY13]. width [LZ11]. Wiener [LSF^{+}17, PBSZ13, WGC14, WBG15]. Wilcoxon [AS15c, DD15, MMR16, PP15]. wild [FMK15, HA10]. Wilk [GEC19]. Wilks [PPGT19]. win [Hay15]. win-probabilities [Hay15]. WinBUGS [CQJ12]. wind [MC16, PF16b]. window [RWCD17, ZST15]. window-based [RWCD17]. window-limited [ZST15]. windowed [LCZ18]. wise [Kim18]. Within [Wil15]. Workbook [Sha19j]. worked [Sha19d]. workers [FF14]. working [Sha13a]. World [Sha19f, Sha15c]. wrapped [ABP16, SRG11]. wrapped-Gaussian [ABP16]. wrong [RSD14]. Wu [Sha19i].

X [TZC^{+}16]. XII [AJM11, ABJR13a, ABJR13b, BAS17, CYRO18, DA16, NK15, PS16, POCdP13, SSMF18, SEAEM13]. xix [Sha19f]. xv [Sha19i].

yield [WW19]. Yule [LRV17].

Zero [CLAH17, HOC^{+}19, PCMMA13, SY19, AWH19, FI17, HK18b, KNM^{+}15, KL17a, KR15a, Li11, LLY19, LLBL14, LWT^{+}17, MH16, MGR15, SB15, SMM19, Sha15b, ZCL19]. Zero-inflated [PCMMA13, SY19, AWH19, FI17, HK18b, KNM^{+}15, KL17a, KR15a, Li11, LLY19, LWT^{+}17, MH16, MGR15, SMM19, ZCL19]. zero-mean [LLBL14]. Zero-modified [CLAH17]. zero-or-one [SB15]. Zero-spiked [HOC^{+}19]. zeros [BBM18, HK16b, LWS19, MH16, Mog11]. ziggurat [McF16]. zigzag [PM10]. Zipf [Zör15]. ZISCMP [SY19].

References

Artiach:2011:EFC

[AA11] Miguel Artiach and Josu Arteche. Estimation of the frequency in cyclical long-memory series. Journal of Statistical Computation and Simulation, 81(11):1627–1639, 2011.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aydogdu:2016:CMV

[AA16] Halil Aydogdu and Ömer Altindag. Computation of the mean value and variance functions in geometric process. Journal of Statistical Computation and Simulation, 86(5):986–995, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

URL http://www.tandfonline.com/doi/abs/10.1080/
00949655.2015.1047778.

Amin:2019:IDG

[AAAQ19] Muhammad Amin, Muhammad Amanullah, Muhammad Aslam, and Muhammad Qasim. Influence diagnostics in gamma ridge regression model. Journal of Statistical Computation and Simulation, 89(3):536–556, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alizadeh:2018:OPC

[AACR18] Morad Alizadeh, Emrah Altun, Gauss M. Cordeiro, and Mahdi Rasekhi. The odd power Cauchy family of distributions: properties, regression models and applications. Journal of Statistical Computation and Simulation, 88(4):785– 807, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alin:2019:RMD

[AAG^{+}19] Aylin Alin, Claudio Agostinelli, Georgi Gergov, Plamen Katsarov, and Yahya Al-Degs. Robust multivariate diagnostics for PLSR and application on high dimensional spectrally overlapped drug systems. Journal of Statistical Computation and Simulation, 89(6):966–984, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Asgharzadeh:2012:RPF

[AAGV12] A. Asgharzadeh, Jafar Ahmadi, Z. Mirzazadeh Ganji, and R. Valiollahi. Reconstruction of the past failure times for the proportional reversed hazard rate model. Journal of Statistical Computation and Simulation, 82(3):475–489, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Albarracin:2019:GAM

[AAH19] Orlando Yesid Esparza Albarracin, Airlane Pereira Alencar, and Linda Lee Ho. Generalized autoregressive and moving average models: multicollinearity, interpretation and a new modified model. Journal of Statistical Computation and Simulation, 89(10):1819–1840, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslam:2016:NVS

[AAJ16] Muhammad Aslam, Osama H. Arif, and Chi-Hyuck Jun. A new variable sample size control chart using MDS sampling. Journal of Statistical Computation and Simulation, 86(18): 3620–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abouammoh:2015:NGL

[AAR15] A. M. Abouammoh, Arwa M. Alshangiti, and I. E. Ragab. A new generalized Lindley distribution. Journal of Statistical Computation and Simulation, 85(18):3662–3678, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akgul:2018:ISS

[AAS18] Fatma Gül Akgül, Sükrü Acitas, and Birdal Senoglu. Inferences on stress-strength reliability based on ranked set sampling data in case of Lindley distribution. Journal of Statistical Computation and Simulation, 88(15):3018–3032, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arrue:2016:BRM

[AAVG16] Jaime Arrué, Reinaldo B. Arellano-Valle, and Héctor W. Gómez. Bias reduction of maximum likelihood estimates for a modified skew-normal distribution. Journal of Statistical Computation and Simulation, 86(15):2967–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aydin:2019:ESR

[AAY19] Dursun Aydin, S. Ejaz Ahmed, and Ersin Yilmaz. Estimation of semiparametric regression model with right-censored high-dimensional data. Journal of Statistical Computation and Simulation, 89(6):985–1004, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alvarez:2010:CPB

[AB10] Susana Alvarez and J. Samuel Baixauli. Coverage properties of beta estimated prediction intervals for multimodal recovery rates. Journal of Statistical Computation and Simulation, 80(1):111–117, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmadi:2011:PMC

[AB11] Jafar Ahmadi and N. Balakrishnan. On Pitman’s measure of closeness of k-records. Journal of Statistical Computation and Simulation, 81(4):497–509, 2011. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

An:2013:NMC

[AB13] Xinming An and Peter M. Bentler. Nesting Monte Carlo EM for high-dimensional item factor analysis. Journal of Statistical Computation and Simulation, 83(1):25–36, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anis:2014:TEA

[AB14] M. Z. Anis and Kinjal Basu. Tests for exponentiality against NBUE alternatives: a Monte Carlo comparison. Journal of Statistical Computation and Simulation, 84(2):231–247, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alin:2015:RCR

[AB15] Aylin Alin and Ayanendranath Basu. Robust confidence regions for multinomial probabilities. Journal of Statistical Computation and Simulation, 85(3):538–551, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alexander:2016:MCS

[AB16a] Blair Alexander and Robert Breunig. A Monte Carlo study of bias corrections for panel probit models. Journal of Statistical Computation and Simulation, 86(1):74–90, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alma:2016:EEV

[AB16b] Ö. Gürünlü Alma and R. Arabi Belaghi. On the estimation of the extreme value and normal distribution parameters based on progressive type-II hybrid-censored data. Journal of Statistical Computation and Simulation, 86(3):569–596, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anvar:2018:SWM

[AB18] P. Muhammed Anvar and N. Balakrishna. Some weighted mixed portmanteau tests for diagnostic checking in linear time series models. Journal of Statistical Computation and Simulation, 88(15):3000–3017, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Azevedo:2012:PRS

[ABA12] Caio L. N. Azevedo, Heleno Bolfarine, and Dalton F. Andrade. Parameter recovery for a skew-normal IRT model under a Bayesian approach: hierarchical framework, prior and kernel sensitivity and sample size. Journal of Statistical Computation and Simulation, 82(11):1679–1699, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslam:2016:IDA

[ABA16] Muhammad Aslam, S. Balamurali, and Touqeer Arif. Improved double acceptance sampling plan based on truncated life test for some popular statistical distributions. Journal of Statistical Computation and Simulation, 86(3):477–493, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alonso:2018:BCT

[ABGM18] Andrés M. Alonso, Guadalupe Bastos, and Carolina García-Martos. Bias correction for time series factor models. Journal of Statistical Computation and Simulation, 88(8):1576– 1602, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslam:2013:ODSa

[ABJR13a] Muhammad Aslam, Saminathan Balamurali, Chi-Hyuck Jun, and Mujahid Rasool. Optimal designing of skip-lot sampling plan of type SkSP-2 with group acceptance sampling plan as reference plan under Burr-Type XII distribution. Journal of Statistical Computation and Simulation, 83 (1):37–51, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslam:2013:ODSb

[ABJR13b] Muhammad Aslam, Saminathan Balamurali, Chi-Hyuck Jun, and Mujahid Rasool. Optimal designing of skip-lot sampling plan of type SkSP-2 with group acceptance sampling plan as reference plan under Burr-Type XII distribution. Journal of Statistical Computation and Simulation, 83 (1):202, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anan:2017:UEH

[ABM17] Orasa Anan, Dankmar Böhning, and Antonello Maruotti. Uncertainty estimation in heterogeneous capture-recapture count data. Journal of Statistical Computation and Simulation, 87(10):2094–2114, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alegria:2016:LBI

[ABP16] Alfredo Alegría, Moreno Bevilacqua, and Emilio Porcu. Likelihood-based inference for multivariate space-time wrapped-Gaussian fields. Journal of Statistical Computation and Simulation, 86(13):2583–2597, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Altieri:2016:BPS

[ACG^{+}16] L. Altieri, D. Cocchi, F. Greco, J. B. Illian, and E. M. Scott. Bayesian P-splines and advanced computing in R for a changepoint analysis on spatio-temporal point processes. Journal of Statistical Computation and Simulation, 86(13):2531–2545, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Acitas:2018:NWD

[Aci18] S. Acitas. A new weighted distribution as an extension of the generalized half-normal distribution with applications. Journal of Statistical Computation and Simulation, 88(12):2325–2341, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alizadeh:2017:OBG

[ACN^{+}17] Morad Alizadeh, Gauss M. Cordeiro, Abraão D. C. Nascimento, Maria do Carmo S. Lima, and Edwin M. M. Ortega. Odd–Burr generalized family of distributions with some applications. Journal of Statistical Computation and Simulation, 87(2):367–389, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akdeniz:2010:LTE

[AD10a] Fikri Akdeniz and Esra Akdeniz Duran. Liu-type estimator in semiparametric regression models. Journal of Statistical Computation and Simulation, 80(8):853–871, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anis:2010:RTE

[AD10b] M. Z. Anis and Somak Dutta. Recent tests of exponentiality against IFR alternatives: a survey. Journal of Statistical Computation and Simulation, 80(12):1373–1387, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Augustyniak:2012:ILS

[AD12] Maciej Augustyniak and Louis G. Doray. Inference for a leptokurtic symmetric family of distributions represented by the difference of two gamma variates. Journal of Statistical Computation and Simulation, 82(11):1621–1634, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amatya:2015:SGM

[AD15] Anup Amatya and Hakan Demirtas. Simultaneous generation of multivariate mixed data with Poisson and normal marginals. Journal of Statistical Computation and Simulation, 85(15):3129–3139, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amatya:2016:CGM

[AD16a] Anup Amatya and Hakan Demirtas. Concurrent generation of multivariate mixed data with variables of dissimilar types.

Journal of Statistical Computation and Simulation, 86(18): 3595–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Atlekhani:2016:IUR

[AD16b] Mohammad Atlekhani and Mahdi Doostparast. Inference under random maintenance policies with a polynomial form as the repair alert function. Journal of Statistical Computation and Simulation, 86(7):1415–1429, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmadi:2018:BCS

[ADA18] Mohammad Vali Ahmadi, Mahdi Doostparast, and Jafar Ahmadi. Block censoring scheme with two-parameter exponential distribution. Journal of Statistical Computation and Simulation, 88(7):1229–1251, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adkins:2012:TPS

[Adk12] Lee C. Adkins. Testing parameter significance in instrumental variables probit estimators: some simulation. Journal of Statistical Computation and Simulation, 82(10):1415–1436, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adragni:2018:MAD

[Adr18] Kofi P. Adragni. Minimum average deviance estimation for sufficient dimension reduction. Journal of Statistical Computation and Simulation, 88(3):411–431, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akdeniz:2015:EGD

[ADRA15] Fikri Akdeniz, Esra Akdeniz Duran, Mahdi Roozbeh, and Mohammad Arashi. Efficiency of the generalized differencebased Liu estimators in semiparametric regression models with correlated errors. Journal of Statistical Computation and Simulation, 85(1):147–165, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abd-Elfattah:2011:GFT

[AE11] A. M. Abd-Elfattah. Goodness of fit test for the generalized Rayleigh distribution with unknown parameters. Journal of Statistical Computation and Simulation, 81(3):357–366, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abd-Elfattah:2017:BST

[AE17] Ehab F. Abd-Elfattah. Bivariate symmetry tests for complete and competing risks data: a saddlepoint approach. Journal of Statistical Computation and Simulation, 87(6): 1269–1275, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Auder:2012:PBC

[AF12] Benjamin Auder and Aurélie Fischer. Projection-based curve clustering. Journal of Statistical Computation and Simulation, 82(8):1145–1168, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abbas:2017:CCM

[AF17] Sermad Abbas and Roland Fried. Control charts for the mean based on robust two-sample tests. Journal of Statistical Computation and Simulation, 87(1):138–155, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmadi:2019:REM

[AG19] Kambiz Ahmadi and Somayeh Ghafouri. Reliability estimation in a multicomponent stress-strength model under generalized half-normal distribution based on progressive type-II censoring. Journal of Statistical Computation and Simulation, 89(13):2505–2548, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arbenz:2018:PLA

[AGA18] Philipp Arbenz and William Guevara-Alarcón. Piecewise linear approximation of empirical distributions under a Wasserstein distance constraint. Journal of Statistical Computation and Simulation, 88(16):3193–3216, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Antoniadis:2015:LCE

[AGM15] A. Antoniadis, I. K. Glad, and H. Mohammed. Local comparison of empirical distributions via nonparametric regression. Journal of Statistical Computation and Simulation, 85 (12):2384–2405, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Andres:2016:CIS

[AH16] A. Martín Andrés and M. Álvarez Hernández. Comment on ‘An improved score interval with a modified midpoint for a binomial proportion’. Journal of Statistical Computation and Simulation, 86(2):388–393, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [YGX14].

Awais:2018:ECM

[AH18a] Muhammad Awais and Abdul Haq. An EWMA chart for monitoring process mean. Journal of Statistical Computation and Simulation, 88(5):1003–1025, 2018. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Azizi:2018:JML

[AH18b] Fariba Azizi and Firoozeh Haghighi. Joint modeling of linear degradation and failure time data with masked causes of failure under simple step-stress test. Journal of Statistical Computation and Simulation, 88(8):1603–1615, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abbasi:2019:EAC

[AH19a] Saba Abbasi and Abdul Haq. Enhanced adaptive CUSUM charts for process mean. Journal of Statistical Computation and Simulation, 89(13):2562–2582, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abbasi:2019:OCA

[AH19b] Saba Abbasi and Abdul Haq. Optimal CUSUM and adaptive CUSUM charts with auxiliary information for process mean. Journal of Statistical Computation and Simulation, 89(2):337–361, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abdel-Hamid:2015:IPS

[AHA15] Alaa H. Abdel-Hamid and Tahani A. Abushal. Inference on progressive-stress model for the exponentiated exponential distribution under type-II progressive hybrid censoring. Journal of Statistical Computation and Simulation, 85 (6):1165–1186, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Hussaini:2010:BTS

[AHAA10] Essam K. Al-Hussaini and Fahimah Al-Awadhi. Bayes twosample prediction of generalized order statistics with fixed and random sample size. Journal of Statistical Computation and Simulation, 80(1):13–28, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abdel-Hamid:2014:BPT

[AHAH14] Alaa H. Abdel-Hamid and Essam K. Al-Hussaini. Bayesian prediction for Type-II progressive-censored data from the Rayleigh distribution under progressive-stress model. Journal of Statistical Computation and Simulation, 84(6):1297– 1312, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmed:2015:EEC

[AHAM15] S. E. Ahmed, A. Hussein, and Marwan Al-Momani. Efficient estimation for the conditional autoregressive model. Journal of Statistical Computation and Simulation, 85(13):2569– 2581, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahn:2015:ECM

[AHC15] Sung K. Ahn, Hanwoom Hong, and Sinsup Cho. Estimation of cointegrated models with exogenous variables. Journal of Statistical Computation and Simulation, 85(10):1945–1962, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abdel-Hamid:2017:NLD

[AHH17] Alaa H. Abdel-Hamid and Atef F. Hashem. A new lifetime distribution for a series-parallel system: properties, applications and estimations under progressive type-II censoring. Journal of Statistical Computation and Simulation, 87(5): 993–1024, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Audigier:2016:MIC

[AHJ16] Vincent Audigier, François Husson, and Julie Josse. Multiple imputation for continuous variables using a Bayesian principal component analysis. Journal of Statistical Computation and Simulation, 86(11):2140–2156, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abuzaid:2013:DOS

[AHM13] A. H. Abuzaid, A. G. Hussin, and I. B. Mohamed. Detection of outliers in simple circular regression models using the mean circular error statistic. Journal of Statistical Computation and Simulation, 83(2):269–277, 2013. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmed:2015:E

[Ahm15] S. Ejaz Ahmed. Editorial. Journal of Statistical Computation and Simulation, 85(13):2547, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmed:2016:CHB

[Ahm16] Jameel Ahmed. A conditionally heteroskedastic binary choice model for macro-financial time series. Journal of Statistical Computation and Simulation, 86(10):2007–2035, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmad:2019:MCM

[Ahm19] M. Rauf Ahmad. Multiple comparisons of mean vectors with large dimension under general conditions. Journal of Statistical Computation and Simulation, 89(6):1044–1059, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Altaher:2012:REB

[AI12] Alsaidi M. Altaher and Mohd Tahir Ismail. Robust estimation for boundary correction in wavelet regression. Journal of Statistical Computation and Simulation, 82(10):1531– 1544, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Asar:2016:FML

[AI16] Özgür Asar and Ozlem Ilk. Forecasting multivariate longitudinal binary data with marginal and marginally specified models. Journal of Statistical Computation and Simulation, 86(2):414–429, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslam:2011:NAS

[AJA11a] Muhammad Aslam, Chi-Hyuck Jun, and Munir Ahmad. New acceptance sampling plans based on life tests for Birnbaum–Saunders distributions. Journal of Statistical Computation and Simulation, 81(4):461–470, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslam:2011:TSG

[AJA11b] Muhammad Aslam, Chi-Hyuck Jun, and Munir Ahmad. A two-stage group sampling plan based on truncated life tests for a general distribution. Journal of Statistical Computation and Simulation, 81(12):1927–1938, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aghamohammadi:2014:PBL

[AJFB14] Seyedeh Zahra Aghamohammadi, Ahad Jamalizadeh, R. Farnoosh, and N. Balakrishnan. Prediction based on linear combinations of order statistics and bivariate concomitants in the case of multivariate elliptical distributions. Journal of Statistical Computation and Simulation, 84(5):1079–1098, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmad:2011:BEU

[AJM11] K. E. Ahmad, Z. F. Jaheen, and Heba S. Mohammed. Bayesian estimation under a mixture of the Burr Type XII distribution and its reciprocal. Journal of Statistical Computation and Simulation, 81(12):2121–2130, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amiri:2016:NTS

[AK16] Masoud Amiri and Baha-Eldin Khaledi. A new test for symmetry against right skewness. Journal of Statistical Computation and Simulation, 86(8):1479–1496, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1071374.

Al-khasawneh:2019:EBS

[Ak19] Mohanad Fayez Al-khasawneh. Empirical Bayes sequential estimation of a finite population mean. Journal of Statistical Computation and Simulation, 89(12):2175–2186, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akdogan:2016:UGD

[AKA^{+}16] Yunus Akdogan, Coskun Kus, Akbar Asgharzadeh, Ismail Kinaci, and Fatemeh Sharafi. Uniform-Geometric distribution. Journal of Statistical Computation and Simulation, 86(9):1754–1770, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amiri:2015:PMG

[AKAW15] Amirhossein Amiri, Mehdi Koosha, Armaghan Azhdari, and Guanghui Wang. Phase I monitoring of generalized linear model-based regression profiles. Journal of Statistical Computation and Simulation, 85(14):2839–2859, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akgul:2019:REM

[Akg19] Fatma Gül Akgül. Reliability estimation in multicomponent stress-strength model for Topp–Leone distribution. Journal of Statistical Computation and Simulation, 89(15):2914– 2929, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslam:2016:CCT

[AKJ16] Muhammad Aslam, Nasrullah Khan, and Chi-Hyuck Jun. A control chart for time truncated life tests using Pareto distribution of second kind. Journal of Statistical Computation and Simulation, 86(11):2113–2122, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aknouche:2013:ROE

[Akn13] Abdelhakim Aknouche. Recursive online EM estimation of mixture autoregressions. Journal of Statistical Computation and Simulation, 83(2):370–383, 2013. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmed:2015:CSB

[AKS^{+}15] S. Ejaz Ahmed, Iskander Kareev, Sujitta Suraphee, Andrei Volodin, and Igor Volodin. Confidence sets based on the positive part James–Stein estimator with the asymptotically constant coverage probability. Journal of Statistical Computation and Simulation, 85(12):2506–2513, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alma:2011:GAO

[AKU11] Özlem Gürünlü Alma, Serdar Kurt, and Aybars Uğur. Genetic algorithms for outlier detection in multiple regression with different information criteria. Journal of Statistical Computation and Simulation, 81(1):29–47, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arashi:2017:RPE

[AKV17] M. Arashi, B. M. Golam Kibria, and T. Valizadeh. On ridge parameter estimators under stochastic subspace hypothesis. Journal of Statistical Computation and Simulation, 87(5): 966–983, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alber:2015:CPD

[AL15] Susan A. Alber and J. Jack Lee. Calibrating the prior distribution for a normal model with conjugate prior. Journal of Statistical Computation and Simulation, 85(15):3108–3128, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Austin:2018:ENC

[AL18] Peter C. Austin and George Leckie. The effect of number of clusters and cluster size on statistical power and Type I error rates when testing random effects variance components in multilevel linear and logistic regression models. Journal of Statistical Computation and Simulation, 88(16):3151–3163, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alin:2012:PPD

[Ali12] Aylin Alin. Penalized power-divergence test statistics for testing linear-by-linear association. Journal of Statistical Computation and Simulation, 82(2):227–239, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alin:2014:APD

[Ali14] Aylin Alin. Adjusted power-divergence test statistics for simple goodness-of-fit under clustered sampling. Journal of Statistical Computation and Simulation, 84(9):1904–1916, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ali:2015:BEW

[Ali15] Sajid Ali. On the Bayesian estimation of the weighted Lindley distribution. Journal of Statistical Computation and Simulation, 85(5):855–880, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Altun:2019:QLO

[Alt19] Gökçen Altun. Quasi local odds symmetry model for square contingency table with ordinal categories. Journal of Statistical Computation and Simulation, 89(15):2899–2913, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alwan:2017:OTG

[Alw17] Saba Mohammed Alwan. Optimal trajectory to the general equilibrium state of the unstable stochastic lattice gas of prey-predator model with pair-approximation. Journal of Statistical Computation and Simulation, 87(9):1887–1900, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abbasi:2013:MCE

[AM13] Saddam Akber Abbasi and Arden Miller. MDEWMA chart: an efficient and robust alternative to monitor process dispersion. Journal of Statistical Computation and Simulation, 83(2):247–268, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abdalla:2019:FMR

[AM19] Abdelbaset Abdalla and Semhar Michael. Finite mixture of regression models for a stratified sample. Journal of Statistical Computation and Simulation, 89(14):2782–2800, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmed:2012:STE

[AMAMS12] S. E. Ahmed, H. A. Muttlak, J. Al-Mutawa, and M. Saheh. Stein-type estimation using ranked set sampling. Journal of Statistical Computation and Simulation, 82(10):1501–1516, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmadi:2012:BPR

[AMB12] Jafar Ahmadi, S. M. T. K. Mirmostafaee, and N. Balakrishnan. Bayesian prediction of k-record values based on progressively censored data from exponential distribution. Journal of Statistical Computation and Simulation, 82(1): 51–62, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amiri:2015:RSS

[AMB15] Saeid Amiri, Reza Modarres, and Dinesh S. Bhoj. Ranked set sampling with random subsamples. Journal of Statistical Computation and Simulation, 85(5):935–946, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alin:2017:SB

[AMBP17] Aylin Alin, Michael A. Martin, Ufuk Beyaztas, and Pramod K. Pathak. Sufficient m-out-of-n (m/n) bootstrap. Journal of Statistical Computation and Simulation, 87(9):1742–1753, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amezziane:2012:BMK

[Ame12] Mohamed Amezziane. A binomial model for the kernel density estimator and related inference. Journal of Statistical Computation and Simulation, 82(2):151–164, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aminzadeh:2011:BER

[Ami11] M. S. Aminzadeh. Bayesian estimation of renewal function for inverse Gaussian renewal process. Journal of Statistical Computation and Simulation, 81(3):331–341, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2010:EB

[Ano10] Anonymous. Editorial Board. Journal of Statistical Computation and Simulation, 80(12):??, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2011:EB

[Ano11] Anonymous. Editorial Board. Journal of Statistical Computation and Simulation, 81(12):ebi, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2012:EB

[Ano12] Anonymous. Editorial Board. Journal of Statistical Computation and Simulation, 82(12):ebi, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2013:EB

[Ano13] Anonymous. Editorial Board. Journal of Statistical Computation and Simulation, 83(12):ebi, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2014:EB

[Ano14a] Anonymous. Editorial Board. Journal of Statistical Computation and Simulation, 84(12):ebi, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2014:E

[Ano14b] Anonymous. Erratum. Journal of Statistical Computation and Simulation, 84(11):i, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2014:RAS

[Ano14c] Anonymous. RETRACTED ARTICLE: Statement of retraction: [Ali A. Ismail and H. M. Aly, Optimal planning of failure-step stress partially accelerated life tests under Type-II censoring]. Journal of Statistical Computation and Simulation, 84(12):2759, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. Retraction [IA10] because a similar article was published in the International Journal of Mathematical Analysis, 3(31) 1509–1523.

Anonymous:2015:C

[Ano15a] Anonymous. Corrigendum. Journal of Statistical Computation and Simulation, 85(14):i, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2015:EBa

[Ano15b] Anonymous. Editorial Board. Journal of Statistical Computation and Simulation, 85(18):ebi, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2015:Ea

[Ano15c] Anonymous. Erratum. Journal of Statistical Computation and Simulation, 85(4):i, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2015:Eb

[Ano15d] Anonymous. Erratum. Journal of Statistical Computation and Simulation, 85(11):i, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2015:Ec

[Ano15e] Anonymous. Erratum. Journal of Statistical Computation and Simulation, 85(12):i, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2016:Ca

[Ano16a] Anonymous. Corrigendum. Journal of Statistical Computation and Simulation, 86(15):i, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2016:Cb

[Ano16b] Anonymous. Corrigendum. Journal of Statistical Computation and Simulation, 86(18):3835–??, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Anonymous:2017:C

[Ano17a] Anonymous. Corrigendum. Journal of Statistical Computation and Simulation, 87(4):i, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2017:E

[Ano17b] Anonymous. Erratum. Journal of Statistical Computation and Simulation, 87(18):i, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2018:C

[Ano18a] Anonymous. Corrigendum. Journal of Statistical Computation and Simulation, 88(9):x, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2018:E

[Ano18b] Anonymous. Erratum. Journal of Statistical Computation and Simulation, 88(14):x, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anonymous:2019:C

[Ano19] Anonymous. Correction. Journal of Statistical Computation and Simulation, 89(17):3330, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Omari:2011:EMB

[AO11] Amer Ibrahim Al-Omari. Estimation of mean based on modified robust extreme ranked set sampling. Journal of Statistical Computation and Simulation, 81(8):1055–1066, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alfaro:2012:RHC

[AO12] José-Luis Alfaro and Juan-Fco Ortega. Robust Hotelling’s T^{2} control charts under non-normality: the case of t-Student distribution. Journal of Statistical Computation and Simulation, 82(10):1437–1447, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ata:2013:SFF

[AÖ13] Nihal Ata and Gamze Özel. Survival functions for the frailty models based on the discrete compound Poisson process. Journal of Statistical Computation and Simulation, 83 (11):2105–2116, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Acar:2016:CVR

[AÖ16] T. Söküt Açar and M. R. Özkale. Cross validation of ridge regression estimator in autocorrelated linear regression models. Journal of Statistical Computation and Simulation, 86(12):2429–2440, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Omari:2016:EEG

[AOH16] Amer Ibrahim Al-Omari and Abdul Haq. Entropy estimation and goodness-of-fit tests for the inverse Gaussian and Laplace distributions using paired ranked set sampling. Journal of Statistical Computation and Simulation, 86(11):2262–2272, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Omari:2019:NSM

[AOH19] Amer Ibrahim Al-Omari and Abdul Haq. A new sampling method for estimating the population mean. Journal of Statistical Computation and Simulation, 89(11):1973–1985, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Omari:2013:EPM

[AOR13] Amer Ibrahim Al-Omari and Mohammad Z. Raqab. Estimation of the population mean and median using truncation-based ranked set samples. Journal of Statistical Computation and Simulation, 83(8):1453–1471, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anselmo:2012:PTH

[AP12] Cézar A. F. Anselmo and Aluísio Pinheiro. Phylogenetic trees via Hamming distance decomposition tests. Journal of Statistical Computation and Simulation, 82(9):1287–1297, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Akhundjanov:2015:MRE

[AP15] Sherzod B. Akhundjanov and Francis Pascual. Moving range EWMA control charts for monitoring the Weibull shape parameter. Journal of Statistical Computation and Simulation, 85(9):1864–1882, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amro:2017:PIP

[AP17] Lubna Amro and Markus Pauly. Permuting incomplete paired data: a novel exact and asymptotic correct randomization test. Journal of Statistical Computation and Simulation, 87(6):1148–1159, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Agraz:2019:ELT

[AP19] Melih Agraz and Vilda Purutçuoğlu. Extended lasso-type MARS (LMARS) model in the description of biological network. Journal of Statistical Computation and Simulation, 89 (1):1–14, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alegria:2018:AMV

[APF18] Alfredo Alegría, Emilio Porcu, and Reinhard Furrer. Asymmetric matrix-valued covariances for multivariate random fields on spheres. Journal of Statistical Computation and Simulation, 88(10):1850–1862, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmadi:2010:NCI

[AR10] Jafar Ahmadi and M. Razmkhah. Nonparametric confidence intervals for quantiles and quantile intervals from inversely sampled record breaking data in a multi-sample plan. Journal of Statistical Computation and Simulation, 80(9):993–1006, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amini:2016:LTS

[AR16a] Morteza Amini and Mahdi Roozbeh. Least trimmed squares ridge estimation in partially linear regression models. Journal of Statistical Computation and Simulation, 86(14):2766– 2780, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amirzadeh:2016:PPT

[AR16b] Vahid Amirzadeh and Mohsen Rezapour. On properties of progressively Type-II censored conditionally N-ordered statistics arising from a non-identical and dependent random vector. Journal of Statistical Computation and Simulation, 86(9):1818–1828, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Andrade:2016:FAB

[AR16c] Bernardo B. Andrade and Pushpa N. Rathie. Fitting asymmetric bimodal data with selected distributions. Journal of Statistical Computation and Simulation, 86(16):3205–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ali:2013:CAM

[ARB13] Irfan Ali, Yashpal Singh Raghav, and Abdul Bari. Compromise allocation in multivariate stratified surveys with stochastic quadratic cost function. Journal of Statistical Computation and Simulation, 83(5):962–976, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Arendacka:2012:AIB

[Are12] Barbora Arendacká. Approximate interval for the betweengroup variance under heteroscedasticity. Journal of Statistical Computation and Simulation, 82(2):209–218, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alka:2019:TSC

[ARQ19] Alka, Piyush Kant Rai, and Muhammad Qasim. Two-step calibration of design weights under two auxiliary variables in sample survey. Journal of Statistical Computation and Simulation, 89(12):2316–2327, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmadi:2015:EGH

[ARY15] K. Ahmadi, M. Rezaei, and F. Yousefzadeh. Estimation for the generalized half-normal distribution based on progressive type-II censoring. Journal of Statistical Computation and Simulation, 85(6):1128–1150, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmadi:2016:PPL

[ARY16] K. Ahmadi, M. Rezaei, and F. Yousefzadeh. Point predictors of the latent failure times of censored units in progressively type-II censored competing risks data from the exponential distributions. Journal of Statistical Computation and Simulation, 86(8):1620–1634, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1079634.

Ahmadi:2017:SAM

[ARY17] K. Ahmadi, M. Rezaei, and F. Yousefzadeh. Statistical analysis of middle censored competing risks data with exponential distribution. Journal of Statistical Computation and Simulation, 87(16):3082–3110, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amin:2012:DAS

[AS12] Zeinab Amin and Maram Salem. On designing an acceptance sampling plan for the Pareto lifetime model. Journal of Statistical Computation and Simulation, 82(8):1115– 1133, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abushal:2015:EPP

[AS15a] Tahani A. Abushal and Ahmed A. Soliman. Estimating the Pareto parameters under progressive censoring data for constant-partially accelerated life tests. Journal of Statistical Computation and Simulation, 85(5):917–934, 2015. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Allison:2015:DDC

[AS15b] J. S. Allison and L. Santana. On a data-dependent choice of the tuning parameter appearing in certain goodness-of-fit tests. Journal of Statistical Computation and Simulation, 85(16):3276–3288, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amezziane:2015:COS

[AS15c] Mohamed Amezziane and Claudia Schmegner. A class of one sample tests based on the Mann–Whitney–Wilcoxon functional. Journal of Statistical Computation and Simulation, 85(3):587–595, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslam:2017:CCC

[ASA^{+}17] Muhammad Aslam, Aamir Saghir, Liaquat Ahmad, Chi-Hyuck Jun, and Jaffer Hussain. A control chart for COM–Poisson distribution using a modified EWMA statistic. Journal of Statistical Computation and Simulation, 87 (18):3491–3502, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Anholeto:2014:APR

[ASB14] Tatiana Anholeto, Mônica C. Sandoval, and Denise A. Botter. Adjusted Pearson residuals in beta regression models. Journal of Statistical Computation and Simulation, 84(5): 999–1014, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abolmohammadi:2019:ELL

[ASBM19] M. Abolmohammadi, A. Seif, M. H. Behzadi, and M. Bameni Moghadam. Effect of Linex loss function on the VSIXν control chart. Journal of Statistical Computation and Simulation, 89(9):1674–1693, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alawadhi:2016:MOE

[ASH16] Fahimah A. Alawadhi, Ammar M. Sarhan, and David C. Hamilton. Marshall–Olkin extended two-parameter bathtubshaped lifetime distribution. Journal of Statistical Computation and Simulation, 86(18):3653–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Abrahantes:2011:CVS

[ASM^{+}11] José Cortiñas Abrahantes, Cristina Sotto, Geert Molenberghs, Geert Vromman, and Bart Bierinckx. A comparison of various software tools for dealing with missing data via imputation. Journal of Statistical Computation and Simulation, 81(11):1653–1675, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Afraei:2017:SAR

[ASM17] Sajjad Afraei, Kourosh Shahriar, and Sayyed Hasan Madani. Statistical analysis of rock-burst events in underground mines and excavations to present reasonable data-driven predictors. Journal of Statistical Computation and Simulation, 87(17):3336–3376, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Sharadqah:2019:SDO

[ASM19] Ali Al-Sharadqah and Majid Mojirsheibani. Semi-doubly optimal concentric circles fitting with presence of heteroscedasticity. Journal of Statistical Computation and Simulation, 89(7):1183–1202, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Allison:2011:TND

[ASS11] James S. Allison, Leonard Santana, and Jan W. H. Swanepoel. Two new data-dependent choices of m when applying the m-out-of-n bootstrap to hypothesis testing. Journal of Statistical Computation and Simulation, 81(12):2107– 2120, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Amiri:2019:MEL

[ATF19] Mohammad Moqaddasi Amiri, Leili Tapak, and Javad Faradmal. A mixed-effects least square support vector regression model for three-level count data. Journal of Statistical Computation and Simulation, 89(15):2801–2812, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Andres:2012:OMM

[ATH12] A. Martín Andrés, I. Herranz Tejedor, and M. Álvarez Hernández. The optimal method to make inferences about a linear combination of proportions. Journal of Statistical Computation and Simulation, 82(1):123–135, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Altun:2018:NGS

[ATON18] Emrah Altun, Huseyin Tatlidil, Gamze Ozel, and Saralees Nadarajah. A new generalization of skew-T distribution with volatility models. Journal of Statistical Computation and Simulation, 88(7):1252–1272, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Austin:2018:SPD

[Aus18] Peter C. Austin. Statistical power to detect violation of the proportional hazards assumption when using the Cox regression model. Journal of Statistical Computation and Simulation, 88(3):533–552, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmed:2019:P

[AV19] S. Ejaz Ahmed and Andrei Volodin. Preface. Journal of Statistical Computation and Simulation, 89(6):949–950, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Acosta:2018:EGS

[AVG18] Jonathan Acosta, Ronny Vallejos, and Daniel Griffth. On the effective geographic sample size. Journal of Statistical Computation and Simulation, 88(10):1958–1975, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Asgharzadeh:2015:PFF

[AVK15] A. Asgharzadeh, R. Valiollahi, and D. Kundu. Prediction for future failures in Weibull distribution under hybrid censoring. Journal of Statistical Computation and Simulation, 85(4):824–838, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmad:2015:THD

[AvR15] M. Rauf Ahmad and D. von Rosen. Tests for highdimensional covariance matrices using the theory of Ustatistics. Journal of Statistical Computation and Simulation, 85(13):2619–2631, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Andersson:2014:SAV

[AW14] Björn Andersson and Ingeborg Waernbaum. Sensitivity analysis of violations of the faithfulness assumption. Journal of Statistical Computation and Simulation, 84(7):1608– 1620, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslam:2017:ASP

[AW17] Muhammad Aslam and Fu-Kwun Wang. Acceptance sampling plans for linear profiles with one-sided specifications. Journal of Statistical Computation and Simulation, 87(4): 806–816, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Wahsh:2019:EZI

[AWH19] H. Al-Wahsh and A. Hussein. Estimation of zero-inflated parameter-driven models via data cloning. Journal of Statistical Computation and Simulation, 89(6):951–965, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslam:2013:DVR

[AWJ^{+}13] Muhammad Aslam, Chien-Wei Wu, Chi-Hyuck Jun, Muhammad Azam, and Negrin Itay. Developing a variables repetitive group sampling plan based on process capability index C_{pk} with unknown mean and variance. Journal of Statistical Computation and Simulation, 83(8):1507–1517, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alkenani:2013:CSR

[AY13] Ali Alkenani and Keming Yu. A comparative study for robust canonical correlation methods. Journal of Statistical Computation and Simulation, 83(4):692–720, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alhamzawi:2014:BLM

[AY14] Rahim Alhamzawi and Keming Yu. Bayesian lasso-mixed quantile regression. Journal of Statistical Computation and Simulation, 84(4):868–880, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Alhamzawi:2015:BTQ

[AY15] Rahim Alhamzawi and Keming Yu. Bayesian Tobit quantile regression using g-prior distribution with ridge parameter. Journal of Statistical Computation and Simulation, 85 (14):2903–2918, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aydin:2018:MES

[AY18] Dursun Aydin and Ersin Yilmaz. Modified estimators in semiparametric regression models with right-censored data. Journal of Statistical Computation and Simulation, 88(8):1470–1498, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Aslam:2011:VRG

[AYJ11] Muhammad Aslam, Ching-Ho Yen, and Chi-Hyuck Jun. Variable repetitive group sampling plans with process loss consideration. Journal of Statistical Computation and Simulation, 81(11):1417–1432, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ahmadi:2016:APT

[AYR16] K. Ahmadi, F. Yousefzadeh, and M. Rezaei. Analysis of progressively Type-I interval censored competing risks data for a class of an exponential distribution. Journal of Statistical Computation and Simulation, 86(18):3629–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Al-Zahrani:2015:SAL

[AZS15] Bander Al-Zahrani and Hanaa Sagor. Statistical analysis of the lomax-logarithmic distribution. Journal of Statistical Computation and Simulation, 85(9):1883–1901, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Basiri:2015:PIG

[BA15] Elham Basiri and Jafar Ahmadi. Prediction intervals for generalized-order statistics with random sample size. Journal of Statistical Computation and Simulation, 85(9):1725– 1741, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhatnagar:2015:CAL

[BAB15] S. R. Bhatnagar, J. Atherton, and A. Benedetti. Comparing alternating logistic regressions to other approaches to modelling correlated binary data. Journal of Statistical Computation and Simulation, 85(10):2059–2071, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bagkavos:2011:LLV

[Bag11] Dimitrios Bagkavos. Local linear variable bandwidth hazard rate estimation. Journal of Statistical Computation and Simulation, 81(11):1517–1531, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baird:2016:GJC

[Bai16] Matthew Baird. Gains from joint cross validation bandwidth selection for derivatives of conditional multidimensional densities. Journal of Statistical Computation and Simulation, 86(4):807–819, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1037766.

Bagheri:2014:ECE

[BAJN14] S. F. Bagheri, M. Alizadeh, E. Baloui Jamkhaneh, and S. Nadarajah. Evaluation and comparison of estimations in the generalized exponential-Poisson distribution. Journal of Statistical Computation and Simulation, 84(11):2345–2360, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baklizi:2014:IES

[Bak14] Ayman Baklizi. Interval estimation of the stress-strength reliability in the two-parameter exponential distribution based on records. Journal of Statistical Computation and Simulation, 84(12):2670–2679, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baklizi:2018:ATP

[Bak18] Ayman Baklizi. Approximating the tail probabilities of the longest run in a sequence of Bernoulli trials. Journal of Statistical Computation and Simulation, 88(14):2751–2760, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Belaid:2016:BLB

[BAKZ16] Nawal Belaid, Smail Adjabi, Célestin C. Kokonendji, and Nabil Zougab. Bayesian local bandwidth selector in multivariate associated kernel estimator for joint probability mass functions. Journal of Statistical Computation and Simulation, 86(18):3667–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Belaghi:2017:EPB

[BAS17] R. Arabi Belaghi, M. Noori Asl, and Sukhdev Singh. On estimating the parameters of the Burr XII model under progressive type-I interval censoring. Journal of Statistical Computation and Simulation, 87(16):3132–3151, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bashtian:2011:UIE

[BAT11] M. Hassanzadeh Bashtian, M. Arashi, and S. M. M. Tabatabaey. Using improved estimation strategies to combat multicollinearity. Journal of Statistical Computation and Simulation, 81(12):1773–1797, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bindele:2019:RES

[BAZ19] Huybrechts F. Bindele, Asheber Abebe, and Peng Zeng. Robust estimation and selection for single-index regression model. Journal of Statistical Computation and Simulation, 89(8):1376–1393, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Basak:2012:ETP

[BB12] Prasanta Basak and N. Balakrishnan. Estimation for the three-parameter inverse Gaussian distribution under progressive Type-II censoring. Journal of Statistical Computation and Simulation, 82(7):1055–1072, 2012. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Basirat:2015:SIS

[BBA15] M. Basirat, S. Baratpour, and Jafar Ahmadi. Statistical inferences for stress-strength in the proportional hazard models based on progressive Type-II censored samples. Journal of Statistical Computation and Simulation, 85(3):431–449, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bisaglia:2010:BAE

[BBC10] Luisa Bisaglia, Silvano Bordignon, and Nedda Cecchinato. Bootstrap approaches for estimation and confidence intervals of long memory processes. Journal of Statistical Computation and Simulation, 80(9):959–978, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Blas:2013:SAC

[BBL13] Betsabé Blas, Heleno Bolfarine, and Victor H. Lachos. Statistical analysis of controlled calibration model with replicates. Journal of Statistical Computation and Simulation, 83(5):941–961, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bourguignon:2018:NGI

[BBM18] Marcelo Bourguignon, Patrick Borges, and Fabio Fajardo Molinares. A new geometric INAR(1) process based on counting series with deflation or inflation of zeros. Journal of Statistical Computation and Simulation, 88(17):3338–3348, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bidram:2013:BWG

[BBT13] H. Bidram, J. Behboodian, and M. Towhidi. The beta Weibull-geometric distribution. Journal of Statistical Computation and Simulation, 83(1):52–67, 2013. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Berta:2017:CSM

[BBV17] Paolo Berta, Matteo Bossi, and Stefano Verzillo. %CEM: a SAS macro to perform coarsened exact matching. Journal of Statistical Computation and Simulation, 87(2):227–238, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Becker:2017:ICU

[BBW17] Devan G. Becker, W. John Braun, and Bethany J. G. White. Interval-censored unimodal kernel density estimation via data sharpening. Journal of Statistical Computation and Simulation, 87(10):2023–2037, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barbeito:2019:SBB

[BC19] Inés Barbeito and Ricardo Cao. Smoothed bootstrap bandwidth selection for nonparametric hazard rate estimation. Journal of Statistical Computation and Simulation, 89(1): 15–37, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bayer:2018:BSA

[BCCN18] Fábio M. Bayer, Renato J. Cintra, and Francisco Cribari-Neto. Beta seasonal autoregressive moving average models. Journal of Statistical Computation and Simulation, 88 (15):2961–2981, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barranco-Chamorro:2012:ARP

[BCJG12] I. Barranco-Chamorro and M. D. Jiménez-Gamero. Asymptotic results in partially non-regular log-exponential distributions. Journal of Statistical Computation and Simulation, 82(3):445–461, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bakouch:2018:NLM

[BCL18] Hassan S. Bakouch, Christophe Chesneau, and Jeremias Leao. A new lifetime model with a periodic hazard rate and an application. Journal of Statistical Computation and Simulation, 88(11):2048–2065, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bertin:2017:SPS

[BCLM17] K. Bertin, X. Collilieux, E. Lebarbier, and C. Meza. Semiparametric segmentation of multiple series using a DP-Lasso strategy. Journal of Statistical Computation and Simulation, 87(6):1255–1268, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Braga:2013:AMG

[BCO13] A. C. Braga, L. Costa, and P. Oliveira. An alternative method for global and partial comparison of two diagnostic systems based on ROC curves. Journal of Statistical Computation and Simulation, 83(2):307–325, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Berrendero:2016:MVS

[BCT16] J. R. Berrendero, A. Cuevas, and J. L. Torrecilla. The mRMR variable selection method: a comparative study for functional data. Journal of Statistical Computation and Simulation, 86(5):891–907, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1042378.

Bai:2016:MLM

[BCY16] Xiuqin Bai, Kun Chen, and Weixin Yao. Mixture of linear mixed models using multivariate t distribution. Journal of Statistical Computation and Simulation, 86(4):771–787, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. URL http://www.tandfonline.

com/doi/abs/10.1080/00949655.2015.1036431.

Brito:2017:TLO

[BCY^{+}17] E. Brito, G. M. Cordeiro, H. M. Yousof, M. Alizadeh, and G. O. Silva. The Topp–Leone odd log-logistic family of distributions. Journal of Statistical Computation and Simulation, 87(15):3040–3058, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Busetti:2012:BLT

[BD12] Fabio Busetti and Silvestro Di Sanzo. Bootstrap LR tests of stationarity, common trends and cointegration. Journal of Statistical Computation and Simulation, 82(9):1343–1355, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bandyopadhyay:2017:CBT

[BD17] Uttam Bandyopadhyay and Radhakanta Das. A comparison between two treatments in a clinical trial with an ethical allocation design. Journal of Statistical Computation and Simulation, 87(17):3377–3397, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2011:PCM

[BDKM11] N. Balakrishnan, Katherine F. Davies, Jerome P. Keating, and Robert L. Mason. Pitman closeness, monotonicity and consistency of best linear unbiased and invariant estimators for exponential distribution under Type II censoring. Journal of Statistical Computation and Simulation, 81(8): 985–999, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bouhaddioui:2014:IEV

[BEBG14] Chafik Bouhaddioui, Mohamed Y. El-Bassiouni, and Kilani Ghoudi. Interval estimation of variance ratio in non-normal unbalanced one-way random models. Journal of Statistical Computation and Simulation, 84(6):1329–1344, 2014. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bergman:2012:GRV

[Ber12] Jakob Bergman. Generating random variates from a bicompositional Dirichlet distribution. Journal of Statistical Computation and Simulation, 82(6):797–805, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Batsidis:2016:TFL

[BET16] A. Batsidis, P. Economou, and G. Tzavelas. Tests of fit for a lognormal distribution. Journal of Statistical Computation and Simulation, 86(2):215–235, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balabdaoui:2012:ECC

[BF12] Fadoua Balabdaoui and Karim Filali. Efficient computation of the cdf of the maximal difference between a Brownian bridge and its concave majorant. Journal of Statistical Computation and Simulation, 82(3):405–418, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balasooriya:2011:LRT

[BG11a] Uditha Balasooriya and Veeresh Gadag. Likelihood ratio tests for testing for multiple contaminants in the shocks and labelled slippage models. Journal of Statistical Computation and Simulation, 81(5):607–618, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bell:2011:SSE

[BG11b] Melanie L. Bell and Gary K. Grunwald. Small sample estimation properties of longitudinal count models. Journal of Statistical Computation and Simulation, 81(9):1067–1079, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Boussaha:2018:PAS

[BH18] Nadia Boussaha and Fayçal Hamdi. On periodic autoregressive stochastic volatility models: structure and estimation. Journal of Statistical Computation and Simulation, 88(9):1637–1668, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baffoun:2019:ESE

[BH19] Hatem Baffoun and Mekki Hajlaoui. Equation-solving estimator based on the general n-step MHDR algorithm. Journal of Statistical Computation and Simulation, 89(14):2733– 2743, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bialek:2015:CCI

[Bia15] Jacek Bialek. Construction of confidence intervals for the Laspeyres price index. Journal of Statistical Computation and Simulation, 85(14):2962–2973, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bhoj:2016:RSS

[BK16] Dinesh S. Bhoj and Debashis Kushary. Ranked set sampling with unequal samples for skew distributions. Journal of Statistical Computation and Simulation, 86(4):676–681, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. URL http://www.tandfonline.

com/doi/abs/10.1080/00949655.2015.1028405.

Bae:2017:BCS

[BK17a] Taehan Bae and Alex Kreinin. A backward construction and simulation of correlated Poisson processes. Journal of Statistical Computation and Simulation, 87(8):1593–1607, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bonat:2017:FTR

[BK17b] Wagner Hugo Bonat and Célestin C. Kokonendji. Flexible Tweedie regression models for continuous data. Journal of Statistical Computation and Simulation, 87(11):2138–2152, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bakar:2016:HSV

[BKJ16] K. Shuvo Bakar, Philip Kokic, and Huidong Jin. Hierarchical spatially varying coefficient and temporal dynamic process models using spTDyn. Journal of Statistical Computation and Simulation, 86(4):820–840, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1038267.

Babaie-Kafaki:2017:RCD

[BKR17] Saman Babaie-Kafaki and Mahdi Roozbeh. A revised Cholesky decomposition to combat multicollinearity in multiple regression models. Journal of Statistical Computation and Simulation, 87(12):2298–2308, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baer:2019:EBM

[BL19] Daniel R. Baer and Andrew B. Lawson. Evaluation of Bayesian multiple stage estimation under spatial CAR model variants. Journal of Statistical Computation and Simulation, 89(1):98–144, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bates:2014:SSM

[BMAW14] Ron A. Bates, Hugo Maruri-Aguilar, and Henry P. Wynn. Smooth supersaturated models. Journal of Statistical Computation and Simulation, 84(11):2453–2464, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Boone:2014:HDA

[BMK14] Edward L. Boone, Jason R. W. Merrick, and Matthew J. Krachey. A Hellinger distance approach to MCMC diagnostics. Journal of Statistical Computation and Simulation, 84(4):833–849, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bahrami:2014:EPN

[BMM14] Mohammad Bahrami, Mohammad Mahdi Maghami, and Yaser Mehrali. Erratum to the paper: A note on reference prior for the scaler skew-normal distribution. Journal of Statistical Computation and Simulation, 84(9):1960–1962, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barbiero:2015:BPP

[BMM15] Alessandro Barbiero, Giancarlo Manzi, and Fulvia Mecatti. Bootstrapping probability-proportional-to-size samples via calibrated empirical population. Journal of Statistical Computation and Simulation, 85(3):608–620, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bispo:2012:SPG

[BMP12] Regina Bispo, Tiago A. Marques, and Dinis Pestana. Statistical power of goodness-of-fit tests based on the empirical distribution function for Type-I right-censored data. Journal of Statistical Computation and Simulation, 82(2): 173–181, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Batsidis:2014:NPD

[BMPZ14] Apostolos Batsidis, Nirian Martin, Leandro Pardo, and Konstantinos Zografos. A necessary power divergence-type family of tests for testing elliptical symmetry. Journal of Statistical Computation and Simulation, 84(1):57–83, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bottai:2015:GSM

[BOG15] Matteo Bottai, Nicola Orsini, and Marco Geraci. A gradient search maximization algorithm for the asymmetric Laplace likelihood. Journal of Statistical Computation and Simulation, 85(10):1919–1925, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Boring:2015:FPE

[Bør15] Pål Børing. Flexible parametric estimation technique for a competing risks model with unobserved heterogeneity: a Monte Carlo study. Journal of Statistical Computation and Simulation, 85(11):2258–2275, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Borges:2017:EAB

[Bor17] Patrick Borges. EM algorithm-based likelihood estimation for a generalized Gompertz regression model in presence of survival data with long-term survivors: an application to uterine cervical cancer data. Journal of Statistical Computation and Simulation, 87(9):1712–1722, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Botts:2011:MAA

[Bot11] Carsten Botts. A modified adaptive accept-reject algorithm for univariate densities with bounded support. Journal of Statistical Computation and Simulation, 81(8):1039–1053, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barabesi:2015:UMG

[BP15] L. Barabesi and L. Pratelli. Universal methods for generating random variables with a given characteristic function. Journal of Statistical Computation and Simulation, 85(8):1679–1691, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baek:2012:PTF

[BPH12] J. S. Baek, J. A. Park, and S. Y. Hwang. Preliminary test of fit in a general class of conditionally heteroscedastic nonlinear time series. Journal of Statistical Computation and Simulation, 82(5):763–781, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Babu:2016:IDS

[BR16] T. K. S. Rathish Babu and N. Sankar Ram. Improvement of the dependency structure of the software architecture model with risk estimation. Journal of Statistical Computation and Simulation, 86(5):908–921, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1042379.

Bruzda:2015:SWE

[Bru15] J. Bruzda. On simple wavelet estimators of random signals and their small-sample properties. Journal of Statistical Computation and Simulation, 85(14):2771–2792, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bruch:2019:ARB

[Bru19] Christian Bruch. Applying the rescaling bootstrap under imputation: a simulation study. Journal of Statistical Computation and Simulation, 89(4):641–659, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bitaraf:2017:TNB

[BRY17] M. Bitaraf, M. Rezaei, and F. Yousefzadeh. Test for normality based on two new estimators of entropy. Journal of Statistical Computation and Simulation, 87(2):280–294, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Blumentritt:2012:MIM

[BS12] Thomas Blumentritt and Friedrich Schmid. Mutual information as a measure of multivariate association: analytical properties and statistical estimation. Journal of Statistical Computation and Simulation, 82(9):1257–1274, 2012. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2013:RRS

[BS13] N. Balakrishnan and H. M. Saleh. Recurrence relations for single and product moments of progressively Type-II censored order statistics from a generalized half-logistic distribution with application to inference. Journal of Statistical Computation and Simulation, 83(9):1704–1721, 2013. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Blumentritt:2014:NEC

[BS14] Thomas Blumentritt and Friedrich Schmid. Nonparametric estimation of copula-based measures of multivariate association from contingency tables. Journal of Statistical Computation and Simulation, 84(4):781–797, 2014. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bello:2016:OAA

[BS16] Ghalib A. Bello and Roy T. Sabo. Outcome-adaptive allocation with natural lead-in for three-group trials with binary outcomes. Journal of Statistical Computation and Simulation, 86(12):2441–2449, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Beyaztas:2019:FFT

[BS19] Ufuk Beyaztas and Han Lin Shang. Forecasting functional time series using weighted likelihood methodology. Journal of Statistical Computation and Simulation, 89(16):3046– 3060, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bourguignon:2014:NCF

[BSC14] Marcelo Bourguignon, Rodrigo B. Silva, and Gauss M. Cordeiro. A new class of fatigue life distributions. Journal of Statistical Computation and Simulation, 84(12):2619–2635, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barreto-Souza:2011:WGD

[BSdMC11] Wagner Barreto-Souza, Alice Lemos de Morais, and Gauss M. Cordeiro. The Weibull-geometric distribution. Journal of Statistical Computation and Simulation, 81(5): 645–657, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barreto-Souza:2015:LRT

[BSS15] Wagner Barreto-Souza and Rodrigo B. Silva. A likelihood ratio test to discriminate exponential-Poisson and gamma distributions. Journal of Statistical Computation and Simulation, 85(4):802–823, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barreto-Souza:2017:IEB

[BSS17] Wagner Barreto-Souza and Alexandre B. Simas. Improving estimation for beta regression models via EM-algorithm and related diagnostic tools. Journal of Statistical Computation and Simulation, 87(14):2847–2867, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Barreto-Souza:2010:BGE

[BSSC10] Wagner Barreto-Souza, Alessandro H. S. Santos, and Gauss M. Cordeiro. The beta generalized exponential distribution. Journal of Statistical Computation and Simulation, 80(2):159–172, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baddeley:2014:BCP

[BT14] Adrian Baddeley and Rolf Turner. Bias correction for parameter estimates of spatial point process models. Journal of Statistical Computation and Simulation, 84(8):1621– 1643, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2016:PTT

[BT16a] Narayanaswamy Balakrishnan and Tao Tan. A parametric test for trend based on moving order statistics. Journal of Statistical Computation and Simulation, 86(4):641–655, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. URL http://www.tandfonline.

com/doi/abs/10.1080/00949655.2015.1027894.

Bee:2016:SAE

[BT16b] M. Bee and L. Trapin. A simple approach to the estimation of Tukey’s gh distribution. Journal of Statistical Computation and Simulation, 86(16):3287–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Becker:2018:SAA

[BTD18] W. E. Becker, S. Tarantola, and G. Deman. Sensitivity analysis approaches to high-dimensional screening problems at low sample size. Journal of Statistical Computation and Simulation, 88(11):2089–2110, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Baddeley:2016:ACL

[BTR16] Adrian Baddeley, Rolf Turner, and Ege Rubak. Adjusted composite likelihood ratio test for spatial Gibbs point processes. Journal of Statistical Computation and Simulation, 86(5):922–941, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

URL http://www.tandfonline.com/doi/abs/10.1080/
00949655.2015.1044530.

Balamurali:2017:DDE

[BU17] S. Balamurali and M. Usha. Developing and designing of an efficient variables sampling system based on the process capability index. Journal of Statistical Computation and Simulation, 87(7):1401–1415, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Burch:2014:EKC

[Bur14] Brent D. Burch. Estimating kurtosis and confidence intervals for the variance under nonnormality. Journal of Statistical Computation and Simulation, 84(12):2710–2720, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bourguignon:2015:IEP

[BV15a] M. Bourguignon and K. L. P. Vasconcellos. Improved estimation for Poisson INAR(1) models. Journal of Statistical Computation and Simulation, 85(12):2425–2441, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Burton:2015:ATT

[BV15b] J. H. Burton and J. Volaufova. Approximate testing in twostage nonlinear mixed models. Journal of Statistical Computation and Simulation, 85(13):2656–2665, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bahamonde:2016:RCF

[BV16] Natalia Bahamonde and Helena Veiga. A robust closed-form estimator for the GARCH(1, 1) model. Journal of Statistical Computation and Simulation, 86(8):1605–1619, 2016.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (elec-

tronic), 1563-5163. URL http://www.tandfonline.com/
doi/abs/10.1080/00949655.2015.1077387.

Bourguignon:2016:PIP

[BVRI16] Marcelo Bourguignon, Klaus L. P. Vasconcellos, Valdério A. Reisen, and Márton Ispány. A Poisson INAR(1) process with a seasonal structure. Journal of Statistical Computation and Simulation, 86(2):373–387, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2014:IME

[BZ14] N. Balakrishnan and Xiaojun Zhu. An improved method of estimation for the parameters of the Birnbaum–Saunders distribution. Journal of Statistical Computation and Simulation, 84(10):2285–2294, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Balakrishnan:2016:ELB

[BZ16a] N. Balakrishnan and Xiaojun Zhu. Exact likelihood-based point and interval estimation for Laplace distribution based on Type-II right censored samples. Journal of Statistical Computation and Simulation, 86(1):29–54, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bindele:2016:SRR

[BZ16b] Huybrechts F. Bindele and Yichuan Zhao. Signed-rank regression inference via empirical likelihood. Journal of Statistical Computation and Simulation, 86(4):729–739, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (elec-

tronic), 1563-5163. URL http://www.tandfonline.com/
doi/abs/10.1080/00949655.2015.1032288.

Balakrishnan:2015:RCS

[BZAZ15] N. Balakrishnan, Xiaojun Zhu, and Bander Al-Zahrani. Recursive computation of the single and product moments of order statistics from the complementary exponentialgeometric distribution. Journal of Statistical Computation and Simulation, 85(11):2187–2201, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bai:2018:MCC

[BZF18] Yu Bai, Shaofu Zhou, and Zhaoyuan Fan. A Monte Carlo comparison of GMM and QMLE estimators for short dynamic panel data models with spatial errors. Journal of Statistical Computation and Simulation, 88(2):376–409, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chitsaz:2012:SER

[CA12] Shabnam Chitsaz and S. Ejaz Ahmed. Shrinkage estimation for the regression parameter matrix in multivariate regression model. Journal of Statistical Computation and Simulation, 82(2):309–323, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Coskun:2019:DMG

[CA19] Burcin Coskun and O. Alpu. Diagnostics of multiple group influential observations for logistic regression models. Journal of Statistical Computation and Simulation, 89(16):3118– 3136, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2017:HDC

[CAC17] Yu-Chuan Chen, Hongshik Ahn, and James J. Chen. Highdimensional canonical forest. Journal of Statistical Computation and Simulation, 87(5):845–854, 2017. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2016:THL

[CAM16] Gauss M. Cordeiro, Morad Alizadeh, and Pedro Rafael Diniz Marinho. The Type I half-logistic family of distributions. Journal of Statistical Computation and Simulation, 86(4):707–728, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

URL http://www.tandfonline.com/doi/abs/10.1080/
00949655.2015.1031233.

Cordeiro:2017:GOL

[CAO^{+}17] Gauss Moutinho Cordeiro, Morad Alizadeh, Gamze Ozel, Bistoon Hosseini, Edwin Moises Marcos Ortega, and Emrah Altun. The generalized odd log-logistic family of distributions: properties, regression models and applications.

Journal of Statistical Computation and Simulation, 87(5): 908–932, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Caron:2016:MCE

[Car16] Pier-Olivier Caron. A Monte Carlo examination of the broken-stick distribution to identify components to retain in principal component analysis. Journal of Statistical Computation and Simulation, 86(12):2405–2410, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Caron:2019:CTE

[Car19] Pier-Olivier Caron. A comparison of the type I error rates of three assessment methods for indirect effects. Journal of Statistical Computation and Simulation, 89(8):1343–1356, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chakraborty:2010:AMO

[CB10] Surupa Chakraborty and Tathagata Banerjee. Analysis of mixed outcomes: misclassified binary responses and measurement error in covariates. Journal of Statistical Computation and Simulation, 80(11):1197–1209, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Campano:2011:RET

[CB11a] Wendell Q. Campano and Erniel B. Barrios. Robust estimation of a time series model with structural change. Journal of Statistical Computation and Simulation, 81(7):909–927, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cunningham:2011:BAA

[CB11b] Caitlin M. Cunningham and James G. Booth. A Bayesian approach to analysis of covariance in balanced randomized block experiments. Journal of Statistical Computation and Simulation, 81(11):1449–1460, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Corzo:2013:MRT

[CB13] J. Corzo and G. Babativa. A modified runs test for symmetry. Journal of Statistical Computation and Simulation, 83(5):984–991, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chatterjee:2019:NAC

[CB19] Suryasish Chatterjee and Uttam Bandyopadhyay. Nonparametric approaches for comparing three-period, twotreatment, four-sequence crossover designs. Journal of Statistical Computation and Simulation, 89(7):1153–1182, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cramer:2016:EDM

[CBG16] E. Cramer, M. Burkschat, and J. Górny. On the exact distribution of the MLEs based on Type-II progressively hybrid censored data from exponential distributions. Journal of Statistical Computation and Simulation, 86(10):2036–2052, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chang:2011:IME

[CC11a] Yu-Mei Chang and Yuh-Ing Chen. Identification of the minimum effective dose based on weighted Kaplan–Meier statistics. Journal of Statistical Computation and Simulation, 81(5):619–634, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheng:2011:AUD

[CC11b] Chuen-Sheng Cheng and Pei-Wen Chen. An ARL-unbiased design of time-between-events control charts with runs rules. Journal of Statistical Computation and Simulation, 81(7): 857–871, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cuevas:2012:TTI

[CC12] Jordan P. Cuevas and Eric Chicken. A trimmed translationinvariant denoising estimator. Journal of Statistical Computation and Simulation, 82(9):1299–1310, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chatterjee:2016:SAC

[CC16a] Moutushi Chatterjee and Ashis Kumar Chakraborty. A simple algorithm for calculating values for folded normal distribution. Journal of Statistical Computation and Simulation, 86(2):293–305, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2016:TSS

[CC16b] Chun-Shu Chen and Kun-Cih Chen. Two-stage signal restoration based on a modified median filter. Journal of Statistical Computation and Simulation, 86(1):122–134, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cerasa:2017:OFM

[CC17a] Andrea Cerasa and Andrea Cerioli. Outlier-free merging of homogeneous groups of pre-classified observations under contamination. Journal of Statistical Computation and Simulation, 87(15):2997–3020, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2017:RSS

[CC17b] Jinyuan Chen and Conghua Cheng. Reliability of stressstrength model for exponentiated Pareto distributions. Journal of Statistical Computation and Simulation, 87(4): 791–805, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Canals:2019:WLE

[CC19] Catalina Canals and Andrea Canals. When is n large enough? Looking for the right sample size to estimate proportions. Journal of Statistical Computation and Simulation, 89(10):1887–1898, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cysneiros:2010:CML

[CCC10] F. J. A. Cysneiros, G. M. Cordeiro, and A. H. M. A. Cysneiros. Corrected maximum likelihood estimators in heteroscedastic symmetric nonlinear models. Journal of Statistical Computation and Simulation, 80(4):451–461, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cafri:2019:HDV

[CCF19] Guy Cafri, Peter Calhoun, and Juanjuan Fan. High dimensional variable selection with clustered data: an application of random multivariate survival forests for detection of outlier medical device components. Journal of Statistical Computation and Simulation, 89(8):1410–1422, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cho:2013:MCM

[CCK13] Sun-Joo Cho, Allan S. Cohen, and Seock-Ho Kim. Markov chain Monte Carlo estimation of a mixture item response theory model. Journal of Statistical Computation and Simulation, 83(2):278–306, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chalise:2012:BAE

[CCM12] Prabhakar Chalise, Eric Chicken, and Daniel McGee. Baseline age effect on parameter estimates in Cox models. Journal of Statistical Computation and Simulation, 82(12):1767– 1774, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cepeda-Cuervo:2014:GLM

[CCMGA14] Edilberto Cepeda-Cuervo, Helio S. Migon, Liliana Garrido, and Jorge A. Achcar. Generalized linear models with random effects in the two-parameter exponential family. Journal of Statistical Computation and Simulation, 84(3):513– 525, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Caiado:2012:TCT

[CCP12] Jorge Caiado, Nuno Crato, and Daniel Peña. Tests for comparing time series of unequal lengths. Journal of Statistical Computation and Simulation, 82(12):1715–1725, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chandra:2012:SAE

[CCS12] Hukum Chandra, Ray Chambers, and Nicola Salvati. Small area estimation of proportions in business surveys. Journal of Statistical Computation and Simulation, 82(6):783–795, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheng:2013:EGEa

[CCZ13a] Conghua Cheng, Jinyuan Chen, and Haiqing Zhao. Estimation of the generalized exponential renewal function. Journal of Statistical Computation and Simulation, 83(8):1570– 1583, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheng:2013:EGEb

[CCZ13b] Conghua Cheng, Jinyuan Chen, and Haiqing Zhao. Estimation of the generalized exponential renewal function. Journal of Statistical Computation and Simulation, 83(9):1781, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cao:2017:BIH

[CCZJ17] Chunzheng Cao, Mengqian Chen, Xiaoxin Zhu, and Shaobo Jin. Bayesian inference in a heteroscedastic replicated measurement error model using heavy-tailed distributions. Journal of Statistical Computation and Simulation, 87(15): 2915–2928, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2011:NFG

[CdC11] Gauss M. Cordeiro and Mário de Castro. A new family of generalized distributions. Journal of Statistical Computation and Simulation, 81(7):883–898, 2011. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cape:2015:EHB

[CDG^{+}15] Joshua Cape, William Dearden, William Gamber, Jeffrey Liebner, Qin Lu, and M. Linh Nguyen. Estimating Heston’s and bates’ models parameters using Markov chain Monte Carlo simulation. Journal of Statistical Computation and Simulation, 85(11):2295–2314, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

CardosoDeOliveira:2010:MEC

[CF10] I. R. Cardoso De Oliveira and D. F. Ferreira. Multivariate extension of chi-squared univariate normality test. Journal of Statistical Computation and Simulation, 80(5):513–526, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Christou:2015:CTS

[CF15] Vasiliki Christou and Konstantinos Fokianos. On count time series prediction. Journal of Statistical Computation and Simulation, 85(2):357–373, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Crackel:2017:BIF

[CF17] Roberto Crackel and James Flegal. Bayesian inference for a flexible class of bivariate beta distributions. Journal of Statistical Computation and Simulation, 87(2):295–312, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Caeiro:2015:BRE

[CG15] Frederico Caeiro and M. Ivette Gomes. Bias reduction in the estimation of a shape second-order parameter of a heavytailed model. Journal of Statistical Computation and Simulation, 85(17):3405–3419, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Caporale:2010:FID

[CGA10] Guglielmo Maria Caporale and Luis A. Gil-Alana. Fractional integration and data frequency. Journal of Statistical Computation and Simulation, 80(2):121–132, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Caporale:2019:LMD

[CGAP19] Guglielmo Maria Caporale, Luis Gil-Alana, and Alex Plastun. Long memory and data frequency in financial markets. Journal of Statistical Computation and Simulation, 89 (10):1763–1779, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2014:AEB

[CGdS14] Gauss M. Cordeiro, Antonio E. Gomes, and Cibele Q. da Silva. Another extended Burr III model: some properties and applications. Journal of Statistical Computation and Simulation, 84(12):2524–2544, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2013:BEW

[CGdSO13] Gauss M. Cordeiro, Antonio Eduardo Gomes, Cibele Queiroz da Silva, and Edwin M. M. Ortega. The beta exponentiated Weibull distribution. Journal of Statistical Computation and Simulation, 83(1):114–138, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chastaing:2015:GSS

[CGP15] G. Chastaing, F. Gamboa, and C. Prieur. Generalized Sobol sensitivity indices for dependent variables: numerical methods. Journal of Statistical Computation and Simulation, 85 (7):1306–1333, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Conover:2018:UCS

[CGSTG18] William Jay Conover, Armando Jesús Guerrero-Serrano, and Víctor Gustavo Tercero-Gómez. An update on ‘a comparative study of tests for homogeneity of variance’. Journal of Statistical Computation and Simulation, 88(8):1454– 1469, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Coia:2014:SME

[CH14] Vincenzo Coia and Mei Ling Huang. A sieve model for extreme values. Journal of Statistical Computation and Simulation, 84(8):1692–1710, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2015:ACA

[CH15] Jen-Hao Chen and Wen-Liang Hung. An automatic clustering algorithm for probability density functions. Journal of Statistical Computation and Simulation, 85(15):3047–3063, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cascone:2019:QML

[CH19] Marcos H. Cascone and Luiz K. Hotta. Quasi-maximum likelihood estimation of GARCH models in the presence of missing values. Journal of Statistical Computation and Simulation, 89(2):292–314, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chang:2014:MSQ

[Cha14] Youngjae Chang. Multi-step quantile regression tree. Journal of Statistical Computation and Simulation, 84(3):663– 682, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chang:2015:DSE

[Cha15] S.-M. Chang. Double shrinkage estimators for large sparse covariance matrices. Journal of Statistical Computation and Simulation, 85(8):1497–1511, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chamroukhi:2016:ULR

[Cha16] Faicel Chamroukhi. Unsupervised learning of regression mixture models with unknown number of components. Journal of Statistical Computation and Simulation, 86(12): 2308–2334, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chatterjee:2017:EEE

[Cha17] Moutushi Chatterjee. Exact expression for the expectation of estimated C_{pk} based on control chart information and the corresponding process capability control charts. Journal of Statistical Computation and Simulation, 87(5):1025–1047, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chakraborty:2018:GWM

[CHB18] Niladri Chakraborty, Schalk W. Human, and Narayanaswamy Balakrishnan. A generally weighted moving average exceedance chart. Journal of Statistical Computation and Simulation, 88(9):1759–1781, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2011:SME

[Che11] Zhenmin Chen. A simple method for estimating parameters of the location-scale distribution family. Journal of Statistical Computation and Simulation, 81(1):49–58, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2014:LBT

[Che14] Lee-Shen Chen. Locally best tests for exponential distributions with interval-censored data. Journal of Statistical Computation and Simulation, 84(9):1946–1959, 2014. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2018:MFC

[Che18] Xiaolin Chen. Model-free conditional feature screening for ultra-high dimensional right censored data. Journal of Statistical Computation and Simulation, 88(12):2425–2446, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheng:2019:ELR

[Che19] Conghua Cheng. Empirical likelihood ratio for two-sample compound Poisson process problems. Journal of Statistical Computation and Simulation, 89(16):3035–3045, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chiu:2010:PBA

[Chi10] Sung Nok Chiu. Parametric bootstrap and approximate tests for two Poisson variates. Journal of Statistical Computation and Simulation, 80(3):263–271, 2010. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2017:ITS

[CHQ17] Zhongxue Chen, Hanwen Huang, and Peihua Qiu. An improved two-stage procedure to compare hazard curves. Journal of Statistical Computation and Simulation, 87(9):1877– 1886, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Christie:2015:EMF

[Chr15] David Christie. Efficient von Mises–Fisher concentration parameter estimation using Taylor series. Journal of Statistical Computation and Simulation, 85(16):3259–3265, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheng:2016:QRM

[CHT16] Jung-Yu Cheng, Shu-Chun Huang, and Shinn-Jia Tzeng. Quantile regression methods for left-truncated and rightcensored data. Journal of Statistical Computation and Simulation, 86(3):443–459, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cai:2014:SMF

[CHTZ14] Yuzhi Cai, Jie Huang, Yu Tang, and Guixia Zhou. A simulation method for finite non-stationary time series. Journal of Statistical Computation and Simulation, 84(7):1563–1579, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2015:BGS

[CI15] Shyh-Huei Chen and Edward H. Ip. Behaviour of the Gibbs sampler when conditional distributions are potentially incompatible. Journal of Statistical Computation and Simulation, 85(16):3266–3275, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ciuperca:2013:ELN

[Ciu13] Gabriela Ciuperca. Empirical likelihood for nonlinear models with missing responses. Journal of Statistical Computation and Simulation, 83(4):739–758, 2013. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chandler:2013:ALA

[CJ13] Gabriel Chandler and Leif T. Johnson. Automatic locally adaptive smoothing for tree-based set estimation. Journal of Statistical Computation and Simulation, 83(2):384–401, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Casals:2010:DSS

[CJS10] José Casals, Miguel Jerez, and Sonia Sotoca. Decomposition of a state-space model with inputs. Journal of Statistical Computation and Simulation, 80(9):979–992, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheon:2014:BSC

[CK14] Sooyoung Cheon and Jaehee Kim. A Bayesian structuralchange analysis via the stochastic approximation Monte Carlo and Gibbs sampler. Journal of Statistical Computation and Simulation, 84(7):1444–1470, 2014. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cetinkaya:2019:ITP

[ÇK19] Merve Kandemir Çetinkaya and Selahattin Kaçiranlar. Improved two-parameter estimators for the negative binomial and Poisson regression models. Journal of Statistical Computation and Simulation, 89(14):2645–2660, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Choi:2015:FLA

[CKP15] Hosik Choi, Ja-Yong Koo, and Changyi Park. Fused least absolute shrinkage and selection operator for credit scoring. Journal of Statistical Computation and Simulation, 85 (11):2135–2147, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chatterjee:2016:CAU

[CKP16] K. Chatterjee, C. Koukouvinos, and C. Parpoula. Computeraided unbalanced supersaturated designs involving interactions. Journal of Statistical Computation and Simulation, 86(4):756–770, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

URL http://www.tandfonline.com/doi/abs/10.1080/
00949655.2015.1035655.

Chongfuangprinya:2011:ISV

[CKPS11] Panitarn Chongfuangprinya, Seoung Bum Kim, Sun-Kyoung Park, and Thuntee Sukchotrat. Integration of support vector machines and control charts for multivariate process monitoring. Journal of Statistical Computation and Simulation, 81(9):1157–1173, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2010:NRL

[CL10] Gauss M. Cordeiro and Sóstenes L. Lins. New results on the likelihood ratio and score tests for the von Mises distribution. Journal of Statistical Computation and Simulation, 80(8):915–926, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cao:2011:DEL

[CL11] Chun-Zheng Cao and Jin-Guan Lin. Diagnostics for elliptical linear mixed models with first-order autoregressive errors. Journal of Statistical Computation and Simulation, 81 (10):1281–1296, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Changpetch:2013:MSL

[CL13] Pannapa Changpetch and Dennis K. J. Lin. Model selection for logistic regression via association rules analysis. Journal of Statistical Computation and Simulation, 83(8):1415– 1428, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2014:SMI

[CL14] Yuh-Ing Chen and Wen-Ming Lin. Statistical model and inference for pharmacokinetic data. Journal of Statistical Computation and Simulation, 84(12):2607–2618, 2014. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chastaing:2015:ADC

[CL15] G. Chastaing and L. Le Gratiet. ANOVA decomposition of conditional Gaussian processes for sensitivity analysis with dependent inputs. Journal of Statistical Computation and Simulation, 85(11):2164–2186, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2016:LUR

[CL16] Cathy W. S. Chen and Sangyeol Lee. A local unit root test in mean for financial time series. Journal of Statistical Computation and Simulation, 86(4):788–806, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1037765.

Conceicao:2017:ZMP

[CLAH17] K. S. Conceição, F. Louzada, M. G. Andrade, and E. S. Helou. Zero-modified power series distribution and its Hurdle distribution version. Journal of Statistical Computation and Simulation, 87(9):1842–1862, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2017:HRF

[CLC17] Dacheng Chen, Jialiang Li, and Juin Kuan Chong. Hazards regression for freemium products and services: a competing risks approach. Journal of Statistical Computation and Simulation, 87(9):1863–1876, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Choi:2019:HDM

[CLC19] Young-Geun Choi, Johan Lim, and Sujung Choi. Highdimensional Markowitz portfolio optimization problem: empirical comparison of covariance matrix estimators. Journal of Statistical Computation and Simulation, 89(7):1278– 1300, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cancho:2016:NLM

[CLDB16] Vicente G. Cancho, Francisco Louzada, Dipak K. Dey, and Gladys D. C. Barriga. A new lifetime model for multivariate survival data with a surviving fraction. Journal of Statistical Computation and Simulation, 86(2):279–292, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2014:AOS

[CLH14] Chun-Shu Chen, Yun-Huan Lee, and Hung-Wei Hsu. Adaptive order selection for autoregressive models. Journal of Statistical Computation and Simulation, 84(9):1963–1974, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Crato:2010:SPD

[CLL10] N. Crato, R. R. Linhares, and S. R. C. Lopes. Statistical properties of detrended fluctuation analysis. Journal of Statistical Computation and Simulation, 80(6):625–641, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2019:CCM

[CLS^{+}19] Zhuang Chen, Chao Lu, Yan Shen, Rui Wang, and Xuejun Wang. On complete and complete moment convergence for weighted sums of ANA random variables and applications. Journal of Statistical Computation and Simulation, 89(15):2871–2898, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cui:2018:APN

[CLYX18] Jiarong Cui, Chunxiao Li, Ke Yang, and Wangli Xu. Adaptive positive and negative runs test. Journal of Statistical Computation and Simulation, 88(7):1314–1335, 2018. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Capizzi:2010:CSE

[CM10] Giovanna Capizzi and Guido Masarotto. Combined Shewhart–EWMA control charts with estimated parameters. Journal of Statistical Computation and Simulation, 80(7):793–807, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Canal:2014:CSC

[CM14] Luisa Canal and Rocco Micciolo. The chi-square controversy: what if Pearson had R? Journal of Statistical Computation and Simulation, 84(5):1015–1021, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chatterjee:2016:EHP

[CM16] Kiranmoy Chatterjee and Diganta Mukherjee. On the estimation of homogeneous population size from a complex dual-record system. Journal of Statistical Computation and Simulation, 86(17):3562–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chacko:2017:EPK

[CM17] Manoj Chacko and Rakhi Mohan. Estimation of parameters of Kumaraswamy–Exponential distribution under progressive type-II censoring. Journal of Statistical Computation and Simulation, 87(10):1951–1963, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2019:GGD

[CM19] Ying-Ju Chen and Tatjana Miljkovic. From grouped to degrouped data: a new approach in distribution fitting for grouped data. Journal of Statistical Computation and Simulation, 89(2):272–291, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Castellares:2013:BLN

[CMC13] F. Castellares, L. C. Montenegro, and G. M. Cordeiro. The beta log-normal distribution. Journal of Statistical Computation and Simulation, 83(2):203–228, 2013. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Castillo:2012:IPE

[CMCH12] Enrique Castillo, Roberto Mínguez, Carmen Castillo, and Ali S. Hadi. Improving parameter estimation using constrained optimization methods. Journal of Statistical Computation and Simulation, 82(10):1471–1499, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cheng:2017:SMF

[CMX17] Yuan Cheng, Amitava Mukherjee, and Min Xie. Simultaneously monitoring frequency and magnitude of events based on bivariate gamma distribution. Journal of Statistical Computation and Simulation, 87(9):1723–1741, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cribari-Neto:2017:NNH

[CNL17] Francisco Cribari-Neto and Sadraque E. F. Lucena. Nonnested hypothesis testing inference for GAMLSS models. Journal of Statistical Computation and Simulation, 87(6): 1189–1205, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2013:GRB

[CNO13] Gauss M. Cordeiro, Saralees Nadarajah, and Edwin M. M. Ortega. General results for the beta Weibull distribution. Journal of Statistical Computation and Simulation, 83(6):1082–1114, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cribari-Neto:2019:TIH

[CNP19] Francisco Cribari-Neto and Inara F. S. Pereira. Testing inference in heteroskedastic linear regressions: a comparison of two alternative approaches. Journal of Statistical Computation and Simulation, 89(8):1437–1465, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cribari-Neto:2014:TIB

[CNQ14] Francisco Cribari-Neto and Marcela P. F. Queiroz. On testing inference in beta regressions. Journal of Statistical Computation and Simulation, 84(1):186–203, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cribari-Neto:2012:TIV

[CNS12] Francisco Cribari-Neto and Tatiene C. Souza. Testing inference in variable dispersion beta regressions. Journal of Statistical Computation and Simulation, 82(12):1827–1843, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cancho:2011:CMP

[COBH11] Vicente G. Cancho, Edwin M. M. Ortega, Gladys D. C. Barriga, and Elizabeth M. Hashimoto. The Conway–Maxwell– Poisson-generalized gamma regression model with longterm survivors. Journal of Statistical Computation and Simulation, 81(11):1461–1481, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Coin:2013:MEP

[Coi13] Daniele Coin. A method to estimate power parameter in exponential power distribution via polynomial regression. Journal of Statistical Computation and Simulation, 83(11):1981–2001, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2014:EWL

[COL14] Gauss M. Cordeiro, Edwin M. M. Ortega, and Artur J. Lemonte. The exponential-Weibull lifetime distribution. Journal of Statistical Computation and Simulation, 84(12): 2592–2606, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Congdon:2010:SSE

[Con10] Peter Congdon. A spatial structural equation model with an application to area health needs. Journal of Statistical Computation and Simulation, 80(4):401–412, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2014:GLF

[COP14] Gauss M. Cordeiro, Edwin M. M. Ortega, and Bozidar V. Popović. The gamma-linear failure rate distribution: theory and applications. Journal of Statistical Computation and Simulation, 84(11):2408–2426, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2015:GLD

[COP15] Gauss M. Cordeiro, Edwin M. M. Ortega, and Bozidar V. Popović. The gamma-Lomax distribution. Journal of Statistical Computation and Simulation, 85(2):305–319, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2013:SFB

[Cor13] Gauss M. Cordeiro. A simple formula based on quantiles for the moments of beta generalized distributions. Journal of Statistical Computation and Simulation, 83(10):1932–1943, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2011:EGG

[COS11] Gauss M. Cordeiro, Edwin M. M. Ortega, and Giovana O. Silva. The exponentiated generalized gamma distribution with application to lifetime data. Journal of Statistical Computation and Simulation, 81(7):827–842, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2014:KMW

[COS14] Gauss M. Cordeiro, Edwin M. M. Ortega, and Giovana O. Silva. The Kumaraswamy modified Weibull distribution: theory and applications. Journal of Statistical Computation and Simulation, 84(7):1387–1411, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cox:2013:ASM

[Cox13] M. A. A. Cox. Analysis of stock market indices through multidimensional scaling. Journal of Statistical Computation and Simulation, 83(11):2015–2029, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Choi:2012:APP

[CP12] Hosik Choi and Changyi Park. Approximate penalization path for smoothly clipped absolute deviation. Journal of Statistical Computation and Simulation, 82(5):643–652, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Charest:2014:UBI

[CP14] Lysiane Charest and Jean-François Plante. Using balanced iterative reducing and clustering hierarchies to compute approximate rank statistics on massive datasets. Journal of Statistical Computation and Simulation, 84(10):2214–2232, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chaturvedi:2019:SIR

[CPK19] Ajit Chaturvedi, Anupam Pathak, and Narendra Kumar. Statistical inferences for the reliability functions in the proportional hazard rate models based on progressive type-II right censoring. Journal of Statistical Computation and Simulation, 89(12):2187–2217, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Choi:2019:RBC

[CPP^{+}19] Hosik Choi, J. C. Poythress, Cheolwoo Park, Jong-June Jeon, and Changyi Park. Regularized boxplot via convex clustering. Journal of Statistical Computation and Simulation, 89(7):1227–1247, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chang:2012:BEC

[CQJ12] Chen-Ye Chang, Xi-Yuan Qian, and Sheng-Yuan Jian. Bayesian estimation and comparison of MGARCH and MSV models via WinBUGS. Journal of Statistical Computation and Simulation, 82(1):19–39, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chatterjee:2018:CBA

[CR18] Moumita Chatterjee and Sugata Sen Roy. A copula-based approach for estimating the survival functions of two alternating recurrent events. Journal of Statistical Computation and Simulation, 88(16):3098–3115, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cabras:2015:HOA

[CRV15] Stefano Cabras, Walter Racugno, and Laura Ventura. Higher order asymptotic computation of Bayesian significance tests for precise null hypotheses in the presence of nuisance parameters. Journal of Statistical Computation and Simulation, 85(15):2989–3001, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2011:PBP

[CS11] Cheng C. Chen and Sarjinder Singh. Pseudo-Bayes and pseudo-empirical Bayes estimators in randomized response sampling. Journal of Statistical Computation and Simulation, 81(6):779–793, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Calzada:2013:SCC

[CS13] Maria E. Calzada and Stephen M. Scariano. A synthetic control chart for the coefficient of variation. Journal of Statistical Computation and Simulation, 83(5):853–867, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cattelan:2016:ESA

[CS16] M. Cattelan and N. Sartori. Empirical and simulated adjustments of composite likelihood ratio statistics. Journal of Statistical Computation and Simulation, 86(5):1056–1067, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. URL http://www.tandfonline.

com/doi/abs/10.1080/00949655.2015.1053091.

Chan:2017:PBC

[CS17] Raymond K. S. Chan and Mike K. P. So. On the performance of the Bayesian composite likelihood estimation of max-stable processes. Journal of Statistical Computation and Simulation, 87(15):2869–2881, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Celik:2018:RET

[ÇŞ18a] Nuri Çelik and Birdal Şenoğlu. Robust estimation and testing in one-way ANOVA for Type II censored samples: skew normal error terms. Journal of Statistical Computation and Simulation, 88(7):1382–1393, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chan:2018:MMS

[CS18b] Raymond K. S. Chan and Mike K. P. So. Multivariate modelling of spatial extremes based on copulas. Journal of Statistical Computation and Simulation, 88(12):2404–2424, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Casals:2015:SME

[CSJ15] José Casals, Sonia Sotoca, and Miguel Jerez. Single and multiple error state-space models for signal extraction. Journal of Statistical Computation and Simulation, 85(5):1053–1069, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chang:2017:ESC

[CSJ17] Yu-Mei Chang, Pao-Sheng Shen, and Yu-Ru Jiang. Estimation of the simultaneous confidence regions for the ratios of quantile residual lifetimes. Journal of Statistical Computation and Simulation, 87(13):2581–2596, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2014:GPB

[CSPO14] Gauss M. Cordeiro, Giovana O. Silva, Rodrigo R. Pescim, and Edwin M. M. Ortega. General properties for the beta extended half-normal model. Journal of Statistical Computation and Simulation, 84(4):881–901, 2014. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chien:2014:PSR

[CT14] Li-Chu Chien and Tsung-Shan Tsou. Parametric simultaneous robust inferences for regression coefficient under generalized linear models. Journal of Statistical Computation and Simulation, 84(4):850–867, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2016:DHH

[CT16] Cathy W. S. Chen and Buu-Chau Truong. On double hysteretic heteroskedastic model. Journal of Statistical Computation and Simulation, 86(13):2684–2705, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chang:2017:FBV

[CTC17] S.-M. Chang, J.-Y. Tzeng, and R.-B. Chen. Fast Bayesian variable screenings for binary response regressions with small sample size. Journal of Statistical Computation and Simulation, 87(14):2708–2723, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2019:HVA

[CTTS19] Cathy W. S. Chen, Hong Than-Thi, and Mike K. P. So. On hysteretic vector autoregressive model with applications. Journal of Statistical Computation and Simulation, 89(2): 191–210, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2018:GMV

[CTX18] Wangxue Chen, Yi Tian, and Minyu Xie. The global minimum variance unbiased estimator of the parameter for a truncated parameter family under the optimal ranked set sampling. Journal of Statistical Computation and Simulation, 88(17):3399–3414, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cybis:2018:CCP

[CVL18] Gabriela B. Cybis, Marcio Valk, and Sílvia R. C. Lopes. Clustering and classification problems in genetics through U-statistics. Journal of Statistical Computation and Simulation, 88(10):1882–1902, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chan:2016:BAC

[CW16] Jennifer So Kuen Chan and Wai Yin Wan. Bayesian analysis of Cannabis offences using generalized Poisson geometric process model with flexible dispersion. Journal of Statistical Computation and Simulation, 86(16):3315–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chowdhury:2017:LBC

[CWM17] Mohammed Chowdhury, Colin Wu, and Reza Modarres. Local Box–Cox transformation on time-varying parametric models for smoothing estimation of conditional CDF with longitudinal data. Journal of Statistical Computation and Simulation, 87(15):2900–2914, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2018:TSS

[CWZ18] Hui Chen, Zhaojun Wang, and Xuemin Zi. Two-sample spatial rank test using projection. Journal of Statistical Computation and Simulation, 88(3):498–510, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2017:CBS

[CYL17] Lee-Shen Chen, Ming-Chung Yang, and TaChen Liang. Curtailed Bayesian sampling plans for exponential distributions based on Type-II censored samples. Journal of Statistical Computation and Simulation, 87(6):1160–1178, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cotos-Yanez:2016:MCN

[CYPGGM16] T. R. Cotos-Yáñez, A. Pérez-González, and W. González-Manteiga. Model checks for nonparametric regression with missing data: a comparative study. Journal of Statistical Computation and Simulation, 86(16):3188–??, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Cordeiro:2018:BXS

[CYRO18] Gauss M. Cordeiro, Haitham M. Yousof, Thiago G. Ramires, and Edwin M. M. Ortega. The Burr XII system of densities: properties, regression model and applications. Journal of Statistical Computation and Simulation, 88(3): 432–456, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Chen:2018:ALA

[CZX18] Yuhui Chen, Jiajia Zhang, and Yangyang Xu. Adaptive lasso for accelerated hazards models. Journal of Statistical Computation and Simulation, 88(15):2948–2960, 2018.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DaVeiga:2015:GSA

[Da 15] Sebastien Da Veiga. Global sensitivity analysis with dependence measures. Journal of Statistical Computation and Simulation, 85(7):1283–1305, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Duran:2013:NDB

[DA13] Esra Akdeniz Duran and Fikri Akdeniz. New differencebased estimator of parameters in semiparametric regression models. Journal of Statistical Computation and Simulation, 83(5):810–824, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Danish:2014:BIR

[DA14] Muhammad Yameen Danish and Muhammad Aslam. Bayesian inference for the randomly censored Weibull distribution. Journal of Statistical Computation and Simulation, 84(1):215–230, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dogru:2016:ORE

[DA16] Fatma Zehra Dogru and Olcay Arslan. Optimal Brobust estimators for the parameters of the Burr XII distribution. Journal of Statistical Computation and Simulation, 86(6):1133–1149, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

URL http://www.tandfonline.com/doi/abs/10.1080/
00949655.2015.1054288.

Doostparast:2011:BAT

[DAB11] M. Doostparast, M. G. Akbari, and N. Balakrishna. Bayesian analysis for the two-parameter Pareto distribution based on record values and times. Journal of Statistical Computation and Simulation, 81(11):1393–1403, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

deAndrade:2016:RNG

[dABS16] B. B. de Andrade, H. Bolfarine, and A. N. Siroky. Random number generation and estimation with the bimodal asymmetric power-normal distribution. Journal of Statistical Computation and Simulation, 86(3):460–476, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Neto:2011:BSR

[dALNCdATdC11] Eufrásio de A. Lima Neto, Gauss M. Cordeiro, and Francisco de A. T. de Carvalho. Bivariate symbolic regression models for interval-valued variables. Journal of Statistical Computation and Simulation, 81(11):1727–1744, 2011. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dey:2015:WED

[DAP15] Sanku Dey, Sajid Ali, and Chanseok Park. Weighted exponential distribution: properties and different methods of estimation. Journal of Statistical Computation and Simulation, 85(18):3641–3661, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doostparast:2010:OSS

[DB10] M. Doostparast and N. Balakrishnan. Optimal sample size for record data and associated cost analysis for exponential distribution. Journal of Statistical Computation and Simulation, 80(12):1389–1401, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doostparast:2011:ORB

[DB11] M. Doostparast and N. Balakrishnan. Optimal record-based statistical procedures for the two-parameter exponential distribution. Journal of Statistical Computation and Simulation, 81(12):2003–2019, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

diBella:2015:DBD

[dB15] Enrico di Bella. Discriminating between distributions using feed-forward neural networks. Journal of Statistical Computation and Simulation, 85(4):711–724, 2015. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Dahmen:2012:OPT

[DBC12] K. Dahmen, M. Burkschat, and E. Cramer. A- and Doptimal progressive Type-II censoring designs based on Fisher information. Journal of Statistical Computation and Simulation, 82(6):879–905, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dumanjug:2010:BPS

[DBL10] C. F. Dumanjug, E. B. Barrios, and J. R. G. Lansangan. Bootstrap procedures in a spatial-temporal model. Journal of Statistical Computation and Simulation, 80(7):809–822, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dovoedo:2013:ODM

[DC13] Y. H. Dovoedo and S. Chakraborti. Outlier detection for multivariate skew-normal data: a comparative study. Journal of Statistical Computation and Simulation, 83(4):773– 783, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dovoedo:2015:PDB

[DC15] Y. H. Dovoedo and S. Chakraborti. Power of depth-based nonparametric tests for multivariate locations. Journal of Statistical Computation and Simulation, 85(10):1987–2006, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dovoedo:2016:RSS

[DC16] Y. H. Dovoedo and S. Chakraborti. On the robustness to symmetry of some nonparametric multivariate one-sample sign-type tests. Journal of Statistical Computation and Simulation, 86(10):1936–1953, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doring:2019:PGF

[DC19] Max Döring and Erhard Cramer. On the power of goodnessof-fit tests for the exponential distribution under progressive Type-II censoring. Journal of Statistical Computation and Simulation, 89(16):2997–3034, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

deCarvalho:2012:DRT

[dCFOM12] M. de Carvalho, M. Fonseca, M. Oliveira, and J. T. Mexia. A dimension reduction technique for estimation in linear mixed models. Journal of Statistical Computation and Simulation, 82(2):219–226, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

daCruz:2016:LOL

[dCOC16] José Nilton da Cruz, Edwin M. M. Ortega, and Gauss M. Cordeiro. The log-odd log-logistic Weibull regression model: modelling, estimation, influence diagnostics and residual analysis. Journal of Statistical Computation and Simulation, 86(8):1516–1538, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1071376.

Dey:2012:BEP

[DD12] Sanku Dey and Tanujit Dey. Bayesian estimation and prediction intervals for a Rayleigh distribution under a conjugate prior. Journal of Statistical Computation and Simulation, 82(11):1651–1660, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DeCapitani:2015:RPE

[DD15] L. De Capitani and D. De Martini. Reproducibility probability estimation and testing for the Wilcoxon rank-sum test. Journal of Statistical Computation and Simulation, 85(3):468–493, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dutta:2017:TPF

[DDD17] Sandipan Dutta, Susmita Datta, and Somnath Datta. Temporal prediction of future state occupation in a multistate model from high-dimensional baseline covariates via pseudovalue regression. Journal of Statistical Computation and Simulation, 87(7):1363–1378, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Demirhan:2010:PSM

[DDDD10] Haydar Demirhan, Nimet Anil Dolgun, Yaprak Parlak Demirhan, and Muhsin Özgür Dolgun. Performance of some multiple comparison tests under heteroscedasticity and dependency. Journal of Statistical Computation and Simulation, 80(10):1083–1100, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lascio:2016:TIC

[DDJ16] F. Marta L. Di Lascio, Fabrizio Durante, and Piotr Jaworski. Truncation invariant copulas and a testing procedure. Journal of Statistical Computation and Simulation, 86(12):2362–2378, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doostparast:2013:ELD

[DDZ13] Mahdi Doostparast, Sanjel Deepak, and Amin Zangoie. Estimation with the lognormal distribution on the basis of records. Journal of Statistical Computation and Simulation, 83(12):2339–2351, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Santo:2015:CEM

[dESM15] A. P. J. do Espirito Santo and J. Mazucheli. Comparison of estimation methods for the Marshall–Olkin extended Lindley distribution. Journal of Statistical Computation and Simulation, 85(17):3437–3450, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Didier:2014:WBS

[DF14] Gustavo Didier and John Fricks. On the wavelet-based simulation of anomalous diffusion. Journal of Statistical Computation and Simulation, 84(4):697–723, 2014. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dotto:2019:RIP

[DF19] Francesco Dotto and Alessio Farcomeni. Robust inference for parsimonious model-based clustering. Journal of Statistical Computation and Simulation, 89(3):414–442, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DiMarzio:2016:NNE

[DFPT16a] M. Di Marzio, S. Fensore, A. Panzera, and C. C. Taylor. A note on nonparametric estimation of circular conditional densities. Journal of Statistical Computation and Simulation, 86(13):2573–2582, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DiMarzio:2016:PPL

[DFPT16b] M. Di Marzio, S. Fensore, A. Panzera, and C. C. Taylor. Practical performance of local likelihood for circular density estimation. Journal of Statistical Computation and Simulation, 86(13):2560–2572, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Duan:2017:BSR

[DFT17] Xingde Duan, Wing Kam Fung, and Niansheng Tang. Bayesian semiparametric reproductive dispersion mixed models for non-normal longitudinal data: estimation and case influence analysis. Journal of Statistical Computation and Simulation, 87(10):1925–1939, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Delord:2016:MIC

[DG16] Marc Delord and Emmanuelle Génin. Multiple imputation for competing risks regression with interval-censored data. Journal of Statistical Computation and Simulation, 86(11):2217–2228, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Duarte:2019:OED

[DG19] Belmiro P. M. Duarte and José F. O. Granjo. Optimal exact design of double acceptance sampling plans by attributes. Journal of Statistical Computation and Simulation, 89(17):3313–3329, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dolgun:2012:CDT

[DGK12] Nimet Anil Dolgun, Harika Gozukara, and Ergun Karaagaoglu. Comparing diagnostic tests: test of hypothesis for likelihood ratios. Journal of Statistical Computation and Simulation, 82(3):369–381, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Director:2017:ESS

[DGLV17] Hannah M. Director, James Gattiker, Earl Lawrence, and Scott Vander Wiel. Efficient sampling on the simplex with a self-adjusting logit transform proposal. Journal of Statistical Computation and Simulation, 87(18):3521–3536, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Diongue:2010:BGM

[DGW10] Abdou Kâ Diongue, Dominique Guégan, and Rodney C. Wolff. BL–GARCH models with elliptical distributed innovations. Journal of Statistical Computation and Simulation, 80(7):775–791, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Demirhan:2014:BCL

[DH14] Haydar Demirhan and Canan Hamurkaroglu. Bayesian X^{−} control limits for exponentially distributed measurements. Journal of Statistical Computation and Simulation, 84(3): 628–643, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deutsch:2014:PAV

[DHP14] Roland C. Deutsch, Brian Habing, and Walter W. Piegorsch. A pool-adjacent-violators-algorithm approach to detect infinite parameter estimates in one-regressor doseresponse models with asymptotes. Journal of Statistical Computation and Simulation, 84(12):2545–2556, 2014. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Durio:2011:HAF

[DI11] A. Durio and E. D. Isaia. The heuristic approach in finding initial values for minimum density power divergence estimators. Journal of Statistical Computation and Simulation, 81(4):471–480, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dias:2010:SGB

[Dia10] J. Rodrigues Dias. A simple generalization of the Box– Muller method for obtaining a pair of correlated standard normal variables. Journal of Statistical Computation and Simulation, 80(9):953–958, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dorugade:2010:VSL

[DK10] A. V. Dorugade and D. N. Kashid. Variable selection in linear regression based on ridge estimator. Journal of Statistical Computation and Simulation, 80(11):1211–1224, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Do:2017:AIC

[DK17] Gipeum Do and Yang-Jin Kim. Analysis of interval censored competing risk data with missing causes of failure using pseudo values approach. Journal of Statistical Computation and Simulation, 87(4):631–639, 2017. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dube:2016:GIE

[DKG16] Madhulika Dube, Hare Krishna, and Renu Garg. Generalized inverted exponential distribution under progressive first-failure censoring. Journal of Statistical Computation and Simulation, 86(6):1095–1114, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1052440.

Dong:2017:DRL

[DKY17] Yuexiao Dong, Bo Kai, and Zhou Yu. Dimension reduction via local rank regression. Journal of Statistical Computation and Simulation, 87(2):239–249, 2017. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dey:2013:CCM

[DL13] Tanujit Dey and Chae Young Lim. Comparisons of computational methods for clustered binary data. Journal of Statistical Computation and Simulation, 83(11):2030–2046, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Davies:2019:ELB

[DL19] Tilman M. Davies and Andrew B. Lawson. An evaluation of likelihood-based bandwidth selectors for spatial and spatiotemporal kernel estimates. Journal of Statistical Computation and Simulation, 89(7):1131–1152, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huayanay:2019:PAL

[dlCHBCD19] Alex de la Cruz Huayanay, Jorge L. Bazán, Vicente G. Cancho, and Dipak K. Dey. Performance of asymmetric links and correction methods for imbalanced data in binary regression. Journal of Statistical Computation and Simulation, 89(9):1694–1714, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

deLara:2017:AME

[dLHT17] Idemauro Antonio Rodrigues de Lara, John Hinde, and Cesar Augusto Taconeli. An alternative method for evaluating stationarity in transition models. Journal of Statistical Computation and Simulation, 87(15):2962–2980, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dong:2018:STC

[DLS18] Ping Dong, Lu Lin, and Yunquan Song. Significance test of clustering under high dimensional setting with applications to cancer data. Journal of Statistical Computation and Simulation, 88(17):3349–3378, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Duan:2019:FAS

[DLZ19] Qihong Duan, Junrong Liu, and Dengfu Zhao. A fast algorithm for short term electric load forecasting by a hidden semi-Markov process. Journal of Statistical Computation and Simulation, 89(5):831–843, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Drews:2012:NAT

[DM12] Kimberly L. Drews and Hossein Mansouri. Nonparametric analysis of treatment effects with missing observations. Journal of Statistical Computation and Simulation, 82(6): 807–824, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DeLozzo:2016:NIU

[DM16] Matthias De Lozzo and Amandine Marrel. New improvements in the use of dependence measures for sensitivity analysis and screening. Journal of Statistical Computation and Simulation, 86(15):3038–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mulder:2017:SAP

[DMV17] Wim De Mulder, Geert Molenberghs, and Geert Verbeke. A sensitivity analysis of probabilistic sensitivity analysis in terms of the density function for the input variables. Journal of Statistical Computation and Simulation, 87(7):1429– 1445, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dewan:2013:EAE

[DN13] Isha Dewan and Swagata Nandi. An EM algorithm for the estimation of parameters of bivariate generalized exponential distribution under random left censoring. Journal of Statistical Computation and Simulation, 83(9):1648–1660, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dey:2018:EPM

[DNMT18] Sanku Dey, Mazen Nassar, Raj Kamal Maurya, and Yogesh Mani Tripathi. Estimation and prediction of Marshall-Olkin extended exponential distribution under progressively type-II censored data. Journal of Statistical Computation and Simulation, 88(12):2287–2308, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dogu:2015:ITS

[Dog15] Eralp Dogu. Identifying the time of a step change with multivariate single control charts. Journal of Statistical Computation and Simulation, 85(8):1529–1543, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Drignei:2013:FRS

[DP13] Dorin Drignei and Dalia Eugenia Popescu. Fast regression surrogates for computer models with time-dependent outputs. Journal of Statistical Computation and Simulation, 83(6):1058–1067, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deldossi:2015:BVS

[DP15] Laura Deldossi and Roberta Paroli. Bayesian variable selection in a class of mixture models for ordinal data: a comparative study. Journal of Statistical Computation and Simulation, 85(10):1926–1944, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dube:2011:PEH

[DPK11] Sulabh Dube, Biswabrata Pradhan, and Debasis Kundu. Parameter estimation of the hybrid censored log-normal distribution. Journal of Statistical Computation and Simulation, 81(3):275–287, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dixit:2017:MDH

[DR17] Anand Dixit and Vivekananda Roy. MCMC diagnostics for higher dimensions using Kullback Leibler divergence. Journal of Statistical Computation and Simulation, 87(13):2622– 2638, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Das:2019:MLM

[DRAR19] Sumonkanti Das, Azizur Rahman, Ashraf Ahamed, and Sabbir Tahmidur Rahman. Multi-level models can benefit from minimizing higher-order variations: an illustration using child malnutrition data. Journal of Statistical Computation and Simulation, 89(6):1090–1110, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dinh:2017:CGG

[DRB17] Vu Dinh, Ann E. Rundell, and Gregery T. Buzzard. Convergence of Griddy Gibbs sampling and other perturbed Markov chains. Journal of Statistical Computation and Simulation, 87(7):1379–1400, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DelSarto:2016:MST

[DRC^{+}16] Simone Del Sarto, Maria Giovanna Ranalli, David Cappelletti, Beatrice Moroni, Stefano Crocchianti, and Silvia Castellini. Modelling spatio-temporal air pollution data from a mobile monitoring station. Journal of Statistical Computation and Simulation, 86(13):2546–2559, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Diniz:2014:BEL

[DRLP14] Carlos Alberto Ribeiro Diniz, Camila Pedrozo Rodrigues, Jose Galvão Leite, and Rubiane Maria Pires. A Bayesian estimation of lag lengths in distributed lag models. Journal of Statistical Computation and Simulation, 84(2):415–427, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dette:2010:FSC

[DS10a] Holger Dette and Regine Scheder. A finite sample comparison of nonparametric estimates of the effective dose in quantal bioassay. Journal of Statistical Computation and Simulation, 80(5):527–544, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Duarte:2010:BBC

[DS10b] D. Duarte and M. A. C. Santos. A block bootstrap comparison for sparse chains. Journal of Statistical Computation and Simulation, 80(8):889–902, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dang:2011:NSM

[DS11] Xin Dang and Robert Serfling. A numerical study of multiple imputation methods using nonparametric multivariate outlier identifiers and depth-based performance criteria with clinical laboratory data. Journal of Statistical Computation and Simulation, 81(5):547–560, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dhumal:2014:MTT

[DS14] B. R. Dhumal and D. T. Shirke. A modified test for testing exponentiality using transformed data. Journal of Statistical Computation and Simulation, 84(2):397–403, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dhumal:2015:MOS

[DS15] B. R. Dhumal and D. T. Shirke. A modified one-sample test for goodness-of-fit. Journal of Statistical Computation and Simulation, 85(2):422–429, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Bin:2018:HCD

[DS18] Riccardo De Bin and Willi Sauerbrei. Handling codependence issues in resampling-based variable selection procedures: a simulation study. Journal of Statistical Computation and Simulation, 88(1):28–55, 2018. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Das:2019:RDI

[DSCS19] Ishapathik Das, Sumen Sen, N. Rao Chaganty, and Pooja Sengupta. Regression for doubly inflated multivariate Poisson distributions. Journal of Statistical Computation and Simulation, 89(13):2549–2561, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

daSilva:2017:DCD

[dSdS17] Paulo H. D. da Silva and Cibele Q. da Silva. Dynamic changepoint detection in count time series: a particle filter approach. Journal of Statistical Computation and Simulation, 87(1):42–68, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dadgar:2016:EID

[DSE16] Aniseh Dadgar, Khalil Shafie, and Mahdi Emadi. Evidential inference for diffusion-type processes. Journal of Statistical Computation and Simulation, 86(1):183–194, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Davis:2015:CIG

[DSP15] Karelyn A. Davis, Sanjoy K. Sinha, and Chul Gyu Park. Constrained inference for generalized linear models with incomplete covariate data. Journal of Statistical Computation and Simulation, 85(4):693–710, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Santos:2019:HTD

[dSSCR19] Daiane de Souza Santos, Vicente Cancho, and Josemar Rodrigues. Hypothesis testing for the dispersion parameter of the hyper-Poisson regression model. Journal of Statistical Computation and Simulation, 89(5):763–775, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dunbar:2014:MEG

[DSVY14] Martin X. Dunbar, Hani M. Samawi, Robert Vogel, and Lili Yu. A more efficient Gibbs sampler estimation using steady-state simulation: applications to public health studies. Journal of Statistical Computation and Simulation, 84 (9):1931–1945, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ding:2013:DAL

[DT13] Chang Ding and Siu-Keung Tse. Design of accelerated life test plans under progressive Type II interval censoring with random removals. Journal of Statistical Computation and Simulation, 83(7):1330–1343, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DeBoeck:2011:STG

[DTRB11] B. De Boeck, O. Thas, J. C. W. Rayner, and D. J. Best. Smooth tests for the gamma distribution. Journal of Statistical Computation and Simulation, 81(7):843–855, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Deng:2012:VIA

[DTZZ12] Aidong Deng, Jianeng Tang, Li Zhao, and Cairong Zou. The variable-interval arithmetic coding using asymptotic deterministic randomness for data compression and encryption. Journal of Statistical Computation and Simulation, 82 (10):1545–1555, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

DeLaHuertaContreras:2015:TEV

[DVA15] Violeta De La Huerta Contreras, Humberto Vaquera Huerta, and Barry C. Arnold. A test for equality of variances with censored samples. Journal of Statistical Computation and Simulation, 85(3):450–467, 2015. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Dorre:2017:BEP

[DW17] Achim Dörre and Rafael Weißbach. Bayesian estimation of a proportional hazards model for double-censored durations. Journal of Statistical Computation and Simulation, 87(3): 493–504, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Duan:2018:OSS

[DW18] Fengjun Duan and Guanjun Wang. Optimal step-stress accelerated degradation test plans for inverse Gaussian process based on proportional degradation rate model. Journal of Statistical Computation and Simulation, 88(2):305–328, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dang:2019:SVF

[DW19a] Yuanchu Dang and Qing Wang. Simultaneous variable and factor selection via sparse group lasso in factor analysis. Journal of Statistical Computation and Simulation, 89 (14):2744–2764, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dong:2019:CMD

[DW19b] Fangli Dong and Xiaozhou Wang. A classifier for multidimensional datasets based on Bayesian multiple kernel grouping learning. Journal of Statistical Computation and Simulation, 89(11):2151–2174, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Desmond:2016:ARS

[DY16] Anthony F. Desmond and Zhenlin Yang. Asymptotically refined score and GOF tests for inverse Gaussian models. Journal of Statistical Computation and Simulation, 86(16): 3243–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dogru:2019:HHT

[DYA19] Fatma Zehra Dogru, Keming Yu, and Olcay Arslan. Heteroscedastic and heavy-tailed regression with mixtures of skew Laplace normal distributions. Journal of Statistical Computation and Simulation, 89(17):3213–3240, 2019. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ding:2010:ALT

[DYT10] Chang Ding, Chunyan Yang, and Siu-Keung Tse. Accelerated life test sampling plans for the Weibull distribution under Type I progressive interval censoring with random removals. Journal of Statistical Computation and Simulation, 80(8):903–914, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Dong:2015:MCP

[DYX15] Zhaochen Dong, Wei Yu, and Wangli Xu. A modified combined p-value multiple test. Journal of Statistical Computation and Simulation, 85(12):2479–2490, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adlouni:2019:BNC

[EB19] Salaheddine El Adlouni and Ismaila Baldé. Bayesian noncrossing quantile regression for regularly varying distributions. Journal of Statistical Computation and Simulation, 89(5):884–898, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

El-Din:2017:EGE

[EDASM17] M. M. M. Mohie El-Din, M. M. Amein, A. R. Shafay, and Samar Mohamed. Estimation of generalized exponential distribution based on an adaptive progressively Type-II censored sample. Journal of Statistical Computation and Simulation, 87(7):1292–1304, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ekiz:2015:SSC

[EE15] O. Ufuk Ekiz and Meltem Ekiz. A small-sample correction factor for S-estimators. Journal of Statistical Computation and Simulation, 85(4):794–801, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Esquivel-Frausto:2013:CAL

[EFGMD13] M. E. Esquivel-Frausto, J. A. Guerrero, and J. E. Macías-Díaz. Computational approximation of the likelihood ratio for testing the existence of change-points in a heteroscedastic series. Journal of Statistical Computation and Simulation, 83(8):1491–1506, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Esemen:2018:PEG

[EG18] Melek Esemen and Selma Gürler. Parameter estimation of generalized Rayleigh distribution based on ranked set sample. Journal of Statistical Computation and Simulation, 88(4):615–628, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

ElMaroufy:2016:BIN

[EKBO16] Hamid El Maroufy, Tewfik Kernane, Sidali Becheket, and Abdellah Ouddadj. Bayesian inference for nonlinear stochastic SIR epidemic model. Journal of Statistical Computation and Simulation, 86(11):2229–2240, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Enders:2018:CMI

[EKE^{+}18] Dirk Enders, Bianca Kollhorst, Susanne Engel, Roland Linder, and Iris Pigeot. Comparison of multiple imputation and two-phase logistic regression to analyse two-phase casecontrol studies with rich phase 1: a simulation study. Journal of Statistical Computation and Simulation, 88(11):2201– 2214, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Emami:2016:LIR

[Ema16] Hadi Emami. Local influence in ridge semiparametric models. Journal of Statistical Computation and Simulation, 86 (17):3357–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

El-Raheem:2019:IOD

[ER19] A. M. Abd El-Raheem. Inference and optimal design of multiple constant-stress testing for generalized half-normal distribution under type-II progressive censoring. Journal of Statistical Computation and Simulation, 89(16):3075–3104, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Erdogan:2013:PBU

[Erd13] Birsen Eygi Erdogan. Prediction of bankruptcy using support vector machines: an application to bank bankruptcy. Journal of Statistical Computation and Simulation, 83(8): 1543–1555, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Erb:2011:AST

[ES11] Jack Erb and Douglas G. Steigerwald. Accurately sized test statistics with misspecified conditional homoskedasticity. Journal of Statistical Computation and Simulation, 81(6):729–747, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ebrahimi:2016:ACR

[EXH16] Nader Ebrahimi, Michelle Xia, and Lei Hua. Assessing component reliability using lifetime data from systems. Journal of Statistical Computation and Simulation, 86(18):3791–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Eklund:2012:SNS

[EZ12] Martin Eklund and Silvelyn Zwanzig. SimSel: a new simulation method for variable selection. Journal of Statistical Computation and Simulation, 82(4):515–527, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fallah:2016:GMA

[Fal16] Afshin Fallah. Gaussian mixture analysis of covariance. Journal of Statistical Computation and Simulation, 86(16): 3158–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Famoye:2012:CSB

[Fam12] Felix Famoye. Comparisons of some bivariate regression models. Journal of Statistical Computation and Simulation, 82(7):937–949, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fan:2017:ABR

[Fan17] Chunpeng Fan. On analysis of binary response data in longitudinal factorial studies. Journal of Statistical Computation and Simulation, 87(1):100–122, 2017. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fan:2019:DRM

[Fan19] Zhipeng Fan. A dimensionality reduction method of continuous dependent variables based supervised Laplacian eigenmaps. Journal of Statistical Computation and Simulation, 89(11):2073–2083, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fotuhi:2019:NAB

[FAT19] Hatef Fotuhi, Amirhossein Amiri, and Ali Reza Taheriyoun. A novel approach based on multiple correspondence analysis for monitoring social networks with categorical attributed data. Journal of Statistical Computation and Simulation, 89(16):3137–3164, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ferreira:2018:EDA

[FAV18] Clécio S. Ferreira and Reinaldo B. Arellano-Valle. Estimation and diagnostic analysis in skew-generalized-normal regression models. Journal of Statistical Computation and Simulation, 88(6):1039–1059, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fernandez:2018:BIM

[FBV18] Renzo Fernández, Cristian L. Bayes, and Luis Valdivieso. A beta-inflated mean regression model with mixed effects for fractional response variables. Journal of Statistical Computation and Simulation, 88(10):1936–1957, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fernandez-Duran:2016:BAC

[FDGD16] Juan José Fernández-Durán and María Mercedes Gregorio-Domínguez. Bayesian analysis of circular distributions based on non-negative trigonometric sums. Journal of Statistical Computation and Simulation, 86(16):3175–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Feng:2014:SDB

[FF14] Yuan Feng and Wenhui Fan. A system dynamics-based simulation model of production line with cross-trained workers. Journal of Statistical Computation and Simulation, 84 (6):1190–1212, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Feng:2016:IPA

[FFP16] Yuanhua Feng, Sarah Forstinger, and Christian Peitz. On the iterative plug-in algorithm for estimating diurnal patterns of financial trade durations. Journal of Statistical Computation and Simulation, 86(12):2291–2307, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Feuer:2013:LQN

[FG13] A. Feuer and G. C. Goodwin. On-line quantization in nonlinear filtering. Journal of Statistical Computation and Simulation, 83(7):1210–1222, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fernando:2014:CEG

[FGH14] W. T. P. S. Fernando, S. Ganesalingam, and M. L. Hazelton. A comparison of estimators of the geographical relative risk function. Journal of Statistical Computation and Simulation, 84(7):1471–1485, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Figueiredo:2012:CSQ

[FGHRM12] Fernanda Figueiredo, M. Ivette Gomes, Lígia Henriques-Rodrigues, and M. Cristina Miranda. A computational study of a quasi-PORT methodology for VaR based on second-order reduced-bias estimation. Journal of Statistical Computation and Simulation, 82(4):587–602, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fonseca:2014:CPD

[FGV14] Giovanni Fonseca, Federica Giummolè, and Paolo Vidoni. Calibrating predictive distributions. Journal of Statistical Computation and Simulation, 84(2):373–383, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fan:2011:NSP

[FH11] Shu-Kai S. Fan and Kuo-Nan Huang. A new search procedure of steepest ascent in response surface exploration. Journal of Statistical Computation and Simulation, 81(6): 661–678, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Feizjavadian:2015:MRW

[FH15] S. H. Feizjavadian and R. Hashemi. Mean residual weighted versus the length-biased Rayleigh distribution. Journal of Statistical Computation and Simulation, 85(14):2823–2838, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fagerland:2016:TGF

[FH16] Morten W. Fagerland and David W. Hosmer. Tests for goodness of fit in ordinal logistic regression models. Journal of Statistical Computation and Simulation, 86(17):3398–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fu:2015:EMS

[FHO15] Chong Yau Fu, Hsin-Yi Huang, and Yueh-Hsing Ou. The effects of missing serial effects and/or heteroscedastic errors on mixed models using repeated growth data. Journal of Statistical Computation and Simulation, 85(16):3367–3382, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Feddag:2012:PML

[FHS12] M.-L. Feddag, J.-B. Hardouin, and V. Sebille. Pairwise- and marginal-likelihood estimation for the mixed Rasch model with binary data. Journal of Statistical Computation and Simulation, 82(3):419–430, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Faraz:2014:DOE

[FHSC14] Alireza Faraz, Cédric Heuchenne, Erwin Saniga, and Antonio F. B. Costa. Double-objective economic statistical design of the VP T^{2} control chart: Wald’s identity approach. Journal of Statistical Computation and Simulation, 84(10):2123–2137, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Faroughi:2017:BZI

[FI17] Pouya Faroughi and Noriszura Ismail. Bivariate zeroinflated negative binomial regression model with applications. Journal of Statistical Computation and Simulation, 87(3):457–477, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fragiadakis:2013:IPV

[FKM13] K. Fragiadakis, D. Karlis, and S. G. Meintanis. Inference procedures for the variance gamma model and applications. Journal of Statistical Computation and Simulation, 83(3): 555–567, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Faraz:2010:EES

[FKS10] Alireza Faraz, R. B. Kazemzadeh, and Erwin Saniga. Economic and economic statistical design of T^{2} control chart with two adaptive sample sizes. Journal of Statistical Computation and Simulation, 80(12):1299–1316, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ferreira:2015:IDS

[FLB15] Clécio S. Ferreira, Víctor H. Lachos, and Heleno Bolfarine. Inference and diagnostics in skew scale mixtures of normal regression models. Journal of Statistical Computation and Simulation, 85(3):517–537, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Forbes:2015:FAS

[FM15a] Peter G. M. Forbes and Kanti V. Mardia. A fast algorithm for sampling from the posterior of a von Mises distribution. Journal of Statistical Computation and Simulation, 85 (13):2693–2701, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fusek:2015:SAT

[FM15b] Michal Fusek and Jaroslav Michálek. Statistical analysis of Type I multiply left-censored samples from exponential distribution. Journal of Statistical Computation and Simulation, 85(11):2148–2163, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fallah:2019:AOP

[FM19] Somayeh Fallah and Farshid Mehrdoust. American option pricing under double Heston stochastic volatility model: simulation and strong convergence analysis. Journal of Statistical Computation and Simulation, 89(7):1322–1339, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ferreira:2016:MSO

[FMB16] Clécio S. Ferreira, Thalita B. Mattos, and Narayanaswamy Balakrishnan. Mean-shift outliers model in skew scalemixtures of normal distributions. Journal of Statistical Computation and Simulation, 86(12):2346–2361, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Furstberger:2015:EPL

[FMK15] Axel Fürstberger, Markus Maucher, and Hans A. Kestler. Extended pairwise local alignment of wild card DNA/RNA sequences using dynamic programming. Journal of Statistical Computation and Simulation, 85(1):3–13, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Florez:2019:FTS

[FMV^{+}19] Alvaro J. Flórez, Geert Molenberghs, Geert Verbeke, Michael G. Kenward, Pavlos Mamouris, and Bert Vaes. Fast two-stage estimator for clustered count data with overdispersion. Journal of Statistical Computation and Simulation, 89(14):2678–2693, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fan:2018:QRC

[FMZ18] Caiyun Fan, Huijuan Ma, and Yong Zhou. Quantile regression for competing risks analysis under case-cohort design. Journal of Statistical Computation and Simulation, 88(6):1060–1080, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ferreira:2017:EPT

[FNRCM17] Guillermo Ferreira, Jean P. Navarrete, Francisco J. Rodríguez-Cortés, and Jorge Mateu. Estimation and prediction of time-varying GARCH models through a statespace representation: a computational approach. Journal of Statistical Computation and Simulation, 87(12):2430–2449, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fachini:2014:BRM

[FOC14] Juliana B. Fachini, Edwin M. M. Ortega, and Gauss M. Cordeiro. A bivariate regression model with cure fraction. Journal of Statistical Computation and Simulation, 84(7):1580–1595, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fenga:2011:BBA

[FP11a] Livio Fenga and Dimitris N. Politis. Bootstrap-based ARMA order selection. Journal of Statistical Computation and Simulation, 81(7):799–814, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ferrari:2011:ILI

[FP11b] Silvia L. P. Ferrari and Eliane C. Pinheiro. Improved likelihood inference in beta regression. Journal of Statistical Computation and Simulation, 81(4):431–443, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ferrari:2014:SSL

[FP14] Silvia Ferrari and Eliane C. Pinheiro. Small-sample likelihood inference in extreme-value regression models. Journal of Statistical Computation and Simulation, 84(3):582–595, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fenga:2015:BOS

[FP15a] Livio Fenga and Dimitris N. Politis. Bootstrap order selection for SETAR models. Journal of Statistical Computation and Simulation, 85(2):235–250, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fraley:2015:MAR

[FP15b] Chris Fraley and Daniel Percival. Model-averaged l_{1} regularization using Markov chain Monte Carlo model composition. Journal of Statistical Computation and Simulation, 85(6):1090–1101, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fenga:2017:LOS

[FP17] Livio Fenga and Dimitris N. Politis. LASSO order selection for sparse autoregression: a bootstrap approach. Journal of Statistical Computation and Simulation, 87(14):2668–2688, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Font:2013:BAF

[FPG13] M. Font, X. Puig, and J. Ginebra. A Bayesian analysis of frequency count data. Journal of Statistical Computation and Simulation, 83(2):229–246, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ferreira:2018:EST

[FPP18] Guillermo Ferreira, Nicolas Piña, and Emilio Porcu. Estimation of slowly time-varying trend function in long memory regression models. Journal of Statistical Computation and Simulation, 88(10):1903–1920, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fedou:2015:EMM

[FR15] J.-M. Fédou and M.-J. Rendas. Extending Morris method: identification of the interaction graph using cycle-equitable designs. Journal of Statistical Computation and Simulation, 85(7):1398–1419, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Frey:2012:EKS

[Fre12] Jesse Frey. An exact Kolmogorov–Smirnov test for the Poisson distribution with unknown mean. Journal of Statistical Computation and Simulation, 82(7):1023–1033, 2012. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Frey:2016:MEM

[Fre16] Jesse Frey. A more efficient mean estimator for judgement post-stratification. Journal of Statistical Computation and Simulation, 86(7):1404–1414, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Feng:2017:IMR

[FRL17] Cindy X. Feng, Mehdi Rostami, and Longhai Li. Impact of misspecified residual correlation structure on the parameter estimates in a shared spatial frailty model. Journal of Statistical Computation and Simulation, 87(12):2384–2410, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Faria:2010:FML

[FS10] Susana Faria and Gilda Soromenho. Fitting mixtures of linear regressions. Journal of Statistical Computation and Simulation, 80(2):201–225, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

From:2013:RHI

[FS13] Steven G. From and Andrew W. Swift. A refinement of Hoeffding’s inequality. Journal of Statistical Computation and Simulation, 83(5):977–983, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fort:2015:BMS

[FS15a] Jean-Claude Fort and Luis A. Salomón. Brownian motion simulation: a quantization approach. Journal of Statistical Computation and Simulation, 85(9):1807–1817, 2015. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Foster:2015:RIS

[FS15b] Dean P. Foster and Robert A. Stine. Risk inflation of sequential tests controlled by alpha investing. Journal of Statistical Computation and Simulation, 85(18):3613–3627, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fu:2010:BEM

[FTS10] Zhi-Hui Fu, Jian Tao, and Ning-Zhong Shi. Bayesian estimation of the multidimensional graded response model with nonignorable missing data. Journal of Statistical Computation and Simulation, 80(11):1237–1252, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fu:2016:HBL

[Fu16] Shuai Fu. Hierarchical Bayesian LASSO for a negative binomial regression. Journal of Statistical Computation and Simulation, 86(11):2182–2203, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fukuda:2011:APC

[Fuk11] Kosei Fukuda. Age-period-cohort decompositions using principal components and partial least squares. Journal of Statistical Computation and Simulation, 81(12):1871–1878, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fukuda:2016:PCB

[Fuk16] Kosei Fukuda. Principal-component-based generalizedleast-squares approach for panel data. Journal of Statistical Computation and Simulation, 86(5):874–890, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1041954.

Fonseca:2013:CRS

[FVB13] Renata Santana Fonseca, Dione Maria Valença, and Heleno Bolfarine. Cure rate survival models with missing covariates: a simulation study. Journal of Statistical Computation and Simulation, 83(1):97–113, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fan:2013:EBF

[FW13] Tsai-Hung Fan and Yi-Fu Wang. An empirical Bayesian forecast in the threshold stochastic volatility models. Journal of Statistical Computation and Simulation, 83(3):486– 500, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fan:2015:BMA

[FW15] Tsai-Hung Fan and Guo-Tzau Wang. Bayesian model averaging in longitudinal regression models with AR(1) errors with application to a myopia data set. Journal of Statistical Computation and Simulation, 85(8):1667–1678, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fang:2016:TPS

[FWF16] Yixin Fang, Binhuan Wang, and Yang Feng. Tuningparameter selection in regularized estimations of large covariance matrices. Journal of Statistical Computation and Simulation, 86(3):494–509, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fang:2015:BLV

[FWZ^{+}15] Kuangnan Fang, Xiaoyan Wang, Shengwei Zhang, Jianping Zhu, and Shuangge Ma. Bi-level variable selection via adaptive sparse group lasso. Journal of Statistical Computation and Simulation, 85(13):2750–2760, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Frey:2018:AAR

[FZ18] Jesse Frey and Yimin Zhang. An algorithm with applications in ranked-set sampling. Journal of Statistical Computation and Simulation, 88(3):471–481, 2018. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gunasekera:2015:GVM

[GA15] Sumith Gunasekera and Malwane M. A. Ananda. Generalized variable method inference for the location parameter of the general half-normal distribution. Journal of Statistical Computation and Simulation, 85(10):2115–2132, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghitany:2014:LEP

[GAB14] M. E. Ghitany, F. Alqallaf, and N. Balakrishnan. On the likelihood estimation of the parameters of Gompertz distribution based on complete and progressively Type-II censored samples. Journal of Statistical Computation and Simulation, 84(8):1803–1812, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gauri:2010:TSP

[Gau10] Susanta Kumar Gauri. A test for suitability of the preliminary samples for constructing control limits of X^{−} chart. Journal of Statistical Computation and Simulation, 80(6): 689–705, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gauri:2011:GAC

[Gau11] Susanta Kumar Gauri. Globally applicable control chart for online monitoring of stability of process mean. Journal of Statistical Computation and Simulation, 81(12):1847–1869, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghosh:2015:SIN

[GB15] Indranil Ghosh and N. Balakrishnan. Study of incompatibility or near compatibility of bivariate discrete conditional probability distributions through divergence measures. Journal of Statistical Computation and Simulation, 85(1):117–130, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghosh:2017:HTB

[GB17a] Indranil Ghosh and N. Balakrishnan. On the hidden truncated bivariate Pareto (IV) model and associated inferential issues. Journal of Statistical Computation and Simulation, 87(7):1467–1487, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:2017:FMO

[GB17b] Neha Gupta and Abdul Bari. Fuzzy multi-objective optimization for optimum allocation in multivariate stratified sampling with quadratic cost and parabolic fuzzy numbers. Journal of Statistical Computation and Simulation, 87 (12):2372–2383, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghosh:2018:ISS

[GB18] Abhik Ghosh and Ayanendranath Basu. Improvements in the small sample efficiency of the minimum S-divergence estimators under discrete models. Journal of Statistical Computation and Simulation, 88(3):511–532, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Giancristofaro:2016:DTF

[GBCS16] Rosa Arboretti Giancristofaro, Stefano Bonnini, Livio Corain, and Luigi Salmaso. Dependency and truncated forms of combinations in multivariate combination-based permutation tests and ordered categorical variables. Journal of Statistical Computation and Simulation, 86(18):3608–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gallardo:2016:EAE

[GBdL16] Diego I. Gallardo, Heleno Bolfarine, and Antonio Carlos Pedroso de Lima. An EM algorithm for estimating the destructive weighted Poisson cure rate model. Journal of Statistical Computation and Simulation, 86(8):1497–1515, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (elec-

tronic), 1563-5163. URL http://www.tandfonline.com/
doi/abs/10.1080/00949655.2015.1071375.

Gerlach:2017:SPE

[GC17a] Richard Gerlach and Cathy W. S. Chen. Semi-parametric expected shortfall forecasting in financial markets. Journal of Statistical Computation and Simulation, 87(6):1084– 1106, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gottard:2017:EMM

[GC17b] Anna Gottard and Giorgio Calzolari. Estimating multiplemembership logit models with mixed effects: indirect inference versus data cloning. Journal of Statistical Computation and Simulation, 87(12):2334–2348, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gocoglu:2019:EPP

[GD19] Aylin Göçoglu and Neslihan Demirel. Estimating the population proportion in modified ranked set sampling methods. Journal of Statistical Computation and Simulation, 89 (14):2694–2710, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Goncalves:2018:BIR

[GdCCDS18] Flávio Bambirra Gonçalves, Bárbara da Costa Campos Dias, and Tufi Machado Soares. Bayesian item response model: a generalized approach for the abilities’ distribution using mixtures. Journal of Statistical Computation and Simulation, 88(5):967–981, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gomez-Deniz:2011:DLD

[GDCO11] Emilio Gómez-Déniz and Enrique Calderín-Ojeda. The discrete Lindley distribution: properties and applications. Journal of Statistical Computation and Simulation, 81(11): 1405–1416, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gomez-Deniz:2018:PAP

[GDCO18] Emilio Gómez-Déniz and Enrique Calderín-Ojeda. Properties and applications of the Poisson-reciprocal inverse Gaussian distribution. Journal of Statistical Computation and Simulation, 88(2):269–289, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gorfine:2012:CME

[GDPH12] Malka Gorfine, Rottem De-Picciotto, and Li Hsu. Conditional and marginal estimates in case-control family data — extensions and sensitivity analyses. Journal of Statistical Computation and Simulation, 82(10):1449–1470, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gazal:2012:AVE

[GDR12] Steven Gazal, Jean-Jacques Daudin, and Stéphane Robin. Accuracy of variational estimates for random graph mixture models. Journal of Statistical Computation and Simulation, 82(6):849–862, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gomes:2014:NLM

[GdSCO14] Antonio E. Gomes, Cibele Q. da Silva, Gauss M. Cordeiro, and Edwin M. M. Ortega. A new lifetime model: the Kumaraswamy generalized Rayleigh distribution. Journal of Statistical Computation and Simulation, 84(2):290–309, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gonzalez-Estrada:2019:SWT

[GEC19] Elizabeth González-Estrada and Waldenia Cosmes. Shapiro– Wilk test for skew normal distributions based on data transformations. Journal of Statistical Computation and Simulation, 89(17):3258–3272, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Genc:2013:ETL

[Gen13] Ali I. Genç. Estimation of P (X > Y) with Topp–Leone distribution. Journal of Statistical Computation and Simulation, 83(2):326–339, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gonzalez-Estrada:2018:RPT

[GEV18] E. González-Estrada and J. A. Villaseñor. An R package for testing goodness of fit: goft. Journal of Statistical Computation and Simulation, 88(4):726–751, 2018. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gallagher:2015:CET

[GFS15] Colin M. Gallagher, Thomas J. Fisher, and Jie Shen. A Cauchy estimator test for autocorrelation. Journal of Statistical Computation and Simulation, 85(6):1264–1276, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gencay:2011:EVE

[GG11] Ramazan Gençay and Nikola Gradojevic. Errors-in-variables estimation with wavelets. Journal of Statistical Computation and Simulation, 81(11):1545–1564, 2011. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gunduz:2016:EFR

[GG16] Fatma F. Gündüz and Ali I. Genç. The exponentiated Fréchet regression: an alternative model for actuarial modelling purposes. Journal of Statistical Computation and Simulation, 86(17):3456–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:2017:MIG

[GG17] Vinay K. Gupta and Gurprit Grover. Multiple imputation for gamma outcome variable using generalized linear model. Journal of Statistical Computation and Simulation, 87(10):1980–1988, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gomez:2018:PPE

[GGA18] Yolanda M. Gómez, Diego I. Gallardo, and Barry C. Arnold. The power piecewise exponential model. Journal of Statistical Computation and Simulation, 88(5):825–840, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:2010:ERM

[GGAM10] Ramesh C. Gupta, M. E. Ghitany, and D. K. Al-Mutairi. Estimation of reliability from Marshall–Olkin extended Lomax distributions. Journal of Statistical Computation and Simulation, 80(8):937–947, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:2013:ERB

[GGAM13] Ramesh C. Gupta, M. E. Ghitany, and D. K. Al-Mutairi. Estimation of reliability from a bivariate log-normal data. Journal of Statistical Computation and Simulation, 83(6): 1068–1081, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gallardo:2017:NEA

[GGdC17] Diego I. Gallardo, Yolanda M. Gómez, and Mário de Castro. A note on the EM algorithm for estimation in the destructive negative binomial cure rate model. Journal of Statistical Computation and Simulation, 87(12):2291–2297, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gallardo:2018:FCR

[GGdC18] Diego I. Gallardo, Yolanda M. Gómez, and Mário de Castro. A flexible cure rate model based on the polylogarithm distribution. Journal of Statistical Computation and Simulation, 88(11):2137–2149, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Garcia:2019:CRF

[GGG19] Claudia García, Román Salmerón Gómez, and Catalina B. García. Choice of the ridge factor from the correlation matrix determinant. Journal of Statistical Computation and Simulation, 89(2):211–231, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gonzalez:2018:NPB

[GGM18] Miguel González, Cristina Gutiérrez, and Rodrigo Martínez. Non-parametric Bayesian inference through MCMC method for Y -linked two-sex branching processes with blind choice. Journal of Statistical Computation and Simulation, 88(18): 3565–3587, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guven:2019:FBA

[GGSS19] Gamze Güven, Özge Gürer, Hatice Samkar, and Birdal Senoglu. A fiducial-based approach to the one-way ANOVA in the presence of nonnormality and heterogeneous error variances. Journal of Statistical Computation and Simulation, 89(9):1715–1729, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghalibaf:2016:DSE

[Gha16] Mohammad Bolbolian Ghalibaf. Dependence structure of extreme order statistics. Journal of Statistical Computation and Simulation, 86(14):2846–2855, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghalibaf:2019:TLT

[Gha19] M. Bolbolian Ghalibaf. Testing for lower tail dependence in extreme value models. Journal of Statistical Computation and Simulation, 89(13):2583–2593, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gorecki:2017:SFP

[GHH17] Jan Górecki, Marius Hofert, and Martin Holena. On structure, family and parameter estimation of hierarchical Archimedean copulas. Journal of Statistical Computation and Simulation, 87(17):3261–3324, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Garcia-Hiernaux:2010:URC

[GHJC10] Alfredo García-Hiernaux, Miguel Jerez, and José Casals. Unit roots and cointegration modelling through a family of flexible information criteria. Journal of Statistical Computation and Simulation, 80(2):173–189, 2010. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gu:2019:BPB

[GHMR19] Xin Gu, Herbert Hoijtink, Joris Mulder, and Yves Rosseel. Bain: a program for Bayesian testing of order constrained hypotheses in structural equation models. Journal of Statistical Computation and Simulation, 89(8):1526–1553, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gomes:2013:APM

[GHRAM13] M. Ivette Gomes, Lígia Henriques-Rodrigues, M. Isabel Fraga Alves, and B. G. Manjunath. Adaptive PORT– MVRB estimation: an empirical comparison of two heuristic algorithms. Journal of Statistical Computation and Simulation, 83(6):1129–1144, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Grilli:2017:FLM

[GI17] Leonardo Grilli and Francesco Innocenti. Fitting logistic multilevel models with crossed random effects via Bayesian Integrated Nested Laplace Approximations: a simulation study. Journal of Statistical Computation and Simulation, 87(14):2689–2707, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guure:2015:BSI

[GIDB15] Chris Bambey Guure, Noor Akma Ibrahim, Duah Dwomoh, and Samuel Bosomprah. Bayesian statistical inference of the loglogistic model with interval-censored lifetime data. Journal of Statistical Computation and Simulation, 85(8):1567– 1583, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ginsbourger:2015:E

[GIP15] David Ginsbourger, Bertrand Iooss, and Luc Pronzato. Editorial. Journal of Statistical Computation and Simulation, 85(7):1281–1282, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ganguly:2016:ASS

[GK16] Ayon Ganguly and Debasis Kundu. Analysis of simple stepstress model in presence of competing risks. Journal of Statistical Computation and Simulation, 86(10):1989–2006, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gani:2016:KDB

[GL16] Walid Gani and Mohamed Limam. A kernel distance-based representative subset selection method. Journal of Statistical Computation and Simulation, 86(1):135–148, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Granzotto:2017:CRT

[GLB17] D. C. T. Granzotto, F. Louzada, and N. Balakrishnan. Cubic rank transmuted distributions: inferential issues and applications. Journal of Statistical Computation and Simulation, 87(14):2760–2778, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Garay:2014:SDN

[GLLO14] Aldo M. Garay, Víctor H. Lachos, Filidor V. Labra, and Edwin M. M. Ortega. Statistical diagnostics for nonlinear regression models based on scale mixtures of skew-normal distributions. Journal of Statistical Computation and Simulation, 84(8):1761–1778, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Golalizadeh:2016:RAE

[GM16] Mousa Golalizadeh and Hashem Mahmoudnejad. A recursive algorithm on estimating the parameters in multilevel models subject to the measurement errors on the covariates. Journal of Statistical Computation and Simulation, 86(2):252–261, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Goyal:2013:MTD

[GMG13] Anju Goyal, Vishal Maurya, and Amar Nath Gill. Mutiple three-decision procedure for comparing several exponential treatments with the best control. Journal of Statistical Computation and Simulation, 83(3):472–485, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Grilli:2015:BEI

[GMR15] L. Grilli, S. Metelli, and C. Rampichini. Bayesian estimation with integrated nested Laplace approximation for binary logit mixed models. Journal of Statistical Computation and Simulation, 85(13):2718–2726, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Goncalves:2018:END

[GOC18] Cátia R. Gonçalves, Cira E. G. Otiniano, and Evelyn C. Cruvinel. Estimation of a nonlinear discriminant function from a mixture of two GEV distributions. Journal of Statistical Computation and Simulation, 88(6):1147–1171, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Goulionis:2011:SSV

[Gou11] J. E. Goulionis. Strategies for sampling with varying probabilities. Journal of Statistical Computation and Simulation, 81(11):1753, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Govindaraju:2017:PDS

[Gov17] K. Govindaraju. Predictive design of single sampling attribute plans. Journal of Statistical Computation and Simulation, 87(3):447–456, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gatto:2015:SAS

[GP15a] Riccardo Gatto and Chantal Peeters. Saddlepoint approximations to sensitivities of tail probabilities of random sums and comparisons with Monte Carlo estimators. Journal of Statistical Computation and Simulation, 85(4):641–659, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Golparvar:2015:BPT

[GP15b] L. Golparvar and A. Parsian. On Bayes predictor of times to failure of Type-II progressively censored sample. Journal of Statistical Computation and Simulation, 85(17):3420–3436, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Geete:2019:NBS

[GP19] Kanu Geete and Manish Pandey. A noise-based stabilizer for convolutional neural networks. Journal of Statistical Computation and Simulation, 89(11):2102–2120, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gomes:2010:TIS

[GRPP10] M. Ivette Gomes, Lígia Henriques Rodrigues, Hugo Pereira, and Dinis Pestana. Tail index and second-order parameters’ semi-parametric estimation based on the log-excesses. Journal of Statistical Computation and Simulation, 80(6): 653–666, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Galati:2014:RNB

[GSL^{+}14] J. C. Galati, K. A. Seaton, K. J. Lee, J. A. Simpson, and J. B. Carlin. Rounding non-binary categorical variables following multivariate normal imputation: evaluation of simple methods and implications for practice. Journal of Statistical Computation and Simulation, 84(4):798–811, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghitany:2017:NMM

[GSW17] M. E. Ghitany, Pingfan Song, and Shaochen Wang. New modified moment estimators for the two-parameter weighted Lindley distribution. Journal of Statistical Computation and Simulation, 87(16):3225–3240, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ghitany:2014:LEG

[GTB14] M. E. Ghitany, V. K. Tuan, and N. Balakrishnan. Likelihood estimation for a general class of inverse exponentiated distributions based on complete and progressively censored data. Journal of Statistical Computation and Simulation, 84(1):96–106, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gulkac:2010:HPM

[Gül10] Vildan Gülkaç. The homotopy perturbation method for the Black–Scholes equation. Journal of Statistical Computation and Simulation, 80(12):1349–1354, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gunasekera:2018:CBG

[Gun18] Sumith Gunasekera. Classical, Bayesian, and generalized inferences of the reliability of a multicomponent system with censored data. Journal of Statistical Computation and Simulation, 88(18):3455–3501, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guo:2017:OPU

[Guo17] Zhidong Guo. Option pricing under the Merton model of the short rate in subdiffusive Brownian motion regime. Journal of Statistical Computation and Simulation, 87(3):519–529, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gustafson:2015:NFR

[Gus15] Karl Gustafson. A new financial risk ratio. Journal of Statistical Computation and Simulation, 85(13):2682–2692, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:2012:SIE

[GV12] Ramesh C. Gupta and Weston Viles. Statistical inference for the extended generalized inverse Gaussian model. Journal of Statistical Computation and Simulation, 82(12):1855–1872, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guevara:2016:EPC

[GV16] Rubén D. Guevara and José A. Vargas. Evaluation of process capability in multivariate nonlinear profiles. Journal of Statistical Computation and Simulation, 86(12):2411–2428, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gurevich:2018:DBE

[GV18] Gregory Gurevich and Albert Vexler. A density based empirical likelihood approach for testing bivariate normality. Journal of Statistical Computation and Simulation, 88 (13):2540–2560, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Garza-Venegas:2018:EAE

[GVTGH^{+}18] J. A. Garza-Venegas, V. G. Tercero-Gómez, L. Lee Ho, P. Castagliola, and G. Celano. Effect of autocorrelation estimators on the performance of the X̄ control chart. Journal of Statistical Computation and Simulation, 88(13):2612– 2630, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Giordan:2017:MLB

[GVW17] M. Giordan, F. Vaggi, and R. Wehrens. On the maximization of likelihoods belonging to the exponential family using a Levenberg–Marquardt approach. Journal of Statistical Computation and Simulation, 87(5):895–907, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gupta:2014:ASD

[GWH14] Ramesh C. Gupta, Qiuyi Wu, and Jie Huang. Analysis of survival data by an exponential-generalized Poisson distribution. Journal of Statistical Computation and Simulation, 84(11):2495–2505, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Guo:2014:AUC

[GWX14] Baocai Guo, Bing Xing Wang, and Min Xie. ARLunbiased control charts for the monitoring of exponentially distributed characteristics based on type-II censored samples. Journal of Statistical Computation and Simulation, 84 (12):2734–2747, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gencturk:2016:MCN

[GY16] Yasemin Gençtürk and Ayten Yigiter. Modelling claim number using a new mixture model: negative binomial gamma distribution. Journal of Statistical Computation and Simulation, 86(10):1829–1839, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

George:2013:URM

[GYV^{+}13] Melissa R. W. George, Na Yang, M. Lee Van Horn, Jessalyn Smith, Thomas Jaki, Daniel J. Feaster, Katherine Masyn, and George Howe. Using regression mixture models with non-normal data: examining an ordered polytomous approach. Journal of Statistical Computation and Simulation, 83(4):759–772, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gan:2013:MME

[GZB13] Jianjun Gan, Hongmei Zhang, and Robert Best. Mixture of measurement errors and their impact on parameter inferences. Journal of Statistical Computation and Simulation, 83(4):613–626, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gao:2018:RIS

[GZL18] Quansheng Gao, Kang Zhou, and Junyong Li. Robust importance sampling for some typical types of utility-based shortfall risk measures using exponential twisting and kernel density techniques. Journal of Statistical Computation and Simulation, 88(2):359–375, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gao:2014:RAS

[GZT14] Xing-Long Gao, Qing-Bin Zhang, and Qian-Gang Tang. Reliability assessment of slot-parachute inflation based on Bayes theory. Journal of Statistical Computation and Simulation, 84(6):1159–1172, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hodoshima:2010:BSR

[HA10] Jiro Hodoshima and Masakazu Ando. Bootstrapping stochastic regression models under homoskedasticity: wild bootstrap vs. pairs bootstrap. Journal of Statistical Computation and Simulation, 80(11):1225–1235, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Holm:2013:EPS

[HA13] Anders Holm and Jacob Nielsen Arendt. Evaluating the performance of simple estimators for probit models with two dummy endogenous regressors. Journal of Statistical Computation and Simulation, 83(6):1156–1178, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hussein:2012:STP

[HAB12] Abdulkadir Hussein, S. Ejaz Ahmed, and S. Bhatti. Sequential testing of process capability indices. Journal of Statistical Computation and Simulation, 82(2):279–292, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hong:2016:ACM

[HAC16] Hanwoom Hong, Sung K. Ahn, and Sinsup Cho. Analysis of cointegrated models with measurement errors. Journal of Statistical Computation and Simulation, 86(3):623–639, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hossain:2014:MSP

[HAH14] Shakhawat Hossain, S. Ejaz Ahmed, and Hatem A. Howlader. Model selection and parameter estimation of a multinomial logistic regression model. Journal of Statistical Computation and Simulation, 84(7):1412–1426, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hajrajabi:2019:MSM

[Haj19] Arezo Hajrajabi. Markov switching model of nonlinear autoregressive with skew-symmetric innovations. Journal of Statistical Computation and Simulation, 89(4):559–575, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Han:2010:APP

[Han10] Chien-Pai Han. Adaptive plotting position and test of normality. Journal of Statistical Computation and Simulation, 80(4):413–420, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Han:2017:BHB

[Han17] Ming Han. The E-Bayesian and hierarchical Bayesian estimations of Pareto distribution parameter under different loss functions. Journal of Statistical Computation and Simulation, 87(3):577–593, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2014:IMD

[Haq14] Abdul Haq. An improved mean deviation exponentially weighted moving average control chart to monitor process dispersion under ranked set sampling. Journal of Statistical Computation and Simulation, 84(9):2011–2024, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2017:EDF

[Haq17] Abdul Haq. Estimation of the distribution function under hybrid ranked set sampling. Journal of Statistical Computation and Simulation, 87(2):313–327, 2017. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2019:OPO

[Haq19a] Abdul Haq. Ordered partially ordered judgment subset sampling with applications to parametric inference. Journal of Statistical Computation and Simulation, 89(18):3354– 3376, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2019:WAM

[Haq19b] Abdul Haq. Weighted adaptive multivariate CUSUM charts with variable sampling intervals. Journal of Statistical Computation and Simulation, 89(3):478–491, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hattab:2016:DPI

[Hat16] Mohammad Wasef Hattab. A derivation of prediction intervals for gamma regression. Journal of Statistical Computation and Simulation, 86(17):3512–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hayter:2015:CBC

[Hay15] A. J. Hayter. Confidence bounds on the coefficient of variation of a normal distribution with applications to winprobabilities. Journal of Statistical Computation and Simulation, 85(18):3778–3791, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hung:2011:EAG

[HBC11] Y. C. Hung, N. Balakrishnan, and C. W. Cheng. Evaluation of algorithms for generating Dirichlet random vectors. Journal of Statistical Computation and Simulation, 81(4): 445–459, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hernandez-Bastida:2011:CRM

[HBFSGD11] A. Hernández-Bastida, M. P. Fernández-Sánchez, and E. Gómez-Déniz. Collective risk model: Poisson–Lindley and exponential distributions for Bayes premium and operational risk. Journal of Statistical Computation and Simulation, 81(6):759–778, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hajjem:2014:MER

[HBL14] Ahlem Hajjem, François Bellavance, and Denis Larocque. Mixed-effects random forest for clustered data. Journal of Statistical Computation and Simulation, 84(6):1313–1328, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2016:HRS

[HBM16] Abdul Haq, Jennifer Brown, and Elena Moltchanova. Hybrid ranked set sampling scheme. Journal of Statistical Computation and Simulation, 86(1):1–28, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2015:EME

[HBMAO15] Abdul Haq, Jennifer Brown, Elena Moltchanova, and Amer Ibrahim Al-Omari. Effect of measurement error on exponentially weighted moving average control charts under ranked set sampling schemes. Journal of Statistical Computation and Simulation, 85(6):1224–1246, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hernandez:2012:NBP

[HBT12] Noslen Hernández, Rolando J. Biscay, and Isneri Talavera. A non-Bayesian predictive approach for statistical calibration. Journal of Statistical Computation and Simulation, 82(4):529–545, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huo:2010:CPL

[HC10] Xiaoming Huo and Jie Chen. Complexity of penalized likelihood estimation. Journal of Statistical Computation and Simulation, 80(7):747–759, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2015:BCQ

[HC15] Hanwen Huang and Zhongxue Chen. Bayesian composite quantile regression. Journal of Statistical Computation and Simulation, 85(18):3744–3754, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hermanns:2017:LIC

[HC17] M. Hermanns and E. Cramer. Likelihood inference for the component lifetime distribution based on progressively censored parallel systems data. Journal of Statistical Computation and Simulation, 87(3):607–630, 2017. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hu:2019:CDC

[HCT19] X. L. Hu, P. Castagliola, and A. A. Tang. Conditional design of the CUSUM median chart for the process position when process parameters are unknown. Journal of Statistical Computation and Simulation, 89(13):2468–2488, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hsu:2016:CCI

[HCY16] Chang-Hsien Hsu, Kuen-Suan Chen, and Chun-Ming Yang. Construction of closed interval for process capability indices C_{pu}, C_{pl}, and S_{pk} based on Boole’s inequality and de Morgan’s laws. Journal of Statistical Computation and Simulation, 86(18):3701–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hassan:2013:MPM

[HEB13] M. Y. Hassan and M. Y. El-Bassiouni. Modelling Poisson marked point processes using bivariate mixture transition distributions. Journal of Statistical Computation and Simulation, 83(8):1440–1452, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Herwartz:2011:KIS

[Her11] Helmut Herwartz. Keep it simple: on specific-to-general predictor selection for time series forecasting in the short, medium and long run. Journal of Statistical Computation and Simulation, 81(8):931–943, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hsieh:2015:QRB

[HH15a] Jin-Jian Hsieh and Ming-Fu Hsiao. Quantile regression based on a weighted approach under semi-competing risks data. Journal of Statistical Computation and Simulation, 85(14):2793–2807, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hwang:2015:SRA

[HH15b] Jing-Shiang Hwang and Tsuey-Hwa Hu. A stepwise regression algorithm for high-dimensional variable selection. Journal of Statistical Computation and Simulation, 85(9):1793– 1806, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hepler:2017:MCA

[HH17] Staci A. Hepler and Radu Herbei. A Monte Carlo approach to quantifying discrepancies between intractable posterior distributions. Journal of Statistical Computation and Simulation, 87(8):1666–1683, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hwang:2015:AEC

[HHC15] Wen-Han Hwang, Yih-Huei Huang, and Fei-Yin Chen. Approximate estimations in capture-recapture model with heteroscedastic measurement errors. Journal of Statistical Computation and Simulation, 85(4):762–776, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hirose:2011:OTD

[Hir11] Hideo Hirose. An optimal test design to evaluate the ability of an examinee by using the stress-strength model. Journal of Statistical Computation and Simulation, 81(1):79–87, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2016:NSC

[HK16a] Abdul Haq and Michael B. C. Khoo. A new synthetic control chart for monitoring process mean using auxiliary information. Journal of Statistical Computation and Simulation, 86(15):3068–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hassanzadeh:2016:ADC

[HK16b] Fatemeh Hassanzadeh and Iraj Kazemi. Analysis of overdispersed count data with extra zeros using the Poisson log-skew-normal distribution. Journal of Statistical Computation and Simulation, 86(13):2644–2662, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2018:NDS

[HK18a] Abdul Haq and Michael B. C. Khoo. A new double sampling control chart for monitoring process mean using auxiliary information. Journal of Statistical Computation and Simulation, 88(5):869–899, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hasan:2018:CIM

[HK18b] Md Sazib Hasan and K. Krishnamoorthy. Confidence intervals for the mean and a percentile based on zero-inflated lognormal data. Journal of Statistical Computation and Simulation, 88(8):1499–1514, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hirose:2016:FIM

[HKK^{+}16] Kei Hirose, Sunyong Kim, Yutaka Kano, Miyuki Imada, Manabu Yoshida, and Masato Matsuo. Full information maximum likelihood estimation in factor analysis with a large number of missing values. Journal of Statistical Computation and Simulation, 86(1):91–104, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huo:2017:RBT

[HKKL17] Lijuan Huo, Tae-Hwan Kim, Yunmi Kim, and Dong Jin Lee. A residual-based test for autocorrelation in quantile regression models. Journal of Statistical Computation and Simulation, 87(7):1305–1322, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huh:2017:MPS

[HKL17] Jaewon Huh, Hanwool Kim, and Sangyeol Lee. Monitoring parameter shift with Poisson integer-valued GARCH models. Journal of Statistical Computation and Simulation, 87 (9):1754–1766, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hughes:2017:ELM

[HKST17] Rachael A. Hughes, Michael G. Kenward, Jonathan A. C. Sterne, and Kate Tilling. Estimation of the linear mixed integrated Ornstein–Uhlenbeck model. Journal of Statistical Computation and Simulation, 87(8):1541–1558, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hayter:2013:ETN

[HL13] A. J. Hayter and Y. Lin. The evaluation of trivariate normal probabilities defined by linear inequalities. Journal of Statistical Computation and Simulation, 83(4):668–676, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2015:CTL

[HL15] Chi-Jui Huang and Shin-Li Lu. Considering Taguchi loss function on statistically constrained economic sum of squares exponentially weighted moving average charts. Journal of Statistical Computation and Simulation, 85(3): 572–586, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huwang:2018:SRB

[HLL18] Longcheen Huwang, Jian-Chi Lin, and Li-Wei Lin. A spatial rank-based EWMA chart for monitoring linear profiles. Journal of Statistical Computation and Simulation, 88(18):3620–3649, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2015:BEV

[HLN^{+}15] Zhipeng Huang, Jialiang Li, David Nott, Lei Feng, Tze-Pin Ng, and Tien-Yin Wong. Bayesian estimation of varyingcoefficient models with missing data, with application to the Singapore Longitudinal Aging Study. Journal of Statistical Computation and Simulation, 85(12):2364–2377, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Holbrook:2018:GLM

[HLVRS18] Andrew Holbrook, Shiwei Lan, Alexander Vandenberg-Rodes, and Babak Shahbaba. Geodesic Lagrangian Monte Carlo over the space of positive definite matrices: with application to Bayesian spectral density estimation. Journal of Statistical Computation and Simulation, 88(5):982–1002, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hobza:2013:SAE

[HM13] Tomás Hobza and Domingo Morales. Small area estimation under random regression coefficient models. Journal of Statistical Computation and Simulation, 83(11):2160–2177, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huda:2017:AAC

[HM17] S. Huda and Rahul Mukerjee. Algorithmic and analytical construction of efficient designs in small blocks for comparing consecutive pairs of treatments. Journal of Statistical Computation and Simulation, 87(16):3195–3207, 2017. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haq:2018:ICC

[HM18a] Abdul Haq and Waqas Munir. Improved CUSUM charts for monitoring process mean. Journal of Statistical Computation and Simulation, 88(9):1684–1701, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Homrighausen:2018:STP

[HM18b] Darren Homrighausen and Daniel J. McDonald. A study on tuning parameter selection for the high-dimensional lasso. Journal of Statistical Computation and Simulation, 88(15):2865–2892, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huttner:2019:SRC

[HM19] Amelie Hüttner and Jan-Frederik Mai. Simulating realistic correlation matrices for financial applications: correlation matrices with the Perron–Frobenius property. Journal of Statistical Computation and Simulation, 89(2):315–336, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2013:RBO

[HMM13] Huaming Huang, Kishan Mehrotra, and Chilukuri K. Mohan. Rank-based outlier detection. Journal of Statistical Computation and Simulation, 83(3):518–531, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hobza:2017:WTT

[HMP17] Tomás Hobza, Nirian Martín, and Leandro Pardo. A Waldtype test statistic based on robust modified median estimator in logistic regression models. Journal of Statistical Computation and Simulation, 87(12):2309–2333, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

He:2011:CED

[HN11] Qinying He and H. N. Nagaraja. Correlation estimation in Downton’s bivariate exponential distribution using incomplete samples. Journal of Statistical Computation and Simulation, 81(5):531–546, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hosseini-Nasab:2013:CVA

[HN13] Mohammad Hosseini-Nasab. Cross-validation approximation in functional linear regression. Journal of Statistical Computation and Simulation, 83(8):1429–1439, 2013. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huong:2018:MTS

[HNPB18] Pham Thi Thu Huong, Darfiana Nur, Hoa Pham, and Alan Branford. A modified two-stage approach for joint modelling of longitudinal and time-to-event data. Journal of Statistical Computation and Simulation, 88(17):3379–3398, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ho:2012:UJP

[Ho12] Kwok-Wah Ho. The use of Jeffreys priors for the studentt distribution. Journal of Statistical Computation and Simulation, 82(7):1015–1021, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ho:2016:STC

[Ho16] Kwok-Wah Ho. Stress testing correlation matrix: a maximum empirical likelihood approach. Journal of Statistical Computation and Simulation, 86(14):2707–2713, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hashimoto:2019:ZSR

[HOC^{+}19] Elizabeth M. Hashimoto, Edwin M. M. Ortega, Gauss M. Cordeiro, Vicente G. Cancho, and Carine Klauberg. Zerospiked regression models generated by gamma random variables with application in the resin oil production. Journal of Statistical Computation and Simulation, 89(1):52–70, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hofert:2012:SRS

[Hof12] Marius Hofert. A stochastic representation and sampling algorithm for nested Archimedean copulas. Journal of Statistical Computation and Simulation, 82(9):1239–1255, 2012.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Holden:2019:MMA

[Hol19] Lars Holden. Mixing of MCMC algorithms. Journal of Statistical Computation and Simulation, 89(12):2261–2279, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hossain:2017:NCC

[HOR17] M. Pear Hossain, M. Hafidz Omar, and Muhammad Riaz. New V control chart for the Maxwell distribution. Journal of Statistical Computation and Simulation, 87(3):594–606, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Howe:2013:SID

[How13] J. Andrew Howe. Simulating from irregular data: Kernel Carlo simulation. Journal of Statistical Computation and Simulation, 83(3):446–457, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2011:AUC

[HP11] Xiaohui Huang and Francis Pascual. ARL-unbiased control charts with alarm and warning lines for monitoring Weibull percentiles using the first-order statistic. Journal of Statistical Computation and Simulation, 81(11):1677–1696, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hanagal:2015:GFM

[HP15] David D. Hanagal and Arvind Pandey. Gamma frailty models for bivariate survival data. Journal of Statistical Computation and Simulation, 85(15):3172–3189, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Horig:2012:MVH

[HR12] M. Hörig and C. Redenbach. The maximum volume hard subset model for Poisson processes: simulation aspects. Journal of Statistical Computation and Simulation, 82(1): 107–121, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Harvey:2017:CSE

[HRR^{+}17] Christine Harvey, Scott Rosen, James Ramsey, Christopher Saunders, and Samar K. Guharay. Computationally and statistically efficient model fitting techniques. Journal of Statistical Computation and Simulation, 87(1):123–137, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huggins:2013:SIO

[HS13] Richard Huggins and Jakub Stoklosa. Semiparametric inference for open populations using the Jolly–Seber model: a penalized spline approach. Journal of Statistical Computation and Simulation, 83(9):1741–1755, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hsu:2011:ELP

[HSC11] Bi-Min Hsu, Ming-Hung Shu, and Bo-Sen Chen. Evaluating lifetime performance for the Pareto model with censored and imprecise information. Journal of Statistical Computation and Simulation, 81(12):1817–1833, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haynes:2016:SDB

[HSC16] Mary E. Haynes, Roy T. Sabo, and N. Rao Chaganty. Simulating dependent binary variables through multinomial sampling. Journal of Statistical Computation and Simulation, 86(3):510–523, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2018:GAE

[HSJ18] Wenpo Huang, Lianjie Shu, and Wei Jiang. A gradient approach to efficient design and analysis of multivariate EWMA control charts. Journal of Statistical Computation and Simulation, 88(14):2707–2725, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Helu:2015:ELP

[HSR15] Amal Helu, Hani Samawi, and Mohammad Z. Raqab. Estimation on Lomax progressive censoring using the EM algorithm. Journal of Statistical Computation and Simulation, 85(5):1035–1052, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hanusz:2012:NTM

[HT12] Zofia Hanusz and Joanna Tarasińska. New tests for multivariate normality based on Small’s and Srivastava’s graphical methods. Journal of Statistical Computation and Simulation, 82(12):1743–1752, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2016:CDC

[HTZ^{+}16] Lu Huang, Li Tang, Bo Zhang, Zhiwei Zhang, and Hui Zhang. Comparison of different computational implementations on fitting generalized linear mixed-effects models for repeated count measures. Journal of Statistical Computation and Simulation, 86(12):2392–2404, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haller:2014:FSC

[HU14] Bernhard Haller and Kurt Ulm. Flexible simulation of competing risks data following prespecified subdistribution hazards. Journal of Statistical Computation and Simulation, 84 (12):2557–2576, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2011:OEP

[Hua11] Mei Ling Huang. Optimal estimation for the Pareto distribution. Journal of Statistical Computation and Simulation, 81(12):2059–2076, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2016:QRB

[Hua16] Yangxin Huang. Quantile regression-based Bayesian semiparametric mixed-effects models for longitudinal data with non-normal, missing and mismeasured covariate. Journal of Statistical Computation and Simulation, 86(6):1183–1202, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. URL http://www.tandfonline.

com/doi/abs/10.1080/00949655.2015.1057732.

Huang:2012:BEP

[HW12] Syuan-Rong Huang and Shuo-Jye Wu. Bayesian estimation and prediction for Weibull model with progressive censoring. Journal of Statistical Computation and Simulation, 82 (11):1607–1620, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2017:OSS

[HW17] Syuan-Rong Huang and Shuo-Jye Wu. Optimal sample size allocation for accelerated life test under progressive type-II censoring with competing risks. Journal of Statistical Computation and Simulation, 87(1):1–16, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hwang:2011:CEN

[Hwa11] Yi-Ting Hwang. Comparisons of estimators of the number of true null hypotheses and adaptive FDR procedures in multiplicity testing. Journal of Statistical Computation and Simulation, 81(2):207–220, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Haridy:2014:DMS

[HWA^{+}14] Salah Haridy, Zhang Wu, Kazem Abhary, Philippe Castagliola, and Mohammad Shamsuzzaman. Development of a multiattribute synthetic-np chart. Journal of Statistical Computation and Simulation, 84(9):1884–1903, 2014. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2016:EPD

[HWWZ16] Jianhua Z. Huang, Xueying Wang, Ximing Wu, and Lan Zhou. Estimation of a probability density function using interval aggregated data. Journal of Statistical Computation and Simulation, 86(15):3093–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2015:WLQ

[HXT15] Mei Ling Huang, Xiaojian Xu, and Dmitry Tashnev. A weighted linear quantile regression. Journal of Statistical Computation and Simulation, 85(13):2596–2618, 2015. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huang:2014:TPE

[HY14] Jiewu Huang and Hu Yang. A two-parameter estimator in the negative binomial regression model. Journal of Statistical Computation and Simulation, 84(1):124–134, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Huan:2016:NIB

[HY16] Haijing Huan and Fengkai Yang. A non-iterative Bayesian sampling algorithm for censored Student-t linear regression models. Journal of Statistical Computation and Simulation, 86(16):3337–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hlawka:2014:MSC

[HZ14] Marcin Hlawka and Adam Zagdański. Model selection criteria for reduced rank multivariate time series: a simulation study. Journal of Statistical Computation and Simulation, 84(9):1851–1883, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hu:2016:BAM

[HZL16] Yuao Hu, Kaifeng Zhao, and Heng Lian. Bayesian Additive Machine: classification with a semiparametric discriminant function. Journal of Statistical Computation and Simulation, 86(4):682–695, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1028937.

Han:2019:NJC

[HZSA19] Shengtong Han, Hongmei Zhang, Wenhui Sheng, and Hasan Arshad. The nested joint clustering via Dirichlet process mixture model. Journal of Statistical Computation and Simulation, 89(5):815–830, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ismail:2010:OPF

[IA10] Ali A. Ismail and H. M. Aly. Optimal planning of failurestep stress partially accelerated life tests under Type-II censoring. Journal of Statistical Computation and Simulation, 80(12):1335–1348, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. This article has been retracted: see [Ano14c].

Ismail:2011:OCS

[IAGEK11] Ali A. Ismail, Abdalla A. Abdel-Ghaly, and Eman H. El-Khodary. Optimum constant-stress life test plans for Pareto distribution under Type-I censoring. Journal of Statistical Computation and Simulation, 81(12):1835–1845, 2011. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Islam:2013:IVR

[IGR13] Md. Ershadul Islam, Ulrike Grote, and Md. Israt Rayhan. Income vulnerability of rural households in Bangladesh: a comparison between Bayesian and classical methods. Journal of Statistical Computation and Simulation, 83(6):1179– 1187, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Irfan:2018:OEP

[IJL18] Muhammad Irfan, Maria Javed, and Zhengyan Lin. Optimized estimation for population mean using conventional and non-conventional measures under the joint influence of measurement error and non-response. Journal of Statistical Computation and Simulation, 88(12):2385–2403, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Iliopoulos:2015:ECI

[Ili15] George Iliopoulos. On exact confidence intervals in a competing risks model with generalized hybrid type-I censored exponential data. Journal of Statistical Computation and Simulation, 85(14):2953–2961, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Iliopoulos:2016:ECI

[Ili16] George Iliopoulos. Exact confidence intervals for the shape parameter of the gamma distribution. Journal of Statistical Computation and Simulation, 86(8):1635–1642, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (elec-

tronic), 1563-5163. URL http://www.tandfonline.com/
doi/abs/10.1080/00949655.2015.1080705.

Islam:2014:BSS

[IP14] A. F. M. Saiful Islam and Lawrence I. Pettit. Bayesian sample size determination for the bounded linex loss function. Journal of Statistical Computation and Simulation, 84(8):1644–1653, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Iachine:2010:RTE

[IPK10] Ivan Iachine, Hans Chr Petersen, and Kirsten O. Kyvik. Robust tests for the equality of variances for clustered data. Journal of Statistical Computation and Simulation, 80(4): 365–377, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ilic:2018:MBM

[IRNB18] Ana V. Miletić Ilić, Miroslav M. Ristić, Aleksandar S. Nastić, and Hassan S. Bakouch. An INAR(1) model based on a mixed dependent and independent counting series. Journal of Statistical Computation and Simulation, 88(2): 290–304, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Islam:2013:IUA

[IS13] Khairul Islam and Tanweer J. Shapla. Inference using attributable risk for a 2×2 case-control study. Journal of Statistical Computation and Simulation, 83(2):355–369, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ismail:2010:BEG

[Ism10] Ali A. Ismail. Bayes estimation of Gompertz distribution parameters and acceleration factor under partially accelerated life tests with Type-I censoring. Journal of Statistical Computation and Simulation, 80(11):1253–1264, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ismail:2014:LIS

[Ism14] Ali A. Ismail. Likelihood inference for a step-stress partially accelerated life test model with Type-I progressively hybrid censored data from Weibull distribution. Journal of Statistical Computation and Simulation, 84(11):2486–2494, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ismkhan:2016:NFH

[Ism16] Hassan Ismkhan. A novel fast heuristic to handle large-scale shape clustering. Journal of Statistical Computation and Simulation, 86(1):160–169, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jin:2019:CBP

[JCW19] Lei Jin, Li Cai, and Suojin Wang. A computational bootstrap procedure to compare two dependent time series. Journal of Statistical Computation and Simulation, 89(15):2831–2847, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jerbi:2013:MSV

[Jer13] Yacin Jerbi. Methodology for stochastic volatility process calibration application to the CAC 40 index. Journal of Statistical Computation and Simulation, 83(3):417–433, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jin:2010:EAW

[JG10] Tongdan Jin and Lakshmana S. Gonigunta. Exponential approximation to Weibull renewal with decreasing failure rate. Journal of Statistical Computation and Simulation, 80(3):273–285, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jia:2016:ANR

[JG16] Xiang Jia and Bo Guo. Analysis of non-repairable coldstandby systems in Bayes theory. Journal of Statistical Computation and Simulation, 86(11):2089–2112, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jimenez-Gamero:2017:ECF

[JGPF17] María Dolores Jiménez-Gamero and Juan Carlos Pardo-Fernández. Empirical characteristic function tests for GARCH innovation distribution using multipliers. Journal of Statistical Computation and Simulation, 87(10):2069– 2093, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jin:2015:CAS

[Jin15] Lei Jin. Comparing autocorrelation structures of multiple time series via the maximum distance between two groups of time series. Journal of Statistical Computation and Simulation, 85(17):3535–3548, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jerbi:2014:CED

[JK14] Yacin Jerbi and Mohamed Kharrat. Conditional expectation determination based on the J-process using Malliavin calculus applied to pricing American options. Journal of Statistical Computation and Simulation, 84(11):2465–2473, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jach:2017:WSP

[JK17] Agnieszka Jach and Piotr Kokoszka. Wavelet semiparametric inference for long memory in volatility in the presence of a trend. Journal of Statistical Computation and Simulation, 87(8):1498–1519, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jana:2019:OCR

[JK19] Nabakumar Jana and Somesh Kumar. Ordered classification rules for inverse Gaussian populations with unknown parameters. Journal of Statistical Computation and Simulation, 89(14):2597–2620, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Juutilainen:2011:BAB

[JKLR11] Ilmari Juutilainen, Heli Koskimäki, Perttu Laurinen, and Juha Röning. BUSDM — an algorithm for the bottomup search of departures from a model. Journal of Statistical Computation and Simulation, 81(5):561–578, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jana:2016:CRT

[JKM16] Nabakumar Jana, Somesh Kumar, and Neeraj Misra. Classification rules for two parameter exponential populations under order restrictions on parameters. Journal of Statistical Computation and Simulation, 86(8):1559–1581, 2016.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (elec-

tronic), 1563-5163. URL http://www.tandfonline.com/
doi/abs/10.1080/00949655.2015.1075540.

Jelsema:2019:TKS

[JKP19] Casey M. Jelsema, Richard K. Kwok, and Shyamal D. Peddada. Threshold knot selection for large-scale spatial models with applications to the Deepwater Horizon disaster. Journal of Statistical Computation and Simulation, 89(11):2121– 2137, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jiang:2016:NMT

[JL16] Xianfeng Jiang and Li Li. Naive method to test the convergence of simulation and its applications in the computation of bankruptcy probability. Journal of Statistical Computation and Simulation, 86(6):1216–1232, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1057822.

Jung:2015:EQR

[JLM15] Yoonsuh Jung, Yoonkyung Lee, and Steven N. MacEachern. Efficient quantile regression for heteroscedastic models. Journal of Statistical Computation and Simulation, 85 (13):2548–2568, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jansen:2010:PMM

[JM10] Ivy Jansen and Geert Molenberghs. Pattern-mixture models for categorical outcomes with non-monotone missingness. Journal of Statistical Computation and Simulation, 80(11):1279–1296, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Janani:2017:CBB

[JMM^{+}17] Leila Janani, Mohammad Ali Mansournia, Kazem Mohammad, Mahmood Mahmoodi, Kamran Mehrabani, and Keramat Nourijelyani. Comparison between Bayesian approach and frequentist methods for estimating relative risk in randomized controlled trials: a simulation study. Journal of Statistical Computation and Simulation, 87(4):640–651, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jacob:2011:RSD

[JO11] Pierre Jacob and Paulo Eduardo Oliveira. Relative smoothing of discrete distributions with sparse observations. Journal of Statistical Computation and Simulation, 81(1):109– 121, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jacob:2012:LSG

[JO12] Pierre Jacob and Paulo Eduardo Oliveira. Local smoothing with given marginals. Journal of Statistical Computation and Simulation, 82(6):915–926, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jones:2016:MWM

[Jon16] Sam Jones. Measuring what’s missing: practical estimates of coverage for stochastic simulations. Journal of Statistical Computation and Simulation, 86(9):1660–1672, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jouini:2015:LBM

[Jou15a] Tarek Jouini. Linear bootstrap methods for vector autoregressive moving-average models. Journal of Statistical Computation and Simulation, 85(11):2214–2257, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jourdan:2015:SFO

[Jou15b] Astrid Jourdan. Space-filling orthogonal arrays of strength two. Journal of Statistical Computation and Simulation, 85 (7):1382–1397, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jiang:2011:IOP

[JP11] Min Jiang and Serge B. Provost. Improved orthogonal polynomial density estimates. Journal of Statistical Computation and Simulation, 81(11):1495–1516, 2011. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Jiang:2014:HBS

[JP14] Min Jiang and Serge B. Provost. A hybrid bandwidth selection methodology for kernel density estimation. Journal of Statistical Computation and Simulation, 84(3):614–627, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jafari:2017:BDO

[JP17] H. Jafari and Sh. Pirmohamadi. The Bayesian D-optimal design for simple beta regression model in two cases (both known and unknown nuisance parameter). Journal of Statistical Computation and Simulation, 87(6):1230–1254, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:2013:ILC

[JR13] D. R. Jensen and D. E. Ramirez. Irregularities in X(Y) from Y (X) in linear calibration. Journal of Statistical Computation and Simulation, 83(10):1807–1828, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:2016:EDE

[JR16] D. R. Jensen and D. E. Ramirez. Enhanced design efficiency through least upper bounds. Journal of Statistical Computation and Simulation, 86(9):1798–1817, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:2017:SMM

[JR17] D. R. Jensen and D. E. Ramirez. Singular majorants and minorants: enhanced design conditioning. Journal of Statistical Computation and Simulation, 87(9):1827–1841, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jensen:2018:MCT

[JR18] D. R. Jensen and D. E. Ramirez. On mitigating collinearity through mixtures. Journal of Statistical Computation and Simulation, 88(8):1437–1453, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jin:2013:PDA

[JSM13] Seoweon Jin, Joan G. Staniswalis, and Indika Mallawaarachchi. Principal differential analysis with a continuous covariate: low-dimensional approximations for functional data. Journal of Statistical Computation and Simulation, 83(10):1964– 1980, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ju:2018:BLI

[JTL18] Yuanyuan Ju, Niansheng Tang, and Xiaoxia Li. Bayesian local influence analysis of skew-normal spatial dynamic panel data models. Journal of Statistical Computation and Simulation, 88(12):2342–2364, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Juhasz:2016:CCS

[Juh16] Ferenc Juhász. Crisis cycles in the slightly unstable block random model. Journal of Statistical Computation and Simulation, 86(16):3270–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jurczyk:2012:ODU

[Jur12] Tomás Jurczyk. Outlier detection under multicollinearity. Journal of Statistical Computation and Simulation, 82(2): 261–278, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Joenssen:2014:PSG

[JV14] Dieter W. Joenssen and Jürgen Vogel. A power study of goodness-of-fit tests for multivariate normality implemented in R. Journal of Statistical Computation and Simulation, 84(5):1055–1078, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jolani:2013:CCD

[JvBF13] S. Jolani, S. van Buuren, and L. E. Frank. Combining the complete-data and nonresponse models for drawing imputations under MAR. Journal of Statistical Computation and Simulation, 83(5):868–879, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jaki:2011:SMK

[JW11] Thomas Jaki and R. Webster West. Symmetric maximum kernel likelihood estimation. Journal of Statistical Computation and Simulation, 81(2):193–206, 2011. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jalilian:2018:FBS

[JW18] Abdollah Jalilian and Rasmus Waagepetersen. Fast bandwidth selection for estimation of the pair correlation function. Journal of Statistical Computation and Simulation, 88 (10):2001–2011, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jiryaie:2016:GCD

[JWWdL16] F. Jiryaie, N. Withanage, B. Wu, and A. R. de Leon. Gaussian copula distributions for mixed data, with application in discrimination. Journal of Statistical Computation and Simulation, 86(9):1643–1659, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jamshidian:2013:DDS

[JY13] Mortaza Jamshidian and Ke-Hai Yuan. Data-driven sensitivity analysis to detect missing data mechanism with applications to structural equation modelling. Journal of Statistical Computation and Simulation, 83(7):1344–1362, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jia:2014:TVN

[JY14] Li Jia and Kai Yuan. Time-varying neuro-fuzzy model using probability density function techniques for batch processes. Journal of Statistical Computation and Simulation, 84(6):1249–1260, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jiang:2013:IBI

[JYML13] Xiaomo Jiang, Yong Yuan, Sankaran Mahadevan, and Xian Liu. An investigation of Bayesian inference approach to model validation with non-normal data. Journal of Statistical Computation and Simulation, 83(10):1829–1851, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jebely:2018:EEP

[JZD18] F. Maleki Jebely, K. Zare, and E. Deiri. Efficient estimation of the PDF and the CDF of the inverse Rayleigh distribution. Journal of Statistical Computation and Simulation, 88 (1):75–88, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Jin:2018:MTC

[JZZH18] Hao Jin, Si Zhang, Jinsuo Zhang, and Han Hao. Modified tests for change points in variance in the possible presence of mean breaks. Journal of Statistical Computation and Simulation, 88(14):2651–2667, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kiani:2013:GMT

[KA13] Kaveh Kiani and Jayanthi Arasan. Gompertz model with time-dependent covariate in the presence of interval-, rightand left-censored data. Journal of Statistical Computation and Simulation, 83(8):1472–1490, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kashif:2018:PSL

[KAA18] Muhammad Kashif, Muhammad Amanullah, and Muhammad Aslam. Pena’s statistic for the Liu regression. Journal of Statistical Computation and Simulation, 88(13):2473– 2488, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kheyri:2019:KDE

[KAJB19] A. Kheyri, M. Amini, H. Jabbari, and A. Bozorgnia. Kernel density estimation under negative superadditive dependence and its application for real data. Journal of Statistical Computation and Simulation, 89(12):2373–2392, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kazemzadeh:2016:MSL

[KAK16] Reza Baradaran Kazemzadeh, Amirhossein Amiri, and Behnoush Kouhestani. Monitoring simple linear profiles using variable sample size schemes. Journal of Statistical Computation and Simulation, 86(15):2923–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kalaylioglu:2014:PBM

[Kal14] Zeynep Kalaylioglu. Performances of Bayesian model selection criteria for generalized linear models with nonignorably missing covariates. Journal of Statistical Computation and Simulation, 84(8):1670–1691, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kampke:2013:DE

[Käm13] Thomas Kämpke. Distance equalizers. Journal of Statistical Computation and Simulation, 83(9):1684–1703, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kang:2017:FSA

[Kan17] Sangwook Kang. Fitting semiparametric accelerated failure time models for nested case-control data. Journal of Statistical Computation and Simulation, 87(4):652–663, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khatib:2013:BRM

[KAR13] B. Khatib, Jafar Ahmadi, and M. Razmkhah. Bayesian reconstruction of the missing failure times in exponential distribution. Journal of Statistical Computation and Simulation, 83(3):501–517, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kara:2015:SIG

[KAT15] Mahmut Kara, Halil Aydogdu, and Özlem Türksen. Statistical inference for geometric process with the inverse Gaussian distribution. Journal of Statistical Computation and Simulation, 85(16):3206–3215, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kayid:2012:TDB

[KAWA12] M. Kayid, I. A. Al-Wasel, and I. A. Ahmad. Testing dominance of biometric functions. Journal of Statistical Computation and Simulation, 82(4):491–502, 2012. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Kirchner:2018:NEP

[KB18] M. Kirchner and A. Bercher. A nonparametric estimation procedure for the Hawkes process: comparison with maximum likelihood estimation. Journal of Statistical Computation and Simulation, 88(6):1106–1116, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kang:2016:HPQ

[KBJ16] Jongkyeong Kang, Sungwan Bang, and Myoungshic Jhun. Hierarchically penalized quantile regression. Journal of Statistical Computation and Simulation, 86(2):340–356, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krey:2015:CET

[KBL^{+}15] Sebastian Krey, Sebastian Brato, Uwe Ligges, Jürgen Götze, and Claus Weihs. Clustering of electrical transmission systems based on network topology and stability. Journal of Statistical Computation and Simulation, 85(1):47–61, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koh:2019:PDS

[KBM19] Y. B. Koh, N. A. Bukhari, and I. Mohamed. Parameterdriven state-space model for integer-valued time series with application. Journal of Statistical Computation and Simulation, 89(8):1394–1409, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kang:2011:LRW

[KBS11] Xiuli Kang, W. John Braun, and James E. Stafford. Local regression when the responses are interval-censored. Journal of Statistical Computation and Simulation, 81(10):1247– 1279, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kiesse:2014:DTA

[KC14] T. Senga Kiessé and H. E. Cuny. Discrete triangular associated kernel and bandwidth choices in semiparametric estimation for count data. Journal of Statistical Computation and Simulation, 84(8):1813–1829, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kulczycki:2016:ACM

[KC16] Piotr Kulczycki and Małgorzata Charytanowicz. An algorithm for conditional multidimensional parameter identification with asymmetric and correlated losses of underand overestimations. Journal of Statistical Computation and Simulation, 86(5):1032–1055, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1049606.

Krishna:2017:EPF

[KDG17] Hare Krishna, Madhulika Dube, and Renu Garg. Estimation of P (Y < X) for progressively first-failure-censored generalized inverted exponential distribution. Journal of Statistical Computation and Simulation, 87(11):2274–2289, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khedhiri:2010:EAO

[KE10] Sami Khedhiri and Ghassen El Montasser. The effects of additive outliers on the seasonal KPSS test: a Monte Carlo analysis. Journal of Statistical Computation and Simulation, 80(6):643–651, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kelly:2016:APV

[Kel16] G. E. Kelly. Approximations to the p-values of tests for a change-point under non-standard conditions. Journal of Statistical Computation and Simulation, 86(7):1430–1449, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koh:2013:ENA

[KEW13] Woon Yuen Koh, Kent M. Eskridge, and Dong Wang. The effects of nonnormality on the analysis of supersaturated designs: a comparison of stepwise, SCAD and permutation test methods. Journal of Statistical Computation and Simulation, 83(1):158–166, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kitromilidou:2016:REM

[KF16] Stella Kitromilidou and Konstantinos Fokianos. Robust estimation methods for a class of log-linear count time series models. Journal of Statistical Computation and Simulation, 86(4):740–755, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1035271.

Khowaja:2012:MOO

[KGA12] Saman Khowaja, Shazia Ghufran, and M. J. Ahsan. Multiobjective optimization for optimum allocation in multivariate stratified sampling with quadratic cost. Journal of Statistical Computation and Simulation, 82(12):1789–1798, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khan:2012:UVD

[Kha12] Naushad Mamode Khan. Use of vector divisions in solving quasi-likelihood equations for a Poisson model. Journal of Statistical Computation and Simulation, 82(9):1337–1342, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kiani:2010:EMG

[Kia10] M. Kiani. Extensions to the modified Gram–Schmidt strategy and its application in the steepest ascent method. Journal of Statistical Computation and Simulation, 80(4):389– 400, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kiani:2012:COD

[Kia12] M. Kiani. On the construction of optimal designs. Journal of Statistical Computation and Simulation, 82(7):1003–1014, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kibria:2012:SLR

[Kib12] B. M. Golam Kibria. Some Liu and ridge-type estimators and their properties under the ill-conditioned Gaussian linear regression model. Journal of Statistical Computation and Simulation, 82(1):1–17, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2018:FAC

[Kim18] Jiwoong Kim. A fast algorithm for the coordinate-wise minimum distance estimation. Journal of Statistical Computation and Simulation, 88(3):482–497, 2018. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kizilaslan:2017:BHB

[Kiz17] Fatih Kizilaslan. The E-Bayesian and hierarchical Bayesian estimations for the proportional reversed hazard rate model based on record values. Journal of Statistical Computation and Simulation, 87(11):2253–2273, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kizilaslan:2018:SRC

[Kiz18] Fatih Kizilaslan. Some reliability characteristics and stochastic ordering of series and parallel systems of bivariate generalized exponential distribution. Journal of Statistical Computation and Simulation, 88(3):553–574, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kozumi:2011:GSM

[KK11] Hideo Kozumi and Genya Kobayashi. Gibbs sampling methods for Bayesian quantile regression. Journal of Statistical Computation and Simulation, 81(11):1565–1578, 2011. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koutrouvelis:2012:CPG

[KK12] Ioannis A. Koutrouvelis and Alex Karagrigoriou. Cumulant plots and goodness-of-fit tests for the inverse Gaussian distribution. Journal of Statistical Computation and Simulation, 82(3):343–358, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krishna:2013:REG

[KK13] Hare Krishna and Kapil Kumar. Reliability estimation in generalized inverted exponential distribution with progressively Type II censored sample. Journal of Statistical Computation and Simulation, 83(6):1007–1019, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2014:ADD

[KK14] Daeyoung Kim and Jong-Min Kim. Analysis of directional dependence using asymmetric copula-based regression models. Journal of Statistical Computation and Simulation, 84 (9):1990–2010, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kobayashi:2015:GMP

[KK15] Genya Kobayashi and Hideo Kozumi. Generalized multiplepoint Metropolis algorithms for approximate Bayesian computation. Journal of Statistical Computation and Simulation, 85(4):675–692, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2018:FLR

[KK18] Hyojoong Kim and Heeyoung Kim. Functional logistic regression with fused lasso penalty. Journal of Statistical Computation and Simulation, 88(15):2982–2999, 2018. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kharrati-Kopaei:2017:SCI

[KKE17] Mahmood Kharrati-Kopaei and Sana Eftekhar. Simultaneous confidence intervals for comparing several inverse Gaussian means under heteroscedasticity. Journal of Statistical Computation and Simulation, 87(4):777–790, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kotthaus:2015:RMC

[KKL^{+}15] Helena Kotthaus, Ingo Korb, Michel Lang, Bernd Bischl, Jörg Rahnenführer, and Peter Marwedel. Runtime and memory consumption analyses for machine learning R programs. Journal of Statistical Computation and Simulation, 85(1):14–29, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kharrati-Kopaei:2016:MTI

[KKM16] Mahmood Kharrati-Kopaei and Arden Miller. A method for testing interaction in unreplicated two-way tables: using all pairwise interaction contrasts. Journal of Statistical Computation and Simulation, 86(6):1203–1215, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1057821.

Karakaya:2015:SIM

[KKY15] Jale Karakaya, Erdem Karabulut, and Recai M. Yucel. Sensitivity to imputation models and assumptions in receiver operating characteristic analysis with incomplete data. Journal of Statistical Computation and Simulation, 85(17):3498–3511, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krishnamoorthy:2010:PBS

[KL10] K. Krishnamoorthy and Fei Lu. A parametric bootstrap solution to the MANOVA under heteroscedasticity. Journal of Statistical Computation and Simulation, 80(8):873–887, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krishnamoorthy:2012:CFA

[KL12] K. Krishnamoorthy and Xiaodong Lian. Closed-form approximate tolerance intervals for some general linear models and comparison studies. Journal of Statistical Computation and Simulation, 82(4):547–563, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krishnamoorthy:2013:NAC

[KL13a] K. Krishnamoorthy and Meesook Lee. New approximate confidence intervals for the difference between two Poisson means and comparison. Journal of Statistical Computation and Simulation, 83(12):2232–2243, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuo:2013:DAC

[KL13b] Tsen-I Kuo and Pei-Hsi Lee. Design of adaptive s control charts. Journal of Statistical Computation and Simulation, 83(11):2002–2014, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2017:REZ

[KL17a] Byungsoo Kim and Sangyeol Lee. Robust estimation for zero-inflated Poisson autoregressive models based on density power divergence. Journal of Statistical Computation and Simulation, 87(15):2981–2996, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2017:FOI

[KL17b] Hanwool Kim and Sangyeol Lee. On first-order integervalued autoregressive process with Katz family innovations. Journal of Statistical Computation and Simulation, 87(3): 546–562, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumar:2017:TMU

[KL17c] Nirpeksh Kumar and Chien-Tai Lin. Testing for multiple upper and lower outliers in an exponential sample. Journal of Statistical Computation and Simulation, 87(5):870–881, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2018:BET

[KL18a] Minjo Kim and Sangyeol Lee. Bootstrap entropy test for general location-scale time series models with heteroscedasticity. Journal of Statistical Computation and Simulation, 88(13):2573–2588, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krakovsky:2018:RIF

[KL18b] Yury Krakovsky and Aleksandr Luzgin. Robust interval forecasting algorithm based on a probabilistic cluster model. Journal of Statistical Computation and Simulation, 88(12):2309–2324, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Klar:2015:NGD

[Kla15] Bernhard Klar. A note on gamma difference distributions. Journal of Statistical Computation and Simulation, 85(18):3708–3715, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2014:NIP

[KLK14] Dal Ho Kim, Woo Dong Lee, and Sang Gil Kang. Noninformative priors for the inverse Weibull distribution. Journal of Statistical Computation and Simulation, 84(5):1039– 1054, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kraus:2015:ENN

[KLK15] Johann M. Kraus, Ludwig Lausser, and Hans A. Kestler. Exhaustive k-nearest-neighbour subspace clustering. Journal of Statistical Computation and Simulation, 85(1):30–46, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kang:2017:OBM

[KLK17] Sang Gil Kang, Woo Dong Lee, and Yongku Kim. Objective Bayesian multiple comparisons for normal variances. Journal of Statistical Computation and Simulation, 87(5): 882–894, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2019:NPB

[KLKK19] Dal Ho Kim, Woo Dong Lee, Sang Gil Kang, and Yongku Kim. Noninformative priors for the between-group variance in the unbalanced one-way random effects model with heterogeneous error variances. Journal of Statistical Computation and Simulation, 89(10):1935–1956, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krishna:2012:REM

[KM12] Hare Krishna and Manish Malik. Reliability estimation in Maxwell distribution with progressively Type-II censored data. Journal of Statistical Computation and Simulation, 82(4):623–641, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kao:2014:FAC

[KM14] Ming-Hung Kao and Hans D. Mittelmann. A fast algorithm for constructing efficient event-related functional magnetic resonance imaging designs. Journal of Statistical Computation and Simulation, 84(11):2391–2407, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2017:RTE

[KM17] Tae-Hwan Kim and Christophe Muller. A robust test of exogeneity based on quantile regressions. Journal of Statistical Computation and Simulation, 87(11):2161–2174, 2017.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2017:SPN

[KMS17] Donghyun Kim, Hyungil Moon, and Hayong Shin. Some properties of nonlinear Lanchester equations with an application in military. Journal of Statistical Computation and Simulation, 87(13):2470–2479, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kizilaslan:2015:EGE

[KN15] Fatih Kizilaslan and Mustafa Nadar. Estimation with the generalized exponential distribution based on record values and inter-record times. Journal of Statistical Computation and Simulation, 85(5):978–999, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kizilaslan:2016:EPK

[KN16] Fatih Kizilaslan and Mustafa Nadar. Estimation and prediction of the Kumaraswamy distribution based on record values and inter-record times. Journal of Statistical Computation and Simulation, 86(12):2471–2493, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kassahun:2015:JMH

[KNM^{+}15] Wondwosen Kassahun, Thomas Neyens, Geert Molenberghs, Christel Faes, and Geert Verbeke. A joint model for hierarchical continuous and zero-inflated overdispersed count data. Journal of Statistical Computation and Simulation, 85(3):552–571, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuran:2019:MSC

[KÖ19] Özge Kuran and M. Revan Özkale. Model selection via conditional conceptual predictive statistic under ridge regression in linear mixed models. Journal of Statistical Computation and Simulation, 89(1):155–187, 2019. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Kharoubi:2019:CCN

[KOM19] Rachid Kharoubi, Karim Oualkacha, and Abdallah Mkhadri. The cluster correlation-network support vector machine for high-dimensional binary classification. Journal of Statistical Computation and Simulation, 89(6):1020–1043, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krishnamoorthy:2015:AOS

[KP15] K. Krishnamoorthy and Jie Peng. Approximate one-sided tolerance limits in random effects model and in some mixed models and comparisons. Journal of Statistical Computation and Simulation, 85(8):1651–1666, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kulkarni:2018:IIS

[KP18] H. V. Kulkarni and K. P. Patil. Improved inference for the shape-scale family of distributions under type-II censoring. Journal of Statistical Computation and Simulation, 88 (12):2259–2272, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Katebi:2019:ESD

[KP19] Mehdi Katebi and Reza Pourtaheri. An economic statistical design of the Poisson EWMA control charts for monitoring nonconformities. Journal of Statistical Computation and Simulation, 89(15):2813–2830, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koo:2013:SKQ

[KPK^{+}13] Ja-Yong Koo, Kwi Wook Park, Byung Won Kim, Kwang-Rae Kim, and Changyi Park. Structured kernel quantile regression. Journal of Statistical Computation and Simulation, 83(1):179–190, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Katebi:2016:EES

[KPM16] Mehdi Katebi, Reza Poortaheri, and Mohammad Bameni Moghadam. Economic and economic statistical designs for three-level control charts. Journal of Statistical Computation and Simulation, 86(8):1463–1478, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1071373.

Kalke:2013:SGG

[KR13] S. Kalke and W.-D. Richter. Simulation of the p-generalized Gaussian distribution. Journal of Statistical Computation and Simulation, 83(4):641–667, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumar:2015:AVZ

[KR15a] C. Satheesh Kumar and A. Riyaz. An alternative version of zero-inflated logarithmic series distribution and some of its applications. Journal of Statistical Computation and Simulation, 85(6):1117–1127, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kundu:2015:ETP

[KR15b] Debasis Kundu and Mohammad Z. Raqab. Estimation of R = P [Y < X] for three-parameter generalized Rayleigh distribution. Journal of Statistical Computation and Simulation, 85(4):725–739, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kohansal:2016:MEE

[KR16] Akram Kohansal and Saeid Rezakhah. Modified entropy estimators for testing normality. Journal of Statistical Computation and Simulation, 86(3):553–568, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khatib:2018:EPP

[KR18] Bahare Khatib and Mostafa Razmkhah. Estimation and prediction in the presence of an outlier under Type-II censoring. Journal of Statistical Computation and Simulation, 88(17):3251–3268, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krutchkoff:2012:EN

[Kru12] Richard G. Krutchkoff. Editor’s note. Journal of Statistical Computation and Simulation, 82(2):149, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kestler:2015:PR

[KS15] Hans A. Kestler and Matthias Schmid. Proceedings of Reisensburg 2013. Journal of Statistical Computation and Simulation, 85(1):1–2, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khan:2016:DCL

[KS16] Md Hasinur Rahaman Khan and J. Ewart H. Shaw. On dealing with censored largest observations under weighted least squares. Journal of Statistical Computation and Simulation, 86(18):3758–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Karageyik:2017:OLB

[KS17a] Basak Bulut Karageyik and Sule Sahin. Optimal lower barrier on modified surplus process. Journal of Statistical Computation and Simulation, 87(8):1520–1540, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kim:2017:MMC

[KS17b] Sungsu Kim and Ashis SenGupta. Multivariate-multiple circular regression. Journal of Statistical Computation and Simulation, 87(7):1277–1291, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumar:2017:MVI

[KS17c] C. Satheesh Kumar and S. Sreejakumari. On a modified version of intervened negative binomial distribution. Journal of Statistical Computation and Simulation, 87(3):437–446, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Khan:2016:MNS

[KSJ16] Naushad Mamode Khan, Yuvraj Sunecher, and Vandna Jowaheer. Modelling a non-stationary BINAR(1) Poisson process. Journal of Statistical Computation and Simulation, 86(15):3106–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Koslovsky:2018:UEA

[KSLN^{+}18] M. D. Koslovsky, M. D. Swartz, L. Leon-Novelo, W. Chan, and A. V. Wilkinson. Using the EM algorithm for Bayesian variable selection in logistic regression models with related covariates. Journal of Statistical Computation and Simulation, 88(3):575–596, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kubilius:2016:EPS

[KSM16] K. Kubilius, V. Skorniakov, and D. Melichov. Estimation of parameters of SDE driven by fractional Brownian motion with polynomial drift. Journal of Statistical Computation and Simulation, 86(10):1954–1969, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kaciranlar:2011:MRR

[KSÖG11] Selahattin Kaçiranlar, Sadullah Sakallioglu, M. Revan Özkale, and Hüseyin Güler. More on the restricted ridge regression estimation. Journal of Statistical Computation and Simulation, 81(11):1433–1448, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumar:2018:SSE

[KTJ18] Sunil Kumar, Menal Trehan, and J. P. Singh Joorel. A simulation study: estimation of population mean using two auxiliary variables in stratified random sampling. Journal of Statistical Computation and Simulation, 88(18):3694–3707, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kayal:2017:EPC

[KTSR17] Tanmay Kayal, Yogesh Mani Tripathi, Devendra Pratap Singh, and Manoj Kumar Rastogi. Estimation and prediction for Chen distribution with bathtub shape under progressive censoring. Journal of Statistical Computation and Simulation, 87(2):348–366, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuk:2018:TSP

[Kuk18] Anthony Kuk. A two-stage procedure to pool information across quantile levels in linear quantile regression. Journal of Statistical Computation and Simulation, 88(14):2852–2864, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuk:2019:ESM

[Kuk19] Anthony Y. C. Kuk. An extended semiparametric model for short-term and long-term hazard ratios. Journal of Statistical Computation and Simulation, 89(11):2138–2150, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kumar:2015:STC

[Kum15] Arun Kumar. Sequential tuning of complex computer models. Journal of Statistical Computation and Simulation, 85(2):393–404, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kuswanto:2011:NST

[Kus11] Heri Kuswanto. A new simple test against spurious long memory using temporal aggregation. Journal of Statistical Computation and Simulation, 81(10):1297–1311, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Krishna:2015:EMD

[KVK15] Hare Krishna, Vivekanand, and Kapil Kumar. Estimation in Maxwell distribution with randomly censored data. Journal of Statistical Computation and Simulation, 85(17):3560– 3578, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Leon:2011:UMM

[LAGM11] T. Leon, G. Ayala, M. Gaston, and F. Mallor. Using mathematical morphology for unsupervised classification of functional data. Journal of Statistical Computation and Simulation, 81(8):1001–1016, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2015:CCU

[LAuaS^{+}15] Hyeseon Lee, Muhammad Aslam, Qurat ul-ain Shakeel, Wonji Lee, and Chi-Hyuck Jun. A control chart using an auxiliary variable and repetitive sampling for monitoring process mean. Journal of Statistical Computation and Simulation, 85(16):3289–3296, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2011:FAB

[LB11] Albert H. Lee and Edward L. Boone. A frequentist assessment of Bayesian inclusion probabilities for screening predictors. Journal of Statistical Computation and Simulation, 81(9):1111–1119, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2019:PTE

[LBN^{+}19] Tao Li, N. Balakrishnan, Hon Keung Tony Ng, Yifan Lu, and Lu An. Precedence tests for equality of two distributions based on early failures of ranked set samples. Journal of Statistical Computation and Simulation, 89(12):2328– 2353, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lyu:2010:EPB

[LC10] JrJung Lyu and MingNan Chen. Evaluating the precision of bivariate attribute measurements. Journal of Statistical Computation and Simulation, 80(1):99–110, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2018:QSU

[LC18] Yi Liu and Xiaolin Chen. Quantile screening for ultra-high-dimensional heterogeneous data conditional on some variables. Journal of Statistical Computation and Simulation, 88(2):329–342, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2019:BVS

[LC19] Kuo-Jung Lee and Ray-Bing Chen. Bayesian variable selection in a finite mixture of linear mixed-effects models. Journal of Statistical Computation and Simulation, 89(13):2434– 2453, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2013:PSS

[LCB13] Chien-Tai Lin, Cheng-Chieh Chou, and N. Balakrishnan. Planning step-stress test plans under Type-I censoring for the log-location-scale case. Journal of Statistical Computation and Simulation, 83(10):1852–1867, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2014:PSS

[LCB14] Chien-Tai Lin, Cheng-Chieh Chou, and N. Balakrishnan. Planning step-stress test under Type-I censoring for the exponential case. Journal of Statistical Computation and Simulation, 84(4):819–832, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2015:EMM

[LCLP15] Seonjoo Lee, Brian S. Caffo, Balaji Lakshmanan, and Dzung L. Pham. Evaluating model misspecification in independent component analysis. Journal of Statistical Computation and Simulation, 85(6):1151–1164, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemonte:2012:BCB

[LCM12] Artur J. Lemonte, Gauss M. Cordeiro, and Germán Moreno. Bartlett corrections in Birnbaum–Saunders nonlinear regression models. Journal of Statistical Computation and Simulation, 82(6):927–935, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2017:PLS

[LCN^{+}17] Zhen Li, Bin Chen, Dandan Ning, Zhichao Song, Gang Guo, and Xiaogang Qiu. Partitioning large-scale artificial society on distributed cluster with statistical movement graph. Journal of Statistical Computation and Simulation, 87(18):3413–3439, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2018:CVM

[LCZ18] Haibo Lin, Li Chen, and Juan Zhao. Critical value model of sample correlation coefficient after windowed moving average. Journal of Statistical Computation and Simulation, 88(18):3681–3693, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Landsittel:2010:EAF

[LDB10] Douglas Landsittel and Nathan Donohue-Babiak. Effect of adding fold-change criteria to significance testing of microarray data. Journal of Statistical Computation and Simulation, 80(1):89–97, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lachos:2017:SML

[LDCL17] Victor H. Lachos, Dipak K. Dey, Vicente G. Cancho, and Francisco Louzada. Scale mixtures log-Birnbaum–Saunders regression models with censored data: a Bayesian approach. Journal of Statistical Computation and Simulation, 87(10): 2002–2022, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lima:2018:RMT

[LdNdSF18] Stênio Rodrigues Lima, Fernando Ferraz do Nascimento, and Valmária Rocha da Silva Ferraz. Regression models for time-varying extremes. Journal of Statistical Computation and Simulation, 88(2):235–249, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Luangkesorn:2016:MCM

[LED16] K. L. Luangkesorn and Z. F. Eren-Dogu. Markov Chain Monte Carlo methods for estimating surgery duration. Journal of Statistical Computation and Simulation, 86(2):262– 278, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2011:MWK

[Lee11] Seung-Hwan Lee. Maximum of the weighted Kaplan–Meier tests for the two-sample censored data. Journal of Statistical Computation and Simulation, 81(8):1017–1026, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2017:ICI

[Lee17] Minjung Lee. Inference for cumulative incidence on discrete failure times with competing risks. Journal of Statistical Computation and Simulation, 87(10):1989–2001, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2019:RBC

[Lee19] Sangyeol Lee. Residual-based CUSUM of squares test for Poisson integer-valued GARCH models. Journal of Statistical Computation and Simulation, 89(17):3182–3195, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemonte:2011:CMF

[Lem11a] Artur J. Lemonte. Covariance matrix formula for Birnbaum–Saunders regression models. Journal of Statistical Computation and Simulation, 81(7):899–908, 2011. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemonte:2011:IML

[Lem11b] Artur J. Lemonte. Improved maximum-likelihood estimation in a regression model with general parametrization. Journal of Statistical Computation and Simulation, 81(8):1027–1037, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemonte:2011:IPE

[Lem11c] Artur J. Lemonte. Improved point estimation for the Kumaraswamy distribution. Journal of Statistical Computation and Simulation, 81(12):1971–1982, 2011. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemonte:2012:LBS

[Lem12] Artur J. Lemonte. A log-Birnbaum–Saunders regression model with asymmetric errors. Journal of Statistical Computation and Simulation, 82(12):1775–1787, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemonte:2013:MBS

[Lem13] Artur J. Lemonte. Multivariate Birnbaum–Saunders regression model. Journal of Statistical Computation and Simulation, 83(12):2244–2257, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemonte:2016:RIB

[Lem16] Artur J. Lemonte. Robust inference for Birnbaum–Saunders regressions. Journal of Statistical Computation and Simulation, 86(3):611–622, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lengyel:2011:APS

[Len11] Tamás Lengyel. On approximating point spread distributions. Journal of Statistical Computation and Simulation, 81(10):1333–1344, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

LaMorgia:2016:PDS

[LF16] Valentina La Morgia and Stefano Focardi. Performance of distance sampling estimators: a simulation study for designs based on footpaths. Journal of Statistical Computation and Simulation, 86(13):2519–2530, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2016:PCR

[LGW16] Chunling Liu, Shuang Guo, and Chuanhua Wei. Principal components regression estimator of the parameters in partially linear models. Journal of Statistical Computation and Simulation, 86(15):3127–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2018:DOI

[LGWZ18] Xuqing Liu, Feng Gao, Yandong Wu, and Zhiguo Zhao. Detecting outliers and influential points: an indirect classical Mahalanobis distance-based method. Journal of Statistical Computation and Simulation, 88(11):2013–2033, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2012:PHC

[LH12] Chien-Tai Lin and Yen-Lung Huang. On progressive hybrid censored exponential distribution. Journal of Statistical Computation and Simulation, 82(5):689–709, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2014:RES

[LH14] Yi-Kuei Lin and Cheng-Fu Huang. Reliability evaluation subject to assured accuracy rate and time for stochastic unreliable-node computer networks. Journal of Statistical Computation and Simulation, 84(7):1530–1542, 2014. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lac:2018:NPS

[LH18] Le An Lac and Shakhawat Hossain. Non-penalty shrinkage estimation of random effect models for longitudinal data with AR(1) errors. Journal of Statistical Computation and Simulation, 88(16):3230–3247, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2011:EBV

[LHB11] Chien-Tai Lin, Yen-Lung Huang, and N. Balakrishnan. Exact Bayesian variable sampling plans for the exponential distribution with progressive hybrid censoring. Journal of Statistical Computation and Simulation, 81(7):873–882, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2013:EBV

[LHB13] Chien-Tai Lin, Yen-Lung Huang, and N. Balakrishnan. Exact Bayesian variable sampling plans for the exponential distribution with progressive hybrid censoring. Journal of Statistical Computation and Simulation, 83(2):402–404, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Lia14].

Lin:2019:ICS

[LHLB19] Chien-Tai Lin, Yao-Yu Hsu, Siao-Yu Lee, and N. Balakrishnan. Inference on constant stress accelerated life tests for log-location-scale lifetime distributions with type-I hybrid censoring. Journal of Statistical Computation and Simulation, 89(4):720–749, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2011:LFT

[Li11] Chin-Shang Li. A lack-of-fit test for parametric zero-inflated Poisson models. Journal of Statistical Computation and Simulation, 81(9):1081–1098, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2014:EPA

[Li14] Mingfei Li. An extension of play against the random past strategy. Choosing the right experts on IBM forecasts. Journal of Statistical Computation and Simulation, 84(12):2680– 2687, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2015:TLN

[Li15a] Chin-Shang Li. Testing the linearity of negative binomial regression models. Journal of Statistical Computation and Simulation, 85(5):1013–1025, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2015:PSC

[Li15b] Dongmei Li. Power and stability comparisons of multiple testing procedures with false discovery rate control. Journal of Statistical Computation and Simulation, 85(14):2808– 2822, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2018:CCB

[Li18] Chung-I Li. Control charts based on quasi-likelihood estimation for monitoring profiles. Journal of Statistical Computation and Simulation, 88(3):457–470, 2018. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liang:2010:DMH

[Lia10] Faming Liang. A double Metropolis–Hastings sampler for spatial models with intractable normalizing constants. Journal of Statistical Computation and Simulation, 80(9):1007– 1022, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liang:2011:CBP

[Lia11] TaChen Liang. Comments on ‘Life prediction under random censorship’. Journal of Statistical Computation and Simulation, 81(6):795–797, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [?].

Liang:2014:CEB

[Lia14] TaChen Liang. Comment on “Exact Bayesian variable sampling plans for the exponential distribution with progressive hybrid censoring”. Journal of Statistical Computation and Simulation, 84(3):468–476, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [LHB13].

Lin:2014:MSI

[Lin14] Ting Hsiang Lin. Model selection information criteria in latent class models with missing data and contingency question. Journal of Statistical Computation and Simulation, 84(1):159–170, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Linhares:2016:SDF

[Lin16] Raquel Romes Linhares. Smoothed detrended fluctuation analysis. Journal of Statistical Computation and Simulation, 86(17):3388–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2018:CAB

[LJ18] Feng Li and Li Jia. Correlation analysis-based error compensation recursive least-square identification method for the Hammerstein model. Journal of Statistical Computation and Simulation, 88(1):56–74, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2015:IDV

[LJL15] Yonghui Liu, Guocheng Ji, and Shuangzhe Liu. Influence diagnostics in a vector autoregressive model. Journal of Statistical Computation and Simulation, 85(13):2632–2655, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lall:2018:AAC

[LJV^{+}18] Shwetank Lall, Seema Jaggi, Eldho Varghese, Cini Varghese, and Arpan Bhowmik. An algorithmic approach to construct D-optimal saturated designs for logistic model. Journal of Statistical Computation and Simulation, 88(6):1191– 1199, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lyhagen:2013:SSP

[LK13] Johan Lyhagen and Katrin Kraus. The small sample performance of estimators of the standard errors of structural equation models. Journal of Statistical Computation and Simulation, 83(3):458–471, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2017:EBG

[LK17] Sangyeol Lee and Minjo Kim. On entropy-based goodnessof-fit test for asymmetric Student-t and exponential power distributions. Journal of Statistical Computation and Simulation, 87(1):187–197, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lang:2015:AMS

[LKM^{+}15] M. Lang, H. Kotthaus, P. Marwedel, C. Weihs, J. Rahnenführer, and B. Bischl. Automatic model selection for high-dimensional survival analysis. Journal of Statistical Computation and Simulation, 85(1):62–76, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2010:REO

[LL10] Sangyeol Lee and Taewook Lee. Robust estimation for order of hidden Markov models based on density power divergences. Journal of Statistical Computation and Simulation, 80(5):503–512, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2011:VRF

[LL11] Sangyeol Lee and Taewook Lee. Value-at-risk forecasting based on Gaussian mixture ARMA–GARCH model. Journal of Statistical Computation and Simulation, 81(9):1131– 1144, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lu:2013:FAN

[LL13] Kewei Lu and Linyuan Li. On Fan’s adaptive Neyman tests for comparing two spectral densities. Journal of Statistical Computation and Simulation, 83(9):1585–1601, 2013. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2015:MPE

[LL15a] Hyejin Lee and Junsoo Lee. More powerful Engle–Granger cointegration tests. Journal of Statistical Computation and Simulation, 85(15):3154–3171, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2015:IAF

[LL15b] Jing Li and Junsoo Lee. Improved autoregressive forecasts in the presence of non-normal errors. Journal of Statistical Computation and Simulation, 85(14):2936–2952, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2016:TCI

[LL16] Weiming Li and Zhi Liu. A test for the complete independence of high-dimensional random vectors. Journal of Statistical Computation and Simulation, 86(16):3135–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2017:GSC

[LL17] Chi-Rong Li and Tsai-Yu Lin. Generalized simultaneous confidence regions for regression coefficients and their ratios. Journal of Statistical Computation and Simulation, 87 (10):1940–1950, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lai:2018:STP

[LL18a] Xin Lai and Liu Liu. A simple test procedure in standardizing the power of Hosmer–Lemeshow test in large data sets. Journal of Statistical Computation and Simulation, 88 (13):2463–2472, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2018:SLB

[LL18b] Xuqing Liu and Xinsheng Liu. Structure learning of Bayesian networks by continuous particle swarm optimization algorithms. Journal of Statistical Computation and Simulation, 88(8):1528–1556, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2019:CMA

[LL19] Yongxin Liu and Lu Lin. Classification with minimum ambiguity under distribution heterogeneity. Journal of Statistical Computation and Simulation, 89(12):2239–2260, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lydersen:2012:EUP

[LLB12] Stian Lydersen, Mette Langaas, and Øyvind Bakke. The exact unconditional z-pooled test for equality of two binomial probabilities: optimal choice of the Berger and Boos confidence coefficient. Journal of Statistical Computation and Simulation, 82(9):1311–1316, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2014:NSA

[LLBL14] Z. H. Liu, Y. Y. Li, J. W. Bian, and H. W. Li. Noise subspace approach for frequency estimation of two-dimensional harmonics in zero-mean multiplicative and additive noise. Journal of Statistical Computation and Simulation, 84(11): 2378–2390, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lim:2017:SAM

[LLC17] Hwa Kyung Lim, Jaejun Lee, and Sooyoung Cheon. Stochastic approximation Monte Carlo EM for change-point analysis. Journal of Statistical Computation and Simulation, 87(1):69–87, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lai:2017:SSC

[LLGP17] Chin-Diew Lai, Gwo Dong Lin, K. Govindaraju, and Sarah Pirikahu. A simulation study on the correlation structure of Marshall–Olkin bivariate Weibull distribution. Journal of Statistical Computation and Simulation, 87(1):156–170, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2016:TEC

[LLM16] Xiaobin Liu, Song Liu, and Chang-Xing Ma. Testing equality of correlation coefficients for paired binary data from multiple groups. Journal of Statistical Computation and Simulation, 86(9):1686–1696, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2013:BEM

[LLN13] Yu-Jau Lin, Y. L. Lio, and Hon Keung Tony Ng. Bayes estimation of Moran–Downton bivariate exponential distribution based on censored samples. Journal of Statistical Computation and Simulation, 83(5):837–852, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lloyd:2010:ETB

[Llo10] Chris J. Lloyd. Exact tests based on pre-estimation and second order pivotals: non-inferiority trials. Journal of Statistical Computation and Simulation, 80(8):841–851, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lim:2014:BME

[LLP^{+}14] Johan Lim, Sungim Lee, Heon-Jin Park, Kyeong Eun Lee, and Shin-Jae Lee. Bootstrap method to evaluate tightness of clusters with application to the Korean standard occlusion study. Journal of Statistical Computation and Simulation, 84(2):360–372, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2016:ASG

[LLQ^{+}16] Lin Li, Yang Li, Yichen Qin, Jiaxu Chen, Limin Wang, and Danhui Yi. Adaptive stochastic gradient boosting tree with composite criterion. Journal of Statistical Computation and Simulation, 86(10):1901–1911, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2013:PTC

[LLS13] Junsoo Lee, Jing Li, and Mark C. Strazicich. Performance of threshold cointegration tests. Journal of Statistical Computation and Simulation, 83(5):793–809, 2013. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Luo:2012:BQR

[LLT12] Youxi Luo, Heng Lian, and Maozai Tian. Bayesian quantile regression for longitudinal data models. Journal of Statistical Computation and Simulation, 82(11):1635–1649, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2014:DSD

[LLV^{+}14] Xiaobo Li, Yonglin Lei, Hans Vangheluwe, Weiping Wang, and Qun Li. Domain-specific decision modelling and statistical analysis for combat system effectiveness simulation. Journal of Statistical Computation and Simulation, 84(6):1261–1279, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2015:FAM

[LLWY15] Juxin Liu, Wei Liu, Lang Wu, and Guohua Yan. A flexible approach for multivariate mixed-effects models with nonignorable missing values. Journal of Statistical Computation and Simulation, 85(18):3727–3743, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2017:NEA

[LLXY17] Wei Li, Lingyun Lu, Xiaotian Xie, and Ming Yang. A novel extension algorithm for optimized latin hypercube sampling. Journal of Statistical Computation and Simulation, 87(13):2549–2559, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2019:TLF

[LLY19] Chin-Shang Li, Shen-Ming Lee, and Ming-Shan Yeh. A test for lack-of-fit of zero-inflated negative binomial models. Journal of Statistical Computation and Simulation, 89 (7):1301–1321, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liseo:2013:OPC

[LM13] Brunero Liseo and Christian Macaro. Objective priors for causal AR(p) with partial autocorrelations. Journal of Statistical Computation and Simulation, 83(9):1613–1628, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2016:SRT

[LM16] Bo Li and Hossein Mansouri. Simultaneous rank tests for detecting differentially expressed genes. Journal of Statistical Computation and Simulation, 86(5):959–972, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1046073.

Li:2018:SIG

[LM18a] Dengkui Li and Changlin Mei. SiZer inference for generalized varying coefficient models. Journal of Statistical Computation and Simulation, 88(18):3544–3564, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lou:2018:ECS

[LM18b] Dongfang Lou and Zhiyuan Ma. Estimating coefficients of single-index regression models by minimizing variation. Journal of Statistical Computation and Simulation, 88(2): 343–358, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemonte:2015:MBS

[LMFMA15] Artur J. Lemonte, Guillermo Martínez-Florez, and Germán Moreno-Arenas. Multivariate Birnbaum–Saunders distribution: properties and associated inference. Journal of Statistical Computation and Simulation, 85(2):374–392, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2017:LCD

[LMRW17] Lina Liu, Shiwei Ma, Ling Rui, and Jiarui Wen. Locality constrained dictionary learning for human behaviour recognition. Journal of Statistical Computation and Simulation, 87(13):2526–2537, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2016:RAE

[LMSX16] Chenglong Li, Amitava Mukherjee, Qin Su, and Min Xie. Robust algorithms for economic designing of a nonparametric control chart for abrupt shift in location. Journal of Statistical Computation and Simulation, 86(2):306–323, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Leuchs:2013:MCT

[LN13] Ann-Kristin Leuchs and Markus Neuhäuser. A modified combination test for the analysis of a clinical trial when a protocol amendment changed the inclusion criteria. Journal of Statistical Computation and Simulation, 83(5):825–836, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lim:2017:RMM

[LNC17] Hwa Kyung Lim, Naveen N. Narisetty, and Sooyoung Cheon. Robust multivariate mixture regression models with incomplete data. Journal of Statistical Computation and Simulation, 87(2):328–347, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2016:SON

[LOK16] Sangin Lee, Miae Oh, and Yongdai Kim. Sparse optimization for nonconvex group penalized estimation. Journal of Statistical Computation and Simulation, 86(3):597–610, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Longford:2012:SSE

[Lon12] Nicholas T. Longford. Small-sample estimators of the quantiles of the normal, log-normal and Pareto distributions. Journal of Statistical Computation and Simulation, 82(9):1383–1395, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lio:2010:BCC

[LP10] Y. L. Lio and Chanseok Park. A bootstrap control chart for inverse Gaussian percentiles. Journal of Statistical Computation and Simulation, 80(3):287–299, 2010. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lagona:2013:MLE

[LP13] Francesco Lagona and Marco Picone. Maximum likelihood estimation of bivariate circular hidden Markov models from incomplete data. Journal of Statistical Computation and Simulation, 83(7):1223–1237, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lagona:2016:MBS

[LP16a] Francesco Lagona and Marco Picone. Model-based segmentation of spatial cylindrical data. Journal of Statistical Computation and Simulation, 86(13):2598–2610, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2016:MLB

[LP16b] Xijia Liu and Daniel Preve. Measure of location-based estimators in simple linear regression. Journal of Statistical Computation and Simulation, 86(9):1771–1784, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2017:TSC

[LP17] Taewook Lee and Cheolwoo Park. Tests for serial correlation in mean and variance of a sequence of time series objects. Journal of Statistical Computation and Simulation, 87(3): 478–492, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2014:EWM

[LPJ14] Sang-Ho Lee, Jang-Ho Park, and Chi-Hyuck Jun. An exponentially weighted moving average chart controlling false discovery rate. Journal of Statistical Computation and Simulation, 84(8):1830–1840, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2018:EGF

[LPK18] Sangyeol Lee, Siyun Park, and Byungsoo Kim. On entropy goodness-of-fit test based on integrated distribution function. Journal of Statistical Computation and Simulation, 88 (12):2447–2461, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2012:SAS

[LPS12] Shuangzhe Liu, Wolfgang Polasek, and Richard Sellner. Sensitivity analysis of SAR estimators: a numerical approximation. Journal of Statistical Computation and Simulation, 82(2):325–342, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lu:2013:EBD

[LPS13] Bo Lu, Juan Peng, and Timothy Sahr. Estimation bias of different design and analytical strategies in dual-frame telephone surveys: an empirical evaluation. Journal of Statistical Computation and Simulation, 83(12):2352–2368, 2013.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Linder:2016:BBL

[LPSR16] Daniel F. Linder, Viral Panchal, Hani Samawi, and Duchwan Ryu. Balanced Bayesian LASSO for heavy tails. Journal of Statistical Computation and Simulation, 86(6):1115–1132, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

URL http://www.tandfonline.com/doi/abs/10.1080/
00949655.2015.1053886.

Lunardon:2013:NEL

[LPV13] Nicola Lunardon, Francesco Pauli, and Laura Ventura. A note on empirical likelihoods derived from pairwise score functions. Journal of Statistical Computation and Simulation, 83(8):1405–1414, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2018:NRM

[LPY18] Yi-Kuei Lin, Chih-Li Pan, and Louis Cheng-Lu Yeng. Network reliability for multipath TCP networks with a retransmission mechanism under the time constraint. Journal of Statistical Computation and Simulation, 88(12):2273–2286, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lloyd:2018:ENA

[LR18a] Chris J. Lloyd and Enrico Ripamonti. An exhaustive numerical assessment of alternative unconditional tests of a binary treatment effect. Journal of Statistical Computation and Simulation, 88(11):2150–2169, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Louzada:2018:ECF

[LR18b] Francisco Louzada and Pedro Luiz Ramos. Efficient closedform maximum a posteriori estimators for the gamma distribution. Journal of Statistical Computation and Simulation, 88(6):1134–1146, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Leisen:2017:NPI

[LRV17] Fabrizio Leisen, Luca Rossini, and Cristiano Villa. A note on the posterior inference for the Yule–Simon distribution. Journal of Statistical Computation and Simulation, 87(6):1179–1188, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2011:LNC

[LS11] Yushu Li and Ghazi Shukur. Linear and nonlinear causality tests in an LSTAR model: wavelet decomposition in a nonlinear environment. Journal of Statistical Computation and Simulation, 81(12):1913–1925, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2014:PEM

[LS14] Jaejun Lee and T. N. Sriram. On the performance of L_{2} E estimation in modelling heterogeneous count responses with extreme values. Journal of Statistical Computation and Simulation, 84(3):564–581, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lee:2019:IBR

[LS19] DongHyuk Lee and Samiran Sinha. Identifiability and bias reduction in the skew-probit model for a binary response. Journal of Statistical Computation and Simulation, 89(9):1621–1648, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lemaitre:2015:DMB

[LSA^{+}15] P. Lemaître, E. Sergienko, A. Arnaud, N. Bousquet, F. Gamboa, and B. Iooss. Density modification-based reliability sensitivity analysis. Journal of Statistical Computation and Simulation, 85(6):1200–1223, 2015. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lisawadi:2016:MSP

[LSA16] Supranee Lisawadi, Muhammad Kashif Ali Shah, and S. Ejaz Ahmed. Model selection and post estimation based on a pretest for logistic regression models. Journal of Statistical Computation and Simulation, 86(17):3495–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2017:RLE

[LSF^{+}17] Tianyu Liu, Quan Sun, Jing Feng, Zhengqiang Pan, and Qizi Huangpeng. Residual life estimation under timevarying conditions based on a Wiener process. Journal of Statistical Computation and Simulation, 87(2):211–226, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Larsen:2010:FSL

[LSL10] Ross A. Larsen, G. Bruce Schaalje, and John S. Lawson. Food shelf life: estimation and optimal design. Journal of Statistical Computation and Simulation, 80(2):143–157, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lin:2016:MSC

[LT16] Ting Hsiang Lin and Min-Hsiao Tsai. Model selection criteria for dual-inflated data. Journal of Statistical Computation and Simulation, 86(13):2663–2672, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lio:2018:DSG

[LTJB18] Y. L. Lio, Tzong-Ru Tsai, Nan Jiang, and Narayanaswamy Balakrishnan. A dynamic system for Gompertz model. Journal of Statistical Computation and Simulation, 88(4): 752–768, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lu:2014:DBV

[Lu14] Yan Lu. Difference-based variance estimator for nonparametric regression in complex surveys. Journal of Statistical Computation and Simulation, 84(2):335–343, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lu:2019:FPB

[Lu19] Jiannan Lu. On finite-population Bayesian inferences for 2^{K} factorial designs with binary outcomes. Journal of Statistical Computation and Simulation, 89(5):927–945, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lue:2019:PDE

[Lue19] Heng-Hui Lue. Pairwise directions estimation for multivariate response regression data. Journal of Statistical Computation and Simulation, 89(5):776–794, 2019. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lutkenoner:2015:FKT

[Lüt15] B. Lütkenöner. A family of kernels and their associated deconvolving kernels for normally distributed measurement errors. Journal of Statistical Computation and Simulation, 85(12):2347–2363, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lovaglio:2017:GSM

[LV17] Pietro Giorgio Lovaglio and Gianmarco Vacca. %Gra: an SAS macro for generalized redundancy analysis. Journal of Statistical Computation and Simulation, 87(5):1048–1060, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Loose:2018:BTI

[LVB18] Laís H. Loose, Dione Maria Valença, and Fábio Mariano Bayer. On bootstrap testing inference in cure rate models. Journal of Statistical Computation and Simulation, 88 (17):3437–3454, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2012:CSB

[LW12] H. Li and L. Wang. A consistent simulation-based estimator in generalized linear mixed models. Journal of Statistical Computation and Simulation, 82(8):1085–1103, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2014:TCM

[LW14] Xiaosong Li and Patricia Pepple Williamson. Testing on the common mean of normal distributions using Bayesian methods. Journal of Statistical Computation and Simulation, 84 (6):1363–1380, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ledwina:2015:DNG

[LW15] Teresa Ledwina and Grzegorz Wylupek. Detection of non-Gaussianity. Journal of Statistical Computation and Simulation, 85(17):3480–3497, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2017:BEG

[LW17] Shengnan Li and Min Wang. Bayesian estimation of the generalized lognormal distribution using objective priors. Journal of Statistical Computation and Simulation, 87(7):1323– 1341, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2019:CCB

[LWS19] Cong Li, Dehui Wang, and Jianguo Sun. Control charts based on dependent count data with deflation or inflation of zeros. Journal of Statistical Computation and Simulation, 89(17):3273–3289, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2017:SCD

[LWT^{+}17] Xueyan Liu, Bryan Winter, Li Tang, Bo Zhang, Zhiwei Zhang, and Hui Zhang. Simulating comparisons of different computing algorithms fitting zero-inflated Poisson models for zero abundant counts. Journal of Statistical Computation and Simulation, 87(13):2609–2621, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2018:FVS

[LWW18] Ruiping Liu, Huiwen Wang, and Shanshan Wang. Functional variable selection via Gram–Schmidt orthogonalization for multiple functional linear regression. Journal of Statistical Computation and Simulation, 88(18):3664–3680, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2017:MFF

[LWZ17] Xiangjie Li, Lei Wang, and Jingxiao Zhang. A model-free feature screening approach based on kernel density estimation. Journal of Statistical Computation and Simulation, 87 (12):2450–2468, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2014:IGE

[LX14] Hui Li and Tianyuan Xiao. Improved generalized energy index method for comprehensive evaluation and prediction of track irregularity. Journal of Statistical Computation and Simulation, 84(6):1213–1231, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2016:FIB

[LX16] Yalan Li and Ancha Xu. Fiducial inference for Birnbaum– Saunders distribution. Journal of Statistical Computation and Simulation, 86(9):1673–1685, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ling:2011:OBS

[LXL11] Li Ling, Wei Xu, and Minghai Li. Optimal bivariate step-stress accelerated life test for Type-I hybrid censored data. Journal of Statistical Computation and Simulation, 81(9):1175–1186, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2017:SAE

[LXL17] Zhaonan Li, Xinyi Xu, and Bo Lu. Small area estimation strategies for large population surveys: a comparison of design and model-based methods. Journal of Statistical Computation and Simulation, 87(4):817–833, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2011:NGV

[LXZ11] Xuhua Liu, Xingzhong Xu, and Jianxin Zhao. A new generalized p-value approach for testing equality of coefficients of variation in k normal populations. Journal of Statistical Computation and Simulation, 81(9):1121–1130, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liang:2013:OBS

[LY13] TaChen Liang and Ming-Chung Yang. Optimal Bayesian sampling plans for exponential distributions based on hybrid censored samples. Journal of Statistical Computation and Simulation, 83(5):922–940, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Latif:2015:ODB

[LY15] Shahid Latif and Mohammad Zafar Yab. D-optimal designs for beta regression models with single predictor. Journal of Statistical Computation and Simulation, 85(9):1709–1724, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2018:FBL

[LY18] Longhai Li and Weixin Yao. Fully Bayesian logistic regression with hyper-LASSO priors for high-dimensional feature selection. Journal of Statistical Computation and Simulation, 88(14):2827–2851, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2017:TEM

[LYL17] Xiaobin Li, Chao Yin, and Fei Liu. A trust estimation method of machine tool resources in the cloud environment. Journal of Statistical Computation and Simulation, 87(13):2572–2580, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2017:EMT

[LYLM17] Xiaobin Liu, Zhengyu Yang, Song Liu, and Chang-Xing Ma. Exact methods of testing the homogeneity of prevalences for correlated binary data. Journal of Statistical Computation and Simulation, 87(15):3021–3039, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2015:RRO

[LYQ^{+}15] Yang Li, Chenqun Yu, Yichen Qin, Limin Wang, Jiaxu Chen, Danhui Yi, Ben-Chang Shia, and Shuangge Ma. Regularized receiver operating characteristic-based logistic regression for grouped variable selection with composite criterion. Journal of Statistical Computation and Simulation, 85(13):2582–2595, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2011:SIV

[LYZ11] Xiao Li Li, Jin Hong You, and Yong Zhou. Statistical inference for varying-coefficient models with error-prone covariates. Journal of Statistical Computation and Simulation, 81 (12):1755–1771, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lu:2010:EBL

[LZ10] Yan Lu and Guoyi Zhang. The equivalence between likelihood ratio test and f-test for testing variance component in a balanced one-way random effects model. Journal of Statistical Computation and Simulation, 80(4):443–450, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2011:CFW

[LZ11] Wei Liu and Sanyu Zhou. Construction of fixed width confidence intervals for a Bernoulli success probability using sequential sampling: a simulation study. Journal of Statistical Computation and Simulation, 81(11):1483–1493, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2014:CAR

[LZGW14] Zhonghua Li, Changliang Zou, Zhen Gong, and Zhaojun Wang. The computation of average run length and average time to signal: an overview. Journal of Statistical Computation and Simulation, 84(8):1779–1802, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2015:MSF

[LZZ^{+}15] Wei Liu, Bo Zhang, Zhiwei Zhang, Jian Tao, and Adam J. Branscum. Model selection in finite mixture of regression models: a Bayesian approach with innovative weighted g priors and reversible jump Markov chain Monte Carlo implementation. Journal of Statistical Computation and Simulation, 85(12):2456–2478, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Liu:2018:EIG

[LZZM18] Yuan Liu, Yan-Yong Zhao, Jian-Qiang Zhao, and Zhang-Xiao Miao. Estimation and inference for generalized semivarying coefficient models. Journal of Statistical Computation and Simulation, 88(11):2215–2231, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Li:2017:NHC

[LZZW17] Shunyong Li, Nahui Zhang, Xiaoqin Zhang, and Guannan Wang. A new heteroskedasticity-consistent covariance matrix estimator and inference under heteroskedasticity. Journal of Statistical Computation and Simulation, 87(1):198– 210, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mahdizadeh:2010:ERS

[MA10] Mahdi Mahdizadeh and N. R. Arghami. Efficiency of ranked set sampling in entropy estimation and goodness-of-fit testing for the inverse Gaussian law. Journal of Statistical Computation and Simulation, 80(7):761–774, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Martella:2017:FMA

[MA17] Francesca Martella and Marco Alfò. A finite mixture approach to joint clustering of individuals and multivariate discrete outcomes. Journal of Statistical Computation and Simulation, 87(11):2186–2206, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Muttlak:2010:SER

[MAAM10] H. A. Muttlak, S. E. Ahmed, and M. Al-Momani. Shrinkage estimation in replicated median ranked set sampling. Journal of Statistical Computation and Simulation, 80(11):1185– 1196, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Muttlak:2011:CIE

[MADASAM11] H. A. Muttlak, W. A. Abu-Dayyeh, E. Al-Sawi, and M. Al-Momani. Confidence interval estimation of the location and scale parameters of the logistic distribution using pivotal method. Journal of Statistical Computation and Simulation, 81(4):391–409, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mukherjee:2014:RTT

[MAEP14] Amitava Mukherjee, A. M. Abd-Elfattah, and Barendra Pukait. A rule of thumb for testing symmetry about an unknown median against a long right tail. Journal of Statistical Computation and Simulation, 84(10):2138–2155, 2014.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Malec:2016:SRF

[Mal16] Lukás Malec. Some remarks on the functional relation between canonical correlation analysis and partial least squares. Journal of Statistical Computation and Simulation, 86(12):2379–2391, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maansson:2013:DLE

[Mån13] Kristofer Månsson. Developing a Liu estimator for the negative binomial regression model: method and application. Journal of Statistical Computation and Simulation, 83(9):1773–1780, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mansouri:2015:SIB

[Man15] Hossein Mansouri. Simultaneous inference based on rank statistics in linear models. Journal of Statistical Computation and Simulation, 85(4):660–674, 2015. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marques:2010:EEF

[Mar10] G. O. L. C. Marques. Evaluating the efficiency of fractional integration parameter estimators. Journal of Statistical Computation and Simulation, 80(3):301–313, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marozzi:2011:LTT

[Mar11] Marco Marozzi. Levene type tests for the ratio of two scales. Journal of Statistical Computation and Simulation, 81(7): 815–826, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marozzi:2014:TCB

[Mar14a] Marco Marozzi. Testing for concordance between several criteria. Journal of Statistical Computation and Simulation, 84(9):1843–1850, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maruotti:2014:RFH

[Mar14b] Antonello Maruotti. Robust fitting of hidden Markov regression models under a longitudinal setting. Journal of Statistical Computation and Simulation, 84(8):1728–1747, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mousavi:2019:MTC

[MARS19] Seyedjavad Ahmadi Mousavi, Vahid Amirzadeh, Mohsen Rezapour, and Ayob Sheikhy. Multivariate tail conditional expectation for scale mixtures of skew-normal distribution. Journal of Statistical Computation and Simulation, 89(17):3167–3181, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maleki:2017:MPE

[MAV17] Mohsen Maleki and Reinaldo B. Arellano-Valle. Maximum a-posteriori estimation of autoregressive processes based on finite mixtures of scale-mixtures of skew-normal distributions. Journal of Statistical Computation and Simulation, 87(6):1061–1083, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Madsen:2013:SDD

[MB13] L. Madsen and D. Birkes. Simulating dependent discrete data. Journal of Statistical Computation and Simulation, 83(4):677–691, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mameli:2016:MLI

[MB16] V. Mameli and A. R. Brazzale. Modern likelihood inference for the maximum/minimum of a bivariate normal vector. Journal of Statistical Computation and Simulation, 86 (10):1869–1890, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maciel:2017:EPF

[MBG17] Leandro Maciel, Rosangela Ballini, and Fernando Gomide. Evolving possibilistic fuzzy modelling. Journal of Statistical Computation and Simulation, 87(7):1446–1466, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maleki:2019:RCH

[MBKW19] Mohsen Maleki, Zahra Barkhordar, Zahra Khodadadi, and Darren Wraith. A robust class of homoscedastic nonlinear regression models. Journal of Statistical Computation and Simulation, 89(14):2765–2781, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mbogning:2015:JML

[MBL15] Cyprien Mbogning, Kevin Bleakley, and Marc Lavielle. Joint modelling of longitudinal and repeated time-to-event data using nonlinear mixed-effects models and the stochastic approximation expectation-maximization algorithm. Journal of Statistical Computation and Simulation, 85(8):1512–1528, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Massa:2013:ECG

[MC13] M. Sofia Massa and Monica Chiogna. Effectiveness of combinations of Gaussian graphical models for model building. Journal of Statistical Computation and Simulation, 83 (9):1602–1612, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mun:2014:ESC

[MC14] Jungwon Mun and Hyonho Chun. Effective simultaneous confidence bands for repeated measurements in linear mixed-effect models. Journal of Statistical Computation and Simulation, 84(8):1748–1760, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mastrantonio:2016:HMM

[MC16] Gianluca Mastrantonio and Gianfranco Calise. Hidden Markov model for discrete circular-linear wind data time series. Journal of Statistical Computation and Simulation, 86(13):2611–2624, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Martinez-Camblor:2016:CDR

[MCBPF16] Pablo Martínez-Camblor, Gustavo F. Bayón, and Sonia Pérez-Fernández. Cumulative/dynamic ROC curve estimation. Journal of Statistical Computation and Simulation, 86(17):3582–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McCullagh:2014:AAP

[McC14] Peter McCullagh. An asymptotic approximation for the permanent of a doubly stochastic matrix. Journal of Statistical Computation and Simulation, 84(2):404–414, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McFarland:2016:MZA

[McF16] Christopher D. McFarland. A modified ziggurat algorithm for generating exponentially and normally distributed pseudorandom numbers. Journal of Statistical Computation and Simulation, 86(7):1281–1294, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

McLeish:2014:SRV

[McL14] Don McLeish. Simulating random variables using momentgenerating functions and the saddlepoint approximation. Journal of Statistical Computation and Simulation, 84(2): 324–334, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Miao:2017:MRM

[MCW17] Shidi Miao, Deyun Chen, and Tengfei Wang. Modelling and research on manufacturer’s capacity of remanufacturing and recycling in closed-loop supply chain. Journal of Statistical Computation and Simulation, 87(17):3325–3335, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ma:2017:FRF

[MCZ17] Xuejun Ma, Xin Chen, and Jingxiao Zhang. Fast robust feature screening for ultrahigh-dimensional varying coefficient models. Journal of Statistical Computation and Simulation, 87(4):724–732, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mazucheli:2018:BCM

[MD18] J. Mazucheli and Sanku Dey. Bias-corrected maximum likelihood estimation of the parameters of the generalized halfnormal distribution. Journal of Statistical Computation and Simulation, 88(6):1027–1038, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meintanis:2015:VTS

[ME15] Simos G. Meintanis and Jochen Einbeck. Validation tests for semi-parametric models. Journal of Statistical Computation and Simulation, 85(1):131–146, 2015. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Medovikov:2016:NPW

[Med16] Ivan Medovikov. Non-parametric weighted tests for independence based on empirical copula process. Journal of Statistical Computation and Simulation, 86(1):105–121, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mehrdoust:2015:NHM

[Meh15] Farshid Mehrdoust. A new hybrid Monte Carlo simulation for Asian options pricing. Journal of Statistical Computation and Simulation, 85(3):507–516, 2015. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meintanis:2011:TNP

[Mei11] Simos G. Meintanis. Testing for normality with panel data. Journal of Statistical Computation and Simulation, 81(11):1745–1752, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Momeni:2019:EPT

[MES19] R. Momeni, J. Etminan, and M. Khanjari Sadegh. Estimation of parameters in the tree order restriction by a randomized decision. Journal of Statistical Computation and Simulation, 89(11):1986–2005, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Micheas:2016:MER

[MFD16] Athanasios Micheas, Neil Fox, and Rima Dey. Modelling the evolution of a random set utilizing covariate information with applications to nowcasting. Journal of Statistical Computation and Simulation, 86(18):3735–??, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Melo:2014:MLR

[MFP14] Tatiane F. N. Melo, Silvia L. P. Ferrari, and Alexandre G. Patriota. Modified likelihood ratio tests in heteroskedastic multivariate regression models with measurement error. Journal of Statistical Computation and Simulation, 84 (10):2233–2247, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Melo:2017:IHT

[MFP17] Tatiane F. N. Melo, Silvia L. P. Ferrari, and Alexandre G. Patriota. Improved hypothesis testing in a general multivariate elliptical model. Journal of Statistical Computation and Simulation, 87(7):1416–1428, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

MaestreliConsulin:2018:PCV

[MFR^{+}18] Cintia Maestreli Consulin, Damiane Ferreira, Idemauro Antonio Rodrigues de Lara, Antonino De Lorenzo, Laura di Renzo, and Cesar Augusto Taconeli. Performance of coefficient of variation estimators in ranked set sampling. Journal of Statistical Computation and Simulation, 88(2): 221–234, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moller:2015:FSS

[MG15] J. Møller and M. Ghorbani. Functional summary statistics for the Johnson–Mehl model. Journal of Statistical Computation and Simulation, 85(5):899–916, 2015. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Martinez:2017:WBA

[MG17] Gustavo I. Rivas Martínez and M. Dolores Jiménez Gamero. A weighted bootstrap approximation for comparing the error distributions in nonparametric regression. Journal of Statistical Computation and Simulation, 87(18):3503–3520, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mersad:2015:SEZ

[MGR15] M. Mersad, M. Ganjali, and F. Rivaz. Some extensions of zero-inflated models and Bayesian tests for them. Journal of Statistical Computation and Simulation, 85(18):3792–3810, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mitha:2012:MMC

[MH12] Faheem Mitha and Mark L. Huber. Monotonic multigamma coupling for perfect sampling. Journal of Statistical Computation and Simulation, 82(4):603–622, 2012. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Martin:2016:RMZ

[MH16] Jacob Martin and Daniel B. Hall. R^{2} measures for zeroinflated regression models for count data with excess zeros. Journal of Statistical Computation and Simulation, 86(18): 3777–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Munir:2017:NCS

[MH17] Waqas Munir and Abdul Haq. New cumulative sum control charts for monitoring process variability. Journal of Statistical Computation and Simulation, 87(15):2882–2899, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Matthews:2010:ERF

[MHA10] Gregory J. Matthews, Ofer Harel, and Robert H. Aseltine. Examining the robustness of fully synthetic data techniques for data with binary variables. Journal of Statistical Computation and Simulation, 80(6):609–624, 2010. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maram:2016:BDO

[MJ16] Parisa Parsa Maram and Habib Jafari. Bayesian D-optimal design for logistic regression model with exponential distribution for random intercept. Journal of Statistical Computation and Simulation, 86(10):1856–1868, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mantalos:2012:BAD

[MK12] P. Mantalos and A. Karagrigoriou. Bootstrapping the augmented Dickey–Fuller test for unit root using the MDIC.

Journal of Statistical Computation and Simulation, 82(3): 431–443, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mejri:2013:EMC

[MKL13] Dhouha Mejri, Riadh Khanchel, and Mohamed Limam. An ensemble method for concept drift in nonstationary environment. Journal of Statistical Computation and Simulation, 83(6):1115–1128, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Muneer:2018:NIR

[MKSN18] Siraj Muneer, Alamgir Khalil, Javid Shabbir, and Ghulam Narjis. A new improved ratio-product type exponential estimator of finite population variance using auxiliary information. Journal of Statistical Computation and Simulation, 88(16):3179–3192, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Men:2016:BIA

[MKW16] Zhongxian Men, Adam W. Kolkiewicz, and Tony S. Wirjanto. Bayesian inference of asymmetric stochastic conditional duration models. Journal of Statistical Computation and Simulation, 86(7):1295–1319, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marchant:2018:RMC

[MLCL18] Carolina Marchant, Víctor Leiva, Francisco José A. Cysneiros, and Shuangzhe Liu. Robust multivariate control charts based on Birnbaum–Saunders distributions. Journal of Statistical Computation and Simulation, 88(1):182–202, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mahmoud:2013:PMC

[MM13a] M. A. Mahmoud and P. E. Maravelakis. The performance of multivariate CUSUM control charts with estimated parameters. Journal of Statistical Computation and Simulation, 83(4):721–738, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maiti:2013:MLE

[MM13b] Saran Ishika Maiti and Sadhan Samar Maiti. Maximum likelihood estimation of the linearly structured correlation matrix by a Jacobi-type iterative scheme. Journal of Statistical Computation and Simulation, 83(5):896–921, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mojirsheibani:2015:ACR

[MM15] Majid Mojirsheibani and Zahra Montazeri. Aggregating classifiers via Rademacher–Walsh polynomials. Journal of Statistical Computation and Simulation, 85(6):1187–1199, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mantalos:2010:FAM

[MMK10] P. Mantalos, K. Mattheou, and A. Karagrigoriou. Forecasting ARMA models: a comparative study of information criteria focusing on MDIC. Journal of Statistical Computation and Simulation, 80(1):61–73, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Matsui:2014:MSC

[MMK14] Hidetoshi Matsui, Toshihiro Misumi, and Shuichi Kawano. Model selection criteria for the varying-coefficient modelling via regularized basis expansions. Journal of Statistical Computation and Simulation, 84(10):2156–2165, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moreira:2012:CLL

[MMP12] Elsa Moreira, João Tiago Mexia, and Luís Santos Pereira. Clustering of log-linear models using LRT p-values to assess homogeneous regions relative to drought class transitions. Journal of Statistical Computation and Simulation, 82(2): 293–308, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Martin:2015:CTT

[MMP15] Nirian Martín, Raquel Mata, and Leandro Pardo. Comparing two treatments in terms of the likelihood ratio order. Journal of Statistical Computation and Simulation, 85 (17):3512–3534, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mahmoudi:2019:CCC

[MMP19] Mohammad Reza Mahmoudi, Marziyeh Mahmoudi, and Abbas Pak. On comparing, classifying and clustering several dependent regression models. Journal of Statistical Computation and Simulation, 89(12):2280–2292, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Malela-Majika:2016:DFC

[MMR16] Jean-Claude Malela-Majika and Eeva Rapoo. Distributionfree cumulative sum and exponentially weighted moving average control charts based on the Wilcoxon rank-sum statistic using ranked set sampling for monitoring mean shifts. Journal of Statistical Computation and Simulation, 86(18): 3715–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Murray:2016:FFM

[MMT16] K. Murray, S. Müller, and B. A. Turlach. Fast and flexible methods for monotone polynomial fitting. Journal of Statistical Computation and Simulation, 86(15):2946–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mateu:2015:MAL

[MN15] Jorge Mateu and Orietta Nicolis. Multiresolution analysis of linearly oriented spatial point patterns. Journal of Statistical Computation and Simulation, 85(3):621–637, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Manouchehri:2019:EPP

[MN19] T. Manouchehri and A. R. Nematollahi. On the estimation problem of periodic autoregressive time series: symmetric and asymmetric innovations. Journal of Statistical Computation and Simulation, 89(1):71–97, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moghimbeigi:2011:STE

[Mog11] Abbas Moghimbeigi. A score test for extra zeros in negative binomial mixed models. Journal of Statistical Computation and Simulation, 81(5):635–644, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mohammadi:2017:PCI

[Moh17] Mohammad Mohammadi. Performance of confidence intervals for the population size in capture-recapture experiment under inverse sampling with replacement. Journal of Statistical Computation and Simulation, 87(5):984–992, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moiseev:2017:FTS

[Moi17a] Nikita A. Moiseev. Forecasting time series of economic processes by model averaging across data frames of various lengths. Journal of Statistical Computation and Simulation, 87(16):3111–3131, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Moiseev:2017:VAC

[Moi17b] Nikita A. Moiseev. p-Value adjustment to control Type I errors in linear regression models. Journal of Statistical Computation and Simulation, 87(9):1701–1711, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mbah:2015:SFT

[MP15] Alfred K. Mbah and Arnut Paothong. Shapiro–Francia test compared to other normality test using expected pvalue. Journal of Statistical Computation and Simulation, 85(15):3002–3016, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Malik:2016:NFS

[MP16] Waqas Ahmed Malik and Hans-Peter Piepho. On a new family of shifted logarithmic transformations. Journal of Statistical Computation and Simulation, 86(9):1697–1708, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mohebbi:2019:NDA

[MPB19] Shima Mohebbi, Esra Pamukcu, and Hamparsum Bozdogan. A new data adaptive elastic net predictive model using hybridized smoothed covariance estimators with information complexity. Journal of Statistical Computation and Simulation, 89(6):1060–1089, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mehmood:2018:PXC

[MQR18] Rashid Mehmood, Miss Sara Qazi, and Muhammad Riaz. On the performance of X̄ control chart for known and unknown parameters supplemented with runs rules under different probability distributions. Journal of Statistical Computation and Simulation, 88(4):675–711, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Muller:2017:MED

[MR17] Klaus Müller and Wolf-Dieter Richter. Modelling extremal data. Journal of Statistical Computation and Simulation, 87(5):933–955, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marin:2015:DCE

[MRBR15] J. M. Marín, M. T. Rodríguez-Bernal, and E. Romero. Data cloning estimation of GARCH and COGARCH models. Journal of Statistical Computation and Simulation, 85 (9):1818–1831, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mai:2011:RMO

[MS11] Jan-Frederik Mai and Matthias Scherer. Reparameterizing Marshall–Olkin copulas with applications to sampling. Journal of Statistical Computation and Simulation, 81(1): 59–78, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mehrdoust:2015:PAA

[MS15] Farshid Mehrdoust and Naghmeh Saber. Pricing arithmetic Asian option under a two-factor stochastic volatility model with jumps. Journal of Statistical Computation and Simulation, 85(18):3811–3819, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mehreyan:2017:SHT

[MS17] Sedigheh Zamani Mehreyan and Abdolreza Sayyareh. Separated hypotheses testing for autoregressive models with nonnegative residuals. Journal of Statistical Computation and Simulation, 87(4):689–711, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Majd:2018:RED

[MS18a] M. H. Hashemi Majd and R. Aliakbari Saba. Robust extreme double ranked set sampling. Journal of Statistical Computation and Simulation, 88(9):1749–1758, 2018. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mousavi:2018:FLR

[MS18b] Seyed Nourollah Mousavi and Helle Sørensen. Functional logistic regression: a comparison of three methods. Journal of Statistical Computation and Simulation, 88(2):250–268, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Morales:2019:SAE

[MS19] Domingo Morales and Laureano Santamaría. Small area estimation under unit-level temporal linear mixed models. Journal of Statistical Computation and Simulation, 89 (9):1592–1620, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Muggeo:2012:QRI

[MSA12] Vito M. R. Muggeo, Mariangela Sciandra, and Luigi Augugliaro. Quantile regression via iterative least squares computations. Journal of Statistical Computation and Simulation, 82(11):1557–1569, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mun:2014:RTA

[MSK14] Hyeong-Ho Mun, Eun-Young Shim, and Tae-Hwan Kim. A robust test for autocorrelation in the presence of a structural break in variance. Journal of Statistical Computation and Simulation, 84(7):1552–1562, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mao:2014:EIC

[MSS14] Song Mao, Yi-Min Shi, and Yu-Dong Sun. Exact inference for competing risks model with generalized type-I hybrid censored exponential data. Journal of Statistical Computation and Simulation, 84(11):2506–2521, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mohamed:2018:FIS

[MSS18] Choukri Mohamed, Stephen A. Sedory, and Sarjinder Singh. A fresh imputing survey methodology using sensible constraints on study and auxiliary variables: dubious random non-response. Journal of Statistical Computation and Simulation, 88(7):1273–1294, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Menardi:2013:RDD

[MT13] Giovanna Menardi and Nicola Torelli. Reducing data dimension for cluster detection. Journal of Statistical Computation and Simulation, 83(11):2047–2063, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meng:2017:BPB

[MT17] Lixin Meng and Jian Tao. Bayesian penalized B-spline estimation approach for epidemic models. Journal of Statistical Computation and Simulation, 87(1):88–99, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meng:2014:IRM

[MTS14] Xiang-Bin Meng, Jian Tao, and Ning-Zhong Shi. An item response model for Likert-type data that incorporates response time in personality measurements. Journal of Statistical Computation and Simulation, 84(1):1–21, 2014. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maurya:2019:PCI

[MTSR19] Raj Kamal Maurya, Yogesh Mani Tripathi, Tanmay Sen, and Manoj Kumar Rastogi. On progressively censored inverted exponentiated Rayleigh distribution. Journal of Statistical Computation and Simulation, 89(3):492–518, 2019.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Muggeo:2016:TNP

[Mug16] Vito M. R. Muggeo. Testing with a nuisance parameter present only under the alternative: a score-based approach with application to segmented modelling. Journal of Statistical Computation and Simulation, 86(15):3059–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Muhammed:2016:BIW

[Muh16] Hiba Z. Muhammed. Bivariate inverse Weibull distribution. Journal of Statistical Computation and Simulation, 86 (12):2335–2345, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Murakami:2012:MBS

[Mur12] Hidetoshi Murakami. Modified Baumgartner statistics for the two-sample and multisample problems: a numerical comparison. Journal of Statistical Computation and Simulation, 82(5):711–728, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Murakami:2015:RBS

[Mur15] Hidetoshi Murakami. A randomized Baumgartner statistic for multivariate two-sample testing hypothesis. Journal of Statistical Computation and Simulation, 85(1):189–201, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Meligkotsidou:2014:DSB

[MV14] Loukia Meligkotsidou and Ioannis D. Vrontos. Detecting structural breaks in multivariate financial time series: evidence from hedge fund investment strategies. Journal of Statistical Computation and Simulation, 84(5):1115–1135, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Micheas:2014:RSM

[MWL14] Athanasios C. Micheas, Christopher K. Wikle, and David R. Larsen. Random set modelling of three-dimensional objects in a hierarchical Bayesian context. Journal of Statistical Computation and Simulation, 84(1):107–123, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mendes:2013:CAA

[MY13] Mehmet Mendes and Soner Yigit. Comparison of ANOVA-F and ANOM tests with regard to Type I error rate and test power. Journal of Statistical Computation and Simulation, 83(11):2093–2104, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marino:2019:HDS

[MZZ^{+}19] Simeone Marino, Nina Zhou, Yi Zhao, Lu Wang, Qiucheng Wu, and Ivo D. Dinov. HDDA: DataSifter: statistical obfuscation of electronic health records and other sensitive datasets. Journal of Statistical Computation and Simulation, 89(2):249–271, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2011:MCCb

[NA11a] Hadi Alizadeh Noughabi and Naser Reza Arghami. Monte Carlo comparison of five exponentiality tests using different entropy estimates. Journal of Statistical Computation and Simulation, 81(11):1579–1592, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2011:MCCa

[NA11b] Hadi Alizadeh Noughabi and Naser Reza Arghami. Monte Carlo comparison of seven normality tests. Journal of Statistical Computation and Simulation, 81(8):965–972, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2011:TEB

[NA11c] Hadi Alizadeh Noughabi and Naser Reza Arghami. Testing exponentiality based on characterizations of the exponential distribution. Journal of Statistical Computation and Simulation, 81(11):1641–1651, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2011:TEU

[NA11d] Hadi Alizadeh Noughabi and Naser Reza Arghami. Testing exponentiality using transformed data. Journal of Statistical Computation and Simulation, 81(4):511–516, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2013:GTG

[NA13] Hadi Alizadeh Noughabi and Naser Reza Arghami. General treatment of goodness-of-fit tests based on Kullback–Leibler information. Journal of Statistical Computation and Simulation, 83(8):1556–1569, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Norouzirad:2017:IRR

[NAA17] M. Norouzirad, M. Arashi, and S. E. Ahmed. Improved robust ridge M-estimation. Journal of Statistical Computation and Simulation, 87(18):3469–3490, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nikzad:2018:ETS

[NAA18] Erfaneh Nikzad, Amirhossein Amiri, and Farzad Amirkhani. Estimating total and specific process capability indices in three-stage processes with measurement errors. Journal of Statistical Computation and Simulation, 88(15):3033–3064, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadarajah:2010:SPR

[Nad10] Saralees Nadarajah. Sum, product and ratio of Pareto and gamma variables. Journal of Statistical Computation and Simulation, 80(10):1071–1082, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nagatsuka:2013:CME

[NB13a] Hideki Nagatsuka and N. Balakrishnan. A consistent method of estimation for the parameters of the threeparameter inverse Gaussian distribution. Journal of Statistical Computation and Simulation, 83(10):1915–1931, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nagatsuka:2013:PQE

[NB13b] Hideki Nagatsuka and N. Balakrishnan. Parameter and quantile estimation for the three-parameter lognormal distribution based on statistics invariant to unknown location. Journal of Statistical Computation and Simulation, 83(9):1629–1647, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nagatsuka:2014:MEP

[NB14] Hideki Nagatsuka and N. Balakrishnan. A method for estimating parameters and quantiles of the three-parameter inverse Gaussian distribution based on statistics invariant to unknown location. Journal of Statistical Computation and Simulation, 84(11):2361–2377, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nagatsuka:2015:CEP

[NB15a] Hideki Nagatsuka and N. Balakrishnan. Consistent estimation of parameters and quantiles of the three-parameter gamma distribution based on Type-II right-censored data. Journal of Statistical Computation and Simulation, 85(12): 2406–2424, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nesterko:2015:BVB

[NB15b] Sergiy Nesterko and Joseph Blitzstein. Bias-variance and breadth-depth tradeoffs in respondent-driven sampling. Journal of Statistical Computation and Simulation, 85(1): 89–102, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nagatsuka:2016:EUC

[NB16] Hideki Nagatsuka and N. Balakrishnan. Existence, uniqueness and consistency of estimation of life characteristics of three-parameter Weibull distribution based on Type-II right censored data. Journal of Statistical Computation and Simulation, 86(6):1248–1279, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1060236.

Neville:2018:MCT

[NB18] Zachariah Neville and Naomi C. Brownstein. Macros to conduct tests of multimodality in SAS. Journal of Statistical Computation and Simulation, 88(17):3269–3290, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nandram:2015:LRT

[NBB15] Balgobin Nandram, Dilli Bhatta, and Dhiman Bhadra. A likelihood ratio test of quasi-independence for sparse twoway contingency tables. Journal of Statistical Computation and Simulation, 85(2):284–304, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadarajah:2011:GRB

[NCO11] Saralees Nadarajah, Gauss M. Cordeiro, and Edwin M. M. Ortega. General results for the beta-modified Weibull distribution. Journal of Statistical Computation and Simulation, 81(10):1211–1232, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadarajah:2012:GRK

[NCO12] Saralees Nadarajah, Gauss M. Cordeiro, and Edwin M. M. Ortega. General results for the Kumaraswamy-G distribution. Journal of Statistical Computation and Simulation, 82(7):951–979, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadarajah:2015:EGF

[NCO15] Saralees Nadarajah, Gauss M. Cordeiro, and Edwin M. M. Ortega. The exponentiated G geometric family of distributions. Journal of Statistical Computation and Simulation, 85(8):1634–1650, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nofuentes:2015:AKC

[NdCOP15] José Antonio Roldán Nofuentes and María del Carmen Olvera Porcel. Average kappa coefficient: a new measure to assess a binary test considering the losses associated with an erroneous classification. Journal of Statistical Computation and Simulation, 85(8):1601–1620, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Neddermeyer:2011:NPF

[Ned11] Jan C. Neddermeyer. Nonparametric particle filtering and smoothing with quasi-Monte Carlo sampling. Journal of Statistical Computation and Simulation, 81(11):1361–1379, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nunes:2012:CTE

[NFFM12] Célia Nunes, Dário Ferreira, Sandra S. Ferreira, and João T. Mexia. Control of the truncation errors for generalized F distributions. Journal of Statistical Computation and Simulation, 82(2):165–171, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nunes:2014:FEA

[NFFM14] Célia Nunes, Dário Ferreira, Sandra S. Ferreira, and João T. Mexia. Fixed effects ANOVA: an extension to samples with random size. Journal of Statistical Computation and Simulation, 84(11):2316–2328, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nabil:2016:CSD

[NG16] M. Nabil and M. Golalizadeh. On clustering shape data. Journal of Statistical Computation and Simulation, 86(15): 2995–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Niu:2014:CSB

[NGXZ14] Cuizhen Niu, Xu Guo, Wangli Xu, and Lixing Zhu. Comparison of several Birnbaum–Saunders distributions. Journal of Statistical Computation and Simulation, 84(12):2721–2733, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Norouzirad:2018:SPE

[NHA18] M. Norouzirad, S. Hossain, and M. Arashi. Shrinkage and penalized estimators in weighted least absolute deviations regression models. Journal of Statistical Computation and Simulation, 88(8):1557–1575, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Neuhauser:2014:RLS

[NHGS14] David Neuhäuser, Christian Hirsch, Catherine Gloaguen, and Volker Schmidt. Ratio limits and simulation algorithms for the Palm version of stationary iterated tessellations. Journal of Statistical Computation and Simulation, 84(7):1486–1504, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nie:2014:SSF

[NHWT14] Ningming Nie, Jianfei Huang, Wenjia Wang, and Yifa Tang. Solving spatial-fractional partial differential diffusion equations by spectral method. Journal of Statistical Computation and Simulation, 84(6):1173–1189, 2014. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nesterkin:2013:RVS

[NJ13] Dmitriy A. Nesterkin and Thomas W. Jones. Random vs. systematic non-response in group-level research. Journal of Statistical Computation and Simulation, 83(1):139–144, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2018:MNP

[NJ18] Hadi Alizadeh Noughabi and Jalil Jarrahiferiz. Moments of nonparametric probability density functions of entropy estimators applied to testing the inverse Gaussian distribution. Journal of Statistical Computation and Simulation, 88 (16):3217–3229, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nofuentes:2014:AHT

[NJdC14] J. A. Roldán Nofuentes, A. E. Marín Jiménez, and J. D. Luna del Castillo. Asymptotic hypothesis test to simultaneously compare the weighted kappa coefficients of multiple binary diagnostic tests in the presence of ignorable missing data. Journal of Statistical Computation and Simulation, 84(2):273–289, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadarajah:2013:NFL

[NJR13] Saralees Nadarajah, K. Jayakumar, and Miroslav M. Ristić. A new family of lifetime models. Journal of Statistical Computation and Simulation, 83(8):1389–1404, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadar:2015:EPB

[NK15] Mustafa Nadar and Fatih Kizilaslan. Estimation and prediction of the Burr type XII distribution based on record values and inter-record times. Journal of Statistical Computation and Simulation, 85(16):3297–3321, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nguelifack:2019:RRB

[NKB19] Brice M. Nguelifack and Isabelle Kemajou-Brown. Robust rank-based variable selection in double generalized linear models with diverging number of parameters under adaptive Lasso. Journal of Statistical Computation and Simulation, 89(11):2051–2072, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadar:2014:CBE

[NKP14] Mustafa Nadar, Fatih Kizilaslan, and Alexander Papadopoulos. Classical and Bayesian estimation of P (Y < X) for Kumaraswamy’s distribution. Journal of Statistical Computation and Simulation, 84(7):1505–1529, 2014. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nandram:2019:PBT

[NKZ19] Balgobin Nandram, Dalho Kim, and Jingran Zhou. A pooled Bayes test of independence for sparse contingency tables from small areas. Journal of Statistical Computation and Simulation, 89(5):899–926, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ng:2012:PET

[NLHD12] H. K. T. Ng, L. Luo, Y. Hu, and F. Duan. Parameter estimation of three-parameter Weibull distribution based on progressively Type-II censored samples. Journal of Statistical Computation and Simulation, 82(11):1661–1678, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noh:2011:RED

[NLK11] Manegseok Noh, Youngjo Lee, and Michael G. Kenward. Robust estimation of dropout models using hierarchical likelihood. Journal of Statistical Computation and Simulation, 81(6):693–706, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noh:2019:HLA

[NLOT19] Maengseok Noh, Youngjo Lee, Johan H. L. Oud, and Toni Toharudin. Hierarchical likelihood approach to non-Gaussian factor analysis. Journal of Statistical Computation and Simulation, 89(9):1555–1573, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nooraee:2018:SHM

[NMOV18] Nazanin Nooraee, Geert Molenberghs, Johan Ormel, and Edwin R. Van den Heuvel. Strategies for handling missing data in longitudinal studies with questionnaires. Journal of Statistical Computation and Simulation, 88(17):3415–3436, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Naranjo:2014:AMB

[NMPR14] L. Naranjo, J. Martín, C. J. Pérez, and M. J. Rufo. Addressing misclassification for binary data: probit and t-link regressions. Journal of Statistical Computation and Simulation, 84(10):2187–2213, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Naderi:2018:ESD

[NMS18] M. H. Naderi, M. Bameni Moghadam, and A. Seif. Economic statistical design of the T^{2} control chart under the Weibull shock model with multiple assignable causes. Journal of Statistical Computation and Simulation, 88(1):1–27, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2013:EE

[NN13] Havva Alizadeh Noughabi and Reza Alizadeh Noughabi. On the entropy estimators. Journal of Statistical Computation and Simulation, 83(4):784–792, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2014:ETR

[NNB14] Reza Alizadeh Noughabi, Havva Alizadeh Noughabi, and Alireza Ebrahimi Moghaddam Behabadi. An entropy test for the Rayleigh distribution and power comparison. Journal of Statistical Computation and Simulation, 84(1):151– 158, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Namba:2010:RPP

[NO10] Akio Namba and Kazuhiro Ohtani. Risk performance of a pre-test ridge regression estimator under the LINEX loss function when each individual regression coefficient is estimated. Journal of Statistical Computation and Simulation, 80(3):255–262, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nomura:2014:CMN

[Nom14] Noboru Nomura. Computation of multivariate normal probabilities with polar coordinate systems. Journal of Statistical Computation and Simulation, 84(3):491–512, 2014. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nooghabi:2019:DOP

[Noo19] Mehdi Jabbari Nooghabi. On detecting outliers in the Pareto distribution. Journal of Statistical Computation and Simulation, 89(8):1466–1481, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2010:NEE

[Nou10] Hadi Alizadeh Noughabi. A new estimator of entropy and its application in testing normality. Journal of Statistical Computation and Simulation, 80(10):1151–1162, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2017:ERS

[Nou17a] Hadi Alizadeh Noughabi. Efficiency of ranked set sampling in tests for normality. Journal of Statistical Computation and Simulation, 87(5):956–965, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2017:GFT

[Nou17b] Hadi Alizadeh Noughabi. Goodness-of-fit tests for lifetime distributions based on Type II censored data. Journal of Statistical Computation and Simulation, 87(9):1787–1798, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2019:NEK

[Nou19] Hadi Alizadeh Noughabi. A new estimator of Kullback– Leibler information and its application in goodness of fit tests. Journal of Statistical Computation and Simulation, 89(10):1914–1934, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Noughabi:2016:TFL

[NP16] Hadi Alizadeh Noughabi and Sangun Park. Tests of fit for the Laplace distribution based on correcting moments of entropy estimators. Journal of Statistical Computation and Simulation, 86(11):2165–2181, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nguyen:2014:FMB

[NPJ14] Thuan Nguyen, Jie Peng, and Jiming Jiang. Fence methods for backcross experiments. Journal of Statistical Computation and Simulation, 84(3):644–662, 2014. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nezhad:2017:RGS

[NS17] Mohammad Saber Fallah Nezhad and Sina Seifi. Repetitive group sampling plan based on the process capability index for the lot acceptance problem. Journal of Statistical Computation and Simulation, 87(1):29–41, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ning:2018:NBA

[NS18] Shaoyang Ning and Neil Shephard. A nonparametric Bayesian approach to copula estimation. Journal of Statistical Computation and Simulation, 88(6):1081–1105, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nematollahi:2015:TDP

[NSMFR15] A. R. Nematollahi, A. R. Soltani, H. A. Mardani-Fard, and E. Razmi. Typical decision problems in the first-order autoregressive time series. Journal of Statistical Computation and Simulation, 85(14):2919–2935, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadarajah:2014:NFP

[NSR14] Saralees Nadarajah, Fatemeh Shahsanaei, and Sadegh Rezaei. A new four-parameter lifetime distribution. Journal of Statistical Computation and Simulation, 84(2):248–263, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nandram:2011:BBS

[NTC11] Balgobin Nandram, Ma. Criselda S. Toto, and Jai Won Choi. A Bayesian benchmarking of the Scott–Smith model for small areas. Journal of Statistical Computation and Simulation, 81(11):1593–1608, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nguyen:2019:OSS

[NTC^{+}19] Quoc Thong Nguyen, Kim Phuc Tran, Philippe Castagliola, Giovanni Celano, and Salim Lardjane. One-sided synthetic control charts for monitoring the multivariate coefficient of variation. Journal of Statistical Computation and Simulation, 89(10):1841–1862, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Niaki:2013:EDV

[NTG13] Seyed Taghi Akhavan Niaki, Moslem Toosheghanian, and Fazlollah Masoumi Gazaneh. Economic design of VSI X^{−} control chart with correlated non-normal data under multiple assignable causes. Journal of Statistical Computation and Simulation, 83(7):1279–1300, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nadeb:2019:SSR

[NTZ19] Hossein Nadeb, Hamzeh Torabi, and Yichuan Zhao. Stressstrength reliability of exponentiated Fréchet distributions based on Type-II censored data. Journal of Statistical Computation and Simulation, 89(10):1863–1876, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Namba:2018:PDP

[NX18a] Akio Namba and Haifeng Xu. PMSE dominance of the positive-part shrinkage estimator in a regression model with proxy variables. Journal of Statistical Computation and Simulation, 88(15):2893–2908, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Namba:2018:SCM

[NX18b] Akio Namba and Haifeng Xu. A sufficient condition for the MSE dominance of the positive-part shrinkage estimator when each individual regression coefficient is estimated in a misspecified linear regression model. Journal of Statistical Computation and Simulation, 88(11):2034–2047, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Nandram:2016:NBP

[NY16] Balgobin Nandram and Jiani Yin. A nonparametric Bayesian prediction interval for a finite population mean. Journal of Statistical Computation and Simulation, 86(16): 3141–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ortega:2012:LEG

[OCPC12] Edwin M. M. Ortega, Gauss M. Cordeiro, Marcelino A. R. Pascoa, and Epaminondas V. Couto. The log-exponentiated generalized gamma regression model for censored data. Journal of Statistical Computation and Simulation, 82(8): 1169–1189, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ong:2018:ESU

[OCT18] Meng Hwee Victor Ong, Sanjay Chaudhuri, and Berwin A. Turlach. Edge selection for undirected graphs. Journal of Statistical Computation and Simulation, 88(17):3291–3322, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ossima:2015:GLR

[ODBT15] Arnaud D. Nze Ossima, Jean-Pierre Daurès, Faiza Bessaoud, and Brigitte Trétarre. The generalized Lehmann ROC curves: Lehmann family of ROC surfaces. Journal of Statistical Computation and Simulation, 85(3):596–607, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ozel:2010:PFG

[ÖI10] Gamze Özel and Ceyhan Inal. The probability function of a geometric Poisson distribution. Journal of Statistical Computation and Simulation, 80(5):479–487, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Olmo-Jimenez:2018:RCD

[OJRACL18] María José Olmo-Jiménez, José Rodríguez-Avi, and Valentina Cueva-López. A review of the CTP distribution: a comparison with other over- and underdispersed count data models. Journal of Statistical Computation and Simulation, 88(14):2684–2706, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ozbay:2017:ATP

[ÖK17] Nimet Özbay and Selahattin Kaçiranlar. The Almon two parameter estimator for the distributed lag models. Journal of Statistical Computation and Simulation, 87(4):834–843, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ozbay:2018:ELR

[ÖK18] Nimet Özbay and Selahattin Kaçiranlar. Estimation in a linear regression model with stochastic linear restrictions: a new two-parameter-weighted mixed estimator. Journal of Statistical Computation and Simulation, 88(9):1669–1683, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ozkale:2019:AJR

[ÖK19] M. Revan Özkale and Özge Kuran. Adaptation of the jackknifed ridge methods to the linear mixed models. Journal of Statistical Computation and Simulation, 89(18):3413–3452, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ozbay:2017:FGR

[ÖKD17] Nimet Özbay, Selahattin Kaçiranlar, and Issam Dawoud. The feasible generalized restricted ridge regression estimator. Journal of Statistical Computation and Simulation, 87(4):753–765, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ounpraseuth:2012:ITI

[OMY12] Songthip Ounpraseuth, Page C. Moore, and Dean M. Young. Imputation techniques for incomplete data in quadratic discriminant analysis. Journal of Statistical Computation and Simulation, 82(6):863–877, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Okeniyi:2012:IKS

[OO12] Joshua Olusegun Okeniyi and Elizabeth Toyin Okeniyi. Implementation of Kolmogorov–Smirnov P -value computation in Visual Basic(R): implication for Microsoft Excel(R) library function. Journal of Statistical Computation and Simulation, 82(12):1727–1741, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Oulhaj:2014:GDI

[OS14] Abderrahim Oulhaj and Ernesto San Martin. Generating data from improper distributions: application to Cox proportional hazards models with cure. Journal of Statistical Computation and Simulation, 84(1):204–214, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ortega-Serrano:2013:SDE

[OSdVM13] I. Ortega-Serrano, M. C. Ruiz de Villa, and A. Miñarro. Searching for differential expression: a non-parametric approach. Journal of Statistical Computation and Simulation, 83(9):1661–1670, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ommen:2017:CMC

[OSN17] Danica M. Ommen, Christopher P. Saunders, and Cedric Neumann. The characterization of Monte Carlo errors for the quantification of the value of forensic evidence. Journal of Statistical Computation and Simulation, 87(8):1608– 1643, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Olieman:2010:EMM

[OvP10] Niels J. Olieman and Bram van Putten. Estimation method of multivariate exponential probabilities based on a simplex coordinates transform. Journal of Statistical Computation and Simulation, 80(3):355–361, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ou:2015:RSP

[OWKC15] Yanjing Ou, Zhang Wu, Michael B. C. Khoo, and Nan Chen. A rational sequential probability ratio test control chart for monitoring process shifts in mean and variance. Journal of Statistical Computation and Simulation, 85(9):1765–1781, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Oluyede:2016:NCC

[OWLP16] Broderick O. Oluyede, Gayan Warahena-Liyanage, and Mavis Pararai. A new compound class of log-logistic Weibull–Poisson distribution: model, properties and applications. Journal of Statistical Computation and Simulation, 86(7):1363–1391, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Oluyede:2015:NCG

[OY15] Broderick O. Oluyede and Tiantian Yang. A new class of generalized Lindley distributions with applications. Journal of Statistical Computation and Simulation, 85(10):2072– 2100, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ozgul:2018:NCE

[Özg18] Nilgün Özgül. New calibration estimator in stratified sampling. Journal of Statistical Computation and Simulation, 88(13):2561–2572, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ozkale:2012:CUE

[Özk12] M. Revan Özkale. Combining the unrestricted estimators into a single estimator and a simulation study on the unrestricted estimators. Journal of Statistical Computation and Simulation, 82(5):653–688, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ozkale:2019:CEG

[Özk19] M. Revan Özkale. The r–d class estimator in generalized linear models: applications on gamma, Poisson and binomial distributed responses. Journal of Statistical Computation and Simulation, 89(4):615–640, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pardo:2014:GRC

[PA14] M. C. Pardo and R. Alonso. GEEs for repeated categorical responses based on generalized residuals. Journal of Statistical Computation and Simulation, 84(2):344–359, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Paul:2015:TET

[PA15] Sudhir Paul and Khurshid Alam. Testing equality of two negative binomial means in presence of unequal overdispersion parameters: a Behrens–Fisher problem analog. Journal of Statistical Computation and Simulation, 85(15): 3140–3153, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pires:2012:CSL

[PAFPM12] J. C. M. Pires, M. C. M. Alvim-Ferraz, M. C. Pereira, and F. G. Martins. Comparison of several linear statistical models to predict tropospheric ozone concentrations. Journal of Statistical Computation and Simulation, 82(2):183–192, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pakyari:2010:DBG

[Pak10] Reza Pakyari. Discriminating between generalized exponential, geometric extreme exponential and Weibull distributions. Journal of Statistical Computation and Simulation, 80(12):1403–1412, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See erratum [Pak11].

Pakyari:2011:EDB

[Pak11] Reza Pakyari. [Erratum]: Discriminating between generalized exponential, geometric extreme exponential and Weibull distributions [J. Statist. Comput. Simul. 80 (2010), pp. 1403–1412]. Journal of Statistical Computation and Simulation, 81(7):929, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Pak10].

Pamukcu:2019:COH

[Pam19] Esra Pamukcu. Choosing the optimal hybrid covariance estimators in adaptive elastic net regression models using information complexity. Journal of Statistical Computation and Simulation, 89(16):2983–2996, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2011:DRT

[Par11] Jin-Hong Park. Dimension reduction transfer function model. Journal of Statistical Computation and Simulation, 81(12):2131–2140, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2012:ANT

[Par12] Jin-Hong Park. Analysing nonlinear time series with central subspace. Journal of Statistical Computation and Simulation, 82(7):1073–1084, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2017:ORH

[Par17] Heewon Park. Outlier-resistant high-dimensional regression modelling based on distribution-free outlier detection and tuning parameter selection. Journal of Statistical Computation and Simulation, 87(9):1799–1812, 2017. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2012:VFH

[PB12] Sangun Park and N. Balakrishnan. A very flexible hybrid censoring scheme and its Fisher information. Journal of Statistical Computation and Simulation, 82(1):41–50, 2012.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pakyari:2013:GFT

[PB13a] Reza Pakyari and N. Balakrishnan. Goodness-of-fit tests for progressively Type-II censored data from location-scale distributions. Journal of Statistical Computation and Simulation, 83(1):167–178, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pakyari:2013:TEB

[PB13b] Reza Pakyari and N. Balakrishnan. Testing exponentiality based on Type-I censored data. Journal of Statistical Computation and Simulation, 83(12):2369–2378, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Price:2015:CTV

[PB15] Robert M. Price, Jr. and Douglas G. Bonett. A comparison of two varying coefficient meta-analysis methods for an average risk difference. Journal of Statistical Computation and Simulation, 85(12):2378–2383, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pal:2017:ETE

[PB17] Suvra Pal and N. Balakrishnan. An EM type estimation procedure for the destructive exponentially weighted Poisson regression cure model under generalized gamma lifetime. Journal of Statistical Computation and Simulation, 87(6):1107–1129, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pokhriyal:2019:BPI

[PB19] H. Pokhriyal and N. Balakrishna. Bootstrap prediction intervals for autoregressive conditional duration models. Journal of Statistical Computation and Simulation, 89(15):2930– 2950, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Punzo:2016:HTM

[PBM16] Antonio Punzo, Ryan P. Browne, and Paul D. McNicholas. Hypothesis testing for mixture model selection. Journal of Statistical Computation and Simulation, 86(14):2797–2818, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pan:2013:BDA

[PBSZ13] Zhengqiang Pan, Narayanaswamy Balakrishnan, Quan Sun, and Jinglun Zhou. Bivariate degradation analysis of products based on Wiener processes and copulas. Journal of Statistical Computation and Simulation, 83(7):1316–1329, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pereira:2010:ADU

[PC10] L. N. Pereira and P. S. Coelho. Assessing different uncertainty measures of EBLUP: a resampling-based approach. Journal of Statistical Computation and Simulation, 80(7): 713–727, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pan:2011:NRE

[PC11] Jeh-Nan Pan and Sheau-Chiann Chen. New robust estimators for detecting non-random patterns in multivariate control charts: a simulation approach. Journal of Statistical Computation and Simulation, 81(3):289–300, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Perumean-Chaney:2013:ZIO

[PCMMA13] Suzanne E. Perumean-Chaney, Charity Morgan, David Mc-Dowall, and Inmaculada Aban. Zero-inflated and overdispersed: what’s one to do? Journal of Statistical Computation and Simulation, 83(9):1671–1683, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pereira:2014:MLR

[PCN14] Tarciana L. Pereira and Francisco Cribari-Neto. Modified likelihood ratio statistics for inflated beta regressions. Journal of Statistical Computation and Simulation, 84(5):982– 998, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Petralias:2013:MMS

[PD13] Athanassios Petralias and Petros Dellaportas. An MCMC model search algorithm for regression problems. Journal of Statistical Computation and Simulation, 83(9):1722–1740, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Perera:2016:OLT

[PDS16] Harsha Perera, Jack Davis, and Tim B. Swartz. Optimal lineups in Twenty20 cricket. Journal of Statistical Computation and Simulation, 86(14):2888–2900, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Perakis:2010:EDB

[Per10] Michael Perakis. Estimation of differences between process capability indices C_{pm} or C_{pmk} for two processes. Journal of Statistical Computation and Simulation, 80(3):315–334, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Picchini:2016:AID

[PF16a] Umberto Picchini and Julie Lyng Forman. Accelerating inference for diffusions observed with measurement error and large sample sizes using approximate Bayesian computation. Journal of Statistical Computation and Simulation, 86(1): 195–213, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pinheiro:2016:CRG

[PF16b] Eliane C. Pinheiro and Silvia L. P. Ferrari. A comparative review of generalizations of the Gumbel extreme value distribution with an application to wind speed data. Journal of Statistical Computation and Simulation, 86(11):2241–2261, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Perez-Gonzalez:2019:ACS

[PGF19] Carlos J. Pérez-González and Arturo J. Fernández. Accurate computation of single and product moments of order statistics under progressive censoring. Journal of Statistical Computation and Simulation, 89(13):2489–2504, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pham-Gia:2011:TES

[PGT11] T. Pham-Gia and N. Turkkan. Testing the equality of several covariance matrices. Journal of Statistical Computation and Simulation, 81(10):1233–1246, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pearn:2011:EPT

[PHCS11] W. L. Pearn, H. N. Hung, Y. S. Chuang, and R. H. Su. An effective powerful test for one-sided supplier selection problem. Journal of Statistical Computation and Simulation, 81 (10):1313–1331, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2015:EBA

[PJ15] Taeyoung Park and Seonghyun Jeong. Efficient Bayesian analysis of multivariate aggregate choices. Journal of Statistical Computation and Simulation, 85(16):3352–3366, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pourmousa:2015:MNM

[PJR15] R. Pourmousa, A. Jamalizadeh, and M. Rezapour. Multivariate normal mean-variance mixture distribution based on Birnbaum–Saunders distribution. Journal of Statistical Computation and Simulation, 85(13):2736–2749, 2015. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pradhan:2011:BEP

[PK11] Biswabrata Pradhan and Debasis Kundu. Bayes estimation and prediction of the two-parameter gamma distribution. Journal of Statistical Computation and Simulation, 81(9):1187–1198, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2016:RLR

[PK16a] Heewon Park and Sadanori Konishi. Robust logistic regression modelling via the elastic net-type regularization and tuning parameter selection. Journal of Statistical Computation and Simulation, 86(7):1450–1461, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

Pradhan:2016:BEB

[PK16b] Biswabrata Pradhan and Debasis Kundu. Bayes estimation for the Block and Basu bivariate and multivariate Weibull distributions. Journal of Statistical Computation and Simulation, 86(1):170–182, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2017:CRI

[PK17] Sangun Park and Ilmun Kim. Cumulative ratio information based on general cumulative entropy. Journal of Statistical Computation and Simulation, 87(3):563–576, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pumi:2015:GGS

[PL15] Guilherme Pumi and Sílvia R. C. Lopes. A generalization of a Gaussian semiparametric estimator on multivariate longrange dependent processes. Journal of Statistical Computation and Simulation, 85(9):1832–1856, 2015. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pollice:2016:I

[PL16a] Alessio Pollice and Giovanna Jona Lasinio. Introduction. Journal of Statistical Computation and Simulation, 86(13): 2495–2496, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Potgieter:2016:NTS

[PL16b] Cornelis J. Potgieter and Fred Lombard. Nonparametric two-sample estimation of location and scale parameters from empirical characteristic functions. Journal of Statistical Computation and Simulation, 86(16):3225–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pordeli:2018:EPL

[PL18] Pooneh Pordeli and Xuewen Lu. Estimation of partially linear single-index additive hazards model with current status data. Journal of Statistical Computation and Simulation, 88(5):935–966, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pereira:2010:CDM

[PM10] Dulce Gamito Pereira and João Tiago Mexia. Comparing double minimization and zigzag algorithms in joint regression analysis: the complete case. Journal of Statistical Computation and Simulation, 80(2):133–141, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Paul:2018:BCS

[PMM18] Erina Paul, Arnab Kumar Maity, and Raju Maiti. Bayesian comparative study on binary time series. Journal of Statistical Computation and Simulation, 88(14):2811–2826, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Perez:2014:ODR

[PMP14] Betsabé Pérez, Isabel Molina, and Daniel Peña. Outlier detection and robust estimation in linear regression models with fixed group effects. Journal of Statistical Computation and Simulation, 84(12):2652–2669, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Patil:2017:CER

[PNN17] Dipak D. Patil and U. V. Naik-Nimbalkar. Computation and estimation of reliability for some bivariate copulas with Pareto marginals. Journal of Statistical Computation and Simulation, 87(18):3563–3589, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pathmanathan:2014:MCS

[PO14] Dharini Pathmanathan and S. H. Ong. A Monte Carlo simulation study of two-sided tolerance intervals in balanced oneway random effects model for non-normal errors. Journal of Statistical Computation and Simulation, 84(11):2329–2344, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Paranaiba:2013:KBX

[POCdP13] Patrícia F. Paranaíba, Edwin M. M. Ortega, Gauss M. Cordeiro, and Marcelino A. R. de Pascoa. The Kumaraswamy Burr XII distribution: theory and practice. Journal of Statistical Computation and Simulation, 83(11):

2117–2143, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Patel:2010:MCC

[PP10] P. A. Patel and Jigna S. Patel. A Monte Carlo comparison of some variance estimators of the Horvitz–Thompson estimator. Journal of Statistical Computation and Simulation, 80(5):489–502, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Philonenko:2015:NTS

[PP15] Petr Philonenko and Sergey Postovalov. A new two-sample test for choosing between log-rank and Wilcoxon tests with right-censored data. Journal of Statistical Computation and Simulation, 85(14):2761–2770, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Philonenko:2019:NRT

[PP19] Petr Philonenko and Sergey Postovalov. The new robust two-sample test for randomly right-censored data. Journal of Statistical Computation and Simulation, 89(8):1357– 1375, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Phong:2019:EDT

[PPGT19] Duong Thanh Phong, T. Pham-Gia, and Dinh Ngoc Thanh. Exact distributions of two non-central generalized Wilks’s statistics. Journal of Statistical Computation and Simulation, 89(10):1798–1818, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Philonenko:2016:LTS

[PPK16] Petr Philonenko, Sergey Postovalov, and Artem Kovalevskii. The limit test statistic distribution of the maximum value test for right-censored data. Journal of Statistical Computation and Simulation, 86(17):3482–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Preisser:2014:CMS

[PQ14] John S. Preisser and Bahjat F. Qaqish. A comparison of methods for simulating correlated binary variables with specified marginal means and correlations. Journal of Statistical Computation and Simulation, 84(11):2441–2452, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2019:CCH

[PQ19] Kayoung Park and Peihua Qiu. Comparing crossing hazard rate functions by joint modelling of survival and longitudinal data. Journal of Statistical Computation and Simulation, 89(18):3391–3412, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pereira:2012:CBJ

[PRMM12] Dulce G. Pereira, Paulo C. Rodrigues, Stanislaw Mejza, and João T. Mexia. A comparison between joint regression analysis and the AMMI model: a case study with barley. Journal of Statistical Computation and Simulation, 82(2):193–207, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Popuri:2018:PCE

[PRNG18] Sai K. Popuri, Andrew M. Raim, Nagaraj K. Neerchal, and Matthias K. Gobbert. Parallelizing computation of expected values in recombinant binomial trees. Journal of Statistical Computation and Simulation, 88(4):657–674, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Potdar:2014:ISP

[PS14] K. G. Potdar and D. T. Shirke. Inference for the scale parameter of lifetime distribution of k-unit parallel system based on progressively censored data. Journal of Statistical Computation and Simulation, 84(1):171–185, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Panahi:2016:EPU

[PS16] Hanieh Panahi and Abdolreza Sayyareh. Estimation and prediction for a unified hybrid-censored Burr Type XII distribution. Journal of Statistical Computation and Simulation, 86(1):55–73, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pan:2018:SVS

[PS18] Juming Pan and Junfeng Shang. A simultaneous variable selection methodology for linear mixed models. Journal of Statistical Computation and Simulation, 88(17):3323–3337, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Paulino:2019:BCD

[PS19a] Carlos Daniel Paulino and Giovani L. Silva. Bayesian comparison of diagnostic tests with largely non-informative missing data. Journal of Statistical Computation and Simulation, 89(10):1877–1886, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pawar:2019:NTM

[PS19b] Somanath D. Pawar and Digambar T. Shirke. Nonparametric tests for multivariate multi-sample locations based on data depth. Journal of Statistical Computation and Simulation, 89(9):1574–1591, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Park:2014:RSR

[PSK14] Heewon Park, Fumitake Sakaori, and Sadanori Konishi. Robust sparse regression and tuning parameter selection via the efficient bootstrap information criteria. Journal of Statistical Computation and Simulation, 84(7):1596–1607, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pal:2018:FEE

[PSKC18] Surya K. Pal, Housila P. Singh, Sunil Kumar, and Kiranmoy Chatterjee. A family of efficient estimators of the finite population mean in simple random sampling. Journal of Statistical Computation and Simulation, 88(5):920–934, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Parvathy:2015:CSB

[PSS15] M. Parvathy, K. Sundarakantham, and S. Mercy Shalinie. Cosine similarity-based clustering and dynamic reputation trust aware key generation scheme for trusted communication on social networking. Journal of Statistical Computation and Simulation, 85(16):3247–3258, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pace:2011:RNS

[PSV11] Luigi Pace, Alessandra Salvan, and Laura Ventura. Remedying the Neyman–Scott phenomenon in model discrimination. Journal of Statistical Computation and Simulation, 81(6):749–757, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Plaza:2018:IEP

[PSY18] Francisco Plaza, Rodrigo Salas, and Eleuterio Yáñez. Identifying ecosystem patterns from time series of anchovy (Engraulis ringens) and sardine (Sardinops sagax) landings in northern Chile. Journal of Statistical Computation and Simulation, 88(10):1863–1881, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Poursina:2014:MOD

[PT14] D. Poursina and H. Talebi. Modified D-optimal design for logistic model. Journal of Statistical Computation and Simulation, 84(2):428–437, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Panza:2017:MPV

[PV17a] C. A. Panza and J. A. Vargas. Monitoring the parameter vector of regression models with time-to-event response in phase II processes. Journal of Statistical Computation and Simulation, 87(14):2779–2798, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Paulauskas:2017:CSP

[PV17b] V. Paulauskas and M. Vaiciulis. Comparison of the several parameterized estimators for the positive extreme value index. Journal of Statistical Computation and Simulation, 87 (7):1342–1362, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Psaradakis:2019:BAT

[PV19] Zacharias Psaradakis and Marián Vávra. Bootstrap-assisted tests of symmetry for dependent data. Journal of Statistical Computation and Simulation, 89(7):1203–1226, 2019.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pearn:2015:EPC

[PWW15] Wen Lea Pearn, Chin Chieh Wu, and Chia Huang Wu. Estimating process capability index C_{pk}: classical approach versus Bayesian approach. Journal of Statistical Computation and Simulation, 85(10):2007–2021, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Paul:2010:TNL

[PZ10] Sudhir R. Paul and Xuemao Zhang. Testing for normality in linear regression models. Journal of Statistical Computation and Simulation, 80(10):1101–1113, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Peng:2014:EAR

[PZY^{+}14] Bin Peng, Dianwen Zhu, Xiaowei Yang, Ling Liu, Wenquan Huang, Xiaohua Zhou, and Dong Yi. An extended association rule mining strategy for gene relationship discovery from microarray data. Journal of Statistical Computation and Simulation, 84(2):384–396, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Pan:2019:VSU

[PZZ19] Jing Pan, Shucong Zhang, and Yong Zhou. Variable screening for ultrahigh dimensional censored quantile regression. Journal of Statistical Computation and Simulation, 89(3): 395–413, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qasim:2018:PSN

[QAA18] Muhammad Qasim, Muhammad Amin, and Muhammad Amanullah. On the performance of some new Liu parameters for the gamma regression model. Journal of Statistical Computation and Simulation, 88(16):3065–3080, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Quinn:2015:GFS

[QHB15] Stephen J. Quinn, David W. Hosmer, and C. Leigh Blizzard. Goodness-of-fit statistics for log-link regression models. Journal of Statistical Computation and Simulation, 85 (12):2533–2545, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qureshi:2018:RCP

[QKUH18] Muhammad Nouman Qureshi, Cem Kadilar, Muhammad Noor Ul Amin, and Muhammad Hanif. Rare and clustered population estimation using the adaptive cluster sampling with some robust measures. Journal of Statistical Computation and Simulation, 88(14):2761–2774, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qomi:2016:ATI

[QKY16] M. Naghizadeh Qomi, A. Kiapour, and Derek S. Young. Approximate tolerance intervals for the discrete Poisson– Lindley distribution. Journal of Statistical Computation and Simulation, 86(4):841–854, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1038715.

Qi:2016:LMD

[QLW16] Dequan Qi, Zhonghua Li, and Zhaojun Wang. On-line monitoring data quality of high-dimensional data streams. Journal of Statistical Computation and Simulation, 86(11):2204– 2216, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qu:2015:OSB

[QMZ15] Yanping Qu, Glen D. Meeden, and Bo Zhang. An objective stepwise Bayes approach to small area estimation. Journal of Statistical Computation and Simulation, 85(7):1474– 1494, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qin:2016:JMC

[QMZ16] Guoyou Qin, Jie Mao, and Zhongyi Zhu. Joint meancovariance model in generalized partially linear varying coefficient models for longitudinal data. Journal of Statistical Computation and Simulation, 86(6):1166–1182, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (elec-

tronic), 1563-5163. URL http://www.tandfonline.com/
doi/abs/10.1080/00949655.2015.1057731.

Quinlan:2018:DRU

[QPQ18] José J. Quinlan, Garritt L. Page, and Fernando A. Quintana. Density regression using repulsive distributions. Journal of Statistical Computation and Simulation, 88(15):2931– 2947, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qureshi:2019:GEE

[QRH19] Muhammad Nouman Qureshi, Naureen Riaz, and Muhammad Hanif. Generalized estimator for the estimation of rare and clustered population variance in adaptive cluster sampling. Journal of Statistical Computation and Simulation, 89(11):2084–2101, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qi:2012:CQS

[QX12] Houduo Qi and Naihua Xiu. A convex quadratic semidefinite programming approach to the partial additive constant problem in multidimensional scaling. Journal of Statistical Computation and Simulation, 82(9):1317–1336, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qiao:2017:TNE

[QYX17] Yunxuan Qiao, Wei Yu, and Wangli Xu. Two new estimators for the proportion of true null hypotheses in multiple test. Journal of Statistical Computation and Simulation, 87(4):712–723, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qin:2016:QRL

[QZZ16] Guoyou Qin, Jiajia Zhang, and Zhongyi Zhu. Quantile regression in longitudinal studies with dropouts and measurement errors. Journal of Statistical Computation and Simulation, 86(17):3527–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rakitzis:2014:CCS

[RA14] Athanasios C. Rakitzis and Demetrios L. Antzoulakos. Control charts with switching and sensitizing runs rules for monitoring process variation. Journal of Statistical Computation and Simulation, 84(1):37–56, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roozbeh:2017:LTS

[RA17] Mahdi Roozbeh and Mohammad Arashi. Least-trimmed squares: asymptotic normality of robust estimator in semiparametric regression models. Journal of Statistical Computation and Simulation, 87(6):1130–1147, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raghav:2014:MON

[RAB14] Yashpal Singh Raghav, Irfan Ali, and Abdul Bari. Multiobjective nonlinear programming problem approach in multivariate stratified sample surveys in the case of nonresponse. Journal of Statistical Computation and Simulation, 84(1):22–36, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raqab:2016:PCP

[RAB16] Mohammad Z. Raqab, Laila A. Alkhalfan, and N. Balakrishnan. Pitman comparisons of predictors of censored observations from progressively censored samples for exponential distribution. Journal of Statistical Computation and Simulation, 86(8):1539–1558, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1071820.

Raissi:2012:CPF

[Rai12] Hamdi Raissi. Comparison of procedures for fitting the autoregressive order of a vector error correction model. Journal of Statistical Computation and Simulation, 82(10):1517– 1529, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raqab:2016:PIF

[RAJ16] Mohammad Z. Raqab and Reem A. Al-Jarallah. Prediction intervals for future Weibull residual data. Journal of Statistical Computation and Simulation, 86(7):1320–1333, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rakitzis:2016:MED

[Rak16a] Athanasios C. Rakitzis. Monitoring exponential data using two-sided control charts with runs rules. Journal of Statistical Computation and Simulation, 86(1):149–159, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rao:2016:BCI

[RAK16b] G. Srinivasa Rao, Muhammad Aslam, and R. R. L. Kantam. Bootstrap confidence intervals of C_{Npk} for inverse Rayleigh and log-logistic distributions. Journal of Statistical Computation and Simulation, 86(5):862–873, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1040799.

Ralph:2017:WSF

[Ral17] William J. Ralph. Warp statistics and financial returns. Journal of Statistical Computation and Simulation, 87(2): 390–405, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roozbeh:2011:RRM

[RAN11] M. Roozbeh, M. Arashi, and H. A. Niroumand. Ridge regression methodology in partial linear models with correlated errors. Journal of Statistical Computation and Simulation, 81(4):517–528, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Renjini:2016:SEL

[RASR16] K. R. Renjini, E. I. Abdul-Sathar, and G. Rajesh. A study of the effect of loss functions on the Bayes estimates of dynamic cumulative residual entropy for Pareto distribution under upper record values. Journal of Statistical Computation and Simulation, 86(2):324–339, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ravichandran:2019:SSX

[Rav19] J. Ravichandran. Self-starting X-bar control chart based on Six Sigma quality and sometimes pooling procedure. Journal of Statistical Computation and Simulation, 89(2): 362–377, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ristic:2012:GEE

[RB12] Miroslav M. Ristić and Narayanaswamy Balakrishnan. The gamma-exponentiated exponential distribution. Journal of Statistical Computation and Simulation, 82(8):1191–1206, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rivers:2017:DAL

[RB17] Derick L. Rivers and Edward L. Boone. Dynamic approach to linear statistical calibration with an application in microwave radiometry. Journal of Statistical Computation and Simulation, 87(1):17–28, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rizzo:2018:SCT

[RB18] Manuel Rizzo and Francesco Battaglia. Statistical and computational tradeoff in genetic algorithm-based estimation. Journal of Statistical Computation and Simulation, 88 (16):3081–3097, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rodrigues:2015:LCR

[RBC^{+}15] Josemar Rodrigues, N. Balakrishnan, Gauss M. Cordeiro, Mário de Castro, and Vicente G. Cancho. Latent cure rate model under repair system and threshold effect. Journal of Statistical Computation and Simulation, 85(14):2860–2873, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramsey:2011:PCM

[RBHSL11] Philip H. Ramsey, Kyrstle Barrera, Pri Hachimine-Semprebom, and Chang-Chia Liu. Pairwise comparisons of means under realistic nonnormality, unequal variances, outliers and equal sample sizes. Journal of Statistical Computation and Simulation, 81(2):125–135, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roozbeh:2016:ELT

[RBK16] M. Roozbeh and S. Babaie-Kafaki. Extended least trimmed squares estimator in semiparametric regression models with correlated errors. Journal of Statistical Computation and Simulation, 86(2):357–372, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rutherford:2015:URC

[RCL15] Mark J. Rutherford, Michael J. Crowther, and Paul C. Lambert. The use of restricted cubic splines to approximate complex hazard functions in the analysis of time-to-event data: a simulation study. Journal of Statistical Computation and Simulation, 85(4):777–793, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Romao:2010:EPC

[RDC10] Xavier Romão, Raimundo Delgado, and Aníbal Costa. An empirical power comparison of univariate goodness-of-fit tests for normality. Journal of Statistical Computation and Simulation, 80(5):545–591, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rampaso:2016:BAS

[RdSF16] Renato Couto Rampaso, Aparecida Doniseti Pires de Souza, and Edilson Ferreira Flores. Bayesian analysis of spatial data using different variance and neighbourhood structures. Journal of Statistical Computation and Simulation, 86(3): 535–552, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramanayake:2010:TCP

[RG10] Asoka Ramanayake and Arjun K. Gupta. Testing for a change point in a sequence of exponential random variables with repeated values. Journal of Statistical Computation and Simulation, 80(2):191–199, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raim:2013:MLE

[RGNM13] Andrew M. Raim, Matthias K. Gobbert, Nagaraj K. Neerchal, and Jorge G. Morel. Maximum-likelihood estimation of the random-clumped multinomial model as a prototype problem for large-scale statistical computing. Journal of Statistical Computation and Simulation, 83(12):2178–2194, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ringrose:2012:BCR

[Rin12] Trevor J. Ringrose. Bootstrap confidence regions for correspondence analysis. Journal of Statistical Computation and Simulation, 82(10):1397–1413, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rozliman:2018:BAE

[RIY18] Nur Aainaa Rozliman, Adriana Irawati Nur Ibrahim, and Rossita Muhamad Yunus. Bayesian approach to errors-in-variables in count data regression models with departures from normality and overdispersion. Journal of Statistical Computation and Simulation, 88(2):203–220, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rikhtehgaran:2017:CSE

[RKV17] Reyhaneh Rikhtehgaran, Iraj Kazemi, and Geert Verbeke. A comparative study on estimation methods to deal with the endogeneity in linear random-intercept models with an extension. Journal of Statistical Computation and Simulation, 87(1):171–186, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Robbertse:2014:MLE

[RL14] Wickes Robbertse and Fred Lombard. On maximum likelihood estimation of the long-memory parameter in fractional Gaussian noise. Journal of Statistical Computation and Simulation, 84(4):902–915, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Reese:2015:TSB

[RL15] Simon Reese and Yushu Li. Testing for structural breaks in the presence of data perturbations: impacts and waveletbased improvements. Journal of Statistical Computation and Simulation, 85(17):3468–3479, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Richter:2019:PCS

[RM19] Scott J. Richter and Melinda H. McCann. Pairwise comparison of scale using deviances. Journal of Statistical Computation and Simulation, 89(9):1730–1739, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rupasinghe:2014:OPI

[RMS14] Maduka Rupasinghe, Purna Mukhopadhyay, and V. A. Samaranayake. Obtaining prediction intervals for FARIMA processes using the sieve bootstrap. Journal of Statistical Computation and Simulation, 84(9):2044–2058, 2014. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ristic:2014:NLD

[RN14] Miroslav M. Ristić and Saralees Nadarajah. A new lifetime distribution. Journal of Statistical Computation and Simulation, 84(1):135–150, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roozbeh:2018:EQC

[RN18] Mahdi Roozbeh and Mohammad Najarian. Efficiency of the QR class estimator in semiparametric regression models to combat multicollinearity. Journal of Statistical Computation and Simulation, 88(9):1804–1825, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roldan-Nofuentes:2017:ACI

[RNA17] J. A. Roldán-Nofuentes and R. M. Amro. Approximate confidence intervals for the weighted kappa coefficient of a binary diagnostic test subject to a case-control design. Journal of Statistical Computation and Simulation, 87(3):530–545, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Raza:2019:DNO

[RNH19] Muhammad Ali Raza, Tahir Nawaz, and Dong Han. On designing new optimal synthetic Tukey’s control charts. Journal of Statistical Computation and Simulation, 89(12):2218– 2238, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roldan-Nofuentes:2019:RMC

[RNSR19] J. A. Roldán-Nofuentes and S. B. Sidaty-Regad. Recommended methods to compare the accuracy of two binary diagnostic tests subject to a paired design. Journal of Statistical Computation and Simulation, 89(14):2621–2644, 2019.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramires:2016:BFD

[ROCH16] Thiago G. Ramires, Edwin M. M. Ortega, Gauss M. Cordeiro, and Niel Hens. A bimodal flexible distribution for lifetime data. Journal of Statistical Computation and Simulation, 86(12):2450–2470, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rodrigues:2018:E

[RP18] Paulo Canas Rodrigues and Emilio Porcu. Editorial. Journal of Statistical Computation and Simulation, 88(10):1847– 1849, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rodriguez-Picon:2017:BDM

[RPFOMGRM17] Luis Alberto Rodríguez-Picón, Víctor Hugo Flores-Ochoa, Luis Carlos Méndez-González, and Manuel Arnoldo Rodríguez-Medina. Bivariate degradation modelling with marginal heterogeneous stochastic processes. Journal of Statistical Computation and Simulation, 87(11):2207–2226, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ristic:2015:LNG

[RPN15] Miroslav M. Ristić, Bozidar V. Popović, and Saralees Nadarajah. Libby and Novick’s generalized beta exponential distribution. Journal of Statistical Computation and Simulation, 85(4):740–761, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramsey:2013:PTG

[RR13a] Philip H. Ramsey and Patricia P. Ramsey. Pairwise testing of group mean vectors in MANOVA with small samples. Journal of Statistical Computation and Simulation, 83 (6):1037–1057, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramsey:2013:UVC

[RR13b] Philip H. Ramsey and Patricia P. Ramsey. Updated version of the critical values of the standardized fourth moment. Journal of Statistical Computation and Simulation, 83(7):1387–1388, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramsey:2010:CBP

[RRB10] Philip H. Ramsey, Patricia P. Ramsey, and Kyrstle Barrera. Choosing the best pairwise comparisons of means from non-normal populations, with unequal variances, but equal sample sizes. Journal of Statistical Computation and Simulation, 80(6):595–608, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rueda:2013:BHR

[RRDU13] Oscar M. Rueda, Cristina Rueda, and Ramon Diaz-Uriarte. A Bayesian HMM with random effects and an unknown number of states for DNA copy number analysis. Journal of Statistical Computation and Simulation, 83(1):82–96, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rodrigues:2013:FSP

[RRV13] Paulo M. M. Rodrigues, Antonio Rubia, and João Valle e Azevedo. Finite sample performance of frequency- and timedomain tests for seasonal fractional integration. Journal of Statistical Computation and Simulation, 83(7):1373–1384, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roozegar:2014:CPS

[RS14] Rasool Roozegar and Ahmad Reza Soltani. Classes of power semicircle laws that are randomly weighted average distributions. Journal of Statistical Computation and Simulation, 84(12):2636–2643, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rublik:2015:BFP

[RS15] F. Rublík and J. Somorcík. On the Behrens–Fisher problem from the spatial median point of view. Journal of Statistical Computation and Simulation, 85(15):3080–3091, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rocha:2017:AAP

[RS17] Andréa V. Rocha and Alexandre B. Simas. Asymptotic adjustments of Pearson residuals in exponential family nonlinear models. Journal of Statistical Computation and Simulation, 87(8):1684–1700, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Roy:2014:ABB

[RSD14] Surupa Roy, Angshuman Sarkar, and Kalyan Das. Analysis of bivariate binary data with possible chances of wrong ascertainment. Journal of Statistical Computation and Simulation, 84(4):724–738, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rastogi:2014:PRE

[RT14] Manoj Kumar Rastogi and Yogesh Mani Tripathi. Parameter and reliability estimation for an exponentiated half-logistic distribution under progressive type II censoring. Journal of Statistical Computation and Simulation, 84 (8):1711–1727, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Romeo:2019:MSH

[RT19] Giovanni Romeo and Magne Thoresen. Model selection in high-dimensional noisy data: a simulation study. Journal of Statistical Computation and Simulation, 89(11):2031–2050, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rodrigues:2018:RSS

[RTM18] Paulo Canas Rodrigues, Pétala G. S. E. Tuy, and Rahim Mahmoudvand. Randomized singular spectrum analysis for long time series. Journal of Statistical Computation and Simulation, 88(10):1921–1935, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ren:2017:SWB

[RWCD17] Lei Ren, Yongchang Wei, Jin Cui, and Yi Du. A sliding window-based multi-stage clustering and probabilistic forecasting approach for large multivariate time series data. Journal of Statistical Computation and Simulation, 87(13):2494–2508, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ramey:2013:CRM

[RY13] John A. Ramey and Phil D. Young. A comparison of regularization methods applied to the linear discriminant function with high-dimensional microarray data. Journal of Statistical Computation and Simulation, 83(3):581–596, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Riggs:2011:IEM

[RYS11] Kent Riggs, Dean Young, and James Stamey. Interval estimation for misclassification rate parameters in a complementary Poisson model. Journal of Statistical Computation and Simulation, 81(9):1145–1156, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rahardja:2012:BCS

[RZ12] Dewi Rahardja and Yan D. Zhao. Bayesian credible sets for the risk difference in binomial data with a fallible classifier and a gold standard. Journal of Statistical Computation and Simulation, 82(1):63–71, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rahardja:2013:OWA

[RZ13a] Dewi Rahardja and Yan D. Zhao. One-way analysis of proportions for misclassified binomial data. Journal of Statistical Computation and Simulation, 83(3):532–541, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Rusch:2013:GIR

[RZ13b] Thomas Rusch and Achim Zeileis. Gaining insight with recursive partitioning of generalized linear models. Journal of Statistical Computation and Simulation, 83(7):1301–1315, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sampath:2012:FPV

[SA12] S. Sampath and S. Ammani. Finite-population variance estimation under systematic sampling schemes with multiple random starts. Journal of Statistical Computation and Simulation, 82(8):1207–1221, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shokripour:2015:WBE

[SA15] Mona Shokripour and Mina Aminghafari. Wavelet-based estimation for multivariate stable laws. Journal of Statistical Computation and Simulation, 85(8):1584–1600, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Simsek:2016:PAA

[SA16] Gülhayat Gölbasi Simsek and Selahattin Aydogdu. Parallel analysis approach for determining dimensionality in canonical correlation analysis. Journal of Statistical Computation and Simulation, 86(17):3419–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Salehi:2015:POS

[SAB15] Mahdi Salehi, Jafar Ahmadi, and N. Balakrishnan. Prediction of order statistics and record values based on ordered ranked set sampling. Journal of Statistical Computation and Simulation, 85(1):77–88, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Safiih:2013:FPS

[Saf13] L. Muhamad Safiih. Fuzzy parametric sample selection model: Monte Carlo simulation approach. Journal of Statistical Computation and Simulation, 83(6):992–1006, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sultan:2014:BML

[SAK14] K. S. Sultan, N. H. Alsadat, and Debasis Kundu. Bayesian and maximum likelihood estimations of the inverse Weibull parameters under progressive type-II censoring. Journal of Statistical Computation and Simulation, 84(10):2248–2265, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sultan:2013:EDF

[SAM13a] K. S. Sultan and A. S. Al-Moisheer. Estimation of a discriminant function from a mixture of two inverse Weibull distributions. Journal of Statistical Computation and Simulation, 83(3):405–416, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sultan:2013:ABE

[SAM13b] Khalaf S. Sultan and A. S. Al-Moisheer. Approximate Bayes estimation of the parameters and reliability function of a mixture of two inverse Weibull distributions under type-2 censoring. Journal of Statistical Computation and Simulation, 83(10):1900–1914, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Santiago:2012:DRP

[San12] Eduardo Santiago. Design of robust parameter experiments in a continuous space using an evolutionary optimization algorithm. Journal of Statistical Computation and Simulation, 82(6):825–847, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sultan:2016:HOM

[SAT16] Khalaf S. Sultan and W. S. Al-Thubyani. Higher order moments of order statistics from the Lindley distribution and associated inference. Journal of Statistical Computation and Simulation, 86(17):3432–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sayyareh:2012:IAS

[Say12] Abdolreza Sayyareh. Inference after separated hypotheses testing: an empirical investigation for linear models. Journal of Statistical Computation and Simulation, 82(9):1275– 1286, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sazak:2019:CDW

[Saz19] Hakan Savas Sazak. The combined dynamically weighted modified maximum likelihood estimators of the location and scale parameters. Journal of Statistical Computation and Simulation, 89(5):751–762, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sutradhar:2011:RQL

[SB11] Brajendra C. Sutradhar and Wasimul Bari. Robust quasilikelihood inference in generalized linear mixed models with outliers. Journal of Statistical Computation and Simulation, 81(2):233–258, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sanjel:2012:JLP

[SB12a] Deepak Sanjel and N. Balakrishnan. Jacobi and Laguerre polynomial approximations for the distributions of statistics useful in testing for outliers in exponential and gamma samples. Journal of Statistical Computation and Simulation, 82(3):463–473, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Song:2012:PFM

[SB12b] Li Song and Pascal Bondon. Piecewise FARIMA models for long-memory time series. Journal of Statistical Computation and Simulation, 82(9):1367–1382, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Song:2013:SCE

[SB13] Li Song and Pascal Bondon. Structural changes estimation for strongly dependent processes. Journal of Statistical Computation and Simulation, 83(10):1783–1806, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Santos:2015:BAZ

[SB15] Bruno Santos and Heleno Bolfarine. Bayesian analysis for zero-or-one inflated proportion data using quantile regression. Journal of Statistical Computation and Simulation, 85 (17):3579–3593, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shafay:2014:BIB

[SBAA14] A. R. Shafay, N. Balakrishnan, and Y. Abdel-Aty. Bayesian inference based on a jointly type-II censored sample from two exponential populations. Journal of Statistical Computation and Simulation, 84(11):2427–2440, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sutradhar:2010:PVL

[SBD10] Brajendra C. Sutradhar, Wasimul Bari, and Kalyan Das. On probit versus logit dynamic mixed models for binary panel data. Journal of Statistical Computation and Simulation, 80(4):421–441, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sharafi:2013:DFC

[SBK13] Maryam Sharafi, N. Balakrishnan, and Baha-Eldin Khaledi. Distribution-free comparison of hazard rates of two distributions under progressive Type-II censoring. Journal of Statistical Computation and Simulation, 83(8):1527–1542, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Solaro:2018:SCI

[SBMF18] N. Solaro, A. Barbiero, G. Manzi, and P. A. Ferrari. A simulation comparison of imputation methods for quantitative data in the presence of multiple data patterns. Journal of Statistical Computation and Simulation, 88(18):3588–3619, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shafay:2014:TSB

[SBS14] A. R. Shafay, N. Balakrishnan, and K. S. Sultan. Twosample Bayesian prediction for sequential order statistics from exponential distribution based on multiply Type-II censored samples. Journal of Statistical Computation and Simulation, 84(3):526–544, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stogiannis:2012:TOI

[SC12] D. Stogiannis and C. Caroni. Tests for outliers in the inverse Gaussian distribution, with application to first hitting time models. Journal of Statistical Computation and Simulation, 82(1):73–80, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shiu:2016:SSU

[SC16] Shang-Ying Shiu and Ting-Li Chen. On the strengths of the self-updating process clustering algorithm. Journal of Statistical Computation and Simulation, 86(5):1010–1031, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. URL http://www.tandfonline.

com/doi/abs/10.1080/00949655.2015.1049605.

Shirke:2019:MMS

[SC19] Digambar Tukaram Shirke and Atul Rajaram Chavan. Multivariate multi-sample tests for location based on data depth. Journal of Statistical Computation and Simulation, 89(18):3377–3390, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Samawi:2018:SAM

[SCL^{+}18] Hani Samawi, Jingxian Cai, Daniel F. Linder, Haresh Rochani, and Jingjing Yin. A simpler approach for mediation analysis for dichotomous mediators in logistic regression. Journal of Statistical Computation and Simulation, 88(6):1211–1227, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2016:RWE

[SCW16] Si-Lian Shen, Jian-Ling Cui, and Chun-Wei Wang. Reweighting estimation of the coefficients in the varying coefficient model with heteroscedastic errors. Journal of Statistical Computation and Simulation, 86(10):1921–1935, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Satman:2015:REV

[SD15] M. Hakan Satman and Erkin Diyarbakirlioglu. Reducing errors-in-variables bias in linear regression using compact genetic algorithms. Journal of Statistical Computation and Simulation, 85(16):3216–3235, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Samawi:2012:SSR

[SDC12] Hani M. Samawi, Martin Dunbar, and Ding-Geng (Din) Chen. Steady-state ranked Gibbs sampler. Journal of Statistical Computation and Simulation, 82(8):1223–1238, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shafiei:2016:IWP

[SDS16] Sobhan Shafiei, Saeed Darijani, and Hadi Saboori. Inverse Weibull power series distributions: properties and applications. Journal of Statistical Computation and Simulation, 86(6):1069–1094, 2016. CODEN JSCSAJ.

ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

URL http://www.tandfonline.com/doi/abs/10.1080/
00949655.2015.1049949.

Song:2017:RPL

[SDWL17] Yunquan Song, Ping Dong, Xiuli Wang, and Lu Lin. Rapid penalized likelihood-based outlier detection via heteroskedasticity test. Journal of Statistical Computation and Simulation, 87(6):1206–1229, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shin:2011:UCC

[SE11] Hyejin Shin and R. L. Eubank. Unit canonical correlations and high-dimensional discriminant analysis. Journal of Statistical Computation and Simulation, 81(2):167–178, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Soliman:2013:EBT

[SEAEM13] Ahmed A. Soliman, Ahmed H. Abd Ellah, Naser A. Abou-Elheggag, and Abdullah A. Modhesh. Estimation from Burr Type XII distribution using progressive first-failure censored data. Journal of Statistical Computation and Simulation, 83(12):2270–2290, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sanchez-Espigares:2019:GCN

[SEGMA19] José A. Sánchez-Espigares, Pere Grima, and Lluís Marco-Almagro. Graphical comparison of normality tests for unimodal distribution data. Journal of Statistical Computation and Simulation, 89(1):145–154, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Seo:2011:GBA

[Seo11] Byungtae Seo. A gradient-based algorithm for semiparametric models with missing covariates. Journal of Statistical Computation and Simulation, 81(4):381–390, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Spiriti:2013:KSL

[SESY13] Steven Spiriti, Randall Eubank, Philip W. Smith, and Dennis Young. Knot selection for least-squares and penalized splines. Journal of Statistical Computation and Simulation, 83(6):1020–1036, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shaby:2012:EBB

[SF12] Benjamin A. Shaby and Daniel Fink. Embedding blackbox regression techniques into hierarchical Bayesian models. Journal of Statistical Computation and Simulation, 82 (12):1753–1766, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Salomon:2013:EHP

[SF13] Luis A. Salomón and Jean-Claude Fort. Estimation of the Hurst parameter in some fractional processes. Journal of Statistical Computation and Simulation, 83(3):542–554, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Seif:2015:EES

[SFS15] Asghar Seif, Alireza Faraz, and Magide Sadeghifar. Evaluation of the economic statistical design of the multivariate T^{2} control chart with multiple variable sampling intervals scheme: NSGA-II approach. Journal of Statistical Computation and Simulation, 85(12):2442–2455, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Su:2015:SSD

[SG15] Nan-Cheng Su and Arjun K. Gupta. On some sampling distributions for skew-normal population. Journal of Statistical Computation and Simulation, 85(17):3549–3559, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shabuz:2018:BCI

[SG18] Zillur R. Shabuz and Paul H. Garthwaite. Bootstrap confidence intervals for the contributions of individual variables to a Mahalanobis distance. Journal of Statistical Computation and Simulation, 88(11):2232–2258, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sysoev:2013:BEV

[SGB13] O. Sysoev, A. Grimvall, and O. Burdakov. Bootstrap estimation of the variance of the error term in monotonic regression models. Journal of Statistical Computation and Simulation, 83(4):627–640, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sevinc:2018:PGR

[SGÇ18] Büsra Sevinç, Selma Gürler, and Bekir Çetintav. Partial groups ranked set sampling and mean estimation. Journal of Statistical Computation and Simulation, 88(14):2799–2810, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schwartz:2013:IML

[SGG13] Jacob Schwartz, Ryan T. Godwin, and David E. Giles. Improved maximum-likelihood estimation of the shape parameter in the Nakagami distribution. Journal of Statistical Computation and Simulation, 83(3):434–445, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Salmeron:2018:VIF

[SGG18] R. Salmerón, C. B. García, and J. García. Variance inflation factor and condition number in multiple linear regression. Journal of Statistical Computation and Simulation, 88 (12):2365–2384, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schettlinger:2010:RBR

[SGGC10] Karen Schettlinger, Sarah Gelper, Ursula Gather, and Christophe Croux. Regression-based, regression-free and model-free approaches for robust online scale estimation. Journal of Statistical Computation and Simulation, 80(9): 1023–1040, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saint-Geours:2015:CFO

[SGTKBL15] N. Saint-Geours, S. Tarantola, V. Kopustinskas, and R. Bolado-Lavin. Computing first-order sensitivity indices with contribution to the sample mean plot. Journal of Statistical Computation and Simulation, 85(7):1334–1357, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shao:2014:SAB

[SGZM14] Wei Shao, Guangbao Guo, Guoqing Zhao, and Fanyu Meng. Simulated annealing for the bounds of Kendall’s τ and Spearman’s ρ. Journal of Statistical Computation and Simulation, 84(12):2688–2699, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Song:2010:BHD

[SH10] Juhee Song and Jeffrey D. Hart. Bootstrapping in a high dimensional but very low-sample size problem. Journal of Statistical Computation and Simulation, 80(8):825–840, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sauter:2016:QCS

[SH16] Rafael Sauter and Leonhard Held. Quasi-complete separation in random effects of binary response mixed models. Journal of Statistical Computation and Simulation, 86 (14):2781–2796, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sinclair:2019:SIC

[SH19] David Sinclair and Giles Hooker. Sparse inverse covariance estimation for high-throughput microRNA sequencing data in the Poisson log-normal graphical model. Journal of Statistical Computation and Simulation, 89(16):3105–3117, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2010:BFS

[Sha10a] Ramalingam Shanmugam. A book on frontiers of statistics errors is reviewed. Journal of Statistical Computation and Simulation, 80(3):363, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2010:BGL

[Sha10b] Ramalingam Shanmugam. Books on general linear models are reviewed. Journal of Statistical Computation and Simulation, 80(2):237–239, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2010:BLL

[Sha10c] Ramalingam Shanmugam. Books on longitudinal and latent growth are reviewed. Journal of Statistical Computation and Simulation, 80(7):823–824, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2010:BPI

[Sha10d] Ramalingam Shanmugam. Books on probability ideas are reviewed. Journal of Statistical Computation and Simulation, 80(6):707–711, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2010:BRV

[Sha10e] Ramalingam Shanmugam. Books on record values are reviewed. Journal of Statistical Computation and Simulation, 80(11):1297–1298, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2010:BRA

[Sha10f] Ramalingam Shanmugam. Books on risk analysis are reviewed. Journal of Statistical Computation and Simulation, 80(8):949–950, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2010:BSS

[Sha10g] Ramalingam Shanmugam. Books on sensory science are reviewed. Journal of Statistical Computation and Simulation, 80(12):1413–1414, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2010:CBI

[Sha10h] Ramalingam Shanmugam. A classic book on information theory is reviewed. Journal of Statistical Computation and Simulation, 80(5):593, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2010:SIB

[Sha10i] Ramalingam Shanmugam. Statistical inference books are reviewed. Journal of Statistical Computation and Simulation, 80(10):1181–1183, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2010:UBE

[Sha10j] Ramalingam Shanmugam. A unique book on environmental statistics is reviewed. Journal of Statistical Computation and Simulation, 80(1):119, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2010:UBL

[Sha10k] Ramalingam Shanmugam. A unique book on long memory time series is reviewed. Journal of Statistical Computation and Simulation, 80(4):475, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2010:UBR

[Sha10l] Ramalingam Shanmugam. A unique book on response surface methodology is reviewed. Journal of Statistical Computation and Simulation, 80(9):1069, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2011:BSS

[Sha11a] Ramalingam Shanmugam. A book on statistics in sports is reviewed. Journal of Statistical Computation and Simulation, 81(4):529, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2011:BOR

[Sha11b] Ramalingam Shanmugam. Books on optimizations are reviewed. Journal of Statistical Computation and Simulation, 81(2):259–260, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2011:ESB

[Sha11c] Ramalingam Shanmugam. Encyclopedia of statistics in behavioural science. Journal of Statistical Computation and Simulation, 81(3):379–380, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2011:MSB

[Sha11d] Ramalingam Shanmugam. Multivariate statistics books are reviewed. Journal of Statistical Computation and Simulation, 81(1):123–124, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2011:TGB

[Sha11e] Ramalingam Shanmugam. Two good books on forensic applications of statistics are reviewed. Journal of Statistical Computation and Simulation, 81(5):659–660, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2013:BSW

[Sha13a] Ramalingam Shanmugam. Books for statisticians working in business, finance and insurance. Journal of Statistical Computation and Simulation, 83(7):1385–1386, 2013. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2013:ORB

[Sha13b] Ramalingam Shanmugam. Operations research books are reviewed. Journal of Statistical Computation and Simulation, 83(6):1188–1189, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2013:SPA

[Sha13c] Ramalingam Shanmugam. Stochastic processes with applications. Journal of Statistical Computation and Simulation, 83(3):597–598, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2014:BSB

[Sha14a] Ramalingam Shanmugam. Basic statistics books are reviewed. Journal of Statistical Computation and Simulation, 84(4):930, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2014:BBRa

[Sha14b] Ramalingam Shanmugam. Bayesian books are reviewed. Journal of Statistical Computation and Simulation, 84(5): 1154–1155, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2014:BBR

[Sha14c] Ramalingam Shanmugam. Biostatistics books are reviewed. Journal of Statistical Computation and Simulation, 84(8): 1841–1842, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2014:BRB

[Sha14d] Ramalingam Shanmugam. Book review: Computational statistics books are reviewed. Journal of Statistical Computation and Simulation, 84(11):2522–2523, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2014:EBB

[Sha14e] Ramalingam Shanmugam. Excellent biostatistics books for clinical trials are reviewed. Journal of Statistical Computation and Simulation, 84(10):2295–2297, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2014:ECD

[Sha14f] Ramalingam Shanmugam. Excellent categorical data analysis books are reviewed. Journal of Statistical Computation and Simulation, 84(9):2071–2072, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shan:2015:EUT

[Sha15a] Guogen Shan. Exact unconditional testing procedures for comparing two independent Poisson rates. Journal of Statistical Computation and Simulation, 85(5):947–955, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2015:BRBb

[Sha15b] Ramalingam Shanmugam. Book review: a book about zero is reviewed. Journal of Statistical Computation and Simulation, 85(10):2133, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2015:BRSb

[Sha15c] Ramalingam Shanmugam. Book review: A simulation book is reviewed. [Handbook of real-world applications in modeling and simulation, ISBN-13 978-1-118-11777-4]. Journal of Statistical Computation and Simulation, 85(4):854, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2015:BRR

[Sha15d] Ramalingam Shanmugam. Book review: Receiver operating characteristic curve book is reviewed: [ROC curves for continuous data, ISBN-13 978-1-4398-0021-8]. Journal of Statistical Computation and Simulation, 85(2):430, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2015:BRSc

[Sha15e] Ramalingam Shanmugam. Book review: SPSS book is reviewed. [Building SPSS graphs to understand data, ISBN-13 978-1-4522-1648-3]. Journal of Statistical Computation and Simulation, 85(5):1070, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2015:BRBa

[Sha15f] Ramalingam Shanmugam. Book review: Survival analysis books are reviewed. Journal of Statistical Computation and Simulation, 85(8):1706–1707, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2015:BRSa

[Sha15g] Ramalingam Shanmugam. Book reviews: SAS books are reviewed. Journal of Statistical Computation and Simulation, 85(3):638–640, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2015:BRSd

[Sha15h] Ramalingam Shanmugam. Book reviews: Statistics and related books are reviewed. Journal of Statistical Computation and Simulation, 85(6):1277–1279, 2015. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2015:BRT

[Sha15i] Ramalingam Shanmugam. Book reviews: Time series modelling books are reviewed. [Time series; modeling, computation and inference, ISBN-13 978-1-4200-9336-0, Bayesian time series models, ISBN 978-0-521-19676-5]. Journal of Statistical Computation and Simulation, 85(7):1495, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2015:URR

[Sha15j] Ramalingam Shanmugam. Uncertainty and risk-related books are reviewed. Journal of Statistical Computation and Simulation, 85(9):1917–1918, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shafay:2016:BEP

[Sha16] A. R. Shafay. Bayesian estimation and prediction based on generalized Type-II hybrid censored sample. Journal of Statistical Computation and Simulation, 86(10):1970–1988, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2018:DSO

[Sha18a] Ramalingam Shanmugam. Data stewardship for open science: implementing FAIR principles / data-driven storytelling. Journal of Statistical Computation and Simulation, 88(16):3249–3250, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2018:ECI

[Sha18b] Ramalingam Shanmugam. Elements of causal inference: foundations and learning algorithms. Journal of Statistical Computation and Simulation, 88(16):3248, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sharma:2018:TLN

[Sha18c] Vikas Kumar Sharma. Topp–Leone normal distribution with application to increasing failure rate data. Journal of Statistical Computation and Simulation, 88(9):1782–1803, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shang:2019:RFT

[Sha19a] Han Lin Shang. A robust functional time series forecasting method. Journal of Statistical Computation and Simulation, 89(5):795–814, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:BCE

[Sha19b] Ramalingam Shanmugam. Bayesian cost-effectiveness analysis of medical treatments. Journal of Statistical Computation and Simulation, 89(18):3497, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:BI

[Sha19c] Ramalingam Shanmugam. Book on inference. Journal of Statistical Computation and Simulation, 89(9):1760, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:BRAa

[Sha19d] Ramalingam Shanmugam. Book review: Applied compositional data analysis: with worked examples in R. Journal of Statistical Computation and Simulation, 89(16):3165, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:BRAb

[Sha19e] Ramalingam Shanmugam. Book review: Applied statistical inference with Minitab. Journal of Statistical Computation and Simulation, 89(16):3165–3166, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:BRD

[Sha19f] Ramalingam Shanmugam. Book review: Discover probability: how to use it, how to avoid misusing it, and how it affects every aspect of your life, by Arieh Ben-Naim, 5 Toh Tuck Link, Singapore 596224, World Scientific Publishing CO. Pte. Ltd. Press, 2015, 318 + xix pp., $38.66 (paperback), ISBN: 978-981-4616-32-4. Journal of Statistical Computation and Simulation, 89(13):2594–2595, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:BRIa

[Sha19g] Ramalingam Shanmugam. Book review: Introduction to engineering statistics and lean six sigma: statistical quality control and design of experiments and systems. Journal of Statistical Computation and Simulation, 89(15):2980, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:BRIb

[Sha19h] Ramalingam Shanmugam. Book review: Introduction to research methods and data analysis in the health sciences. Journal of Statistical Computation and Simulation, 89(15): 2981, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:BRB

[Sha19i] Ramalingam Shanmugam. Book review: The beauty of mathematics in computer science, by Jun Wu, Boca Raton, FL, CRC Press, 2019, pp. 268+xv, $39.95 (Paperback), ISBN: 978-1-138-04960-4. Journal of Statistical Computation and Simulation, 89(13):2594, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:IAM

[Sha19j] Ramalingam Shanmugam. Book reviews: An Information Approach to Mitochondrial Dysfunction: Extending Swerdlow’s Hypothesis / Essentials of Health Information Management: Principles and Practices / Statistics for Health Care Research: A Practical Workbook. Journal of Statistical Computation and Simulation, 89(1):189–190, 2019. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:BDS

[Sha19k] Ramalingam Shanmugam. Books on data science. Journal of Statistical Computation and Simulation, 89(7):1340– 1341, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:BPO

[Sha19l] Ramalingam Shanmugam. Books on programming-oriented books. Journal of Statistical Computation and Simulation, 89(5):946–947, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:BRMb

[Sha19m] Ramalingam Shanmugam. Books on reservoir modeling. Journal of Statistical Computation and Simulation, 89(7): 1340, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:BRMa

[Sha19n] Ramalingam Shanmugam. Books on reservoir modelling. Journal of Statistical Computation and Simulation, 89(1): 188, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:CIG

[Sha19o] Ramalingam Shanmugam. Combinatorial inference in geometric data analysis. Journal of Statistical Computation and Simulation, 89(18):3497–3498, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:COB

[Sha19p] Ramalingam Shanmugam. Computation oriented books. Journal of Statistical Computation and Simulation, 89(9): 1761–1762, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:DOS

[Sha19q] Ramalingam Shanmugam. Data-oriented statistics books. Journal of Statistical Computation and Simulation, 89(10): 1971–1972, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:DMH

[Sha19r] Ramalingam Shanmugam. Decision making in health and medicine. Journal of Statistical Computation and Simulation, 89(12):2393, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:MRW

[Sha19s] Ramalingam Shanmugam. Measuring research: what everyone needs to know. Journal of Statistical Computation and Simulation, 89(3):557–558, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:MSS

[Sha19t] Ramalingam Shanmugam. Modern survey sampling. Journal of Statistical Computation and Simulation, 89(5):948, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2019:RRS

[Sha19u] Ramalingam Shanmugam. Robust response surfaces, regression, and positive data analyses. Journal of Statistical Computation and Simulation, 89(3):557, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2010:IEJ

[She10] Pao-Sheng Shen. Interval estimation of the joint survival function for successive duration times under left truncation and right censoring. Journal of Statistical Computation and Simulation, 80(11):1265–1277, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2011:PHR

[She11] Pao-Sheng Shen. Proportional hazards regression for cancer screening data. Journal of Statistical Computation and Simulation, 81(3):367–377, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2012:NAI

[She12] Pao-Sheng Shen. Nonparametric analysis of interval censored and doubly truncated data. Journal of Statistical Computation and Simulation, 82(12):1845–1854, 2012. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2014:ATD

[She14a] Pao-Sheng Shen. Analysing truncated data with semiparametric transformation models. Journal of Statistical Computation and Simulation, 84(11):2474–2485, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2014:PHR

[She14b] Pao-Sheng Shen. Proportional hazards regression with interval-censored and left-truncated data. Journal of Statistical Computation and Simulation, 84(2):264–272, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2014:WQR

[She14c] Pao-Sheng Shen. A weighted quantile regression for lefttruncated and right-censored data. Journal of Statistical Computation and Simulation, 84(3):596–604, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2015:NTL

[She15a] Pao-Sheng Shen. Nonparametric tests for left-truncated and interval-censored data. Journal of Statistical Computation and Simulation, 85(8):1544–1553, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2015:STM

[She15b] Pao-Sheng Shen. Semiparametric transformation models with random effects for clustered doubly censored data. Journal of Statistical Computation and Simulation, 85(8): 1554–1566, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shih:2015:MIB

[Shi15] Nenghui Shih. The model identification of beta distribution based on quantiles. Journal of Statistical Computation and Simulation, 85(10):2022–2032, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sanagawa:2017:SHP

[SHLT17] Ryoji Sanagawa, Kyoko Hasegawa, Liang Li, and Satoshi Tanaka. The stochastic highlighting of polygon edges in the transparent visualization of large-scale polygon meshes: application to visualizing a high-energy elementary particle detector. Journal of Statistical Computation and Simulation, 87(13):2560–2571, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shitan:2012:NAT

[SHP12] Mahendran Shitan, Lai Choo Heng, and A. R. Padmanabhan. A new adaptive test for paired data for small to moderate sample sizes. Journal of Statistical Computation and Simulation, 82(6):907–914, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shu:2013:CRL

[SHST13] Lianjie Shu, Wenpo Huang, Yan Su, and Kwok-Leung Tsui. Computation of the run-length percentiles of CUSUM control charts under changes in variances. Journal of Statistical Computation and Simulation, 83(7):1238–1251, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shutoh:2012:AED

[Shu12] Nobumichi Shutoh. An asymptotic expansion for the distribution of the linear discriminant function based on monotone missing data. Journal of Statistical Computation and Simulation, 82(2):241–259, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Siray:2019:LEA

[Şir19] Gülesen Üstünda Şiray. Liu estimation approach to the measurement error models. Journal of Statistical Computation and Simulation, 89(18):3453–3496, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sutradhar:2010:TDS

[SJ10] Brajendra C. Sutradhar and Vandna Jowaheer. Treatment design selection effects on parameter estimation in dynamic logistic models for longitudinal binary data. Journal of Statistical Computation and Simulation, 80(9):1053–1067, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Safavinejad:2015:DBE

[SJN15] M. Safavinejad, S. Jomhoori, and H. Alizadeh Noughabi. A density-based empirical likelihood ratio goodness-of-fit test for the Rayleigh distribution and power comparison. Journal of Statistical Computation and Simulation, 85(16):3322– 3334, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sen:2017:SCS

[SJZ17] Deborshee Sen, Ajay Jasra, and Yan Zhou. Some contributions to sequential Monte Carlo methods for option pricing. Journal of Statistical Computation and Simulation, 87(4):

733–752, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2011:NEB

[SK11] Pao-Sheng Shen and Jai-Chi Kang. Nonparametric estimation of the bivariate survival function for one modified form of current-status data. Journal of Statistical Computation and Simulation, 81(9):1199–1209, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sharma:2013:STS

[SK13] Dinesh Sharma and B. M. Golam Kibria. On some test statistics for testing homogeneity of variances: a comparative study. Journal of Statistical Computation and Simulation, 83(10):1944–1963, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shoaee:2016:SIW

[SK16] Shirin Shoaee and Esmaile Khorram. Statistical inference of for Weibull distribution under Type-II progressively hybrid censored data. Journal of Statistical Computation and Simulation, 86(18):3815–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shirke:2017:PCD

[SK17] D. T. Shirke and S. D. Khorate. Power comparison of data depth-based nonparametric tests for testing equality of locations. Journal of Statistical Computation and Simulation, 87(8):1489–1497, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Siju:2018:BER

[SK18] K. C. Siju and M. Kumar. Bayesian estimation of reliability using time-to-failure distribution of parametric degradation models. Journal of Statistical Computation and Simulation, 88(9):1717–1748, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sunecher:2017:GEA

[SKJ17] Yuvraj Sunecher, Naushad Mamode Khan, and Vandna Jowaheer. A GQL estimation approach for analysing nonstationary over-dispersed BINAR(1) time series. Journal of Statistical Computation and Simulation, 87(10):1911–1924, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saracoglu:2012:EED

[SKK12] Bugra Saraçoglu, Ismail Kinaci, and Debasis Kundu. On estimation of R = P (Y < X) for exponential distribution under progressive Type-II censoring. Journal of Statistical Computation and Simulation, 82(5):729–744, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Strzalkowska-Kominiak:2014:KME

[SKM14] E. Strzalkowska-Kominiak and M. Mahdizadeh. On the Kaplan–Meier estimator based on ranked set samples. Journal of Statistical Computation and Simulation, 84(12):2577– 2591, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sukchotrat:2011:ICA

[SKTC11] Thuntee Sukchotrat, Seoung Bum Kim, Kwok-Leung Tsui, and Victoria C. P. Chen. Integration of classification algorithms and control chart techniques for monitoring multivariate processes. Journal of Statistical Computation and Simulation, 81(12):1897–1911, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shan:2018:AUV

[SKX^{+}18] Guogen Shan, Le Kang, Min Xiao, Hua Zhang, and Tao Jiang. Accurate unconditional p-values for a two-arm study with binary endpoints. Journal of Statistical Computation and Simulation, 88(6):1200–1210, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Seo:2015:NAM

[SL15] Byungtae Seo and Taewook Lee. A new algorithm for maximum likelihood estimation in normal scale-mixture generalized autoregressive conditional heteroskedastic models. Journal of Statistical Computation and Simulation, 85(1): 202–215, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Santos:2017:ETA

[SL17] Cristiano C. Santos and Rosangela H. Loschi. EM-Type algorithms for heavy-tailed logistic mixed models. Journal of Statistical Computation and Simulation, 87(15):2940–2961, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shah:2017:MDM

[SLA17] Muhammad Kashif Ali Shah, Supranee Lisawadi, and S. Ejaz Ahmed. Merging data from multiple sources: pretest and shrinkage perspectives. Journal of Statistical Computation and Simulation, 87(8):1577–1592, 2017. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stange:2016:CAM

[SLD16] J. Stange, N. Loginova, and T. Dickhaus. Computing and approximating multivariate chi-square probabilities. Journal of Statistical Computation and Simulation, 86(6):1233–1247, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

URL http://www.tandfonline.com/doi/abs/10.1080/
00949655.2015.1058798.

Song:2018:WLR

[SLLZ18] Zhi Song, Yanchun Liu, Zhonghua Li, and Jiujun Zhang. A weighted likelihood ratio test-based chart for monitoring process mean and variability. Journal of Statistical Computation and Simulation, 88(7):1415–1436, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stein:2016:CPD

[SLM16] J. Stein, S. R. C. Lopes, and A. V. Medino. Continuous processes derived from the solution of generalized Langevin equation: theoretical properties and estimation. Journal of Statistical Computation and Simulation, 86(14):2819–2845, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schkoda:2015:CMS

[SLSW15] Ryan F. Schkoda, Robert Lund, Nan Su, and John Wagner. A comparison of multivariate signal discrimination techniques. Journal of Statistical Computation and Simulation, 85(3):494–506, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Schouten:2018:GMV

[SLV18] Rianne Margaretha Schouten, Peter Lugtig, and Gerko Vink. Generating missing values for simulation purposes: a multivariate amputation procedure. Journal of Statistical Computation and Simulation, 88(15):2909–2930, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Scaccia:2011:MBT

[SM11] L. Scaccia and R. J. Martin. Model-based tests for simplification of lattice processes. Journal of Statistical Computation and Simulation, 81(1):89–107, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sampaio:2015:IER

[SM15] J. M. Sampaio and P. A. Morettin. Indirect estimation of randomized generalized autoregressive conditional heteroskedastic models. Journal of Statistical Computation and Simulation, 85(13):2702–2717, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shirozhan:2018:NCI

[SM18] M. Shirozhan and M. Mohammadpour. A new class of INAR(1) model for count time series. Journal of Statistical Computation and Simulation, 88(7):1348–1368, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Salehi:2010:EEA

[SMBS10] Mohammad Salehi, Mohammad Moradi, Jennifer A. Brown, and David Smith. Efficient estimators for adaptive stratified sequential sampling. Journal of Statistical Computation and Simulation, 80(10):1163–1179, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sutton:2019:HSS

[SML19] Matthew Sutton, Kerrie Mengersen, and Benoit Liquet. [HDDA] sparse subspace constrained partial least squares.

Journal of Statistical Computation and Simulation, 89(6): 1005–1019, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sarvi:2019:GBZ

[SMM19] Fatemeh Sarvi, Abbas Moghimbeigi, and Hossein Mahjub. GEE-based zero-inflated generalized Poisson model for clustered over or under-dispersed count data. Journal of Statistical Computation and Simulation, 89(14):2711–2732, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sha:2017:BIB

[SN17] Naijun Sha and Tun Lee Ng. Bayesian inference for Birnbaum–Saunders distribution and its generalization. Journal of Statistical Computation and Simulation, 87(12): 2411–2429, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sanqui:2012:LOT

[SNG12] Jose Almer T. Sanqui, Truc T. Nguyen, and Arjun K. Gupta. Locally optimal test of normality against skewnormality. Journal of Statistical Computation and Simulation, 82(3):359–368, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saavedra-Nieves:2016:CSS

[SNGMRC16] Paula Saavedra-Nieves, Wenceslao González-Manteiga, and Alberto Rodríguez-Casal. A comparative simulation study of data-driven methods for estimating density level sets. Journal of Statistical Computation and Simulation, 86(2): 236–251, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sutradhar:2010:MSS

[SO10] Brajendra C. Sutradhar and Alwell J. Oyet. Maximum Studentized score tests for the detection of outliers in time series regression models. Journal of Statistical Computation and Simulation, 80(12):1355–1372, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sheu:2013:PIS

[SOH13] Shey-Huei Sheu, Chih-Wen Ouyoung, and Tsung-Shin Hsu. Phase II statistical process control for functional data. Journal of Statistical Computation and Simulation, 83(11):2144– 2159, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saha:2011:IEN

[SP11] Krishna K. Saha and Sudhir R. Paul. Interval estimation of the negative binomial dispersion parameter. Journal of Statistical Computation and Simulation, 81(12):1993–2002, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saleh:2016:CBC

[SP16] Moein Saleh and Rong Pan. A clustering-based coordinate exchange algorithm for generating G-optimal experimental designs. Journal of Statistical Computation and Simulation, 86(8):1582–1604, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1077252.

Saraiva:2019:DDS

[SPS19] E. F. Saraiva, C. A. B. Pereira, and A. K. Suzuki. A datadriven selection of the number of clusters in the Dirichlet allocation model via Bayesian mixture modelling. Journal of Statistical Computation and Simulation, 89(15):2848–2870, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sutradhar:2011:EIF

[SR11] Brajendra Sutradhar and R. Prabhakar Rao. On efficient inferences in familial-longitudinal binary models with two variance components. Journal of Statistical Computation and Simulation, 81(11):1609–1626, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Steward:2016:BWA

[SR16] Robert M. Steward and Steven E. Rigdon. A Bayesian wavelet approach to estimation of a change-point in a nonlinear multivariate time series. Journal of Statistical Computation and Simulation, 86(13):2625–2643, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shadid:2011:MBB

[SRAO11] Maisa R. Shadid, Mohammad Z. Raqab, and Amer I. Al-Omari. Modified BLUEs and BLIEs of the location and scale parameters and the population mean using ranked set sampling. Journal of Statistical Computation and Simulation, 81(3):261–274, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sarma:2011:CFW

[SRG11] Ramabhadra Sarma, A. V. Dattatreya Rao, and S. V. S. Girija. On characteristic functions of the wrapped lognormal and the wrapped Weibull distributions. Journal of Statistical Computation and Simulation, 81(5):579–589, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Singh:2013:ISD

[SRK13] Bhupendra Singh, Shubhi Rathi, and Sachin Kumar. Inferential statistics on the dynamic system model with timedependent failure-rate. Journal of Statistical Computation and Simulation, 83(1):1–24, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Salvati:2011:SAE

[SRP11] N. Salvati, M. G. Ranalli, and M. Pratesi. Small area estimation of the mean using non-parametric M-quantile regression: a comparison when a linear mixed model does not hold. Journal of Statistical Computation and Simulation, 81(8):945–964, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2011:ECP

[sS11] Pao sheng Shen. Estimation in the Cox proportional hazards model with doubly censored and truncated data. Journal of Statistical Computation and Simulation, 81(11):1717–1725, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sajesh:2012:ODH

[SS12a] T. A. Sajesh and M. R. Srinivasan. Outlier detection for high dimensional data using the Comedian approach. Journal of Statistical Computation and Simulation, 82(5):745– 757, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2012:EML

[sS12b] Pao sheng Shen. Estimation in multiple linear regression model with doubly censored data. Journal of Statistical Computation and Simulation, 82(4):503–514, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sun:2013:GEL

[SS13] Bingrui Sun and Brajendra C. Sutradhar. GQL estimation in linear dynamic models for panel data. Journal of Statistical Computation and Simulation, 83(3):568–580, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sugaya:2015:PIA

[SS15] Yuki Sugaya and Ritei Shibata. Probability inheritance algorithm and its implementation. Journal of Statistical Computation and Simulation, 85(10):2033–2048, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2016:EJS

[sS16] Pao sheng Shen. Estimation of the joint survival function for successive duration times under double-truncation and right-censoring. Journal of Statistical Computation and Simulation, 86(13):2673–2683, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2019:EJS

[sS19] Pao sheng Shen. Estimation of the joint survival function for three successive duration times under double truncation and right censoring. Journal of Statistical Computation and Simulation, 89(4):576–590, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shafiei:2017:GIP

[SSD17] Sobhan Shafiei, Hadi Saboori, and Mahdi Doostparast. Generalized inferential procedures for generalized Lorenz curves under the Pareto distribution. Journal of Statistical Computation and Simulation, 87(2):267–279, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Singh:2013:BER

[SSK13] Sanjay Kumar Singh, Umesh Singh, and Dinesh Kumar. Bayes estimators of the reliability function and parameter of inverted exponential distribution using informative and non-informative priors. Journal of Statistical Computation and Simulation, 83(12):2258–2269, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saadatmelli:2018:ESD

[SSMF18] Aitin Saadatmelli, A. Seif, M. Bameni Moghadam, and A. Faraz. Economic statistical design of a multivariate control chart under a Burr XII shock model with multiple assignable causes. Journal of Statistical Computation and Simulation, 88(11):2111–2136, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Saracli:2013:CLR

[ST13] Sinan Saraçli and Ayça Hatice Türkan. A comparison of linear regression techniques in method comparison studies. Journal of Statistical Computation and Simulation, 83 (10):1890–1899, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stapleton:2010:ADW

[Sta10] Robert D. Stapleton. Algorithm for determining whether various two-level fractional factorial split-plot row-column designs are non-isomorphic. Journal of Statistical Computation and Simulation, 80(1):43–59, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Staahl:2016:PEU

[Stå16] Olivia Ståhl. Point estimation using tail modelling for right skew populations. Journal of Statistical Computation and Simulation, 86(11):2073–2088, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shalabh:2016:IIE

[STL16] Shalabh, Jia-Ren Tsai, and Pen-Hwang Liau. Immaculating the inconsistent estimator of slope parameter in measurement error model with replicated data. Journal of Statistical Computation and Simulation, 86(17):3371–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Singh:2014:GPE

[STS14] Housila P. Singh, Rajesh Tailor, and Sarjinder Singh. General procedure for estimating the population mean using ranked set sampling. Journal of Statistical Computation and Simulation, 84(5):931–945, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Singh:2015:EPP

[STW15] Sukhdev Singh, Yogesh Mani Tripathi, and Shuo-Jye Wu. On estimating parameters of a progressively censored lognormal distribution. Journal of Statistical Computation and Simulation, 85(6):1071–1089, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Steel:2011:VSP

[SU11] S. J. Steel and D. W. Uys. A variable selection proposal for multiple linear regression analysis. Journal of Statistical Computation and Simulation, 81(12):2095–2105, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Su:2016:SSD

[Su16] Pei-Fang Su. Sample size determination for clustered rightcensored data using copula models. Journal of Statistical Computation and Simulation, 86(4):696–706, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1031232.

Susam:2018:TIA

[SU18] Selim Orhun Susam and Burcu Hudaverdi Ucer. Testing independence for Archimedean copula based on Bernstein estimate of Kendall distribution function. Journal of Statistical Computation and Simulation, 88(13):2589–2599, 2018.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sun:2011:HRP

[Sun11] Hokeun Sun. Hybrid regions for post-change mean in a sequence of normal variables. Journal of Statistical Computation and Simulation, 81(2):221–231, 2011. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shi:2018:ATE

[SW18] X. Shi and A. Wong. Accurate tests for the equality of coefficients of variation. Journal of Statistical Computation and Simulation, 88(18):3529–3543, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Su:2015:PEI

[SWLZ15] Zhi-Gang Su, Pei-Hong Wang, Yi-Guo Li, and Ze-Kun Zhou. Parameter estimation from interval-valued data using the expectation-maximization algorithm. Journal of Statistical Computation and Simulation, 85(2):320–338, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Samura:2019:CPE

[SWW19] Sallieu Kabay Samura, Xuejun Wang, and Yi Wu. Consistency properties for the estimators of partially linear regression model under dependent errors. Journal of Statistical Computation and Simulation, 89(13):2410–2433, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shi:2018:ACA

[SWXJ18] Yueyong Shi, Yuanshan Wu, Deyi Xu, and Yuling Jiao. An ADMM with continuation algorithm for non-convex SICApenalized regression in high dimensions. Journal of Statistical Computation and Simulation, 88(9):1826–1846, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Smith:2015:HOI

[SWZ15] J. B. Smith, A. Wong, and X. Zhou. Higher order inference for stress-strength reliability with independent Burr-type X distributions. Journal of Statistical Computation and Simulation, 85(15):3092–3107, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sun:2017:BAU

[SXTJ17] Liangrui Sun, Michelle Xia, Yuanyuan Tang, and Philip G. Jones. Bayesian adjustment for unidirectional misclassification in ordinal covariates. Journal of Statistical Computation and Simulation, 87(18):3440–3468, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2015:BCW

[SY15] Yan Shen and Zhenlin Yang. Bias-correction for Weibull common shape estimation. Journal of Statistical Computation and Simulation, 85(15):3017–3046, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sevil:2017:PCK

[SY17a] Yusuf Can Sevil and Tugba Ozkal Yildiz. Power comparison of the Kolmogorov–Smirnov test under ranked set sampling and simple random sampling. Journal of Statistical Computation and Simulation, 87(11):2175–2185, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shen:2017:ILI

[SY17b] Yan Shen and Zhenlin Yang. Improved likelihood inferences for Weibull regression model. Journal of Statistical Computation and Simulation, 87(12):2349–2371, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sellers:2019:ZIS

[SY19] Kimberly F. Sellers and Derek S. Young. Zero-inflated sum of Conway–Maxwell–Poissons (ZISCMP) regression. Journal of Statistical Computation and Simulation, 89(9):1649– 1673, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Si:2014:EBF

[SYL^{+}14] Weixin Si, Zhiyong Yuan, Xiangyun Liao, Zhaoliang Duan, Mingui Sun, Pengfei Hu, and Jianhui Zhao. An energybased free boundary asynchronous diffusion model for 3D warping of tissue dynamics. Journal of Statistical Computation and Simulation, 84(6):1280–1296, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Su:2016:ELI

[SZ16] Feng Su and Xiaojun Zhu. Exact likelihood inference for two exponential populations based on a joint generalized Type-I hybrid censored sample. Journal of Statistical Computation and Simulation, 86(7):1342–1362, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shi:2019:PDA

[SZJW19] Yueyong Shi, Zehui Zhou, Yuling Jiao, and Ju Wang. A primal dual active set with continuation algorithm for highdimensional nonconvex SICA-penalized regression. Journal of Statistical Computation and Simulation, 89(5):864–883, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Suesse:2017:CAE

[SZM17] Thomas Suesse and Andrew Zammit-Mangion. Computational aspects of the EM algorithm for spatial econometric models with missing data. Journal of Statistical Computation and Simulation, 87(9):1767–1786, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Song:2014:MNC

[SZS14] Xiao Song, Shaoyun Zhang, and Xuecheng Shi. Measurement of network complexity and capability in command and control system. Journal of Statistical Computation and Simulation, 84(6):1232–1248, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shu:2016:SAW

[SZS16] Lianjie Shu, Ruoyu Zhou, and Yan Su. A self-adjusted weighted likelihood ratio test for global clustering of disease. Journal of Statistical Computation and Simulation, 86(5):996–1009, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

URL http://www.tandfonline.com/doi/abs/10.1080/
00949655.2015.1049604.

Song:2019:RIE

[SZWL19] Yunquan Song, Yanji Zhu, Xiuli Wang, and Lu Lin. Robust inference for estimating equations with nonignorably missing data based on SIR algorithm. Journal of Statistical Computation and Simulation, 89(17):3196–3212, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Toribio:2011:DMI

[TA11] S. G. Toribio and J. H. Albert. Discrepancy measures for item fit analysis in item response theory. Journal of Statistical Computation and Simulation, 81(10):1345–1360, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tak:2017:FCP

[Tak17] H. Tak. Frequency coverage properties of a uniform shrinkage prior distribution. Journal of Statistical Computation and Simulation, 87(15):2929–2939, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tarr:2012:SSP

[Tar12] Garth Tarr. Small sample performance of quantile regression confidence intervals. Journal of Statistical Computation and Simulation, 82(1):81–94, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Trivez:2010:ELS

[TC10] F. Javier Trívez and Beatriz Catalán. Effects of level shifts and temporary changes on the estimation of GARCH models. Journal of Statistical Computation and Simulation, 80(6):667–688, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tahir:2018:INH

[TCA^{+}18] M. H. Tahir, Gauss M. Cordeiro, Sajid Ali, Sanku Dey, and Aroosa Manzoor. The inverted Nadarajah–Haghighi distribution: estimation methods and applications. Journal of Statistical Computation and Simulation, 88(14):2775–2798, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tang:2019:AEM

[TCHS19] Anan Tang, Philippe Castagliola, Xuelong Hu, and Jinsheng Sun. The adaptive EWMA median chart for known and estimated parameters. Journal of Statistical Computation and Simulation, 89(5):844–863, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsai:2014:EBS

[TCLY14] Tzong-Ru Tsai, Jyun-You Chiang, TaChen Liang, and Ming-Chung Yang. Efficient Bayesian sampling plans for exponential distributions with Type-I-censored samples. Journal of Statistical Computation and Simulation, 84(5):964– 981, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Terpstra:2011:USC

[TCM11] Jeff T. Terpstra, Chia-Hao Chang, and Rhonda C. Magel. On the use of Spearman’s correlation coefficient for testing ordered alternatives. Journal of Statistical Computation and Simulation, 81(11):1381–1392, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tchahou:2013:TCV

[TD13] Herbert Nkwimi Tchahou and Pierre Duchesne. On testing for causality in variance between two multivariate time series. Journal of Statistical Computation and Simulation, 83 (11):2064–2092, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taylor:2014:IMT

[TD14] Benjamin M. Taylor and Peter J. Diggle. INLA or MCMC? A tutorial and comparative evaluation for spatial prediction in log-Gaussian Cox processes. Journal of Statistical Computation and Simulation, 84(10):2266–2284, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tang:2019:CDS

[TD19] Fengqin Tang and Wenwen Ding. Community detection with structural and attribute similarities. Journal of Statistical Computation and Simulation, 89(4):668–685, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tomaya:2018:HME

[TdC18] Lorena Cáceres Tomaya and Mário de Castro. A heteroscedastic measurement error model based on skew and heavy-tailed distributions with known error variances. Journal of Statistical Computation and Simulation, 88(11):2185– 2200, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Taconeli:2019:NTS

[TdSC19] Cesar Augusto Taconeli and Angelo da Silva Cabral. New two-stage sampling designs based on neoteric ranked set sampling. Journal of Statistical Computation and Simulation, 89(2):232–248, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tenreiro:2019:AST

[Ten19] Carlos Tenreiro. On the automatic selection of the tuning parameter appearing in certain families of goodness-of-fit tests. Journal of Statistical Computation and Simulation, 89(10):1780–1797, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Toribio:2012:EBA

[TGL12a] S. G. Toribio, B. R. Gray, and S. Liang. An evaluation of the Bayesian approach to fitting the N-mixture model for use with pseudo-replicated count data. Journal of Statistical Computation and Simulation, 82(8):1135–1143, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tunno:2012:ALT

[TGL12b] Ferebee Tunno, Colin Gallagher, and Robert Lund. Arc length tests for equivalent autocovariances. Journal of Statistical Computation and Simulation, 82(12):1799–1812, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tian:2018:RMO

[THCZ18] Haitao Tian, Lei Huang, Ching-Yu Cheng, and Liang Zhang. Regression models with ordered multiple categorical predictors. Journal of Statistical Computation and Simulation, 88(16):3164–3178, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Thomson:2015:ASE

[THG15] T. Thomson, S. Hossain, and M. Ghahramani. Application of shrinkage estimation in linear regression models with autoregressive errors. Journal of Statistical Computation and Simulation, 85(16):3335–3351, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Trucios:2017:RBF

[THR17] Carlos Trucíos, Luiz K. Hotta, and Esther Ruiz. Robust bootstrap forecast densities for GARCH returns and volatilities. Journal of Statistical Computation and Simulation, 87 (16):3152–3174, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Trucios:2018:RBD

[THR18] Carlos Trucíos, Luiz K. Hotta, and Esther Ruiz. Robust bootstrap densities for dynamic conditional correlations: implications for portfolio selection and value-at-risk. Journal of Statistical Computation and Simulation, 88(10):1976– 2000, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tu:2013:CCA

[TJK13] Chunhao Tu, Shuo Jiao, and Woon Yuen Koh. Comparison of clustering algorithms on generalized propensity score in observational studies: a simulation study. Journal of Statistical Computation and Simulation, 83(12):2206–2218, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tu:2014:CAU

[TK14] Chunhao Tu and Woon Yuen Koh. Comparison of adjusted and unadjusted marginal probabilities for estimating binomial parameters in observational studies. Journal of Statistical Computation and Simulation, 84(12):2644–2651, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tripathy:2015:EEC

[TK15] Manas Ranjan Tripathy and Somesh Kumar. Equivariant estimation of common mean of several normal populations. Journal of Statistical Computation and Simulation, 85(18):3679–3699, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tu:2016:CIA

[TK16] Chunhao Tu and Woon Yuen Koh. Causal inference for average treatment effects of multiple treatments with nonnormally distributed outcome variables. Journal of Statistical Computation and Simulation, 86(5):855–861, 2016. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (elec-

tronic), 1563-5163. URL http://www.tandfonline.com/
doi/abs/10.1080/00949655.2015.1040019.

Tanju:2018:CTB

[TK18] Özge Tanju and Zeynep Kalaylioğlu. A cluster tree based model selection approach for logistic regression classifier. Journal of Statistical Computation and Simulation, 88(7): 1394–1414, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tumlinson:2016:LEE

[TKB16] S. E. Tumlinson, J. P. Keating, and N. Balakrishnan. Linear estimation for the extended exponential power distribution. Journal of Statistical Computation and Simulation, 86(7):1392–1403, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Toker:2018:TSL

[TKG18] Selma Toker, Selahattin Kaçiranlar, and Hüseyin Güler. Two-stage Liu estimator in a simultaneous equations model. Journal of Statistical Computation and Simulation, 88(11): 2066–2088, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tu:2013:UGD

[TKJ13] Chunhao Tu, Woon Yuen Koh, and Shuo Jiao. Using generalized doubly robust estimator to estimate average treatment effects of multiple treatments in observational studies. Journal of Statistical Computation and Simulation, 83 (8):1518–1526, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tutkun:2019:DTS

[TKK19] Nihal Ata Tutkun, Nursel Koyuncu, and Ugur Karabey. Discrete-time survival analysis under ranked set sampling: an application to Turkish motor insurance data. Journal of Statistical Computation and Simulation, 89(4):660–667, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tekeli:2019:ODP

[TKÖ19] Erkut Tekeli, Selahattin Kaçıranlar, and Nimet Özbay. Optimal determination of the parameters of some biased estimators using genetic algorithm. Journal of Statistical Computation and Simulation, 89(18):3331–3353, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tojeiro:2014:CWG

[TLRB14] Cynthia Tojeiro, Francisco Louzada, Mari Roman, and Patrick Borges. The complementary Weibull geometric distribution. Journal of Statistical Computation and Simulation, 84(6):1345–1362, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Torabi:2018:WRE

[TMG18] Hamzeh Torabi, Narges H. Montazeri, and Aurea Grané. A wide review on exponentiality tests and two competitive proposals with application on reliability. Journal of Statistical Computation and Simulation, 88(1):108–139, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Teimouri:2016:BCM

[TN16] Mahdi Teimouri and Saralees Nadarajah. Bias corrected MLEs under progressive type-II censoring scheme. Journal of Statistical Computation and Simulation, 86(14):2714– 2726, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Teng:2017:BCL

[TNJ17] Ming Teng, Farouk Nathoo, and Timothy D. Johnson. Bayesian computation for log-Gaussian Cox processes: a comparative analysis of methods. Journal of Statistical Computation and Simulation, 87(11):2227–2252, 2017. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Totterdell:2017:BHM

[TNM17] James A. Totterdell, Darfiana Nur, and Kerrie L. Mengersen. Bayesian hidden Markov models in DNA sequence segmentation using R: the case of Simian Vacuolating virus (SV40). Journal of Statistical Computation and Simulation, 87(14):2799–2827, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Teimouri:2014:EAB

[TNS14] Mahdi Teimouri, Saralees Nadarajah, and Shou Hsing Shih. EM algorithms for beta kernel distributions. Journal of Statistical Computation and Simulation, 84(2):451–467, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tzavelas:2013:SIS

[TP13] George Tzavelas and Demosthenes Panagiotakos. Statistical inference for the size-biased Weibull distribution. Journal of Statistical Computation and Simulation, 83(7):1252–1265, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tissot:2015:ROA

[TP15a] J.-Y. Tissot and C. Prieur. A randomized orthogonal arraybased procedure for the estimation of first- and second-order Sobol’ indices. Journal of Statistical Computation and Simulation, 85(7):1358–1381, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tomer:2015:EPM

[TP15b] Sanjeev K. Tomer and M. S. Panwar. Estimation procedures for Maxwell distribution under type-I progressive hybrid censoring scheme. Journal of Statistical Computation and Simulation, 85(2):339–356, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tavakoli:2017:EES

[TPM17] Masoud Tavakoli, Reza Pourtaheri, and Mohamad B. Moghadam. Economic and economic-statistical designs of VSI Bayesian control chart using Monte Carlo method and ABC algorithm. Journal of Statistical Computation and Simulation, 87(4):766–776, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsagris:2017:NHT

[TPW17] Michail Tsagris, Simon Preston, and Andrew T. A. Wood. Nonparametric hypothesis testing for equality of means on the simplex. Journal of Statistical Computation and Simulation, 87(2):406–422, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Torres:2010:PRT

[TQP10] Patricia S. Torres, Marta B. Quaglino, and Valério D. Pillar. Properties of a randomization test for multifactor comparisons of groups. Journal of Statistical Computation and Simulation, 80(10):1131–1150, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Turley:2018:CTE

[TRC^{+}18] Falynn C. Turley, David Redden, Janice L. Case, Charles Katholi, Jeff Szychowski, and Derek DuBay. Comparison of Type I error rates and statistical power of different propensity score methods. Journal of Statistical Computation and Simulation, 88(4):769–784, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Torabi:2014:HBE

[TS14] Mahmoud Torabi and Farhad Shokoohi. Hierarchical Bayes estimation in small area estimation using cross-sectional and time-series data. Journal of Statistical Computation and Simulation, 84(3):605–613, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Takahashi:2016:TPF

[TS16] Sho Takahashi and Nobumichi Shutoh. Tests for parallelism and flatness hypotheses of two mean vectors in high-dimensional settings. Journal of Statistical Computation and Simulation, 86(6):1150–1165, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1055269.

Tafakori:2017:NCT

[TS17] Laleh Tafakori and A. R. Soltani. A note on the Cauchytype mixture distributions. Journal of Statistical Computation and Simulation, 87(9):1901–1910, 2017. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsai:2010:GCI

[Tsa10a] Jia-Ren Tsai. Generalized confidence interval for the slope in linear measurement error model. Journal of Statistical Computation and Simulation, 80(8):927–936, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsay:2010:MLE

[Tsa10b] Wen-Jen Tsay. Maximum likelihood estimation of stationary multivariate ARFIMA processes. Journal of Statistical Computation and Simulation, 80(7):729–745, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsai:2011:CVM

[Tsa11] Shing Chih Tsai. Control-variate methods for comparison with a standard. Journal of Statistical Computation and Simulation, 81(11):1703–1716, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsai:2018:AEI

[Tsa18] Min-Hsiao Tsai. Analytical expression for the integrated squared density partial derivative of a multivariate normal mixture distribution. Journal of Statistical Computation and Simulation, 88(14):2726–2750, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsiotas:2015:QBM

[Tsi15] Georgios Tsiotas. A quasi-Bayesian model averaging approach for conditional quantile models. Journal of Statistical Computation and Simulation, 85(10):1963–1986, 2015.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsou:2015:URM

[Tso15] Tsung-Shan Tsou. A universal robust method for analysing bivariate continuous and proportion data. Journal of Statistical Computation and Simulation, 85(18):3700–3707, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tsou:2016:EUR

[Tso16] Tsung-Shan Tsou. Expedient universal robust likelihood method for general dispersed count data. Journal of Statistical Computation and Simulation, 86(15):2985–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tian:2019:BBR

[TTWT19] Yuzhu Tian, Manlai Tang, Liyong Wang, and Maozai Tian. Bayesian bridge-randomized penalized quantile regression estimation for linear regression model with AP(q) perturbation. Journal of Statistical Computation and Simulation, 89(15):2951–2979, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tarassenko:2015:SBR

[TTZS15] P. F. Tarassenko, S. S. Tarima, A. V. Zhuravlev, and S. Singh. On sign-based regression quantiles. Journal of Statistical Computation and Simulation, 85(7):1420–1441, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tu:2019:CVM

[Tu19] Chunhao Tu. Comparison of various machine learning algorithms for estimating generalized propensity score. Journal of Statistical Computation and Simulation, 89(4):708–719, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tang:2014:NCD

[TW14] Cheng Yong Tang and Tong Tong Wu. Nested coordinate descent algorithms for empirical likelihood. Journal of Statistical Computation and Simulation, 84(9):1917–1930, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tian:2014:VSS

[TX14] Ruiqin Tian and Liugen Xue. Variable selection for semiparametric errors-in-variables regression model with longitudinal data. Journal of Statistical Computation and Simulation, 84(8):1654–1669, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Trindade:2010:TSM

[TZA10] A. Alexandre Trindade, Yun Zhu, and Beth Andrews. Time series models with asymmetric Laplace innovations. Journal of Statistical Computation and Simulation, 80(12):1317– 1333, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tzavelas:2011:PES

[Tza11] George Tzavelas. Pooled estimators for the shape parameter of the three parameter gamma distribution. Journal of Statistical Computation and Simulation, 81(9):1099–1109, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tahir:2016:PXF

[TZC^{+}16] M. H. Tahir, M. Zubair, Gauss M. Cordeiro, Ayman Alzaatreh, and M. Mansoor. The Poisson-X family of distributions. Journal of Statistical Computation and Simulation, 86(14):2901–2921, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Tabar:2019:LDA

[TZSP19] Vahid Rezaei Tabar, Hamid Zareifard, Selva Salimi, and Dariusz Plewczynski. Learning directed acyclic graphs by determination of candidate causes for discrete variables. Journal of Statistical Computation and Simulation, 89(10): 1957–1970, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ullah:2016:MOG

[UA16] Shafi Ullah and Ravi P. Agarwal. Multi-objective geometric programming approach in multivariate stratified sample survey with quadratic cost function. Journal of Statistical Computation and Simulation, 86(10):1840–1855, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Upadhyay:2010:BAM

[UG10] S. K. Upadhyay and Ashutosh Gupta. A Bayes analysis of modified Weibull distribution via Markov chain Monte Carlo simulation. Journal of Statistical Computation and Simulation, 80(3):241–254, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Upadhyay:2013:BCW

[UGMK13] S. K. Upadhyay, Ashutosh Gupta, Bhaswati Mukherjee, and Ankit Khokhar. A Bayes comparison of Weibull extension and modified Weibull models for data showing bathtub hazard rate. Journal of Statistical Computation and Simulation, 83(1):68–81, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Uranga:2014:LCM

[UM14] R. Uranga and G. Molenberghs. Longitudinal conditional models with intermittent missingness: SAS code and applications. Journal of Statistical Computation and Simulation, 84(4):753–780, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ucer:2012:EGF

[UY12] Burcu H. Ucer and Tugba Ozkal Yildiz. Estimation and goodness-of-fit procedures for Farlie–Gumbel–Morgenstern bivariate copula of order statistics. Journal of Statistical Computation and Simulation, 82(1):137–147, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vujacic:2015:CFA

[VAW15] Ivan Vujacić, Antonino Abbruzzo, and Ernst Wit. A computationally fast alternative to cross-validation in penalized Gaussian graphical models. Journal of Statistical Computation and Simulation, 85(18):3628–3640, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Villejo:2017:RED

[VBL17] Stephen Jun V. Villejo, Erniel B. Barrios, and Joseph Ryan G. Lansangan. Robust estimation of a dynamic spatiotemporal model with structural change. Journal of Statistical Computation and Simulation, 87(3):505–518, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Veerakumari:2015:EAS

[VC15] K. Pradeepa Veerakumari and Vivian Clements. Endorsement of acceptance sampling techniques by implementing neural networks. Journal of Statistical Computation and Simulation, 85(9):1857–1863, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

vanderArk:2015:CCI

[vdAvA15] L. Andries van der Ark and Robbie C. M. van Aert. Comparing confidence intervals for Goodman and Kruskal’s gamma coefficient. Journal of Statistical Computation and Simulation, 85(12):2491–2505, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Volterman:2014:NPF

[VDBA14] William Volterman, Katherine F. Davies, N. Balakrishnan, and Jafar Ahmadi. Nonparametric prediction of future order statistics. Journal of Statistical Computation and Simulation, 84(3):683–695, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vallejo:2010:AUF

[VFLR10] G. Vallejo, M. P. Fernández, and P. E. Livacic-Rojas. Analysis of unbalanced factorial designs with heteroscedastic data. Journal of Statistical Computation and Simulation, 80(1): 75–88, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

vanGinkel:2017:EMI

[vGK17] Joost R. van Ginkel and Pieter M. Kroonenberg. Evaluation of multiple-imputation procedures for three-mode component models. Journal of Statistical Computation and Simulation, 87(16):3059–3081, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Villanueva-Guerra:2017:CCV

[VGTGCFC17] Elena Cristina Villanueva-Guerra, Vìctor Gustavo Tercero-Gómez, Alvaro Eduardo Cordero-Franco, and William Jay Conover. A control chart for variance based on squared ranks. Journal of Statistical Computation and Simulation, 87(18):3537–3562, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

VanderElst:2016:UCS

[VHV^{+}16] W. Van der Elst, L. Hermans, G. Verbeke, M. G. Kenward, V. Nassiri, and G. Molenberghs. Unbalanced cluster sizes and rates of convergence in mixed-effects models for clustered data. Journal of Statistical Computation and Simulation, 86(11):2123–2139, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vonta:2014:GFT

[VK14] Ilia Vonta and Alex Karagrigoriou. Goodness-of-fit tests via φ-measures of divergence for censored data. Journal of Statistical Computation and Simulation, 84(5):946–963, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

VanGinkel:2014:MDP

[VKK14] Joost R. Van Ginkel, Pieter M. Kroonenberg, and Henk A. L. Kiers. Missing data in principal component analysis of questionnaire data: a comparison of methods. Journal of Statistical Computation and Simulation, 84(11):2298–2315, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Varshney:2019:OAM

[VM19] R. Varshney and Mradula. Optimum allocation in multivariate stratified sampling design in the presence of nonresponse with Gamma cost function. Journal of Statistical Computation and Simulation, 89(13):2454–2467, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vitoratou:2014:MLE

[VNM14] Silia Vitoratou, Ioannis Ntzoufras, and Irini Moustaki. Marginal likelihood estimation from the Metropolis output: tips and tricks for efficient implementation in generalized linear latent variable models. Journal of Statistical Computation and Simulation, 84(10):2091–2105, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vanegas:2016:ELS

[VP16] Luis Hernando Vanegas and Gilberto A. Paula. An extension of log-symmetric regression models: R codes and applications. Journal of Statistical Computation and Simulation, 86(9):1709–1735, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vanegas:2013:DTG

[VRC13] Luis Hernando Vanegas, Luz Marina Rondón, and Gauss M. Cordeiro. Diagnostic tools in generalized Weibull linear regression models. Journal of Statistical Computation and Simulation, 83(12):2315–2338, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vehkalahti:2015:SSP

[VS15] Kimmo Vehkalahti and Reijo Sund. Solving Survo puzzles using matrix combinatorial products. Journal of Statistical Computation and Simulation, 85(13):2666–2681, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Venegas:2018:BBG

[VSG^{+}18] Osvaldo Venegas, Hugo S. Salinas, Diego I. Gallardo, Heleno Bolfarine, and Héctor W. Gómez. Bimodality based on the generalized skew-normal distribution. Journal of Statistical Computation and Simulation, 88(1):156–181, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Vo-Van:2017:MGA

[VVNTVD^{+}17] Tai Vo-Van, Trung Nguyen-Thoi, Trung Vo-Duy, Vinh Ho-Huu, and Thao Nguyen-Trang. Modified genetic algorithmbased clustering for probability density functions. Journal of Statistical Computation and Simulation, 87(10):1964–1979, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

VanAelst:2011:SDE

[VVW11] S. Van Aelst, E. Vandervieren, and G. Willems. Stahel– Donoho estimators with cellwise weights. Journal of Statistical Computation and Simulation, 81(1):1–27, 2011. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2015:MPM

[Wan15] Chun-Chao Wang. A MATLAB package for multivariate normality test. Journal of Statistical Computation and Simulation, 85(1):166–188, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2018:BMP

[Wan18a] Hai-Bin Wang. A Bayesian multivariate partially linear single-index probit model for ordinal responses. Journal of Statistical Computation and Simulation, 88(8):1616–1636, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2018:EPG

[Wan18b] Liang Wang. Estimation for the parameters of the generalized half-normal distribution. Journal of Statistical Computation and Simulation, 88(4):629–645, 2018. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Waterman:2015:SNS

[WBAS15] Megan J. Waterman, Jeffrey B. Birch, and Abdel-Salam G. Abdel-Salam. Several nonparametric and semiparametric approaches to linear mixed model regression. Journal of Statistical Computation and Simulation, 85(5):956–977, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2015:RLEb

[WBG15] Xiaolin Wang, Narayanaswamy Balakrishnan, and Bo Guo. Residual life estimation based on nonlinear-multivariate Wiener processes. Journal of Statistical Computation and Simulation, 85(9):1742–1764, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2015:RLEa

[WBGJ15] Xiaolin Wang, Narayanaswamy Balakrishnan, Bo Guo, and Ping Jiang. Residual life estimation based on bivariate nonstationary gamma degradation process. Journal of Statistical Computation and Simulation, 85(2):405–421, 2015. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2014:CTA

[WC14] Shu-Fei Wu and Chien-Ju Chiu. Computational testing algorithmic procedure of assessment for lifetime performance index of products with two-parameter exponential distribution based on the multiply type II censored sample. Journal of Statistical Computation and Simulation, 84(10):2106– 2122, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2017:SDS

[WC17] Xiaodi Wang and Xueping Chen. Symmetrical design for symmetrical global sensitivity analysis of model output. Journal of Statistical Computation and Simulation, 87(14):2738–2759, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2013:MSV

[WCC13] Joanna J. J. Wang, S. T. Boris Choy, and Jennifer S. K. Chan. Modelling stochastic volatility using generalized t distribution. Journal of Statistical Computation and Simulation, 83(2):340–354, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2011:VEB

[WCK11] Yishi Wang, Cuixian Chen, and Fanhui Kong. Variance estimation of the Buckley–James estimator under discrete assumptions. Journal of Statistical Computation and Simulation, 81(4):481–496, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2015:FSS

[WCW15] Chien-Wei Wu, James C. Chen, and Tai-Hsi Wu. A flexible sampling scheme for variables inspection with loss consideration. Journal of Statistical Computation and Simulation, 85(18):3766–3777, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

White:2016:MMC

[WD16] Josie A. White and Katherine F. Davies. Minimal-maximal correlation-type goodness-of-fit tests. Journal of Statistical Computation and Simulation, 86(11):2157–2164, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wan:2015:VSM

[WDCK15] Y. Wan, S. Datta, D. J. Conklin, and M. Kong. Variable selection models based on multiple imputation with an application for predicting median effective dose and maximum effect. Journal of Statistical Computation and Simulation, 85(9):1902–1916, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wong:2018:PPH

[WDY18] George Y. C. Wong, Qinggang Diao, and Qiqing Yu. Piecewise proportional hazards models with interval-censored data. Journal of Statistical Computation and Simulation, 88(1):140–155, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wilcox:2014:CTI

[WEHCC14] Rand R. Wilcox, David M. Erceg-Hurn, Florence Clark, and Michael Carlson. Comparing two independent groups via the lower and upper quantiles. Journal of Statistical Computation and Simulation, 84(7):1543–1551, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Weiss:2011:EMS

[Wei11] Christian H. Weiß. Empirical measures of signed serial dependence in categorical time series. Journal of Statistical Computation and Simulation, 81(4):411–429, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Weiss:2012:PCA

[Wei12] Christian H. Weiß. Process capability analysis for serially dependent processes of Poisson counts. Journal of Statistical Computation and Simulation, 82(3):383–404, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wellek:2016:TNB

[Wel16] S. Wellek. Testing for noninferiority of binomial distributions referring to a modified equivalence region with piecewise linear boundary. Journal of Statistical Computation and Simulation, 86(9):1736–1753, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Westgate:2016:CCA

[Wes16] Philip M. Westgate. A covariance correction that accounts for correlation estimation to improve finite-sample inference with generalized estimating equations: a study on its applicability with structured correlation matrices. Journal of Statistical Computation and Simulation, 86(10):1891–1900, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2018:FBN

[WFC^{+}18] Yuh-Jenn Wu, Wei-Quan Fang, Li-Hsueh Cheng, Kai-Chi Chu, Yin-Tzer Shih, and Li-Chu Chien. A flexible Bayesian non-parametric approach for fitting the odds to case II interval-censored data. Journal of Statistical Computation and Simulation, 88(16):3132–3150, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2014:RLE

[WGC14] Xiaolin Wang, Bo Guo, and Zhijun Cheng. Residual life estimation based on bivariate Wiener degradation process with time-scale transformations. Journal of Statistical Computation and Simulation, 84(3):545–563, 2014. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2014:IPS

[WH14] Fu-Kwun Wang and Ping-Ren Huang. Implementing particle swarm optimization algorithm to estimate the mixture of two Weibull parameters with censored data. Journal of Statistical Computation and Simulation, 84(9):1975–1989, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2017:GPA

[WH17] Tsaipei Wang and Wen-Liang Hung. A generalized possibilistic approach to shell clustering of template-based shapes. Journal of Statistical Computation and Simulation, 87(3):423–436, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2019:OPI

[WH19] Shuo-Jye Wu and Syuan-Rong Huang. Optimal progressive interval censoring plan under accelerated life test with limited budget. Journal of Statistical Computation and Simulation, 89(17):3241–3257, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2017:NSA

[WHX^{+}17] Yi Wang, Liang Han, Song Xiao, Jiangyun Wang, and Xiang Zhai. A novel statistical approach to remove salt-and-pepper noise. Journal of Statistical Computation and Simulation, 87(13):2538–2548, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2019:SES

[WHX19] Wenbo Wu, Haileab Hilafu, and Yuan Xue. Simultaneous estimation for semi-parametric multi-index models. Journal of Statistical Computation and Simulation, 89(12):2354–2372, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wilcox:2010:CNP

[Wil10] Rand R. Wilcox. Comparing non-parametric regression lines via regression depth. Journal of Statistical Computation and Simulation, 80(4):379–387, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Williams:2011:FSC

[Wil11] Dennis C. Williams. Finite sample correction factors for several simple robust estimators of normal standard deviation. Journal of Statistical Computation and Simulation, 81 (11):1697–1702, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wilcox:2015:WGA

[Wil15] Rand R. Wilcox. Within groups analysis of covariance: multiple comparisons at specified design points using a robust measure location when there is curvature. Journal of Statistical Computation and Simulation, 85(16):3236–3246, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wei:2017:MCC

[WJ17] Wei Wei and Wenxin Jiang. On Monte Carlo computation of posterior expectations with uncertainty. Journal of Statistical Computation and Simulation, 87(10):2038–2049, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Weiss:2019:BBB

[WJ19] C. H. Weiß and C. Jentsch. Bootstrap-based bias corrections for INAR count time series. Journal of Statistical Computation and Simulation, 89(7):1248–1264, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2019:WBL

[WKJ^{+}19] Yafei Wang, Linglong Kong, Bei Jiang, Xingcai Zhou, Shimei Yu, Li Zhang, and Giseon Heo. Wavelet-based LASSO in functional linear quantile regression. Journal of Statistical Computation and Simulation, 89(6):1111–1130, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2010:NAD

[WL10] Shu-Fei Wu and Mei-Chu Liang. A note on the asymptotic distribution of the process capability index C_{pmk}. Journal of Statistical Computation and Simulation, 80(2):227–235, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2014:OAC

[WL14] Shu-Fei Wu and Ying-Po Lin. The optimal allocation combination for the two-sided sequential screening procedure based on the individual misclassification error. Journal of Statistical Computation and Simulation, 84(9):2025–2043, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2015:SSE

[WL15a] Kangning Wang and Lu Lin. Simultaneous structure estimation and variable selection in partial linear varying coefficient models for longitudinal data. Journal of Statistical Computation and Simulation, 85(7):1459–1473, 2015. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2015:BAM

[WL15b] Wan-Lun Wang and Tsung-I Lin. Bayesian analysis of multivariate t linear mixed models with missing responses at random. Journal of Statistical Computation and Simulation, 85(17):3594–3612, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2015:BIP

[WL15c] Yan-Ping Wang and Zhen-Zhou Lu. Bayesian inference and prediction analysis of the power law process based on a natural conjugate prior. Journal of Statistical Computation and Simulation, 85(5):881–898, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2016:EVR

[WL16] Pan Wang and Zhenzhou Lu. Efficient variance reduction approach based on the variation of the input importance. Journal of Statistical Computation and Simulation, 86(14):2856–2873, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2017:SIV

[WL17] Wentao Wang and Dengkui Li. Structure identification and variable selection in geographically weighted regression models. Journal of Statistical Computation and Simulation, 87(10):2050–2068, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2019:MAG

[WL19] Yi-Hua Tina Wang and Yunru Lai. Monitoring of autocorrelated general linear profiles. Journal of Statistical Computation and Simulation, 89(3):519–535, 2019. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2018:MOD

[WLCL18] Tao Wang, Qun Li, Bin Chen, and Zhonghua Li. Multiple outliers detection in sparse high-dimensional regression. Journal of Statistical Computation and Simulation, 88(1): 89–107, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2012:CCG

[WLL12] Shu-Fei Wu, Kuo-Ren Lou, and Ying-Po Lin. Computational comparison of the general weighted moments estimators of the scale parameter of a Pareto distribution with a known shape parameter with other estimators based on a multiply Type II-censored sample. Journal of Statistical Computation and Simulation, 82(12):1813–1825, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2018:TIH

[WLSZ18] Hao Wang, Baisen Liu, Ning-Zhong Shi, and Shurong Zheng. Testing identity of high-dimensional covariance matrix. Journal of Statistical Computation and Simulation, 88 (13):2600–2611, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Weng:2012:PCI

[WM12] Shih-Feng Weng and Leann Myers. Performance of confidence interval tests for the ratio of two lognormal means applied to Weibull and gamma distribution data. Journal of Statistical Computation and Simulation, 82(7):1035–1054, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wagler:2015:ISI

[WM15] Amy Wagler and Melinda McCann. Improved simultaneous intervals for linear combinations of parameters from generalized linear models. Journal of Statistical Computation and Simulation, 85(12):2514–2532, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Walker:2011:SSA

[WMDO11] C. G. Walker, M. L. Mackenzie, C. R. Donovan, and M. J. O’Sullivan. SALSA — a spatially adaptive local smoothing algorithm. Journal of Statistical Computation and Simulation, 81(2):179–191, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Withers:2011:AOS

[WN11a] Christopher S. Withers and Saralees Nadarajah. Asymptotically optimal shrinkage estimates for non-normal data. Journal of Statistical Computation and Simulation, 81(12): 2021–2037, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Withers:2011:BRR

[WN11b] Christopher S. Withers and Saralees Nadarajah. Bias reduction for the ratio of means. Journal of Statistical Computation and Simulation, 81(12):1799–1816, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Withers:2011:CIL

[WN11c] Christopher S. Withers and Saralees Nadarajah. Confidence intervals for the length of a vector mean. Journal of Statistical Computation and Simulation, 81(5):591–605, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Withers:2012:CIC

[WN12] Christopher S. Withers and Saralees Nadarajah. Confidence intervals for the correlation from a bivariate normal. Journal of Statistical Computation and Simulation, 82(11):1591– 1606, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Withers:2013:CIL

[WN13] Christopher S. Withers and Saralees Nadarajah. Confidence intervals for lognormal regression and a non-parametric alternative. Journal of Statistical Computation and Simulation, 83(5):880–895, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wiegand:2018:AML

[WN18] Martin Wiegand and Saralees Nadarajah. Approximation methods for lognormal characteristic functions. Journal of Statistical Computation and Simulation, 88(18):3650–3663, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Woods:2010:RDB

[Woo10] D. C. Woods. Robust designs for binary data: applications of simulated annealing. Journal of Statistical Computation and Simulation, 80(1):29–41, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2015:PTT

[WPC15] C. H. Wu, W. L. Pearn, and C. C. Chuang. A powerful test for two-sided multiple independent characteristics supplier selection problem. Journal of Statistical Computation and Simulation, 85(4):839–853, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2014:AOD

[WPXL14] Lei Wang, Xiaolong Pu, Dongdong Xiang, and Yan Li. Asymptotic optimality of double sequential mixture likelihood ratio test. Journal of Statistical Computation and Simulation, 84(4):916–929, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Welz:2018:NPM

[WRN18] Anke Welz, Graeme D. Ruxton, and Markus Neuhäuser. A non-parametric maximum test for the Behrens–Fisher problem. Journal of Statistical Computation and Simulation, 88 (7):1336–1347, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wilcox:2018:IMM

[WRP18] Rand R. Wilcox, Guillaume A. Rousselet, and Cyril R. Pernet. Improved methods for making inferences about multiple skipped correlations. Journal of Statistical Computation and Simulation, 88(16):3116–3131, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wilhelm:2016:GSR

[WS16] Matthieu Wilhelm and Laura M. Sangalli. Generalized spatial regression with differential regularization. Journal of Statistical Computation and Simulation, 86(13):2497–2518, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2018:TSN

[WSC18] Yishi Wang, Ann E. Stapleton, and Cuixian Chen. Twosample nonparametric stochastic order inference with an application in plant physiology. Journal of Statistical Computation and Simulation, 88(14):2668–2683, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2017:NMB

[WSLX17] Ronghua Wang, Naijun Sha, Fangfang Liu, and Xiaoling Xu. A new mixed bivariate geometric-exponential life distribution with applications to series systems. Journal of Statistical Computation and Simulation, 87(2):250–266, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2018:ENA

[WT18] Yan Wang and Lu Tian. An efficient numerical algorithm for exact inference in meta analysis. Journal of Statistical Computation and Simulation, 88(4):646–656, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2017:MOF

[WTJW17] Chun Wang, Na Tian, Zhicheng Ji, and Yan Wang. Multiobjective fuzzy flexible job shop scheduling using memetic algorithm. Journal of Statistical Computation and Simulation, 87(14):2828–2846, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2010:IEP

[Wu10] Shu-Fei Wu. Interval estimation for the Pareto distribution based on the progressive Type II censored sample. Journal of Statistical Computation and Simulation, 80(4):463–474, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2011:UNA

[Wu11] Chien-Wei Wu. Using a novel approach to assess process performance in the presence of measurement errors. Journal of Statistical Computation and Simulation, 81(3):301–314, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2012:NMA

[Wu12] Zhengxiao Wu. A note on the monotonicity of the ES algorithm. Journal of Statistical Computation and Simulation, 82(5):759–761, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2013:LPS

[Wu13] Tong Tong Wu. Lasso penalized semiparametric regression on high-dimensional recurrent event data via coordinate descent. Journal of Statistical Computation and Simulation, 83(6):1145–1155, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2016:PDB

[Wu16a] Jibo Wu. Performance of the difference-based almost unbiased Liu estimator in partial linear model. Journal of Statistical Computation and Simulation, 86(14):2874–2887, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2016:NOS

[Wu16b] Shu-Fei Wu. New one-stage multiple comparison procedures with the average for exponential location parameters under heteroscedasticity. Journal of Statistical Computation and Simulation, 86(14):2740–2748, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2016:QHM

[Wu16c] Zhengxiao Wu. Quasi-hidden Markov model and its applications in change-point problems. Journal of Statistical Computation and Simulation, 86(12):2273–2290, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2012:PRT

[WW12] Chien-Hua Wu and Shu-Mei Wan. The precision of regression-type estimator for incremental cost-effectiveness ratio. Journal of Statistical Computation and Simulation, 82(8):1105–1114, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2016:FGL

[WW16a] Bin Wang and Xiao-Feng Wang. Fitting the generalized lambda distribution to pre-binned data. Journal of Statistical Computation and Simulation, 86(9):1785–1797, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2016:MMS

[WW16b] Yi-Hua Tina Wang and Hsiuying Wang. On the monitoring of mixture simple linear profiles. Journal of Statistical Computation and Simulation, 86(15):3009–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2019:ISP

[WW19] Zih-Huei Wang and Chien-Wei Wu. An improved sampling plan by variables inspection with consideration of process yield and quality loss. Journal of Statistical Computation and Simulation, 89(13):2395–2409, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2018:VBI

[WWB18] Xinyang Wu, Xiaoyue Wu, and Narayanaswamy Balakrishnan. Variance-based importance analysis measure for mission reliability of phased mission system. Journal of Statistical Computation and Simulation, 88(5):841–868, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2011:SIT

[WWCL11] Shu-Fei Wu, Chin-Chuan Wu, Chi-Hsiang Chou, and Huan-Min Lin. Statistical inferences of a two-parameter distribution with the bathtub shape based on progressive censored sample. Journal of Statistical Computation and Simulation, 81(3):315–329, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2015:BNE

[WWW15] Juan Wu, Xue Wang, and Stephen G. Walker. Bayesian nonparametric estimation of a copula. Journal of Statistical Computation and Simulation, 85(1):103–116, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wei:2017:IVD

[WWW17] Yuan Wei, Shanshan Wang, and Huiwen Wang. Intervalvalued data regression using partial linear model. Journal of Statistical Computation and Simulation, 87(16):3175–3194, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2011:HLA

[WXF11] Ningning Wang, Shuyi Xu, and Jiqian Fang. Hierarchical likelihood approach for the Weibull frailty model. Journal of Statistical Computation and Simulation, 81(3):343–356, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2011:SCI

[WY11] Shu-Fei Wu and Yuh-Ru Yu. The simultaneous confidence intervals for all distances from the extreme populations for two-parameter exponential populations based on the multiply Type II censored samples. Journal of Statistical Computation and Simulation, 81(2):137–165, 2011. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2019:IEG

[WY19] Bing Xing Wang and Qian Yu. Interval estimation for the generalized inverted exponential distribution under progressive first failure censoring. Journal of Statistical Computation and Simulation, 89(7):1265–1277, 2019. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wei:2017:IMO

[WyT17] Wei Wei and Zhen yu Tian. An improved multi-objective optimization method based on adaptive mutation particle swarm optimization and fuzzy statistics algorithm. Journal of Statistical Computation and Simulation, 87(13):2480– 2493, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2012:RPM

[WYW12] Jiaping Wang, Qiqing Yu, and George Y. C. Wong. The random partition masking model for interval-censored and masked competing risks data. Journal of Statistical Computation and Simulation, 82(7):981–1002, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2017:CER

[WYX17] Jingjing Wu, Weixin Yao, and Sijia Xiang. Computation of an efficient and robust estimator in a semiparametric mixture model. Journal of Statistical Computation and Simulation, 87(11):2128–2137, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2016:SAN

[WYZK16] Xiaodi Wang, Ming Yang, Yingshan Zhang, and Melody Kiang. A screening approach for non-parametric global sensitivity analysis. Journal of Statistical Computation and Simulation, 86(4):656–675, 2016. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-

5163. URL http://www.tandfonline.com/doi/abs/10.
1080/00949655.2015.1027895.

Wan:2013:SRS

[WZ13] Shuwen Wan and Biao Zhang. Semiparametric ROC surface estimation for continuous diagnostic tests via polytomous logistic regression procedures. Journal of Statistical Computation and Simulation, 83(12):2195–2205, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2018:MHD

[WZ18] Jingjing Wu and Xiaofan Zhou. Minimum Hellinger distance estimation for a semiparametric location-shifted mixture model. Journal of Statistical Computation and Simulation, 88(13):2507–2527, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2019:SEB

[WZ19] Qing Wang and Adriano Z. Zambom. Subsamplingextrapolation bandwidth selection in bivariate kernel density estimation. Journal of Statistical Computation and Simulation, 89(9):1740–1759, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wang:2017:DMA

[WZS17] Xiaofeng Wang, Yingzhi Zhang, and Guixiang Shen. Degradation model analysis of two-component series and parallel systems. Journal of Statistical Computation and Simulation, 87(13):2639–2656, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wu:2013:VSJ

[WZX13] Liu-Cang Wu, Zhong-Zhan Zhang, and Deng-Ke Xu. Variable selection in joint location and scale models of the skewnormal distribution. Journal of Statistical Computation and Simulation, 83(7):1266–1278, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2018:FRA

[XBL18] Xingjian Xu, Tian Ban, and Yuehua Li. Flexible reliability assessment of digital circuits based on signal probability. Journal of Statistical Computation and Simulation, 88 (13):2528–2539, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2011:RGS

[XG11] Chonggang Xu and George Zdzislaw Gertner. Reliability of global sensitivity indices. Journal of Statistical Computation and Simulation, 81(12):1939–1969, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xie:2017:GAC

[XHEM17] Yimeng Xie, Yili Hong, Luis A. Escobar, and William Q. Meeker. A general algorithm for computing simultaneous prediction intervals for the (log)-location-scale family of distributions. Journal of Statistical Computation and Simulation, 87(8):1559–1576, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xiao:2014:SIA

[XHYX14] Xun Xiao, Qingpei Hu, Dan Yu, and Min Xie. Study of an imputation algorithm for the analysis of interval-censored data. Journal of Statistical Computation and Simulation, 84(3):477–490, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xie:2014:RAT

[Xie14] Changchun Xie. Relations among three parametric multiple testing methods for correlated tests. Journal of Statistical Computation and Simulation, 84(4):812–818, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xiao:2012:RAC

[XLB12] Yuanhui Xiao, Jiawei Liu, and Madhusudan Bhandary. Resampling approaches for common intraclass correlation coefficients. Journal of Statistical Computation and Simulation, 82(9):1357–1366, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2019:CBF

[XLMX19] X. Xu, P. Lu, S. N. MacEachern, and R. Xu. Calibrated Bayes factors for model comparison. Journal of Statistical Computation and Simulation, 89(4):591–614, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xie:2010:IAA

[XLW10] Feng-Chang Xie, Jin-Guan Lin, and Bo-Cheng Wei. Influence analysis of additive mixed-effects nonlinear regression models via EM algorithm. Journal of Statistical Computation and Simulation, 80(10):1115–1129, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2018:SDR

[XLZ18] Xinyi Xu, Xiangjie Li, and Jingxiao Zhang. Sufficient dimension reduction and prediction through cumulative slicing PFC. Journal of Statistical Computation and Simulation, 88(6):1172–1190, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2014:PBT

[XMC^{+}14] Li-Wen Xu, Bo Mei, Ran-Ran Chen, Hong-Xia Guo, and Jia jie Wang. Parametric bootstrap tests for unbalanced nested designs under heteroscedasticity. Journal of Statistical Computation and Simulation, 84(9):2059–2070, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xiong:2017:AMI

[XSF17] Xiaoyu Xiong, Václav Smídl, and Maurizio Filippone. Adaptive multiple importance sampling for Gaussian processes. Journal of Statistical Computation and Simulation, 87(8):1644–1665, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2017:NNT

[Xu17a] Kai Xu. A new nonparametric test for high-dimensional regression coefficients. Journal of Statistical Computation and Simulation, 87(5):855–869, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2017:TDH

[Xu17b] Kai Xu. Testing diagonality of high-dimensional covariance matrix under non-normality. Journal of Statistical Computation and Simulation, 87(16):3208–3224, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2018:RGE

[XWMA18] Yaqing Xu, Mengyun Wu, Shuangge Ma, and Syed Ejaz Ahmed. Robust gene-environment interaction analysis using penalized trimmed regression. Journal of Statistical Computation and Simulation, 88(18):3502–3528, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2015:MSM

[XWZS15] Lin Xu, You-Gan Wang, Shurong Zheng, and Ning-Zhong Shi. Model selection with misspecified spatial covariance structure. Journal of Statistical Computation and Simulation, 85(11):2276–2294, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xiao:2015:MSA

[XX15] Nan Xiao and Qing-Song Xu. Multi-step adaptive elasticnet: reducing false positives in high-dimensional variable selection. Journal of Statistical Computation and Simulation, 85(18):3755–3765, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2011:RCE

[XY11] Jianwen Xu and Hu Yang. On the restricted r–k class estimator and the restricted r–d class estimator in linear regression. Journal of Statistical Computation and Simulation, 81(6):679–691, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xiang:2016:NIC

[XY16] Sijia Xiang and Weixin Yao. A new information criterionbased bandwidth selection method for non-parametric regressions. Journal of Statistical Computation and Simulation, 86(17):3446–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xiuyun:2019:RAB

[XYZ19] Peng Xiuyun, Xiao Yan, and Yan Zaizai. Reliability analysis of Birnbaum–Saunders model based on progressive type-II censoring. Journal of Statistical Computation and Simulation, 89(3):461–477, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2019:LCP

[XZ19] Da Xu and Yong Zhou. Local composite partial likelihood estimation for length-biased and right-censored data. Journal of Statistical Computation and Simulation, 89(14):2661– 2677, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2015:SNS

[XZD15] Dengke Xu, Zhongzhan Zhang, and Jiang Du. Skew-normal semiparametric varying coefficient model and score test. Journal of Statistical Computation and Simulation, 85(2): 216–234, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xu:2013:CCG

[XZY13] Jinfeng Xu, Gang Zheng, and Ao Yuan. Case-control genome-wide joint association study using semiparametric empirical model and approximate Bayes factor. Journal of Statistical Computation and Simulation, 83(7):1191–1209, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Xiao:2015:PID

[XZZ15] Weilin Xiao, Weiguo Zhang, and Xili Zhang. Parameter identification for the discretely observed geometric fractional Brownian motion. Journal of Statistical Computation and Simulation, 85(2):269–283, 2015. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2016:NAM

[YA16a] Su-Fen Yang and Barry C. Arnold. A new approach for monitoring process variance. Journal of Statistical Computation and Simulation, 86(14):2749–2765, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yuzbasi:2016:SPE

[YA16b] Bahadir Yüzbasi and S. Ejaz Ahmed. Shrinkage and penalized estimation in semi-parametric models with multicollinear data. Journal of Statistical Computation and Simulation, 86(17):3543–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yolcu:2013:TSF

[YAEU13] Ufuk Yolcu, Cagdas Hakan Aladag, Erol Egrioglu, and Vedide R. Uslu. Time-series forecasting with a novel fuzzy time-series approach: an example for Istanbul stock market. Journal of Statistical Computation and Simulation, 83(4):599–612, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yamazaki:2019:TSM

[Yam19] Daisuke Yamazaki. Testing for shifts in mean with monotonic power against multiple structural changes. Journal of Statistical Computation and Simulation, 89(11):2006–2030, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2010:DFC

[Yan10] James J. Yang. Distribution of Fisher’s combination statistic when the tests are dependent. Journal of Statistical Computation and Simulation, 80(1):1–12, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yao:2015:LSS

[Yao15] Weixin Yao. Label switching and its solutions for frequentist mixture models. Journal of Statistical Computation and Simulation, 85(5):1000–1012, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yousfi:2015:OBM

[YBAA15] Smail Yousfi, Rachid Boumaza, Djamil Aissani, and Smail Adjabi. Optimal bandwidth matrices in functional principal component analysis of density functions. Journal of Statistical Computation and Simulation, 85(11):2315–2330, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yen:2015:RVA

[YCA15] Ching-Ho Yen, Chia-Hao Chang, and Muhammad Aslam. Repetitive variable acceptance sampling plan for one-sided specification. Journal of Statistical Computation and Simulation, 85(6):1102–1116, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2015:SBA

[YCD15] Mingan Yang, Lihua Chen, and Guanghui Dong. Semiparametric Bayesian accelerated failure time model with interval-censored data. Journal of Statistical Computation and Simulation, 85(10):2049–2058, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ye:2014:SIE

[YCXN14] Zhi-Sheng Ye, Ping-Shing Chan, Min Xie, and Hon Keung Tony Ng. Statistical inference for the extreme value distribution under adaptive Type-II progressive censoring schemes. Journal of Statistical Computation and Simulation, 84(5):1099–1114, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2016:GPR

[YD16] Qiqing Yu and Junyi Dong. Generation of pseudo random numbers and estimation under Cox models with timedependent covariates. Journal of Statistical Computation and Simulation, 86(14):2727–2739, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ye:2016:SAS

[Ye16] Lu Ye. Saddlepoint approximations for subordinator processes. Journal of Statistical Computation and Simulation, 86(11):2053–2072, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yurdusev:2010:GRN

[YFT10] Mehmet Ali Yurdusev, Mahmut Firat, and Mustafa Erkan Turan. Generalized regression neural networks for municipal water consumption prediction. Journal of Statistical Computation and Simulation, 80(4):477–478, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2014:ISI

[YGX14] Wei Yu, Xu Guo, and Wangli Xu. An improved score interval with a modified midpoint for a binomial proportion. Journal of Statistical Computation and Simulation, 84 (5):1022–1038, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [AH16].

Yang:2018:FSU

[YHWS18] Guangren Yang, Sumin Hou, Luheng Wang, and Yanqing Sun. Feature screening in ultrahigh-dimensional additive Cox model. Journal of Statistical Computation and Simulation, 88(6):1117–1133, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yamazaki:2015:TPC

[YK15] Daisuke Yamazaki and Eiji Kurozumi. Testing for parameter constancy in the time series direction in panel data models. Journal of Statistical Computation and Simulation, 85(14):2874–2902, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yao:2014:OBM

[YL14] Weixin Yao and Longhai Li. An online Bayesian mixture labelling method by minimizing deviance of classification probabilities to reference labels. Journal of Statistical Computation and Simulation, 84(2):310–323, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2015:CDR

[YL15] Ye Yang and Roderick Little. A comparison of doubly robust estimators of the mean with missing data. Journal of Statistical Computation and Simulation, 85(16):3383–3403, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yan:2018:GFI

[YL18] Liang Yan and Xuhua Liu. Generalized fiducial inference for generalized exponential distribution. Journal of Statistical Computation and Simulation, 88(7):1369–1381, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2015:RVS

[YLG15] Hu Yang, Jing Lv, and Chaohui Guo. Robust variable selection in modal varying-coefficient models with longitudinal. Journal of Statistical Computation and Simulation, 85 (15):3064–3079, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2017:FGE

[YLL17] Ming Yang, Zhizhao Liu, and Wei Li. A fast general extension algorithm of latin hypercube sampling. Journal of Statistical Computation and Simulation, 87(17):3398–3411, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ye:2014:IRB

[YML14] Ren-Dao Ye, Tie-Feng Ma, and Kun Luo. Inferences on the reliability in balanced and unbalanced one-way random models. Journal of Statistical Computation and Simulation, 84(5):1136–1153, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yoo:2013:CST

[Yoo13] Jae Keun Yoo. Chi-squared tests in kth-moment sufficient dimension reduction. Journal of Statistical Computation and Simulation, 83(1):191–201, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yoshiba:2018:MLE

[Yos18] Toshinao Yoshiba. Maximum likelihood estimation of skewt copulas with its applications to stock returns. Journal of Statistical Computation and Simulation, 88(13):2489–2506, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Young:2014:PAN

[You14] Derek S. Young. A procedure for approximate negative binomial tolerance intervals. Journal of Statistical Computation and Simulation, 84(2):438–450, 2014. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2010:EPS

[YP10] Xing Yang and Nabendu Pal. Estimation of a population size through capture-mark-recapture method: a comparison of various point and interval estimators. Journal of Statistical Computation and Simulation, 80(3):335–354, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yoder:2017:SSK

[YP17] Jordan Yoder and Carey E. Priebe. Semi-supervised kmeans++. Journal of Statistical Computation and Simulation, 87(13):2597–2608, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2011:CSG

[YPAC11] Xing Yang, Nabendu Pal, Azmy S. Ackleh, and Jacoby Carter. A case study of green tree frog population size estimation by repeated capture-mark-recapture method with individual tagging: a parametric bootstrap method vs.

Jolly–Seber method. Journal of Statistical Computation and Simulation, 81(12):1879–1895, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yoon:2013:PRM

[YPL13] Young Joo Yoon, Cheolwoo Park, and Taewook Lee. Penalized regression models with autoregressive error terms. Journal of Statistical Computation and Simulation, 83(9): 1756–1772, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

You:2019:FCD

[YQ19] Lu You and Peihua Qiu. Fast computing for dynamic screening systems when analyzing correlated data. Journal of Statistical Computation and Simulation, 89(3):379–394, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yenigun:2015:VSR

[YR15] C. Deniz Yenigün and Maria L. Rizzo. Variable selection in regression using maximal correlation and distance correlation. Journal of Statistical Computation and Simulation, 85(8):1692–1705, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yap:2011:CVT

[YS11] B. W. Yap and C. H. Sim. Comparisons of various types of normality tests. Journal of Statistical Computation and Simulation, 81(12):2141–2155, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2013:RPS

[YS13] Dan Yang and Dylan S. Small. An R package and a study of methods for computing empirical likelihood. Journal of Statistical Computation and Simulation, 83(7):1363–1372, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yildiz:2018:PSG

[YS18] Tugba Ozkal Yildiz and Yusuf Can Sevil. Performances of some goodness-of-fit tests for sampling designs in ranked set sampling. Journal of Statistical Computation and Simulation, 88(9):1702–1716, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Younis:2019:GRT

[YS19] Faryal Younis and Javid Shabbir. Generalized ratio-type and ratio-exponential-type estimators for population mean under modified Horvitz–Thompson estimator in adaptive cluster sampling. Journal of Statistical Computation and Simulation, 89(8):1505–1515, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yura:2014:RDS

[YTNT14] Yoshihiro Yura, Hideki Takayasu, Kazuyuki Nakamura, and Misako Takayasu. Rapid detection of the switching point in a financial market structure using the particle filter. Journal of Statistical Computation and Simulation, 84(10):2073– 2090, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2011:EAS

[Yu11] Binbing Yu. Estimating age-specific incidence of dementia using prevalent cohort data. Journal of Statistical Computation and Simulation, 81(8):973–983, 2011. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2015:BQR

[Yu15] Yang Yu. Bayesian quantile regression for hierarchical linear models. Journal of Statistical Computation and Simulation, 85(17):3451–3467, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yuan:2015:RGB

[Yua15] Sen Yuan. Random gradient boosting for predicting conditional quantiles. Journal of Statistical Computation and Simulation, 85(18):3716–3726, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yucel:2017:IND

[Yuc17] Recai Yucel. Impact of the non-distinctness and nonignorability on the inference by multiple imputation in multivariate multilevel data: a simulation assessment. Journal of Statistical Computation and Simulation, 87(9):1813– 1826, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yuan:2014:EMC

[YW14] J. Yuan and G. Wei. An efficient Monte Carlo EM algorithm for Bayesian lasso. Journal of Statistical Computation and Simulation, 84(10):2166–2186, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2018:TAA

[YWL18] Kai Yang, Dehui Wang, and Han Li. Threshold autoregression analysis for finite-range time series of counts with an application on measles data. Journal of Statistical Computation and Simulation, 88(3):597–614, 2018. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2018:TBO

[YWZ18] Xiaona Yang, Zhaojun Wang, and Xuemin Zi. Thresholdingbased outlier detection for high-dimensional data. Journal of Statistical Computation and Simulation, 88(11):2170– 2184, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2016:CIE

[YXZ16] Wei Yu, Wangli Xu, and Lixing Zhu. Confidence interval estimation for negative binomial group distribution. Journal of Statistical Computation and Simulation, 86(3):524–534, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2019:MPT

[YXZ19] Wei Yu, Wangli Xu, and Lixing Zhu. Multiple permutation test for high-dimensional data: a components-combined algorithm. Journal of Statistical Computation and Simulation, 89(4):686–707, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2016:JMM

[YYG16] Lili Yang, Menggang Yu, and Sujuan Gao. Joint models for multiple longitudinal processes and time-to-event outcome. Journal of Statistical Computation and Simulation, 86(18):

3682–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2015:SPM

[YYW15] Xiaobin Yang, Keying Ye, and Yanping Wang. A study of the probit model with latent variables in Phase I clinical trials. Journal of Statistical Computation and Simulation, 85(8):1621–1633, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2014:CMD

[YZ14] Yi Yang and Hui Zou. A coordinate majorization descent algorithm for l_{1} penalized learning. Journal of Statistical Computation and Simulation, 84(1):84–95, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2015:NME

[YZ15] Yi Yang and Hui Zou. Nonparametric multiple expectile regression via ER-Boost. Journal of Statistical Computation and Simulation, 85(7):1442–1458, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2017:MLC

[YZL17] Su-Fen Yang, Ruoyu Zhou, and Shan-Wen Lu. A median loss control chart for monitoring quality loss under skewed distributions. Journal of Statistical Computation and Simulation, 87(17):3241–3260, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2019:BNM

[YZWM19] Yang Yu, Zhihong Zou, Shanshan Wang, and Renate Meyer. Bayesian nonparametric modelling of the link function in the single-index model using a Bernstein–Dirichlet process prior. Journal of Statistical Computation and Simulation, 89(17):3290–3312, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yang:2018:ACQ

[YZX18] Ke Yang, Liping Zhu, and Wangli Xu. Adaptive composite quantile regressions and their asymptotic relative efficiency. Journal of Statistical Computation and Simulation, 88(5):

900–919, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zamanzade:2011:GFT

[ZA11] Ehsan Zamanzade and Nasser Reza Arghami. Goodnessof-fit test based on correcting moments of modified entropy estimator. Journal of Statistical Computation and Simulation, 81(12):2077–2093, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zamanzade:2012:TNB

[ZA12a] Ehsan Zamanzade and Nasser Reza Arghami. Testing normality based on new entropy estimators. Journal of Statistical Computation and Simulation, 82(11):1701–1713, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2012:ACE

[ZA12b] Jian Zhang and Weilian An. Assessing circular error probable when the errors are elliptical normal. Journal of Statistical Computation and Simulation, 82(4):565–586, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zougab:2013:ASA

[ZAK13] N. Zougab, S. Adjabi, and C. C. Kokonendji. Adaptive smoothing in associated kernel discrete functions estimation using Bayesian approach. Journal of Statistical Computation and Simulation, 83(12):2219–2231, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zamanzade:2015:TUB

[Zam15] Ehsan Zamanzade. Testing uniformity based on new entropy estimators. Journal of Statistical Computation and Simulation, 85(16):3191–3205, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zardasht:2017:CIM

[Zar17] Vali Zardasht. On confidence intervals for the mean past lifetime function under random censorship. Journal of Statistical Computation and Simulation, 87(11):2153–2160, 2017.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zardasht:2019:TDO

[Zar19] V. Zardasht. Testing the dilation order by using cumulative residual Tsallis entropy. Journal of Statistical Computation and Simulation, 89(8):1516–1525, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhen:2010:EBM

[ZB10] X. Zhen and I. V. Basawa. Estimation for binary models generated by Gaussian autoregressive processes. Journal of Statistical Computation and Simulation, 80(9):1041–1051, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhu:2013:MCB

[ZB13] Qi Zhu and Tomasz Burzykowski. A Markov-chain-based regression model with random effects for the analysis of ^{18}Olabelled mass spectra. Journal of Statistical Computation and Simulation, 83(1):145–157, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhao:2018:SSP

[ZBG18] Qian Zhao, Vytaras Brazauskas, and Jugal Ghorai. Smallsample performance of the MTM and MWM estimators for the parameters of log-location-scale families. Journal of Statistical Computation and Simulation, 88(4):808–824, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2013:AFL

[ZC13] Guangxiang Zhang and John J. Chen. Adaptive fitting of linear mixed-effects models with correlated random effects. Journal of Statistical Computation and Simulation, 83(12):2291–2314, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zavaleta:2019:LBT

[ZCL19] Katherine E. C. Zavaleta, Vicente G. Cancho, and Artur J. Lemonte. Likelihood-based tests in zero-inflated power series models. Journal of Statistical Computation and Simulation, 89(3):443–460, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2017:SMO

[ZCW^{+}17] Yidu Zhang, Long Chen, Qiong Wu, Zhengsheng Chen, and Youyun Peng. A statistical multivariable optimization method using improved orthogonal algorithm based on large data. Journal of Statistical Computation and Simulation, 87 (14):2657–2667, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2015:NRE

[ZCZ15] Guoyi Zhang, Fletcher Christensen, and Wei Zheng. Nonparametric regression estimators in complex surveys. Journal of Statistical Computation and Simulation, 85(5):1026– 1034, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2017:CAA

[ZCZ17] Jun Zhang, Qian Chen, and Nanguang Zhou. Correlation analysis with additive distortion measurement errors. Journal of Statistical Computation and Simulation, 87(4):664– 688, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhu:2011:VSE

[ZF11] Mu Zhu and Guangzhe Fan. Variable selection by ensembles for the Cox model. Journal of Statistical Computation and Simulation, 81(12):1983–1992, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhu:2016:RAM

[ZF16] W. Zhu and Y. Fan. Relabelling algorithms for mixture models with applications for large data sets. Journal of Statistical Computation and Simulation, 86(2):394–413, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zheng:2018:VSL

[ZFZQ18] Xueying Zheng, Bo Fu, Jiajia Zhang, and Guoyou Qin. Variable selection for longitudinal data with high-dimensional covariates and dropouts. Journal of Statistical Computation and Simulation, 88(4):712–725, 2018. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhou:2014:LRB

[ZGW14] Maoyuan Zhou, Wei Geng, and Zhaojun Wang. Likelihood ratio-based distribution-free sequential change-point detection. Journal of Statistical Computation and Simulation, 84 (12):2748–2758, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zevallos:2012:IOG

[ZH12] Mauricio Zevallos and Luiz K. Hotta. Influential observations in GARCH models. Journal of Statistical Computation and Simulation, 82(11):1571–1589, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2011:EST

[Zha11] Shibin Zhang. Exact simulation of tempered stable Ornstein–Uhlenbeck processes. Journal of Statistical Computation and Simulation, 81(11):1533–1544, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2014:EBE

[Zha14] Jin Zhang. Empirical Bayesian estimation of location of the Cauchy distribution. Journal of Statistical Computation and Simulation, 84(6):1381–1385, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2015:SCI

[Zha15] Guoyi Zhang. Simultaneous confidence intervals for pairwise multiple comparisons in a two-way unbalanced design with unequal variances. Journal of Statistical Computation and Simulation, 85(13):2727–2735, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2016:ELM

[Zha16] Shibin Zhang. Empirical likelihood methods for discretely observed Gaussian moving averages. Journal of Statistical Computation and Simulation, 86(5):942–958, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1046071.

Zhang:2017:IAS

[Zha17] Lin Zhang. Introduction to AsiaSim/SCS AutumnSim special issue. Journal of Statistical Computation and Simulation, 87(13):2469, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2018:GPB

[ZHB18] Man Zhang, Yili Hong, and Narayanaswamy Balakrishnan. The generalized Poisson-binomial distribution and the computation of its distribution function. Journal of Statistical Computation and Simulation, 88(8):1515–1527, 2018. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zamani:2019:IFP

[ZHH19] Atefeh Zamani, Maryam Hashemi, and Hossein Haghbin. Improved functional portmanteau tests. Journal of Statistical Computation and Simulation, 89(8):1423–1436, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ziegelmann:2011:SEV

[Zie11] Flavio A. Ziegelmann. Semiparametric estimation of volatility: some models and complexity choice in the adaptive functional-coefficient class. Journal of Statistical Computation and Simulation, 81(6):707–728, 2011. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2011:CSD

[ZL11] Guangyu Zhang and Roderick Little. A comparative study of doubly robust estimators of the mean with missing data. Journal of Statistical Computation and Simulation, 81(12):2039–2058, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhao:2015:BTQ

[ZL15] Kaifeng Zhao and Heng Lian. Bayesian Tobit quantile regression with single-index models. Journal of Statistical Computation and Simulation, 85(6):1247–1263, 2015. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhao:2016:NEC

[ZL16] Kaifeng Zhao and Heng Lian. A note on the efficiency of composite quantile regression. Journal of Statistical Computation and Simulation, 86(7):1334–1341, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2019:NRM

[ZLL19] Jun Zhang, Bingqing Lin, and Gaorong Li. Nonlinear regression models with general distortion measurement errors. Journal of Statistical Computation and Simulation, 89(8):1482–1504, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2017:SMC

[ZLQ^{+}17] Lianyi Zhang, Kueiming Lo, Duzheng Qing, Weijing Wang, and Lixin Yu. Statistical model checking of stochastic component-based systems. Journal of Statistical Computation and Simulation, 87(13):2509–2525, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2016:ERC

[ZLW16] Jing Zhang, Yanyan Liu, and Yuanshan Wu. Exponential regression for censored data with outliers. Journal of Statistical Computation and Simulation, 86(3):431–442, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2018:DRR

[ZLZZ18] Junhua Zhang, Bingqing Lin, Yan Zhou, and Jun Zhang. Dimension reduction regressions with measurement errors subject to additive distortion. Journal of Statistical Computation and Simulation, 88(13):2631–2649, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zarrin:2019:TSM

[ZMKAV19] Parisa Zarrin, Mohsen Maleki, Zahra Khodadai, and Reinaldo B. Arellano-Valle. Time series models based on the unrestricted skew-normal process. Journal of Statistical Computation and Simulation, 89(1):38–51, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zuo:2013:SSC

[ZMW13] Jianying Zuo, William Q. Meeker, and Huaiqing Wu. A simulation study on the confidence interval procedures of some mean cumulative function estimators. Journal of Statistical Computation and Simulation, 83(10):1868–1889, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhu:2019:TMT

[ZNW19] Xiaojie Zhu, Hon Keung Tony Ng, and Wayne A. Woodward. Testing for monotonic trend in time series based on resampling methods. Journal of Statistical Computation and Simulation, 89(10):1899–1913, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zornig:2015:ZLR

[Zör15] Peter Zörnig. Zipf’s law for randomly generated frequencies: explicit tests for the goodness-of-fit. Journal of Statistical Computation and Simulation, 85(11):2202–2213, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2014:VFR

[ZP14] Xuemao Zhang and Sudhir Paul. Variance function in regression analysis of longitudinal data using the generalized estimating equation approach. Journal of Statistical Computation and Simulation, 84(12):2700–2709, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2016:MHL

[ZPL16] X. Zhang, S. Paul, and D. Li. Modelling heterogeneity in longitudinal binomial responses by generalized estimating equations. Journal of Statistical Computation and Simulation, 86(10):1912–1920, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zare:2014:RLL

[ZR14] Karim Zare and Abdolrahman Rasekh. Residuals and leverages in the linear mixed measurement error models. Journal of Statistical Computation and Simulation, 84(7):1427– 1443, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zahid:2015:RPO

[ZRH15] Faisal Maqbool Zahid, Shahla Ramzan, and Christian Heumann. Regularized proportional odds models. Journal of Statistical Computation and Simulation, 85(2):251–268, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2016:EEW

[ZS16a] Chunfang Zhang and Yimin Shi. Estimation of the extended Weibull parameters and acceleration factors in the stepstress accelerated life tests under an adaptive progressively hybrid censoring data. Journal of Statistical Computation and Simulation, 86(16):3303–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhong:2016:BBU

[ZS16b] Xiao Zhong and V. A. Samaranayake. Bootstrap-based unit root tests for higher order autoregressive models with GARCH(1, 1) errors. Journal of Statistical Computation and Simulation, 86(15):3025–??, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zheng:2018:IDL

[ZS18] Nan Zheng and Brajendra C. Sutradhar. Inferences in dynamic logit models in semi-parametric setup for repeated binary data. Journal of Statistical Computation and Simulation, 88(7):1295–1313, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhao:2015:WLG

[ZST15] Honghao Zhao, Lianjie Shu, and Kwok-Leung Tsui. A window-limited generalized likelihood ratio test for monitoring Poisson processes with linear drifts. Journal of Statistical Computation and Simulation, 85(15):2975–2988, 2015.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2014:E

[ZT14] Lin Zhang and Fei Tao. Editorial. Journal of Statistical Computation and Simulation, 84(6):1156–1158, 2014. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2016:HRA

[ZWCK16] Huaiye Zhang, Yuefeng Wu, Lulu Cheng, and Inyoung Kim. Hit and run ARMS: adaptive rejection Metropolis sampling with hit and run random direction. Journal of Statistical Computation and Simulation, 86(5):973–985, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic),

1563-5163. URL http://www.tandfonline.com/doi/abs/
10.1080/00949655.2015.1046074.

Zhang:2019:EBE

[ZWDM19] Ying-Ying Zhang, Ze-Yu Wang, Zheng-Min Duan, and Wen Mi. The empirical Bayes estimators of the parameter of the Poisson distribution with a conjugate gamma prior under Stein’s loss function. Journal of Statistical Computation and Simulation, 89(16):3061–3074, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zeng:2012:GVS

[ZX12] Lingmin Zeng and Jun Xie. Group variable selection for data with dependent structures. Journal of Statistical Computation and Simulation, 82(1):95–106, 2012. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhao:2014:GFT

[ZX14] Jianxin Zhao and Xingzhong Xu. Goodness-of-fit tests for location-scale families based on random distance. Journal of Statistical Computation and Simulation, 84(4):739–752, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zheng:2015:LRA

[ZY15] Ming Zheng and Wen Yu. Linear regression analysis with inequality constraints on the regression parameters via empirical likelihood. Journal of Statistical Computation and Simulation, 85(9):1782–1792, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2019:RPA

[ZY19] Qihuang Zhang and Grace Y. Yi. R package for analysis of data with mixed measurement error and misclassification in covariates: augSIMEX. Journal of Statistical Computation and Simulation, 89(12):2293–2315, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhong:2015:IAD

[ZZ15a] Wei Zhong and Liping Zhu. An iterative approach to distance correlation-based sure independence screening. Journal of Statistical Computation and Simulation, 85(11):2331– 2345, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhou:2015:CMA

[ZZ15b] Tingyou Zhou and Liping Zhu. Conditional median absolute deviation. Journal of Statistical Computation and Simulation, 85(10):2101–2114, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2017:SEM

[ZZ17] Yaowu Zhang and Liping Zhu. Semiparametric estimation of multivariate partially linear models. Journal of Statistical Computation and Simulation, 87(11):2115–2127, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Zhang:2017:BRP

[ZZXS17] Ying-Ying Zhang, Ming-Qin Zhou, Yu-Han Xie, and Wen-He Song. The Bayes rule of the parameter in (0, 1) under the power-log loss function with an application to the beta-binomial model. Journal of Statistical Computation and Simulation, 87(14):2724–2737, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.